A Bibliography of Publications about the Maple Symbolic Algebra Language

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 May 2026 Version 2.247

Title word cross-reference

(1 + 1) [Ikl94]. 1 + 1 [WF90]. \frac{1}{2} [Bre88]. 2 [BCG^{+}05, DB15, HS92]. 2(3n)!/((n + 1)!(2n + 1)!) [Zei92]. $20.95 [Ano96a]. $24.00 [Ano96c]. $29.00 [Ano96e]. $34.00 [Ano96d]. $34.50/£ [Ano95b]. $49.00 [Ano98a]. $49.95 [Ano99a, Ano99b, Ano03a, Ano04]. 6 [Chi93a]. $64.95 [Ano98b]. $69.00 [Ano96f]. $695 [Ano90]. 7 [Chi93a]. 8 [Alu91]. ^{(}R) [LS04]. ^{S} [Lav91]. C^{1} [GHC92]. K [Tar92]. D [DVL89, PFG06]. d/dx [Ano95c, Str94a]. D_{4}(q) [GP92]. e [Oka92]. E_{6} [FJ93]. E_{7} [FJ93]. F [GH94]. F E [Zir94]. G_{2}(q) [HS92]. III [Car90]. ∫_{0}^{∞} y^{′2} ≤ K(∫_{0}^{∞} |y|) max_{(0,∞)} |y′′| [BK94]. j [Alu91]. jj [GF02]. k(2, q) [KU03]. K(X) [AGR94]. λ [VMRA14]. LU [CES94]. m(λ) [BEEK94]. N [VČ88, McD97, ENS92, Zei92]. O(3) [LLLL94]. ∥ [Sie93]. PGL(2, k) [Sid94]. PGL_{2}(p) [Mal93]. PSL_{2}(p) [Mal93]. q [Déc93, Gar99b, GP92, Koo93]. Q(t) [Mal93]. R^{2} [LLLL94]. R^{n} [PG03]. S [Vol93]. ∥ [JKP98, PS95]. x^{5} + ax + b [SW94b]. z [WyW93].

-analogue [Koo93]. -blocks [HS92]. -dimensional [Ikl94, VČ88]. -function [BEEK94]. -functions [PFG06]. -invariant [Alu91]. -Matrices [Zir94]. -modular [HS92]. -product [Gar99b]. -series [Déc93, Déc93, Gar99b]. -singularities [Tar92]. -symmetries [VMRA14]. -test [GH94]. -transformation [Vol93]. -transformations [WyW93].

0 [Ano95b, CT96, Ryb98]. 0-387-94512-1 [CT96]. 0-387-94630-6 [Ryb98]. 0-8176-3724-9 [Ano95b].

1 [CT96]. 10th [Ano00a, Gar99a, Glo91, Mad92]. 125 [Con92]. 13th [Ano00a, HR92]. 14 [MP10]. 15 [Fos11]. 17 [ECTB^{+}14, MP12, Pos13]. 17-21 [GW00]. 1992 [CvGV92]. 1993 [BBST94]. 1994 [Lop94a, SW94a]. 1er [AL98].

2 [CW05b, TTLM94]. 2-D [TTLM94]. 2000 [Bod01]. 2007 [Tod07]. 2012 [KRR12]. 2019 [GK20]. 21st [Jef08]. 23rd [Gra93]. 26th [Ano95a]. 27th [HA94]. 28-30 [vH83]. 2e [Ano05]. 2ODEs [ADdM14].

3 [Ano96f]. 3-j [Lai94]. 3.0 [Bra96]. 33rd [Tod07]. 38.50 [Ano04]. 386/486 [Wat95a, Wat95b]. 3D [RLvLRM10]. 3M [CW05d]. 3Ms [CW05b, CW05c, CW05a]. 3rd [Ano00a, Ska95].

4 [B^{+}98, Bra96, CT97, RM97, Ano96e]. 48.00 [Ano96b]. 4th [Ano02a, HY14, Lum93, dN93].

5} [Tan99]. 5th [Car96, Ano97d].

6 [Her01, Ryb98, vRFS^{+}00]. 6-j [Lai94]. 6.0 [EHR^{+}96]. 60.00 [Ano97a]. 6th [Mal01, Bod01].

7 [Ano01, Cor02, MGH^{+}01, Wat01a, Wat01b]. 7.0 [Was02]. 7th [S^{+}01, SvADS07, Cha90b].

8 [Fos11, MGH^{+}02a, MGH^{+}02b, Sch03b, Wat02a, Wat02b, Wat02c]. ’83 [vH83]. ’86 [Cha86]. ’88 [Gia89]. ’89 [Gon89]. 895 [BC09]. 8th [Cha91].

9 [Ano95b, Map03a, Map03b, MGH^{+}03a, MGH^{+}03b, Was04]. 9-j [Lai94]. 9.5 [Map04]. ’90 [CC91, WN90, DG99]. 90g [Con92]. ’91 [Wat91]. ’92 [Fit93, dN93]. 926 [Tho13]. ’93 [Bro93, IEE93b]. ’94 [ACM94, SW94a]. ’95 [Lev95]. ’96 [Lak96]. ’97 [KH97]. 98 [Glo98].

= [Gom90].

Abell [Mea96, Cor96a, Wri96]. Aberration [Atz94]. abstract [KSS00, KSS07]. accelerated [FIT03]. accompany [CcMMK96, Kir95]. According [EZ23]. accuracy [Shi05a, Shi05b]. Accurate [ABF14]. ACM [SW94a, ACM82, Gon89]. ACM-SIGSAM [Gon89]. Acoustics [IEE93a]. Active [Ano00b]. Activeworlds.com [Ano00b]. activity [VAS06]. adapted [Lav91]. Adaptive [CR94b, Dou93, DP95]. Addendum [Con92]. addition [RW92]. addition/contraction [RW92]. ADM [CZL11]. ADM-Padé [CZL11]. admitting [HK89]. ADrien [VM99]. Advanced [EM94, KN94, MGH^{+}02a, MGH^{+}03a, MGH^{+}05a, MGH^{+}07a, MGH^{+}08a, MGH^{+}09a, MGH^{+}10a, Ric02, EM07]. Advances [Fat90]. Adventures [Res92]. Affordably [Put94]. AG [Car97a, Car97b]. again [Fos01]. AI [RSL^{+}26]. Aid [CP93]. Aide [Lev97, JLP99]. Aide-mémoire [Lev97]. Aided [LEGS93, DJ89, KK91, KD90]. Aided-Design [LEGS93]. aiding [Sil04]. Airy [CJR92a, CJR92b]. AISMC [CC92]. AISMC-1 [CC92]. Akopia [Ano00b]. Albuquerque [ACM82]. Algcurves [DP11]. Algebra [ADG^{+}01, AV02, Ano97d, Ano03b, Asl96, Aue90, Aue91, Bag95, BEJ95, BMW94, BB92b, C^{+}84a, Con92, DST88, DPV14, Edw95, E^{+}92, EJLS96, FS95, GMV00, GCL92, HWH91, Hog07, HH95, Joh93, Kei01, Kne94, M^{+}94, Mal06, Mur94, Par97a, Pro00, Sch99b, Sch01a, Sch03a, Sco5 , Sid94, SVV95, SCL95, Vei97, Abł96, Akr89, Ano99a, Ano99b, BSS03, Bay96, Bee05, Ben98, Ber93, BS99, Bre88, Bro97a, BCLA83, CJ97, CGGG83, CGG^{+}86, CvGV92, CLO92, CSS94, CLO07, CC98, DST93, DG98, Die92, Ell97a, EM07, FP89, FP95, FOW87, GK91, GKW03, Her01, HPS05, Kla10, KSS00, KSS07, Koe92b, KSK96, LHJ92, LL96, LM91a, Leo86, Mat89, Mat90, MBC92, Pav85, PF95, SMB03, SF10, Sny95, Ste94, SH10, Sza02, TH94, Tsa93, LM91b, Vul02, Wan06]. algebra [Wes99, YSH96, vH83, Car96]. Algebraic [ACM94, Ada94, Boy93, BG90, Bro93, Bro07, CM91, Cha86, DST88, DP11, Doo99, Gia89, Glo98, Gon89, Gut04, Har98a, Har00, JLL91, Kal93, Lai94, Lak96, LEGS93, M^{+}94, Mit00, MSW00, MCG13, SE11, Sch99a, SMW00a, SMW00c, SMW00b, SMW01, Wan92, WN90, Wat91, vdHvH12, AV11, AY03, BS03, Ben89, BDE^{+}13, Bro91b, BCLA83, CG95, CDF92, Cha90a, ČPV90, CM01, CLO92, CLO07, DST93, DM05, Fat90, Kau93, KT91, Lev95, Ose96, PW95, YA07b, van94b, van94c]. algebras [ALP96, AF05, Gec94, IS00]. algèbre [LBP98, PR98]. Algorithm [Gru93, Kal93, UW95, AV11, AGR95, AF92, Bar93, Bar95, Ber94, CS89, CG86, DGS08, Die92, GG12, Hen89, Koo93, LM88, LTFN01, McG84, NS05, PG06, PG08, PG10, PGG13, Shi93, jZzFbL06, van94b, van95, BC09, Sch01b, Tho13]. Algorithmen [tHB91]. Algorithmic [Koe92a, LRB95, MN93]. Algorithmischen [Ger92]. Algorithms [BHC95, Bia00, CM11, DST88, DP11, GCL92, Mit00, Mon94, Mul97, Mul06, Mul16, WF90, AFK^{+}94, CLO92, CLO07, DST93, FF88, GR92, NS05, Sei92, SSKT91, Sie93, tHB91]. Aligned [Car90]. Alltag [EHR^{+}96]. Alternative [ES10, Gar95, Ste08, Ste13]. America [Gra93]. American [Gon88]. Amsterdam [CvGV92]. analogue [Koo93]. Analysis [CS93, GKLM91, Kol00, Nes93, Rhe93, Van94a, Ano96k, Bel95, Bet98, BG90, Bra93, CW92, Cor94, Dou93, DP95, EG97, FF11, GW89, GW94, GS12b, Hed92, Hen94, Hen89, IEE89, IOP21, JKP96, KK91, MBC92, SSKT91, VAS06, Vol93, WH92, YB00a, Ano98b]. Analytic [BS99, BC94b, PN98, AH92]. Analytical [MG94, Ano96k, GS12b]. Analytically [Zir94, Bro95]. Analyzing [Mur94, SMW00a, Sch00a]. Ananda [Ano99a, Ano99b]. André [Ano97a, Ano04]. Andrew [CcMMK96]. Andrews [Gar99a]. angles [GW20]. angular [Fri97, FVGF98, FIBT01, FIT03, Fri09, GFF01, GF02, GSF05, IFF01, PFG06]. animation [Put03]. Ann [Lee93]. Annotated [Kir95]. Annual [ACM82, Cha90b, Gra93, Gra94, Lum94, Jef08, Cha91, Mad92]. Ansatz [JWK^{+}21]. Anton [AH96]. Anwendung [EJLS94, Ben98, Ell97b]. Any [ARL20]. Anymore [Sim92]. anyone [BWB97]. Appell [AAFMJ01]. Applets [And04]. Application [Ano00b, Ano01, CSS94, IA94a, JKP98, Ell97b, GK91, GQ93, IOP21, KU03, Li00a, LL04b, Lop94a, PS95, SCMS96]. Applications [Ada94, Aue91, Corxx, KSS00, KSS07, Kol00, Lee93, Ler95, Pav85, PS97, Rav94, SZ94b, Sny95, Wan94, Akr89, Ano94a, Bet01, BGR91, CC91, GKW03, Gro89, Har14, Har15, JKP96, JS00, LLLL94, Lyn01, Lyn09, McL94b, PG03, SF90, SMFC92, SSKT91, Tay98b, Tsa93, Aue90, Tod07, Ano95b]. Applied [And97, GW89, GW94, Glo91, LW88, Lop93, Pin94, Rao02, TA96, WA96, Gon88, GH94, Mat90]. Apprenticeship [Lop93]. Approach [Ada94, BM94a, CDK94, FKM95, HT98, Lop94b, MSW00, Pet00b, Pet00c, Red95, Red96b, ADdM14, BF02, BF09, Dav01, Gre99, KSK96, MBC92, PC02, RLHAL11, SLC07, Shi09, Sid94, Ano03a]. Approaches [MG94, PW95]. approach^{T M} [Ano96f]. Approx [CT96]. Approximants [ČPRV87, BG90]. Approximate [LSYL11, AV11, CZL11, ČPV90]. Approximating [Gro90]. Approximation [Qua18, BC94b]. approximations [Ano95c, CB02, Oka92, Str94a]. April [Ano03b, Fit93, IEE93a]. Arbeitsbuch [Wer96]. Arbitrary [Atz94, Sei92, Sei93]. Arbor [Lee93]. Architectures [Ano00a]. Arguments [Tho13, CJR92a, CJR92b]. arising [Con92]. Aristoteles [vRFS^{+}00]. Arithmetic [CC93, Corxx, GK98, LMRS08b, LMRS11, PM07, SE22]. arithmétique [PR98]. ARMA [JKP03]. ARMA-representations [JKP03]. arrangements [AF92]. Arrays [WG83]. article [Bar92]. Artificial [CC92, HR92]. Arts [Put94]. Ascend [Ano01]. ASCM [GW00]. Asian [GW00]. aspects [JWK^{+}21]. assertion [Rei96]. assessments [Bia00]. assisted [Rei96, VAS06]. assume [AB05b]. astro [DJ89]. astro-geophysics [DJ89]. Asymptotic [GH94, Gru93, Rhe93, San96, BB92a, BS05b, Die92, Sal89, Sal91]. Atipa [Ano00b]. atom [VČ88]. Atomic [ÖSZ94, BS99, CSS94]. Atoms [EHR^{+}96, McD97, Pie93]. attractor [Rei96]. August [Bod01, Bro07, CC92, Glo98, HY14, Kol00, Lop94a, S^{+}01, SSKT91, WN90, dN93]. Austria [Ano02a, Jef08]. Austrian [Ano00a, Ano02a]. Austrian-Hungarian [Ano00a, Ano02a]. Automata [Car97a, Car97b]. Automated [ADG^{+}01, Col05, PW95]. Automatic [ARL20, Bac96, CR94a, Gro89, Hed92, JL94, Nes93, WW91, WyW93, CMC92, Sal89, Sal91, VM99]. Automatically [BR14, Cam06]. Automating [Pet94]. Automaton [Car98, Car00]. automorphism [BSS03]. aux [KL96]. averaged [BB92a]. AVS [LW94]. AXIOM [Ben98, Ben98, JS92, Ken01, Grä96].

B [Ano96d, Hea96]. Bäcklund [MWCY14]. balanced [OP95]. Balaton [Ano00a]. Balatonfured [Ano00a]. Band [vRFS^{+}00]. Banded [CES94]. bang [SZ03]. bang-bang [SZ03]. Based [Bac96, Kun02, Sch99b, SKB01a, Ano95d, IA93, LL04b, VMRA14, WEM98]. Bases [Arn95, Gar95, Cor96b, CG86, IA93, IA94a, IA94b, OS95, Shi93, Stu89]. basic [Par97b]. Basile [LM91a, LM91b]. Basis [BCG^{+}05, GHC92, van94b]. Bath [Fit93]. BC [Doo99]. BDF [GW20]. Beach [HA94, ÖSZ94]. Beauty [vRFS^{+}00]. Begriff [EHR^{+}96]. Begriff/Vogel [EHR^{+}96]. Behind [TT94]. Beijing [SvADS07]. Beispiele [JKM04, Vet95]. benchmarks [MP16]. benzene [ČPV90, Vin92]. Bergman [JT91]. Berkeley [Wan92]. Berlekamp [Sch01b]. Berlin [Ano96b, CT96, Ryb98]. Bernoulli [Wil98]. Bessel [FS92, GQ93]. best [BK94]. Bestimmter [Sch93a]. Bestimmung [tHB91]. beta [PG06]. Bethe [JWK^{+}21]. Betounes [Ano98b]. Between [Pie93, BHC95, Nay02]. Bézier [PQ23]. Bézier-Spline [PQ23]. Bias [GU93, SC09]. Bifurcation [GKLM91, SSKT91, EG97, Nag93, PNG91, Xu95]. bifurcations [BC94a, QBH06]. Big [Sch01b]. BIGEBRA [AF05]. bilinear [MWCY14, dYyR13, jZzFbL06, jZzFbL10]. binary [ACS91, Lev92]. Biochemistry [GDS94a, GDS94b]. BioconX [Ano01]. biological [YC07]. biology [SH09, YSH96]. Biometric [Ano01]. Birkhäuser [Ano95b]. Birmingham [BBST94]. bis [vRFS^{+}00]. bivariate [BLPR13, DET07]. Black [CM22, CM24, DK95]. Block [Liu19, Bar93]. blocks [HS92]. Body [Pie93, Ron14]. Boltzmann [SD10]. bon [DG97]. Bonn [Wat91]. Book [Ano95b, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano97b, Ano97c, Ano98a, Ano98b, Ano99a, Ano99b, Ano03a, Ano04, Bag95, CT96, Gan03, Gre97, Mea96, Ryb98, Ste96, Sto03, SMC97, Tan99, Wri96, Gar02, RC96, RC97, Ano97c]. boolean [BDE^{+}13]. Born [BS03, BS99]. Boson [Mur07b]. Bosons [Mur07a]. Boston [Ano95b, MCG13]. Boundary [EP96b, EP04a, Hoh01, Wic95, Art98, KRR10, KRR12]. bounds [Lah92, Vin91]. box [PG03, PGG13]. Boxes [CM22, CM24, DK95]. brackets [Ano96j, Hol96]. Branch [ECTB^{+}14, Har97]. Braselton [Cor96a, Mea96, Wri96]. Brazil [M^{+}94]. Brazilian [M^{+}94]. breaking [LLLL94]. Brian [Ano96a]. Bringing [LMRS11, LMRS08b]. Brown [McG84]. Bubble [BBST94, Sha94]. Buchberger [CG86, Stu89]. Buchbesprechungen [EHR^{+}96]. Bücher [vRFS^{+}00]. Built [TMH94]. Bulgaria [Kol00, Tod07]. BVP [Kwo90].

C [Ano96b, Gre97, CM22, LS04, Phi04, Ste94, SH10, WEM98, Zac96, Ano97b, Ryb98]. CA [SE11, Wou86]. Cabri [Kla10]. CACSD [LM91a, LM91b]. Calabi [Rab96]. CalcLabs [BBRM95]. Calcul [GSZ95, RGM95, ZSG5 , KL96, Lem97, RM97]. Calculating [Zir94]. Calculation [GU93, OS95, Vul03, WY93, ABF14, JL94, SD10, Vin91, dJ91]. Calculations [Dav88, KMG96, Bay96, CV01, CD01, CFFM97, CSS94, CM87, Van88, Xu95]. Calculo [BLM97]. calculs [MNP97]. Calculus [AH96, Ano95f, AH92, AB94d, AKEJ94, BF91, BF96, BDS96, BD93, C^{+}87, CH94, Dec94, Dev93b, EP94, EL89, E^{+}92, EJLS96, FV94, Fat92, Fat96, Ful94, GC^{+}88, GMM^{+}89, HC95, HC96, Har94, Har92, HEBS93, HL94, Isr96, Jud96, Kir95, KR03, Lop94b, McL93, Put94, Roy93, SSTW94, Ano95d, CJ97, Cor94, Ell97a, Gre99, HL95, Ken01, LHJ92, Par97b, Str99, SCMS96]. California [IEE89, Wan92]. Calques3D [RLvLRM10]. Campusline [PM5 ]. Can [Ful94, Mon94, Ste08, DPV14]. Canada [Doo99, Lev95, Bro07]. canonical [DM86]. canonicalization [Liu19]. Cantabria [Gut04]. capable [Bro95]. Carlo [YC07]. Carmichael [Arn95]. Carolina [Ano03b]. Cartesian [IA94c]. CAS [CA07, Dur19, RLvLRM10]. CASC [GMV00]. CASC-2000 [GMV00]. Case [Mav93, BF02, BF09, CA07, PS95, Sal93]. Cat [Gra94]. Catalan [Fee90]. catalytic [Bay05]. caustics [IA94c]. CD [Ano98b, Ano99a, Ano99b, Ano03a]. CD-ROM [Ano98b, Ano99a, Ano99b, Ano03a]. Celestial [Gro96]. Cell [DS94b]. Center [AHK93, Ano02b, IEE93a, IEE93b, QBH06]. Central [Ska95, Bre88]. certain [JM98, Koe93, Sid94]. certifying [ME18]. Chains [Zim95, CDC^{+}02, FF88]. Challenge [Mon93]. Champaign [Tho94]. Chandler [Ano97c]. change [Str99]. Chaos [EHR^{+}96, SSKT91]. character [Gec94]. characteristic [Wan95b]. characteristics [TP94]. Characterization [Car98, SW94b, Car00]. characters [GP92, HS92]. Charafi [Ano03a]. charge [Bre88]. Chebyshev [Boy93]. Checking [BS02a]. Chemical [TA94, TA95, Whi10]. Chemistry [Ano88, CVW93, GDS94a, GDS94b]. chercheur [CT95, CT97]. Chern [Wu92]. Chevalley [GP92, HS92]. Chiang [GW00]. Chicago [Ano02b, Tan89]. China [SvADS07]. cientifico [RGM95]. ciphers [TT00]. circuits [Bra93, KSK96]. Class [Dev93a, Ron14, dv96, GQ93, QBH06]. classes [AL98, FJ93, Pet94, Ram98, Wu92]. Classical [Gre95, CT96]. Classification [Chi93a, HK89, LM88, Chi93a]. Classify [BR14]. classroom [CGG^{+}86]. Clean [EZ23]. Clemson [Ano03b]. Clifford [Abł96, ALP96, AF05]. closed [BCEW12, MNEA95]. closed-form [MNEA95]. Cluster [Har98b, SMW00a, SMW00c, SKB01a, SMW01, ČPV90, Vin92]. Clusters [LH93, ČPV90]. coagulation [Ikl94]. coagulation-fragmentation [Ikl94]. Coast [Ano03b]. Code [Bac96, Gom90, GS98, PG03, PG06, PG08, PG10, PGG13, SD10]. codes [ACS91, JS00]. Coding [Joy00, TW06, CC91]. Coefficients [Atz94, SVV95, AV11, AFK^{+}94, Cam06, FIBT01, GSF05]. Coffman [Kwo90]. College [Ano95d, Lum94, BWB97, Bro97a, Put94]. Collegiate [Lum94, Lum93]. collision [SD10]. Colloquium [Ano96g, Kol00]. Colombia [LW88]. Colorado [Har94]. column [CVG94]. combinations [BDE^{+}13]. Combinatorial [Dev93a, FS95, Hen89, LL95, JWK^{+}21]. Combining [HT98, SW98]. Combustion [GKLM91, EG97, GKLM92]. Coming [CW05e]. Common [DK95, AV11, ENS92]. commutative [CLO92, CLO07]. Compact [C^{+}84a, CGGG83, PG06, PG08]. CompactPCI [Ano00b]. Companion [DHM95a, Bar93]. comparative [CVG94]. compared [BFS93]. Comparison [AFK^{+}94, Zot07]. compatibles [Ano95f]. Competition [Gra94]. competitor [Ano90]. Compiled [Dub00, LMRS08a]. Compiled-Interpreted [LMRS08a]. Compiler [DG99]. complete [Boy15, HJOP03, WF90]. completely [FF88]. Complex [Asl96, Bil94, CCM91, CJR92a, CJR92b, Les94]. Complexity [vRFS^{+}00, DET07, LTFN01]. complicated [CTE95, ECT93]. Component [KMG96]. Components [Ano01, AFK^{+}94]. composition [Lah92]. Comprehensive [NWF97, NW96]. comprise [CVG94]. Computable [Ngu93]. Computation [ACM94, Ano95b, Ano96g, Ano01, BSS03, Bee17, Bro93, Bro07, Cha86, DST88, Doo99, DB15, Fee90, Gia89, Glo98, Gon89, Jef08, JS92, JKP98, KH97, KMG96, Lai94, Lak96, Lee93, LMS07, LMRS08a, McD97, MCG13, Ngu93, Ngu94, OP95, Pie93, SE11, Var95, Wan92, Was02, WN90, Wat91, vdHvH12, AAFMJ01, AFK^{+}94, Ber94, BEEK94, BCLA83, CGG83a, Cha90a, CV88, DST93, Fit93, Fit89, Fiu95, GK91, Gro90, GDS94b, HT85, Jon96, KRR12, KZ94, Lev95, LS96, PS95, SMF^{+}90, SMM^{+}90, SF90, Sie93, Tay98a, VAS06, Van94a, VČ88, VC89, Vin91, WW91, WyW93, qXbL04, de 93, de 95, Gut04]. Computational [AHK93, Ano88, Ano97b, DS94a, Faj19, ÖSZ94, PQ23, Qua18, Ryb98, Top99, Whi10, Ano98b, Bet98, CLO92, CLO07, Mic90, Sid94, ZJEP95, Zac96, SvADS07]. Computations [CP93, GDS94a, WWK92, Abł96, ALP96, Bog16, Fri97, FIT03, Ioa94, Wan90]. compute [Boy15, Die92, Shi93]. Computer [ADG^{+}01, AV02, AHK93, Ano95a, Ano97d, Ano03b, Asl96, BMW94, BFS93, BCLA83, BD93, C^{+}84a, CGG^{+}86, DST88, DST93, DJ89, DPV14, EM07, FS95, FOW87, FF88, GMV00, GW00, GCL92, GKW03, HWH91, Har98b, HSW94, HSW98, HH95, IA93, Kol00, KN94, KD90, LEGS93, M^{+}94, Mur94, Pro00, Rie94, Sch99b, Sch01a, Sch03a, Sco5 , Ska95, SVV95, TH94, Vei97, Wes99, vH83, Akr89, Ben98, Ber93, BS99, Car96, CGGG83, CvGV92, CSS94, Die92, FP89, FP95, KK91, Koe92b, KSK96, LHJ92, LM91a, Mat89, Mat90, MBC92, Pav85, PF95, SMB03, SH98, Sny95, Ste94, SH10, Tsa93, LM91b, Vul02, Wan06, YSH96, CvGV92, Bag95]. Computer-aided [DJ89, KD90, KK91]. Computer-Algebra [Sch03a, FOW87]. Computer-based [IA93]. Computeralgebra [Ben98, Kam93]. Computeralgebra-Systemen [Ben98]. Computers [CHM94, Dev92, Dev93c, Dev94a, Dev94b, Gre93, Tan89, Wat95a, Wat95b]. Computing [AGR94, ALL04, BR09, BCEW12, BP92, BKR^{+}91, Car97a, CDL^{+}11, Cor95b, DP11, DK95, EP96a, EP96b, EP96c, EP04a, GH93, GH95, GH97, GMV00, Glo91, Gru93, Har97, HR92, Kal93, LP12, Le01, LTFN01, SMW00a, SMW00c, SKB01a, SMW01, Ste96, Stu89, Top99, Wri01, van94c, Bar93, BS03, Ben89, BR02, BS11, CC92, CDF92, CM11, Dom03, EP03, GH04, GLHR19, GHL^{+}00, GKaM01, Har14, Har15, JLL91, LP05, LMJ93, M^{+}94, Mal01, Möl07, QBH06, Sei92, Vân94d, Wou86, jZzFbL10, dN93, van94b, van95]. con [BLM97, RGM95, SV96, SCL95]. Concepts [Ano00a]. Concise [Str94b]. Condensed [GKLM91, ÖSZ94, GKLM92]. Condensed-Phase [GKLM91, GKLM92]. conditional [Bia00]. conditions [Cam06]. Cone [Har98a, Har00]. Conference [ACM82, AHK93, Ano02b, Bod01, CC92, GK20, Glo91, GH84, Gra94, HY14, IEE93a, IEE93b, Lum94, Mal01, Nes93, S^{+}01, SvADS07, SW94a, Tan89, TRC94, Tod07, vH83, Gra93, HA94, JLL91, Lum93, SSKT91, Ska95, Top99, dN93]. configurations [GSF05]. Conformal [JT91]. conformally [Wil89]. conformance [Bee04]. Congress [HR92, IEE93b]. conjecture [Col05, Yan94, Zei92]. Conjectures [Col02, Koe91]. conjugacy [FJ93]. conjugate [BV93]. conjugates [BV97]. conley [BR09]. Connected [Pet01b]. Connecting [RLvLRM10]. connection [BR09, LW94]. connections [BCEW12]. conquer [LTFN01]. conservation [Bay05, Fit89, MWCY14, YL06, YB00b]. Considered [Pro00]. consistent [BCG^{+}05]. CONSLAW [YL06]. constant [BK94, Fee90]. Constants [RSL^{+}26]. constrained [CS89]. constraint [AB01, AB05b]. constraints [AY03, PG03, PGG13, YA07b]. construct [dYyR13, YL06]. ConstructibleSetTools [CML^{+}08]. Constructing [Arn95, YZ07]. construction [Fat90, Kau93, Pet94]. constructions [Rya99]. Constructive [Mit00]. Content [Man02]. context [ACGL04, WVG90]. contextual [AB01, AB05b]. continued [BC09]. continuous [Sei92, Sei93]. contraction [RW92]. Control [Ano96g, SZ94b, JLL91, Van94a, dJ91, de 93, de 95]. controlled [Lie94]. controller [LS96]. controllers [SRRB91]. Controversy [Wic95]. Convention [IEE93a]. Convergence [Rhe93, FS92]. conversion [CMC92]. converting [Gro89]. convex [AF92]. Coombes [Ano96a]. Cooperative [Pet00b, Pet00c]. Coordinates [DB15, IA94c]. Copenhagen [KT91]. Coq [DM05]. Corner [Cor96b]. Corporation [Ano00b]. corrections [SC09]. Correlated [Dav88, Dou93]. Correlations [vRFS^{+}00]. corrigés [LM98a, LM98b]. Cosmetic [Was04]. Counting [EZ23]. Coupled [Har98b, MG94, ČPV90, FH04, pLbL03, Vin92]. Coupled-Cluster [Har98b]. coupling [Fri97, FVGF98, FIBT01, FIT03, Fri09, GFF01, GF02, GSF05, IFF01, PFG06, YZ07]. cours [AL98, LM98a, LM98b, MNP97]. Course [CM94b, Kne94, AR06, LP05, Run97, ZJEP95]. courses [EM94, TH94]. Coxeter [BB92b]. Create [ES10, Ste08]. creating [Ste13]. criteria [GH94]. cross [BS03]. Crow [CVG94]. Cryptography [G´ 13, May09, TW06, Cos00, Joy00]. crystal [JM98, KOW93]. CSG [TTL93, TTLM94]. CSV [Ano00b]. cubic [CHF94, ES92, PGG13]. cubics [Alu91]. Currents [ČPRV87]. Curriculum [FGH94, Gre93]. Curvature [PQ23, AFK^{+}94]. Curve [MSW00, PQ23, Bro91b, Vol93]. Curved [Mur07b, Mur07a, PMS20, Vul03]. Curves [DP11, SMW00a, SMW00c, Ano98a, Bog16, Con92, ES92, JS00, KRTZ20, KK97, LSYL11, Rov00a, Rov00b, Sei92, Sei93, YJ99, van94c]. cuts [ECTB^{+}14]. Cyclades [Ano00b]. cycle [AL98]. cycles [Bay05]. cyclic [Vin91]. Cyganowski [Gan03, Sto03].

Cylindrical [BDE^{+}13]. Czech [Ska95]. Czechoslovakia [dN93].

D [BCG^{+}05, DB15, TTLM94]. D-Eigenvalue [BCG^{+}05]. DAEs [CM94a]. D’Alembertian [AZ96]. dans [Gom90]. Dapsys [Ano02a]. DAPSYS’2000 [Ano00a]. Darboux [ADdM14]. Darren [Ano96e, Ano96f, Ano97c]. Data [Ano95f, Fri97, MP14, SW94a, Dou93, FP89, MP12]. Dataflow [Sch00b, Sch01a]. Date [ARL20, Fri09]. David [Ano98b]. Day [Ano03b]. DC [Gra93]. Dealing [DM05]. Deborah [CcMMK96]. Dec [CvGV92]. Decades [Wic95]. December [GW00]. decide [SW98]. decimal [OS95]. Decision [SZ94b, RLHAL11]. Decomposition [BS93, LP94, AGR95, GR92, JM98]. decompositions [BDE^{+}13, CM11, GLHR19]. Deeba [Ano99a, Ano99b]. defects [KOW93]. defined [LSYL11]. definite [AZ91]. deformations [Sha94]. degenerations [Wu92]. degré [Chi93a]. degree [Sei92, Sei93]. Degrees [Zir94, Chi93a]. delay [Pec91]. delay-differential [Pec91]. delayed [BC94a]. Delft [Kne94]. delivers [Was04]. Delivery [Ano01]. Demonstration [Kun02]. demonstrations [KRR12]. d’Emploi [GSZ95]. Denmark [KT91]. Denotational [Pin94]. Deployment [Man02]. Derivation [Bay05, GHC92, IA94c, YB00b, DJ89, Hed92, Vin92]. Derivative [Ano05, Pri00]. derivatives [BG11, CT94a, Dom03]. DERIVE [Ben98, SC09, Ben98]. Descent [BB92b]. Design [Ano01, Bil94, C^{+}84a, CFG^{+}84, C^{+}84b, EM94, Fit93, LEGS93, MP14, PPD05, Bra93, CGGG83, Fat90, Hel10, Ioa94, LHJ92, Leo86, LS96, Run97, Van94a]. designs [Kla10]. DESIR [Pfl97]. DESIR-II [Pfl97]. DESOLVII [VJC12]. d’espèces [Chi93b]. detect [WF90]. detecting [IS00]. detection [FF88]. determination [FFB99, LRB95]. Determining [tHB91, Rab96]. Developers [Sch93b]. Developing [Sch99a]. Development [Ano01, GWV93, Koe92a]. Developmental [WGV90, GWV93, WVG90]. Devlin [CHM94]. Devmol [ALL04]. DGS [RLvLRM10]. diagonal [Bar93]. diagonalization [BEEK94]. Diagrammatic [ARL20]. Diamantina [Sid94]. Dick [Ano95f]. Dick/Patton [Ano95f]. dictionary [BB90, BWB97]. didn’t [BWB97]. difference [AZ96, Ano95c, Str94a]. differentiable [LSYL11]. Differential [AB94a, AB00, Ban01, Bet01, Bug95, Con96, Con98, C^{+}96, Coo97, Cor96a, Dav01, DiB95, EP96a, EP96b, EP96c, EP03, EP04a, E^{+}92, EJLS96, Gan03, GH03, Hea96, Koe94, PBA06, Sch5 , Sta07, ST99, Sto03, Xie10, dv96, AZ96, AZ91, Ano98b, Art98, ADdM12, ADdM14, AH97, Bar01, Bar93, BF02, BF09, B^{+}98, BC94a, Bet98, BC94b, Bro91b, CG95, CJ97, CDF92, Cyg96, CGK01, CKO02, Die92, EP04b, Ell97a, FH04, FF11, GS12b, GKaM01, HK89, KOC02, Koe93, KRR10, LS04, Mat89, Pec91, QBH06, Rod92, Rya99, SH98, ST04, VMRA14, VJC12, qXbL04, ZF91, Ano96a]. Differentialgeometrie [Rec98]. Differentialrechnung [Wer96]. Differentiation [MN93, Wan94, JL94, VM99]. differentsial’ [ZF91]. diffusion [Tal94]. Digest [Ano96g]. DIMACS [DS94a]. dimension [Bry91, GHL^{+}00, WF90]. dimensional [BSS03, CM91, Ikl94, Li00b, Rab96, VČ88]. dioid [GK91]. Dirac [Vul02, Vul03]. Direct [IA92]. directed [Vân94d]. DISCO [Fit93]. Discovering [Har92, HL95, KK97, SKH99, Ano98a]. discovery [Sil04]. Discrete [Ano01, Har98a, DS94a, GK91, Har00, Ikl94, JKP03, Kau93, LSYL11, Ros97, ZF91]. Discrete-Event [Ano01]. Discrete-group [ZF91]. discussion [McG84]. diskette [Ano96f, Ano97b, Ano97c, Ano98a]. Diskretno [ZF91]. Diskretno-gruppovye [ZF91]. Diskussion [Ger92]. Dispersions [LH93]. dispersive [Li00b]. dissipative [Li00b]. dissipative-dispersive [Li00b]. distance [ACS91, Liu19]. Distributed [Ano00a, Ano02a, CDK94, Pet00b, Pet00c, Sch99a, Sch00a, Sch01a, Sch03a, Top99, Ber97a, BS02a, BS02b, BS02c, BS05a, Pet00a, Pet01b, Sch98a, Sch00a, Sch00b, SKB01b, Sch01b, Sch03a, SMB03]. distributions [SC09]. Disturbing [MH93]. Diverging [DVL89]. divide [LTFN01]. divide-and-conquer [LTFN01]. division [MP10]. Divisors [DK95]. DM [Ano96b, Ano97a, Ano97b, Ano97c, Ano98a, Ano98b, Ano99a, Ano99b, CT96]. documents [YF04]. Dodson [Ano96b]. Domain [Bac96, LP94]. domains [Jef93]. dot [TT94]. double [Nag93]. drive [MNEA95]. Driven [BFKS93]. Dublin [HR92]. durch [EHR^{+}96]. dyad [CMC92]. Dynamic [Ano05, Zie07, KK91, KZ94, LS96, Pri00]. Dynamical [Lyn01, Lyn09]. dynamics [BBST94, Kim97, SIE98, dJ91].

e-mails [Bra96]. early [KR03]. East [Ano03b]. Easy [Ano00b]. econometricians [SKH99]. Economic [Ngu94, Zie07]. economics [Tod07]. economy [PG10]. Econophysics [vRFS^{+}00]. Edgar [Ano97c]. edited [Ano95b]. Edition [Ano04, BBRM95, DHM95b, Kir95, KN94, Wat95f, Wat95j, Wat95l, Wat95h, Wat95k, Wat95d, Wat95g, Wat95i, Wat95e, Wat95a, Wat95b, Wat95c, Ano97a, Bro97b, Bro97c, CcMMK96, Ell96]. Editor [Cor96b]. Educating [Gra94]. Education [Ano95a, FLC93, Gra94, EM94, Gra93]. effect [SC94]. Efficient [Sal93, GS98, GKW94]. eigenproblems [Cor96b]. Eigenschaften [tHB91]. Eigenvalue [BCG^{+}05]. Eighth [Kol00, MH93]. Einführung [EHR^{+}96, HHR96a]. Einiger [tHB91]. Einstein [DVL89]. Elasticity [Hoh01, IA93]. Electrochemical [ČPRV87]. Électrocinétique [LM98a]. Electrodynamics [Har98a, Har00]. Electronic [Ano95f]. electronics [Bra93]. Électronique [LM98b]. Elektrodynamik [vRFS^{+}00]. element [QBH06]. Elementare [Rec98]. Elementarer [Ger92]. Elementary [EP04b, Gan03, Mul97, Mul06, Mul16, Pav94, Sto03, CKO02, FP95, KOW93, KOC02, PF95]. Elements [Akr89, MH93, Ano03a, GFF01, PC02, SD10, SC94]. Elias [Ano99a, Ano99b]. Elimination [Wan95a, LL04b]. ELISE [Die92]. Elliptic [Gar95, SVV95, Boy15, Con92, CA07, YJ99]. else [BWB97]. Embedded [Ano01]. Emden [Kwo90]. Enable [May09]. Energy [LH93, TRC94, KOW93, VČ88, Vin91]. Energy-Sources [TRC94]. Engineer [MB98]. Engineering [AB02, Bel95, FLC93, Glo91, Gra93, KN94, RW95, Rob96, Roe94, TA94, TA95, TMH94, EM94, Har14, Har15, HSW94, Mat90, SF90, SMFC92, TRC94, Tod07, TH94, Whi10, ZJEP95, Ben98]. Engineering-Physical [AB02]. Engineers [Com96, EM97a, EM97c, Gra94, Rao02, BWB97, EM00, MB98, Ros96, Sch98b, Sch99c, SKH99, TA96, Xie10, Gre97]. engines [VAS06]. England [ACM94, vH83]. enhance [TH94]. Enigma [Joy00]. Enlivening [FGH94]. Enns [Gre97]. enseignement [Ziz93]. entry [ZJEP95]. entry-level [ZJEP95]. enumeration [AF92]. enumerations [Sad96]. Enumerative [KT91]. Environment [Ano01, Ano05, Dev93a, LMRS08a, Nay02, Lav91, LM91a, Pri00, LM91b]. environmental [Sco01]. Environments [Sch03a, SMB03, Wou86]. enzymatic [YB00a]. enzyme [CR94a]. equation [BF02, BC94a, Bro91b, CTE95, GKaM01, IA94c, LS04, LM88, LW94, Roa94, SW90, Vul03, YZ07, jZzFbL10]. Equations [AB94a, Bug95, CES94, Cor96a, DiB95, DS94b, EP96a, EP96b, EP96c, EP04a, E^{+}92, EJLS96, Gan03, GH03, Har98b, Izu14, MG94, PBA06, Sch5 , Sta07, ST99, Sto03, dv96, AB00, AZ96, AZ91, ABF14, Ano96a, Ano98b, Art98, ADdM12, AH97, Ban01, Bar93, BPS12, BF09, B^{+}98, Bet98, Bet01, BC94b, Bro95, CG95, CJ97, CDF92, ČPV90, Con96, Con98, C^{+}96, Coo97, Cyg96, CGK01, CKO02, CMC92, Dav01, Die92, DJ89, EP03, EP04b, ECT93, Ell97a, FH04, FF11, FOW87, GS12b, Hea96, HK89, IA92, IA93, Kim97, KBG01, KOC02, Koe93, Koe94, bLpL02, LL04a, Lie94, pLbL03, Mat89, MWCY14, Ose96, Pec91, QBH06, Rod92, Rya99, Sid94, SC94, SH98, ST04, VMRA14, Vin92, VJC12, XYXZ16, Xie10, qXbL04, dYyR13, YL06, YC92]. equations [ZF91, jZzFbL10]. Equator [Ano01]. equilibria [BC94a]. Equilibrium [Ngu93, CT94b, SIE98]. Equivalent [BMW94]. Erice [BGR91]. Erratum [Ano95c]. error [Cam06, MBC92]. Erste [Bur94a]. Erweiterung [GK98]. Erwin [KN94]. ESP [Ano00b]. espèces [Chi93a]. Essays [Yau92]. Essential [Cor95a, Cor02, Man02, BWB97]. estimation [CR94b]. estimators [BS05b]. étudiant [CT95, CT97, Ell96]. Euclidean [BV93, Kal93]. Euler [Bor89, RSL^{+}26]. EUR49.95 [Ano04]. Euro [Bod01, Top99, Bod01, S^{+}01]. Euro-conference [Top99]. Euro-Par [Bod01, Bod01, S^{+}01]. EUROCAL [vH83]. Eurocode [CC91]. Europe [Ska95]. European [JLL91, vH83]. Evaluating [Ron14, GFF01]. Evaluation [LF95, TP94, AB01, CK11, CJR92a, CJR92b, FVGF98, GK91, McL94a]. Evaluations [DK95]. Evaluator [Kun02]. Event [Ano01, GK91]. evolution [bLpL02, LL04a, pLbL03, XYXZ16, YL06, jZzFbL10]. Exact [Dou93, DP95, GW20, GU93, Ikl94, LL04a, pLbL03, LL04b, LTFN01]. Exakten [Ger92]. Example [AB94b, Zie07, AB99, AB05a, Wur96, Wri96]. Examples [CVW93, Gon88, Sal89, Sal91, CV88, Möl07]. Excel [Ano01, PM12]. exceptional [Gec94]. exchange [GKW94]. Excursion [Gar95, Lev92]. Exemples [GSZ95, ZSG5 ]. Exercices [JLP99, MNP97, LM98a, LM98b]. exotic [Bry91]. Expansion [MH93, ALL04, VČ88]. Expansions [McD97, Dom03, GH94, Gro89, Sal89, Sal91]. expectation [Dou93, DP95]. Experience [DHM95a, DHM95b, Man02]. Experiences [FLC93, CVG94]. Experiment [Bil94]. Experimental [CC93, EJ17, Viv01]. expérimentales [KL96]. Experiments [Ano95d, DG95, EM97b, MS94b, Ano96b]. Expert [HR92]. Explicit [ČPV90]. exploited [Vin91]. exploration [HJR^{+}03, SKH99]. Explorations [Bug95, HC95, May09, Jud96, KT95, KT99, Opr00]. Explorer [RLAH14]. Exploring [AKEJ94, CcMMK96, HEBS93, Ros97]. expressionals [MP16]. Expressions [Pin94, CB01, DM05, GKW94, Jef93, YF04]. extended [SRRB91, FIT03]. Extending [Dub00]. Extension [Ron14, AB05b, Bia00, DDdMS02, GK98, PN98, PPD05]. Extensions [Faj19]. extent [Jef93]. external [VČ88]. Extracting [YF04]. extrapolation [Gro90]. EZTerm [Ano00b].

facility [AB05b, Grä96]. Factor [CM24]. Factoring [CM22]. Factorization [CES94, Mon93, Rie94, Ber97b]. factors [AV11, Lah92]. Failures [BS02b, BS02c, BS05a]. falling [CT94b]. falling-leaves [CT94b]. FAQ [Bee05]. Fast [YJ99, JL94, LMRS08b, LMRS11]. Fault [Ano01, SKB01a, SKB01b]. Fearful [vRFS^{+}00]. Features [Vin91]. Feb [LW88]. February [Nes93, ÖSZ94, Ska95]. feedback [CR94b]. Feedforward [CP93]. Feigenbaum [SW90]. Fenchel [BV97].

Fermat [MP16]. Fermi [ABF14]. few [BP92, Kla10]. fibre [HK89]. fibre-preserving [HK89]. Field [AGR94, Ano96j, Ber97b, DM05, Hol96, VČ88, Vul02, van94b]. Fields [EHR^{+}96, Mur07b, DVL89, Wil89]. Fields/Herrmann [EHR^{+}96]. Fifth [Lum94]. figures [CT96]. Files [Ano95f, EM97b]. films [Opr00]. filter [Dou93, DP95]. final [Stu89]. Finance [vRFS^{+}00, PG10]. Finance/Maple [vRFS^{+}00]. find [ADdM14]. Finding [Ges95, PQ23, VJC12, AV11, bLpL02, LL04a, pLbL03, LL04b, MWCY14, Ose96, XYXZ16]. finds [AZ91]. Finite [BB92b, PC02, Ano95c, BSS03, Ber97b, FJ93, Sil04, SC94, Str94a, Ano03a]. finite-dimensional [BSS03]. Firewall [Ano00b]. First [BBRM95, Bur94b, C^{+}85a, C^{+}85b, CGG^{+}88a, C^{+}89, CGG^{+}89, CGG^{+}90, CGG^{+}92, M^{+}94, Roe94, ADdM12, ADdM14, BP92, BS99, Boy15, JLL91, LP05]. Fits [PM12]. Fitting [PM12]. Five [Shi05a]. flat [Wil89]. flexible [Lie94]. Flight [E^{+}92, EJLS96, Ell97a]. floating [MBC92, OS95, Shi93]. floating-point [MBC92]. Florida [ÖSZ94]. flow [CA07, Fit89]. flows [CDS95]. fluids [Car90]. fonctionnalités [Lem97]. Forces [Pie93]. forecast [CDdM14]. form [Bar93, BCEW12, ČPV90, CS89, CMC92, DM86, GS12a, Lie94, MNEA95, dYyR13, jZzFbL06, van95, SW98]. Formal [Izu14, KS12a, KS12b, Pec91, Bar01]. formalism [CV01, CM87, Van88]. formalisms [CC98]. Formel [GSZ95, RM97, ZSG5 ]. Forms [San96, BLPR13, Bro91a, MWCY14, Sei93]. formula [Fee90, Kim97]. formulae [PNG91]. Formulas [Lai94, RSL^{+}26, HSW94, HSW98]. Formulation [Bac96]. Forschung [EHR^{+}96]. Forschung/Omnès [EHR^{+}96]. Forte [Ano01]. Fortran [DG99, Gom90, GS98, Gro89, DG99, LP05, LS04]. Foundations [SZ94b, GKW03]. Four [Boy15, Bry91, CM91, RLAH14, Shi05b]. four-dimensional [CM91]. four-valued [RLAH14]. Fourier [DB15, Mat90]. Fourth [Gra94, GW00]. Fowler [Kwo90]. FRAC [AGR94]. Fraction [BS93]. fractional [BG11]. fractions [BC09]. fragmentation [Ikl94]. framework [KS13]. France [vdHvH12]. Frank [KOW93]. Fraser [Doo99, vRFS^{+}00]. fraud [BB92c]. Free [ES10, BLD96, Ber97b, Car90, Ste08, Ste13]. Freedom [Zir94]. French [Chi93a, Chi93b, Déc93, Las91]. Friendly [Kai94, Was04]. Frontiers [Gra94, Gra93]. Full [BS93]. Function [AGR94, Bee17, Dav88, MH93, Rav94, Roa96, AGR95, BP92, BEEK94, GQ93, GR92, YZ07, van94b]. Functional [Izu14, SKB01a, BPS12, CT94a, QBH06]. Functions [Asl96, BS93, tHB91, Col02, Dev93a, FMM94, Gar95, GHC92, Mul97, Mul06, Mul16, Pro00, Qua18, Ste95, Tho13, WL95, BC09, CD01, CJR92a, CJR92b, CA07, Dom03, FS92, Gar99a, Gro89, Gro90, Gro91, PFG06, PG08, PG10, SZ94a, Tar92, VHC17]. Funktionen [Sch93a, tHB91, Ger92]. fuzzy [Jon96].

G [Ano96d, CcMMK96]. Galois

[Koh94, Mal93]. Galoisgruppen [Mal93]. gamma [PG08, PG10, Har97, Tho13]. Gander [Ste96]. Garrett [Ano96a]. gas [EG97]. Gathen [Mon93]. Gauss [GM94]. Gaussian [Dou93]. Gaussians [PM12]. GBP [Ano98a, Ano99a, Ano99b, Ano04]. GCD [LF95]. Geddes [Ano96d]. Gegen [EHR^{+}96]. Geheimschreiber [Joy00]. General [CFFM97, McD97, Ngu93, Alu91, Ano94a, CM94a, CG95, Car90, Cha90a, CC98, ES92, Gro90, M^{+}94, McL94b, PG06, PS97]. general-purpose [Cha90a]. généralisés [Las91]. Generalized [JKP98, Kal93, GKaM01, Koe94, Las91]. generate [Bis91]. générateur [Gom90]. Generating [Dev93a, PNG91, Rav94, SVV95, BP92, Cam06, GS98, SZ94a]. Generatingfunctionology [Wil06]. Generation [ARL20, Bac96, Har98b, Gro89, Rod92, SC94]. generator [Gom90, Shi05a]. Generic [LMS07]. genus [Bog16]. Geometric [BR14]. Geometrical [Atz94]. geometrically [Sei92, Sei93]. geometries [Bry91]. Geometry [Mit00, Rov00a, Rov00b, Sch99a, SMW00b, SMW01, AV11, AH92, Bar01, CCM91, CTE95, CLO92, CLO07, ECT93, KT91, Tan89, Wan95a]. geophysics [DJ89]. George [Gre97]. German [Ben98, tHB91, Ell97b, GK98, Mal93]. Germany [AHK93, Bod01, CC92, Glo98, Top99, Wat91]. Geroch [CV01]. Gerthsen [EHR^{+}96]. Gets [Was02]. Getting [CKM98, CKM04, Map03a, Map04, Map05a, Map07a, Map08a, MMCK09, Put94, Wat01a, Wat02a]. Geyer [EHR^{+}96]. GFUN [SZ94a]. GHP [CV01]. Given [CM24, DK95, Alu91, BP92]. Glanz [EHR^{+}96]. Gleason [CcMMK96]. GLie [AH97]. Global [PG03, PG06, Was06, CDdM14, PG08, PG10]. goes [Fos06]. Goldbach [Yan94]. Gonzalez [Ano96b]. gradient [LLLL94, MN93]. Gramians [OP95]. graph [KU03, Sad96]. Graphical [Fos06, Ban01, Fiu95, LP05, Tay98b, WA96]. Graphics [LM94, Ska95, Ste25]. graphs [JM98, RW92]. Grassmann [AH97, AF05]. Grassmann-valued [AH97]. Gravity [Vul02]. Grazyna [Ano98a]. Greatest [DK95]. Greene [CT96]. Grenoble [vdHvH12]. Grid [PPD05]. Gripenberg [CVG94]. Gröbner [Cor96b, CG86, IA93, IA94a, IA94b, OS95, Shi93, Stu89]. Große [vRFS^{+}00]. großer [vRFS^{+}00]. Group [Koh94, BSS03, Li00a, Mic90, Pos13, Sid94, TP94, VAS06, YJ99, ZF91]. Groups [BB92b, EHR^{+}96, FJ93, GP92, HS92, HK89, Mal93]. Groups/Schmidt [EHR^{+}96]. Groups/Schmidt-Rohr [EHR^{+}96]. growth [CB01]. Grundlagen [JKM04, Wer96]. Grundstudium [Wer96]. gruppovye [ZF91]. Guide [BCD^{+}11, C^{+}85a, C^{+}85b, DHM95a, HWH91, HHR95, HHR96b, HHR98, Kai94, Kam95, Map03a, Map03b, Map04, Map05a, Map07a, Map08a, MGH^{+}96, MGH^{+}98, MGH^{+}01, MGH^{+}02a, MGH^{+}02b, MGH^{+}03a, MGH^{+}03b, MGH^{+}05a, MGH^{+}05b, MGH^{+}07a, MGH^{+}07b, MGH^{+}08a, MGH^{+}08b, MGH^{+}09a, MGH^{+}09b, MGH^{+}10a, MGH^{+}10b, Wat01a, Wat01b, Wat02a, Wat02b, Wat02c, EM07, Kam99, Wes99, BCD^{+}12, Ano96c, Ano96d]. guided [SH98]. Gunawardena [Ano99a, Ano99b]. Gustaf [CVG94].

H [Gre97]. Hagenberg [Jef08]. Hallett [CcMMK96]. Hamel [CA07]. hand [Sco01]. Handbook [AB94c, Bee17, Hog07, RW95, Red93, Red94, Red96a, Ano96e, Con96, Con98, GKW03, Hea96, Ros96, Mea96]. Handling [Koo91]. Hansen [Ano96c, SVV95]. Harada [Con92]. Hard [Ano00b]. Hardcover [CT96]. Hardware [Mav93]. harmonic [DJ89]. harmonics [IFF01]. Harper [Lum94]. HBFGen [dYyR13]. Heal [Ano96c, Ano96d]. Heck [Ano97a, Ano04]. Hecke [Gec94]. Heisenberg [EHR^{+}96, vRFS^{+}00]. Heisenberg/Zee [vRFS^{+}00]. Held [CV01, Nes93, AHK93, BBST94, CvGV92, EM94, Fit93, ÖSZ94, SSKT91, Tan89, Wou86]. Helium [Dav88]. Help [Dur19, Ful94]. helping [Sco01]. hereditary [FOW87]. Hermitian [KD90]. Herold [EHR^{+}96]. Herrmann [EHR^{+}96]. Heuristic [Mon92, LF95]. HI [KH97]. Hickory [Ano02b]. Hierarchical [SZ94b]. hierarchy [YZ07]. High [LMRS08a, BB92c, BWB97, PM07]. High-Performance [LMRS08a]. higher [CM94a, FOW87, VJC12]. Highlighting [Phi04]. Hilbert [Bro91a]. Hirota [dYyR13, jZzFbL06, jZzFbL10]. Hodge [Rab96]. HOL [HT93, HT98]. holonomic [SZ94a]. holonomies [Bry91]. Homogeneous [Wil89]. homotopy [HJR^{+}03]. Hopf [AF05, Koh94, Nag93, PNG91, QBH06]. Hopf-Galois [Koh94]. Horbatsch [SMC97]. HotGAMS [Bra96]. HR [Col05]. Hrebícek [Ste96]. Hubbard [Vin92]. Hughes [CcMMK96]. Hughes-Hallett [CcMMK96]. hulls [AF92]. Hungarian [Ano00a, Ano02a]. Hungary [Ano00a]. Hunt [Ano96a]. Huygens [CCM02]. Hybrid [GF92, LMRS08a]. hydrodynamic [CJ90]. hydrogen [VČ88]. hyperbolicity [Fit89]. hyperelliptic [Ber94]. hyperexponential [AZ91]. Hyperfine [GSF06]. Hyperformix [Ano01]. HYPERG [Gau99]. Hypergeometric [Le01, Roa96, Gau99, Koo91, PN98, Zei91]. hypersurfaces [Wu92]. hypervirial [VČ88].

IBFI [AHK93]. ICASSP [IEE93a]. ICCS [SvADS07]. ICMS [HY14]. IDE [Ano01]. Ideals [CLO92, CLO07]. Ideas [Ano95b, Lee93, Sch93b]. identifiability [IOP21]. Identification [BMW94, CR94a, MNEA95]. Identities [Ges95, ACS91, Sil04, Zei91]. IEE [Ano96g]. II [Sid94, CM94a, Car90, FVGF98, Pfl97, SvADS07, Tay98b, WL95]. III [GFF01]. Illinois [Lum94, Tan89]. IMACS [HR92]. Image [MS94b]. Impact [Run97]. Implementation [CM22, LP94, Mul97, Mul06, Mul16, ADdM12, Ber94, Bia00, DDdMS02, Fit93, GS89, GR92, McG84, PPD05, VM99, Wan95b, jZzFbL06]. implementing [CG86]. implicit [NS05]. imprecise [Bia00]. improved [CDdM14, LM88]. improving [CS08]. in-plane [BLD96]. InCert [Ano01]. included [Ano97b, Ano97c, Ano98a, Ano98b, Ano99a, Ano99b, Ano03a]. includes [Bar92, Ste94]. including [Ano96f, SH10]. inclusion [ČPV90]. Incomplete [Tho13]. Indefinite [Wan94, PS95]. index [CM94a, HF92]. Indexed [Wan94]. Indicator [Déc93]. indicatrices [Déc93]. Indicial [KMG96]. Individual [MH93]. induced [Lah92]. Induction [Pie93]. industry [CvGV92]. inequalities [KZ91]. inequality [BK94]. Infamous [Wic95]. infinite [LRB95]. Infinity [Vil95]. Influence [Kne94]. Informal [VV13]. Informant [Bra96]. information [BD87]. informatique [Ziz93]. ingénieur [CT95, CT97]. Ingenieure [Wes97]. Ingenieurmathematik [Ben98]. Ingenieurstudium [Kra97]. Ingredients [Man02]. inhomogeneous [AZ96]. Initial [Pet01a, FH04]. input [Dou93]. INRIA [Glo91]. ins [vRFS^{+}00]. installment [CVG94]. Institute [Lop94a]. Instruction [CW05b, CW05c, Zie07, Run97]. Instructor [Har15, Kir95]. Insurance [AST96]. integrability [SW98, WF90]. integrable [BCEW12, FF88, YZ07]. Integral [Kun02, Ron14, Yan93, Boy15, IA92, SC94, YC92, van94b]. Integrals [Ron14, AZ91, Ber94, Bog16, GQ93]. Integrating [ADG^{+}01, Gre93]. Integration [Jef93, Kun02, Pav94, Str94b, Tef02, Wan94, CA07, GS89, GF92, ZF91]. integrator [Hel10]. integrators [SIE98]. integrirovaniya [ZF91]. integro [KRR10]. integro-differential [KRR10]. Intelligence [HR92, CC92]. interaction [XYXZ16]. Interactions [Nay02]. Interactive [AB02, Ano05, CGG83b, CGG84, CGG85, DG98, HL94, Par00, Sch93b, BWB97, Dav01, Her01, Pri00, SKH99, Ano99a, Ano99b]. Interface [BHC95, BBST94, LHJ92, Leo86]. Interfaces [BFKS93]. interlocking [RLHAL11]. intermediate [CB01].

Intermolecular [Pie93]. internal [WVG90]. International [ACM94, AHK93, Ano88, Ano02b, Bod01, BGR91, Bro93, Bro07, CC92, Doo99, Fit93, GMV00, Gia89, Glo98, Glo91, Gon89, Gut04, HY14, IEE93a, IEE93b, Jef08, KH97, Kol00, Lak96, Lev95, LW88, Lum93, Lum94, Mal01, MCG13, Nes93, ÖSZ94, S^{+}01, SE11, SvADS07, Ska95, SW94a, Tod07, Wan92, WN90, Wat91, vdHvH12, CC91, dN93]. Internet [Fos99]. interpolation [RW92]. Interpretation [EHR^{+}96]. Interpreted [LMRS08a]. Interval [CDC^{+}02, CC93, Corxx, OS95, GK98]. Intervallrechnung [GK98]. intpak [GK98, Corxx, GK98]. intpakX [GK98, Krä07, GK98]. Introducing [Roe94, SKB01b]. Introduction [ASV04, Ano94a, Bau94, CGG83b, CGG84, C^{+}85a, C^{+}85b, C^{+}85c, CFG^{+}86, CGG^{+}88a, C^{+}89, CGG^{+}89, CGG^{+}90, CGG^{+}92, Cor95a, Corxx, CT95, EJ17, FM94, G´ 13, GM94, HC92, Hec93, HK95, Hec96, Hec03, Krä07, KL96, Lev98, NWF97, Sch98b, Sch99c, Sch03b, ST99, Str94b, TW06, Var95, VV13, Zac96, Zie07, vRFS^{+}00, Bay94, BR02, BS11, Bro97a, Cor02, CT97, CT01, CLO92, CLO07, Fiu95, Kof97, Lem97, NW96, SH09, ST04, Sza02, YSH96, Ano97a, Ano97b, Ano04, Ryb98]. Introductory [Abe93, MGH^{+}02b, MGH^{+}03b, MGH^{+}05b, MGH^{+}07b, MGH^{+}08b, MGH^{+}09b, MGH^{+}10b, LHJ92, Sny95]. Invariant [IS00, Alu91, Liu19, Nag93]. Invariants [CHF94, ADdM14, CM91, FFB99]. Inverse [JKP98, Boy15, Dom03, GKaM01, Ioa94, WW91, WyW93]. inverse/design [Ioa94]. investigate [Wil98]. Investigating [CB01]. investigation [FOW87]. Investigations [MS94a]. invited [SMW00b]. ionisation [BS03]. Iopsis [Ano01]. Ireland [HR92]. Iris [Leo86]. Irreducibility [Mon92]. Irvine [IEE89]. ISAAC [Gut04]. Isabelle [BHC95]. ISBN [Ano95b, CT96, Ryb98].

Ising [FMM94]. Isogroups [CDF92]. isolation [GG12]. ISOLDE [BPS12]. isotropic [BLD96]. ISSAC [Gia89, Jef08, ACM94, Bro93, Doo99, Glo98, Gon89, KRR12, Lak96, Lev95, MCG13, SE11, WN90, Wat91, vdHvH12]. Issue [Mon96]. issues [Phi04]. Itô [Ken01]. Italy [CC91, Gia89]. iterated [Bog16]. Iterative [Izu14]. itsol [Izu14]. IUTAM [BBST94]. IV [IFF01]. Iwahori [Gec94]. Iwahori-Hecke [Gec94]. IX [PFG06]. IZIC [FKM95]. Izmir [EM94].

J [Ano96a, Ano96b, Con92, Sto03, Lai94]. Jacobi [Las91]. Jacobian [JL94]. James [Mea96, Wri96]. Janeiro [M^{+}94]. January [ACM82, TRC94]. Japan [IEE93b, WN90]. Java [Ano01, And04, Kun02, LP05, LS04, Sch99b]. JCNN [IEE93b]. Jeffery [CA07]. Jerzy [Gan03]. Jets [Bar01]. Jírí [Ste96]. John [Ano96a, CVG94]. Joint [IEE93b, LM91a, LM91b]. Jose [SE11]. Joseph [Ano97b, Ryb98]. Joy [Gro96]. Julian [Zei92]. July [ACM94, Bro93, Bro07, Cha86, Doo99, EM94, Gia89, Gon89, Gut04, HR92, Jef08, KH97, Lak96, Lev95, M^{+}94, Wan92, Wat91, vdHvH12, GH84]. June [Ano02b, BGR91, HA94, MCG13, SE11, Tod07, Lee93]. Just [Sim92].

K. [Con92]. Kafka [EHR^{+}96]. Kaplan [GU93]. Karlsruhe [CC92]. Kassel [EHR^{+}96]. KdV [MWCY14]. KdV-type [MWCY14]. Keith [CHM94]. kernel [BV93]. Kernels [Pet01b]. Kevin [Ano96a]. Kiev [Bro93]. kinamatic [Kim97]. kind [Boy15]. kinematics [AHK93]. Kinetic [ČPRV87]. kinetics [CR94a, YB00a]. Kingdom [ACM94]. Klimek [Ano98a]. Kloeden [Gan03, Sto03]. knot [KU03]. Kofler [EHR^{+}96]. Kolodner [Kwo90]. Kolodner-Coffman [Kwo90]. Kopenhagen [EHR^{+}96].

Kopenhagen/Luchner [EHR^{+}96]. Korea [HY14]. Korteweg [Lev92]. Kosmologie [vRFS^{+}00]. Kosmologie/Meschede [vRFS^{+}00]. Kovacic [UW95]. Kovalevskaya [KBG01]. Krafft [vRFS^{+}00]. Kramers [FMM94]. Kreyszig [KN94]. Kurvendiskussion [Sch93a]. Kutta [Cam06].

L [Ano96a, Ano96c, CT96, Mea96, Ryb98, Wri96]. L. [Con92]. l986 [Tan89]. Laar [Lev92]. Lab [Wat95f, Wat95d, Wat95g, Wat95e, Ano97c, RC96, RC97]. Labahn [Ano96d]. Laboratories [BF91]. Laboratory [EM97a, Con96, Con97, Con98, Hea96, SCMS96]. Labs [FV94, Fat92, Fat96, Har94, BDS96, LL96, KR03]. Lafayette [Nes93]. Lagrange [Kim97]. Lake [Ano00a]. Lang [Con92]. LANGAGE [Car98, Car00]. Language [Ano01, C^{+}91, CGG^{+}91b, CVW93, Mav93, Vod82, Bra93, Phi04, Str99, TP94]. Languages [ACM82, Boy93, Car98, Car00]. Laplace [LW94]. Large [Pin93, KSK96, Yan95]. Laser [vRFS^{+}00]. Laser/Mantegna [vRFS^{+}00]. latest [Bro95, Pfl97]. lattice [Con92]. laws [Fit89, MWCY14, YL06]. Lax [MWCY14]. leading [LLLL94]. leaf [Shi98]. Learn [Dur19]. Learned [Sch03a]. Learning [Ano96c, HHR95, HHR96b, HHR98, Map03b, Pav94, Wat01b, Wat02b, Her01, Möl07]. Least [PM12]. Least-Squares [PM12]. Leaves [C^{+}85a, C^{+}85b, CGG^{+}88a, C^{+}89, CGG^{+}89, CGG^{+}90, CGG^{+}92, CT94b]. lecons [LBQTV97]. lecture [M^{+}94]. Lee [Ano95b]. Leech [Con92]. Legendre [Boy93]. Lehr [Wer96]. Lehr- [Wer96]. Leibniz [BWB97]. Leipzig [EHR^{+}96]. Leipzig/Kofler [EHR^{+}96]. Leitfaden [EHR^{+}96]. length [CB01, Zei92]. lernen [Wer93a, Wer96]. Lessons [Sch03a]. level [ZJEP95]. Levin [Gro89]. Lexikon [vRFS^{+}00]. Liberal [Put94]. Libraries [FS95]. Library [AV02, Bee17, BR14, BKR^{+}91, CGG^{+}91a, Faj19, GK98, SMW00b, CML^{+}08, CDL^{+}11, JM17, LMX05, LMRS08b, LMRS11, PM07, DG99, GK98]. Licht [vRFS^{+}00]. Lie [CCM91, AH97, FJ93, HK89, LRB95, Rod92]. Lie-series [Rod92]. Life [AST96]. Light [Har98a, Har00]. Light-Cone [Har98a, Har00]. likelihood [BS05b]. Lillestol [vRFS^{+}00]. Limited [Ful94]. Limits [VHC17]. linéaire [LBP98]. lineal [SCL95]. Linear [Aue91, Bac96, BEJ95, CES94, Edw95, E^{+}92, EJLS96, Her01, Hog07, Joh93, Kei01, Kne94, LMJ93, Mal06, Pie93, SF10, Sza02, WY93, WL95, AZ96, Ano99a, Ano99b, Bar93, BPS12, BLPR13, CJ97, DG98, Die92, DU97, Ell97a, HSW98, HPS05, JKP96, Kla10, LL96, Rod92, Rya99, SE22, YC92, Aue90]. linearization [SRRB91]. lines [Alu91, Wu92]. link [Kim97]. linkages [Shi05a, Shi05b]. Links [BKR^{+}91]. Linux [Ano00b, Ano01]. Linz [Ano02a]. Liouvillian [ADdM12]. Lipsman [Ano96a]. liquid [KOW93]. LLC [Ano00b]. LMS [Dou93, DP95]. local [Vol93]. Loci [BR14]. log [Har97]. log-Gamma [Har97]. Logging [BS02a]. logic [RLHAL11]. logiciel [Lem97]. logics [RLAH14, RLAH14]. Lommel [FS92]. London [vH83, TR94]. Long [Pie93]. Long-Range [Pie93]. Look [Pav94, CVG94]. Lorentzian [CM91]. lösen [Wer93b]. Lösungen [Vet95]. Louisiana [Nes93, TRC94]. Lower [Lah92, Vin91]. Lozi [Rei96]. LS [GF02]. lubrication [CJ90]. Luchner [EHR^{+}96].

M [Ano96c, Ano96d, CcMMK96]. M. [Con92, Koe91]. M.-L. [Con92]. M0 [Bau94]. MA [MCG13]. Macaulay [Min04]. machine [Ber97a]. machines [GS98]. Maciej [Ano98a]. Macintosh [C^{+}89, Wat95c, Bro97b]. MACROFORT [Gom90]. MACSYMA [Ben98, Ben98, GH84, CVG94, Grä96, PZ96]. Magic [Sco5 ]. Magma [ES10, Ste08, Ste13]. Magnetic [EHR^{+}96]. Mai [GW00]. mails [Bra96]. Mainframes [Sim92]. Making [Col02, SZ94b, Col05]. Management [Ano01, SW94a, BD87]. Manager [Sch00b, Sch01a]. Manager-Worker [Sch00b, Sch01a]. Manchester [S^{+}01]. Manhattan [Ano95d]. manifold [QBH06]. Manifolds [Mur07a, Yau92]. Manipulating [Car97b, Gau99]. Manipulation [BB92b, Boy93, Fat90, IA93, SZ94a, TP94]. manipulator [MNEA95]. Mantegna [vRFS^{+}00]. Manual [Aue90, C^{+}85a, C^{+}85b, CGG^{+}88b, C^{+}90, C^{+}91, CGG^{+}91a, CGG^{+}91b, E^{+}92, EJLS96, EM97a, GGC81a, GGC81b, GGC83, Har15, KN94, Map05b, Map07b, Map08b, Map09, Map10, Map11, Map12, Sch98a, WG83, Ell97a, SCMS96]. Manufacturing [SZ94b]. Many [Pie93, Alu91, GFF01]. Many-Body [Pie93]. many-particle [GFF01]. MAPinsure [AST96]. MAPLE [Ben98, JKP98, PS95, Ste96, Ano95b, Ano95f, Ano96e, Ano96f, Ano96h, Ano97b, Ano99a, Ano99b, Ano00b, Ano03a, AB01, Bag95, Cha90b, Cha91, Cor96a, EHR^{+}96, GK98, Gre97, Har15, HHR96a, Hea96, Lop94a, Mad92, Mea96, MD96, NWF97, Ryb98, Tan99, Vet95, Vod82, Wer93a, Wer93b, Wer96, Wri96, vRFS^{+}00, AZ91, Ano94a, AST96, AH97, Bar01, Bay96, Bay05, BCG^{+}05, Ben98, Bor89, Bra93, Bro97a, C^{+}90, CB02, CCM02, CZL11, Cyg96, CMC92, Die92, Dou96, EG97, FFB99, Fit89, GH93, GKLM91, Gar99b, GGC83, GS89, GF92, GLHR19, Gom90, Gro89, Her01, Hrě95, IS00, IA92, JT91, JKM04, JKP03, JS00, KRTZ20, KBG01, KOC02, LMX05, Lev92, LMRS08b, Lyn01, McD97, McL94b, Mur07b, Mur07a, Ose96, PN98, Pie93, PG03, PG06, PG10, PGG13]. MAPLE [PNG91, Qua18, Rab96, Rei96, Rod92, Rov00a, Rov00b, Rya99, Sta07, ST04, Ste94, TP94, Tef02, Tho94, Van94a, Van88, Vei97, VJC12, Vul02, Vul03, WEM98, Wan90, WW91, Wur96, Xu95, YB00a, YB00b, YZ07, Kir95, Abe93, AB94a, AB94b, AB94c, AB99, AB00, AB05a, Abł96, AF05, ACGL04, Ada97, ADG^{+}01, ASV04, AV11, AV02, AB02, ABP02, AL98, AAFMJ01, ARL20, AGR94, AY03, And04, And97, Ano90, Ano92a, Ano94b, Ano95d, Ano95e, Ano96i, Ano96j, Ano01, Ano05, AR06, AB94d, AB05b, Art98, AKEJ94, Atz94, Aue90, ALL04, ADdM12, ADdM14, BHC95, BR09, Bar92, BSS03, BPS12, BF02, BF09, BLM97, Bar96, B^{+}98, BS03, BF91, Bau94, BEJ95, BF96, Bau96, BV97, Bay94, Bee05, Bel95, BFS93, Ber97a, BMH02, BCD^{+}11]. Maple [BCD^{+}12, Bet98, Bet01, BR02, Bia00, Bil94, Bis91, BM94a, BM94b, Bla5 , BS99, BBRM95, BS95, BS11, BS02a, BS02b, BS02c, BS05a, BR14, BS05b, Bra96, BDS96, Bro97b, Bro97c, Bro95, Bug95, BD93, Bur94a, Bur94b, CB01, CDC^{+}02, Cap93, CK11, CL95, CDdM14, CJ97, CV01, Car97a, Car97b, Car98, Car00, CS93, CD01, CDK94, Cha96, CGGG83, CGG83b, CGG83a, C^{+}84a, CFG^{+}84, C^{+}84b, CGG84, CGG85, C^{+}85a, C^{+}85b, C^{+}85c, CFG^{+}86, C^{+}87, CGG^{+}88a, CGG^{+}88b, C^{+}89, CGG^{+}89, CGG^{+}90, C^{+}91, CGG^{+}91a, CGG^{+}91b, CGG^{+}92, CT94a, CML^{+}08, CDL^{+}11, CM22, CM24, CH94, CcMMK96, CKM98, CKM04, CM96, CW05b, CW05c, CW05d, CW05e, CFFM97, CVW93, CM99, Col02, Col05, CM01, CC93, Con96, Con97, Con98, CP93, C^{+}96, Coo97, CS08, CJ90, CES94, Cor95a, Cor95b]. Maple [Cor02, CA07, Corxx, CT95, CT97, CT01, Cos00, CSS94, CVG94, CM94b, CC98, CGK01, CKO02, CM87, Dav88, Dav01, Dec94, DP11, DG98, DM05, Dev93b, Dev93a, DGS08, DU97, DG95, Don97, DS94b, DB15, DDdMS02, Dub00, DG97, Dur98, Dur19, EP94, EL89, E^{+}92, EJLS94, Ell96, EJLS96, Ell97a, Ell97b, ECTB^{+}14, EM97a, EM97c, EM97b, EM00, ES10, FH04, Faj19, FV94, Fat92, Fat96, Fer98, FF11, FGH94, FLC93, Fos01, Fos06, Fos11, Fri97, FVGF98, FIBT01, FIT03, Fri09, Ful94, GFF01, GF02, GSF05, GSF06, GH95, GH97, GH04, GW20, GI09, Gar97, Gar99a, Gar02, GS12a, Gau99, GGC81a, GGC81b, GC^{+}88, GMM^{+}89, GK20, GK98, GH03, GU93, GHL^{+}00, Glo93, GSZ95, GS98, G´ 13, Gon88, Grä96, Gre95, Gre99, GS12b, Gro91]. Maple [GHC92, GKaM01, HC95, HC96, Har94, HDC94, HC92, Har92, HL95, Har98a, Har00, Har14, HT93, HT98, HJR^{+}03, HHR95, HHR96b, HHR98, Hec93, HK95, Hec96, Hec03, Hel10, Hol96, HEBS93, Hor95a, Hör95b, HJOP03, HH95, IOP21, IFF01, IA94c, IA94b, Ioa94, Isr96, Izu14, JWK^{+}21, JLP99, JM17, Joh93, Jon96, JKP96, JM98, Jud96, Kam95, Kam99, Kam93, KT95, KT99, Kei01, KS12a, KS12b, Kim97, Kla10, KSS00, KSS07, KK97, Kne94, Kof94a, Kof94b, Kof97, Koh94, Kom95, Koo91, KU03, KRR10, KRR12, Krä07, KR03, Kra97, KN94, KL96, KSK96, Kun02, Lai94, LP05, LL96, LM91a, Lee93, LS04, LBQTV97, Lem97, LBP98, Ler95, LM98a, LM98b, Lev97, Lev98, LP94, Li00a, Li00b, bLpL02, LL04a, LMRS11, LEGS93, pLbL03, LL04b, Lop93, LW94, Lop94b, Lop94a]. Maple [LTFN01, LM94, Lyn09, ME18, MG94, MNEA95, Mal06, Map03a, Map03b, Map04, Map05a, Map05b, Map07a, Map07b, Map08a, Map08b, Map09, Map10, Map11, Map12, MS94a, Mav93, May09, McL93, MB98, MMCK09, MNP97, MWCY14, Min04, Mol94, Möl07, MN93, MGH^{+}96, MP97, MGH^{+}98, MGH^{+}01, MGH^{+}02a, MGH^{+}02b, MGH^{+}03a, MGH^{+}03b, MGH^{+}05a, MGH^{+}05b, MGH^{+}07a, MGH^{+}07b, MGH^{+}08a, MGH^{+}08b, MGH^{+}09a, MGH^{+}09b, MGH^{+}10a, MGH^{+}10b, MP10, MP12, MP14, MP16, NS05, Nay02, Ngu94, NW96, OP95, Opr00, OS95, PFG06, Par97a, Par97b, Par97c, Par00, PM07, PR98, Pet00a, Pet00b, Pet00c, Pet01a, Pet01b, PPD05, Pet92, Phi04, PQ23, Pin93, Pin94, PG08, PMS20, PF95, PC02, PS97, PZ96, Pos13, PM5 , Pri00, Pro00, Put94, Put03, PM12, QBH06, RAB03, Ram98, Rav94, Rec98]. Maple [Red93, Red94, Red95, RC96, Red96a, Red96b, RC97, Rhe93, Ric02, RGM95, RLvLRM10, Rob96, Roe94, Ron14, Ros97, Rot97, Roy93, Run97, RM97, SE22, Sad96, Sal93, SZ94a, SW98, SCMS96, Sch98a, Sch99b, Sch00a, Sch00b, SKB01b, Sch01b, Sch03a, SMB03, Sch5 , Sch98b, Sch99c, Sch03b, MAP87, SF90, SMFC92, Sco5 , Sco01, SIE98, SLC07, Shi09, Shi05a, Shi05b, Shi98, SF10, SD10, SH09, Sie93, Sil04, SSTW94, SV96, SCL95, Spi94, ST99, SH10, SZ03, Ste08, Ste13, Ste95, Ste25, SC09, SKH99, Sym88, Sza02, TT00, TR94, TA94, TA95, Tay98b, Tay98a, TTL93, TTLM94, Tho17, TA96, TMH94, VAS06, Vân94d, LM91b, VMRA14, VM99, Viv01, VV13, Wat01a, Wat01b, Wat02a, Wat02b, Wat02c, Wat96, WH92, WWK92, WyW93, Wan95b, Wan06, Was02]. Maple [Was04, Was06, Wat93, Wat95f, Wat95j, Wat95l, Wat95h, Wat95k, Wat95d, Wat95g, Wat95i, Wat95e, Wat95a, Wat95b, Wat95c, WG83, Wes95, Wes97, Whi10, Wil98, Wri01, XYXZ16, qXbL04, Yan95, YJ99, YA07a, YA07b, Yan93, dYyR13, YL06, YC92, Yua97, Zac96, Zei91, jZzFbL06, jZzFbL10, ZSG5 , Zir94, Zot07, Gan03, Ano96b, Ano97a, Ano04, Bee04, CT96, RLHAL11, RLAH14, SMC97, Ano96c, Ano96d]. Maple-assisted [VAS06]. Maple-AVS [LW94]. MAPLE-package [Bar01, AY03]. Maple-V [Shi05a, Shi05b]. Maple/Mathematica [Bel95]. MAPLE/MATLAB/C [WEM98]. MapleNet [Man02]. Maple(R) [Ano97c, Ano98b, Ano98a, Sto03]. maplet [YC07]. MapleTech [Mon96]. MapleTensor [KMG96]. Maple^{T M} [Ano96f, Ano96a]. Mapping [JT91, LP94, CDdM14]. mappings [BB92a]. Maps [ARL20, LLLL94]. Mar [Tan89].

March [DS94a, IEE89, JLL91, Nes93, Top99, vH83]. Mario [CVG94]. Marko [SMC97]. Markov [CDC^{+}02]. Markovketten [JKM04]. Marriage [HT93]. MARS [WEM98]. Martha [Mea96, Wri96]. massively [Ber97a]. matches [FM89]. Matematicas [SV96]. Material [Roy93]. Math [Ano90, Bau96, Fos99, Sim92, BWB97, Fos91]. MATHCAD [Ben98, EHR^{+}96, Ben98, Bar92, Möl07, BFS93]. MathCW [Bee17]. MATHEMATICA [Ben98, Ben98, Ste94, Tho94, Ano90, Bar92, BLM97, Bel95, BFS93, Bra96, CW05b, CW05c, CW05d, CW05e, Dur98, ES10, Fos11, Grä96, Har14, JLP99, LP05, LC90, Möl07, PZ96, SLC07, Shi09, ST99, ST04, SH10, Ste08, Ste13, SC09, Zot07, Bag95, Har15]. Mathematical [AB02, ABP02, Ano95e, Ano01, BF02, BF09, Bee17, BR02, CM94b, DHM95a, DHM95b, ES10, EM94, FKM95, FM94, Gon88, HY14, Lee93, Man02, Mon94, MP14, RSL^{+}26, SH09, Was04, AR06, BS11, CC92, Ell97b, GKW94, Hen94, Jon96, Ric02, Run97, SMFC92, Wan06, YF04, Ano95b]. Mathematical-Function [Bee17]. Mathematicians [DPV14, SKH99]. Mathematics [CHM94, CVG94, Dev92, Dev93c, Dev94a, Dev94b, DG99, EJ17, FGH94, Har14, HHR95, KN94, LW88, Lop93, Lum94, RW95, Rob96, Sch93b, ASV04, And97, Ano96b, Ben98, Ber93, BWB97, CJ97, CGG^{+}86, DS94a, DG95, GW00, Lop94a, Lum93, Mat90, Opr00, Ros97, SLC07, Shi09, SKH99, Tod07, Viv01, YSH96, Har15]. Mathematik [Wer93a, Wer96, Wes97]. Mathematikaufgaben [Wer93b]. Mathematiker [Glo93]. Mathématiques [Rot97, Ziz93, Don97, DG97, Dur98, JLP99, KL96]. mathematischen [EJLS94, Ell97b]. MATHLAP [Ben98, Ben98]. MathML [Ano02b, BMH02, Nay02]. MathResource [BWB97]. Maths [Dou96, Fer98].

MathView [Wat97a, Wat97b]. MATLAB [WEM98, BF09, ES10, GH93, GH95, GH97, GH04, GS12b, KSS07, Ste13, Ano05, Cam06, CW05b, CW05c, CW05d, CW05e, Möl07, Pri00, SH09, SCL95, Ste08, LS04, Ste96]. Matrices [Pin93, Zir94, BR09, JL94, PFG06, FMM94]. Matrix [WY93, Cor96b, DM86, GFF01, IS00, KU03, KD90, KZ91, SD10]. Matter [ÖSZ94]. Maui [KH97]. MAX [CVG94]. Maxima [SH10]. Maximum [PQ23, BS05b, Jef93]. Maxwell [DVL89, ECT93]. May [SvADS07, SW94a]. McGuire [Gre97]. Means [SVV95, Gro90, PN98]. mechanical [Les94, SMF^{+}90, SMM^{+}90, VC89, Wan95a]. Mechanics [Abe93, EHR^{+}96, Gro96, SMC97, TMH94, Top99, CT96, CV88, Gre95, Hor95a, Ste94, SH10]. Mechanics/Neuser [EHR^{+}96]. Mechanik [vRFS^{+}00]. meeting [Jef08]. Meets [ADG^{+}01]. Meier [GU93]. mémoire [Lev97]. memory [Ber97a]. Meschede [vRFS^{+}00]. mesh [GS12a]. metabolic [YB00b]. Method [BCG^{+}05, Ges95, JT91, PQ23, Qua18, San96, Tef02, BG90, CM94a, Dom03, DDdMS02, GS89, Gro90, GKaM01, JL94, KU03, Kwo90, LL04b, Liu19, Stu89, Wan95b, Yan94]. Methoden [Wer96]. Methods [Ano88, Bay94, Boy93, Glo91, MS94b, ÖSZ94, Rao02, Rie94, AR06, Ban01, FOW87, GS12b, Par00, Pec91, Ric02, Run97, Wan06, WA96, Whi10, ZF91]. metody [ZF91]. Mexico [ACM82]. Michigan [Lee93]. Micro [CW92]. Micro-analysis [CW92]. Microsoft [PM12]. Minimal [Le01]. MiniMaple [KS13]. minimum [ACS91]. Minneapolis [IEE93a, SW94a]. Minnesota [IEE93a, SW94a]. mirror [Yau92]. Misfortunes [DPV14]. Misiurewicz [Rei96]. Mixed [Hoh01, Phi04]. Mixture [Bil94]. Mixtures [BH93, Lev92, SD10]. Mode [GSZ95, LS96, Sha94, SRRB91]. Model [Bro95, FMM94, ČPV90, Ikl94, MNEA95, Vin92]. Modeler [Ano01]. Modeling [Ano01, CCN06, Cap93, EP96a, EP96b, EP96c, EP04a, Ngu93, Ngu94, EP03, EM07, Les94, SCMS96, VAS06, Yua97]. Modelling [CM94b, FM94, BF02, BF09, Bro95, CDS95, EM94]. Models [BMW94, CR94a, GWV93, IOP21, Par00, Vin91, WVG90, YSH96]. Modern [vRFS^{+}00, BS11]. Moderne [Kom95]. modpn [LMRS08b, LMRS11]. Modular [Gar99a, HS92, JS00]. module [Fri09]. Modules [CM99, CML^{+}08, Her01, Lah92]. moduli [Bog16]. moléculaires [Chi93a, Chi93b]. Molecular [LH93, ÖSZ94, ALL04, Chi93a, Chi93b, CSS94, SIE98, VAS06]. Molecule [Pie93]. molecules [VAS06]. Moment [McD97]. momenta [Fri97, FVGF98, FIBT01, FIT03, Fri09, GFF01, GF02, GSF05, IFF01, PFG06]. moments [CB02]. Monagan [Ano96d]. Monolith [Ano00b]. monoterm [Liu19]. MontaVista [Ano00b]. Monte [YC07]. month [CVG94]. Monthly [Gon88]. Montréal [Lev95]. mosaic [CT94b]. Moser [Bar95]. Motif [Ano00b]. Motion [FM89, IEE89, Lie94, Shi05a]. movement [IA94a]. MP [GKW94]. MPL [Bog16]. MR [Con92, Min04]. much [Bro95]. Multi [Kim97, YZ07]. multi-function [YZ07]. Multi-link [Kim97]. Multibody [Cap93, Lie94]. Multidimensional [EHR^{+}96, LL95]. Multilevel [GWV93]. MultInt [Tef02]. multiobjective [PGG13]. Multiple [Asl96, Pet01b, San96, Tef02, BC94a, Pet94]. multiple-delayed [BC94a]. multiple-point [Pet94]. Multiple-Valued [Asl96]. multiplication [MP10]. multiplicities [Lah92]. multiplicity [FM89]. Multipole [McD97]. multisoliton [jZzFbL10]. Multivariable [CJ97, CH94, CcMMK96]. Multivariate [CM22, CM24, AV11, Ber97b, DGS08, LMJ93]. Munich [Bod01]. MuPAD [Ben98, Ben98, Grä96, PZ96].

Myrtle [HA94].

nach [vRFS^{+}00]. NAG [BKR^{+}91]. Nagoya [IEE93b]. NATO [EM94]. Natur [EHR^{+}96]. Naturwissenschaft [Hör95b]. Naturwissenschaftler [vRFS^{+}00]. Naturwissenschaftler/Weizsäcker [vRFS^{+}00]. Navigator [Ano00b]. Near [CW05e, AGR95]. near-separated [AGR95]. Negative [Tho13]. negativity [ME18]. Neighbors [EZ23]. Nested [Dom03]. Net [Ano00b]. Netherlands [CvGV92]. Network [Ano00b, Ano96k, Tsa93, Vân94d]. networked [SMB03]. Networks [HDC94, IEE93b, YB00b]. Neural [IEE93b, Tsa93]. Neuser [EHR^{+}96]. Newman [CM87]. News [Bra96]. Newsletter [MAP87]. Newton [BWB97, ACS91]. NG [PM12]. Ninth [ACM82]. NIST [Bra96]. NJ [DS94a]. NMR [EHR^{+}96]. No.1996 [Ano96g]. No.1996/078 [Ano96g]. No.94CH35723 [Gra94]. Noether [GHL^{+}00, Vol93]. Non [Bac96, SIE98, CDS95, ME18, Rod92, SE22]. Non-equilibrium [SIE98]. Non-Linear [Bac96, Rod92, SE22]. non-negativity [ME18]. non-uniqueness [CDS95]. noncommutative [HJOP03]. nonexpansive [BB92a]. Nonhyperbolic [Nag93]. Nonlinear [CS93, EM97a, EM97c, EM97b, EM00, Les94, MG94, Mur94, San96, CM94a, CZL11, GLHR19, LW88, Li00b, bLpL02, LL04a, pLbL03, Ose96, Roa94, SRRB91, Van94a, WH92, XYXZ16, qXbL04, dYyR13, YL06, jZzFbL10, dJ91, de 93, de 95, Gre97]. nonrelativistic [VC89]. Norm [KZ91]. Normal [San96, CS89, van95]. Norman [GS89]. North [HA94, TR94]. Northern [Lum94]. Note [UW95, Oka92]. Notebook [Ano95f]. notes [M^{+}94]. November [CC91, Gra93, Lum93, Lum94, Wou86]. Novosibirsk [Mal01]. NP [AFK^{+}94, CM87]. NPspinor [CCM02]. NPTOOLS [CC98]. NQL [Ano01]. NSWC [DG99]. null [DVL89]. Number [Cos00, Dur19, EZ23, Col05, ES92, ENS92, Rab96, Zei92]. Numbers [Arn95, Rie94, BB90, Wil98, Yan95]. Numeric [BCG^{+}05, BKR^{+}91, GDS94a, ALP96, GF92, GDS94b, LP05, Roa94]. Numerical [BFKS93, CJR92a, CJR92b, FH04, Kol00, LP12, McL94a, MSW00, Nes93, Pet00a, Rao02, SW90, Ban01, BEEK94, CM94a, Fri97, GW89, GW94, GS12b, Gro89, WA96, Xu95, Tay98b]. numerische [Wer96]. NY [Ano04]. nykh [ZF91].

O [Ano96d]. O.D.E [Ano97c]. O.D.E. [RC96, RC97]. Object [KSK96, Ano96j, Hol96]. object-oriented [Ano96j, Hol96]. Objects [Dev93a, MP14]. obtain [Jef93]. obyknovennykh [ZF91]. October [AHK93, Ano02a, GMV00, IEE93b]. odd [GP92]. ODE [Cha96, IOP21, PZ96]. ODEs [Bar96, LSYL11, Pet00a]. OEM [TMH94]. Ombach [Gan03, Sto03]. Omnès [EHR^{+}96]. One [GS12a, SZ94a]. One-peak [GS12a]. ongoing [Ken01]. Only [Wat95h, Wat95i]. Ontario [Bro07, Cha86]. open [Ste08, Ste13]. OpenMath [Nay02]. Operating [Vod82]. operations [Par00, YJ99]. Operators [Pro00, FOW87, KRR10, SD10]. Optical [Atz94, YC07]. Optics [Atz94, Möl07]. Optik [vRFS^{+}00]. Optimal [BB92a]. optimality [SZ03]. Optimization [Bac96, Was06, Zim95, PG03, PG06, PG08, PG10, PGG13]. Oracle [BF91]. Orbital [MH93]. Orbitals [PM12]. Order [EHR^{+}96, MH93, ADdM12, ADdM14, Cam06, CB02, FOW87, HK89, LLLL94, Mat89, VMRA14]. Order/Kafka [EHR^{+}96]. Ordinary [LS04, ADdM12, HK89, KRR12, Rod92, Rya99, VMRA14, ZF91]. Oregon [Gon89]. oriented [Ano96j, Hol96, KSK96]. Orleans [TRC94]. ORTHOFRAME [Van88]. Orthogonal [BGR91, Koe93]. orthonormal [Bis91, Van88]. Osborn [Ano96a]. oscillating [Sha94]. oscillations [WH92]. Oscillatory [San96]. Oseen [KOW93]. Oseen-Frank [KOW93]. other [AZ96]. Otter [Col05]. overview [Ano94a]. Oxford [ACM94].

P [Mea96, Sto03, Wri96]. PA [Pav85]. Package [ABP02, AGR94, AST96, Car98, CC93, DP11, Izu14, Mur07a, PMS20, Pro00, Ste95, Tef02, Vei97, AF05, AY03, Ano96j, ALL04, ADdM14, BC09, Bar01, BPS12, Bro95, Cam06, CDdM14, CV01, Car00, CCM02, CZL11, Cyg96, CC98, CM87, FFB99, FF11, GI09, Gar99b, Gau99, GLHR19, GK98, GHL^{+}00, Gro91, Hol96, HJOP03, JWK^{+}21, Jon96, JM98, KRTZ20, KBG01, KRR10, bLpL02, LL04a, pLbL03, LL04b, ME18, MWCY14, Min04, OS95, Pos13, RLAH14, Rya99, Sad96, SZ94a, Sil04, Tal94, Van88, VJC12, XYXZ16, YA07a, dYyR13, YL06, Yua97]. Packages [And04, Car97a, Car97b, BFS93, jZzFbL10]. PaCT [Mal01]. Padé [CZL11, ČPRV87]. Page [HF92]. pages [Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano97b, Ano97c, Ano98a, Ano99a, Ano99b, Ano03a, Ano04, CT96]. Painlevé [KBG01, qXbL04]. Paipa [LW88]. pairs [ENS92, MWCY14]. Pak [Wat95f, Wat95d, Wat95g, Wat95e]. Paket [GK98]. paper [Con92]. Papers [ACM82, HR92, Wou86]. Parallel [Ano00a, Ano02a, KH97, Mal01, Mit00, MSW00, Pet01a, PS95, Sch99b, Sch00a, Sch03a, Top99, Ber97a, Bod01, Cha90a, GWV93, S^{+}01, SMB03, Wou86]. Parallelism [Sch00b, Sch01a]. Parallelizing [SMW00b, Sch01b, Sie93]. parameter [SC09]. parameterizations [van94c]. parameters [KZ94]. parametric [KRTZ20]. ParametricSystemTools [CML^{+}08]. parametrisation [GSF06]. Paris [JLL91]. Part [Con97, Mic90, SvADS07, Sid94, CVG94, CW05b]. Partial [Art98, Bet98, BS93, CK11, DiB95, ST99, ST04, dv96, CT94a, GS12b, LS04, qXbL04, Ano98b]. particle [GFF01, GQ93, IA94a]. Partition [FMM94]. partitions [ENS92]. parts [Ano95c, Str94a]. PASCO [KH97]. path [Bry91]. pattern [Sad96]. Patton [Ano95f]. Pauli [Bay96]. Pauli-algebra [Bay96]. PBW [Faj19]. PC [Ano95f, Bau94, Lav91, RM97]. PC/Windows [RM97]. PC300 [Ano00b]. PDE [jZzFbL06]. PDEBellII [MWCY14]. PDEs [GLHR19, LP94, LRB95]. peak [GS12a]. pedagogical [Kla10]. Peer [BS02a]. Penrose [CV01, CM87]. perfect [Car90]. Performance [CFG^{+}84, C^{+}84b, LMRS08a, Sch00a, Zot07, GK91, PM07, Run97]. Performing [KMG96, CV01, CM87, Van88]. periodic [GKLM92]. permutations [Zei92]. Perturbation [Ada94, VČ88]. Perturbative [SMF^{+}90, SMM^{+}90]. Pervasive [Ano00b]. Peter [Gan03]. Petrov [Car90, DVL89, LM88]. Phase [GKLM91, Fit89, GKLM92]. Phenomena [GKLM91, BBST94, EG97, LW88]. Philadelphia [Pav85]. Photoeffekt [EHR^{+}96]. Photoeffekt/Kleint [EHR^{+}96]. Physical [AB02, Bay94, CS93, GDS94a, GDS94b, Har14, Har15]. Physicists [CM96]. Physics [CVW93, Com96, EM97a, EM97c, EM97b, Gre93, Gre97, Wan06, Wic95, dN93, vRFS^{+}00, EM00, GQ93, Run97]. Physics/Fraser [vRFS^{+}00]. Physics/Schroeder [vRFS^{+}00]. Physik [EHR^{+}96, Kom95, vRFS^{+}00, BS95]. Physik/Kassel [EHR^{+}96]. Physiker [vRFS^{+}00]. Pi [Bor89]. Piecewise [Qua18]. Pilsen [Ska95]. Piscataway [DS94a]. pitchfork [Xu95]. pivoting [AF92]. Plane [BR14, DP11, Zir94, Alu91, BLD96]. Planetary [MH93]. plasma [CTE95, ECT93]. plates [BLD96]. plot [Bee04]. Plotting [MSW00, SMW00c]. Plovdiv [Kol00]. Plus [EHR^{+}96]. Point [ČPRV87, FM89, HK89, MBC92, Pet94, Shi93, VAS06]. points [Shi05a, Shi05b]. Poisson [Ano96j, BV93, Hol96]. Polar [DB15, Sei93]. POLY [MP12, MP14]. polyene [Vin91]. polyhedra [AF92, JM17]. polykays [DGS08]. Polymers [EHR^{+}96]. Polymers/Ruder [EHR^{+}96]. Polynômes [Las91]. Polynome [Mal93]. Polynomial [Qua18, BG11, CB01, CM11, CDL^{+}11, FFB99, Grä96, LMRS08b, LMRS11, LF95, MP10, MP12, PM07, Vân94d, VHC17]. Polynomials [CM22, CM24, DK95, Mon92, Pro00, Vei97, AV11, AAFMJ01, AGR95, Ber97b, Bis91, BGR91, FS92, GG12, Koe93, Koe94, Las91, Liu19, LMJ93, Stu89, Mal93]. Polytechnic [Lop94a]. Ponte [ÖSZ94]. Popkin [Ano01]. Portable [Bee17, C^{+}84a, CGGG83]. Portela [Ano03a]. Portland [Gon89, Lum93]. Portuguese [Sid94]. posets [GS12a]. position [Alu91]. positive [GS12a]. possible [Vul03]. PostScript [YF04, Bee04]. Potentials [Pie93]. pour [Lem97, LBP98, ZSG5 ]. Power [Koe92b, Krä07, Was02, Bra93, HJOP03, JL94, Koe92a, Rya99]. Powerful [C^{+}84a, CGGG83]. powers [KZ91, Pak90]. pp [Ano95b]. PPP [ČPV90]. Practical [BM94a, Red95, Red96b, Ano96f, GR92, PW95, Wes99]. practicas [BLM97]. practice [CvGV92]. Prague [dN93]. Praktiker [EHR^{+}96]. Praktiker/Mathcad [EHR^{+}96]. Precalculus [MS94a, Bau96]. precise [Bia00]. precision [BB92c]. Preliminary [C^{+}87, Sch99a, CcMMK96, ES92]. Prelle [DDdMS02]. Prepackaged [Fos91]. préparatoires [AL98]. prepas [Ram98]. preprocessing [AV11]. presence [IS00]. Presented [ACM82, TRC94]. presents [CVG94]. preserving [HK89]. Presto [Tal94]. price [CT96]. Pricing [Ano05, Pri00]. Primality [Yan95]. Prime [Rie94, Sch01b]. Primer [Gar97]. Principal [Har97]. principle [CCM02]. Principles [ACM82]. Print [Ano00b]. Pro [Ano00b]. Probability [Gan03, KT95, KT99, Sto03, Bia00, CDC^{+}02, CKO02, KOC02]. Problem [BCG^{+}05, Bro95, CDK94, CvGV92, GDS94a, SMW01, Ano97b, Bay94, CM01, CJ90, Gar99a, GDS94b, KD90, Kwo90, LP05, Ryb98, SLC07, Shi09, Zac96]. Problem-Solving [GDS94a, GDS94b]. Probleme [Vet95]. Problèmes [Dur98, AL98, LM98a, LM98b, MNP97]. Problems [AB02, Ano95f, C^{+}87, EP96b, EP04a, GH93, GH95, GH97, GC^{+}88, GMM^{+}89, Hoh01, Pet01a, San96, Ste96, Art98, CZL11, GH04, Gon88, IA93, IA94a, IA94b, Ioa94, KRR10, KRR12, Mic90, SMF^{+}90, SMM^{+}90, SF90, SMFC92, VC89]. Procedure [LP94, VMRA14]. Procedures [Mur07b, CT94a, Fri97, FVGF98, FIBT01, FIT03, Fri09, GFF01, GF02, GSF05, IS00, IFF01, Ose96, PFG06, Wan95a]. Proceedings [ACM94, AHK93, Ano00a, Cha86, CvGV92, EM94, Fit93, Glo98, Glo91, GH84, Gon89, Gra94, IEE89, KH97, Lak96, Lee93, Lum94, MCG13, Nes93, ÖSZ94, SE11, SSKT91, SvADS07, Tan89, vdHvH12, BBST94, Bod01, BGR91, Bro93, CC92, Doo99, GW00, Gut04, IEE93b, JLL91, KT91, LW88, Lop94a, Mal01, S^{+}01, WN90, Wat91, dN93, Bro07, Car96, Gia89, HY14, Jef08, SW94a]. process [AB01]. Processes [Pet00b, Pet00c, Pet01a, Res92, VV13, Tal94]. Processing [Ano02a, IEE93a, Mol94, MS94b, Bod01, S^{+}01]. Product [Ano94c, Gar99b]. Products [Ano00b, Ano01, MP14, Bre88, JM98]. Program [BCG^{+}05, CM24, AZ91, AH97, BS03, Bis91, Bog16, BWB97, CB01, CD01, KSK96, Leo86, QBH06, Rab96, Rod92, Vân94d, Zei91]. programmation [MNP97]. Programmer [MP97]. Programmers [Cor95a, Cor02]. Programmieren [MD96]. Programming [ACM82, AB02, Ano96d, Bee17, BCD^{+}11, BCD^{+}12, Dub00, LMS07, LMRS08a, Mav93, MGH^{+}96, MGH^{+}98, MGH^{+}01, MGH^{+}02a, MGH^{+}02b, MGH^{+}03a, MGH^{+}03b, MGH^{+}05a, MGH^{+}05b, MGH^{+}07a, MGH^{+}07b, MGH^{+}08a, MGH^{+}08b, MGH^{+}09a, MGH^{+}09b, MGH^{+}10a, MGH^{+}10b, Vod82, Zie07, Ano97b, Ano03a, BR02, Bra93, GS12a, GWV93, Phi04, PC02, Ryb98, Sal93, Zac96]. Programs [CW92, FV94, KS12a, KS12b, GS98, Gro89, KS13, Ste94, SH10]. Progress [Ano00b, CM94a, Cha90a, KMG96, LP12]. Project [Roe94, Ano95d, ES10]. projective [GHL^{+}00]. Projects [BF96, EP96a, GH03, Kir95, Str99]. proof [Rei96, Sil04, Zei92]. Properties [tHB91, CT94b, YC07]. Prospects [LP12]. protocol [GKW94]. proves [AZ91]. Proving [ADG^{+}01, Wan95a, Zei91]. pseudogroups [LRB95]. Pseudoprimes [Arn95]. Pspice [LM98a, LM98b]. PSsolver [ADdM12]. PTOPO [KRTZ20]. Publication [CW05e]. pure [Wil89]. purpose [Cha90a]. PVMaple [Pet00b, Pet00c, Pet01b]. PVS [ADG^{+}01]. PWS [Ano95f]. Python [GS12a].

QPSI [FFB99]. QR [NS05]. Quadratec [Ano00b]. quadratic [GS12a]. quadtree [HF92]. quadtrees [LL95]. quality [Bee04]. quantities [GFF01]. Quantized [Har98a, Har00]. Quantum [Abe93, Ada94, Ano88, EHR^{+}96, Hor95a, Ste94, SH10, Wic95, CV88, FF88, Joy00, SMF^{+}90, SMM^{+}90, VC89, SMC97]. quartic [LM88]. Quasi [GKLM92, FFB99]. Quasi-periodic [GKLM92]. quasi-polynomial [FFB99]. Quasispin [Vin92]. quaternion [Li00a]. quaternions [Li00a]. query [CR94b]. questions [SW98]. queueing [FP95, PF95]. Quick [Ano00b, BM94b]. Quicksort [Hen89]. quintic [ES92]. quintics [SW94b]. quinze [Don97]. quivers [GS12a]. quotients [VHC17].

R [Ano96a, CT96]. Racah [Fri09]. radiation [Wil89]. RAEEM [LL04a]. railway [RLHAL11]. Rainey [Lum94]. raisonnée [CT95, CT97]. Ramanujan [Bor89, Fee90, Gar95, Sil04]. random [TT94]. Range [Pie93]. rappels [LM98a, LM98b]. rate [FS92]. rates [BB92a]. RATH [bLpL02]. Rational [AGR94, BS93, GK91, Rav94, AGR95, Bar95, DM86, Gro89, GR92, HJOP03, LSYL11, MP16, Oka92, PS95, van94c]. rationally [Vol93]. Real [tHB91, AY03, BG11, BB90, CDL^{+}11, DET07, SE22, YA07b]. RealCertify [ME18]. realizations [OP95]. Realized [Vei97, Die92]. Reals [HT93]. Reasoning [HT93, PW95]. Rebhan [vRFS^{+}00]. Recipes [Cor95b, EM07]. reciprocal [Pak90]. Recognition [Tar92]. reconstruction [AB05b]. Record [ACM82]. Recoupling [FIBT01]. rectangle [LW94]. rectangular [BLD96]. Recurrence [SVV95]. Recurrences [Zim95, AZ91]. recurring [LMJ93]. recursive [Kim97]. Redfern [Ano96e, Ano96f, Ano97c]. REDUCE [Ste94, Vul02, Vul03, Grä96]. Reduction [GHC92, Har98a, Har00]. Reellen [tHB91]. Reference [BM94b, C^{+}85a, C^{+}85b, CGG^{+}88b, C^{+}90, C^{+}91, CGG^{+}91a, CGG^{+}91b, Sch98a, Kof97]. Referenzhandbuch [Spi94]. reflection [Nag93]. Reform [Roy93]. regions [Fit89]. Regular [Car98, Car00]. RegularChains [BR14, CML^{+}08, CDL^{+}11, LMX05]. regularity [BB92a]. related [Bar92]. Relations [SVV95]. relationships [Bay05, Hed92, YB00b]. Relativitätstheorie [vRFS^{+}00]. relativity [Ano94a, Car90, CFFM97, CC98, M^{+}94, McL94b, PS97]. Release [Bra96, CT97, EHR^{+}96, EJLS94, Gar97, Kof94a, Kof94b, Red96a, Spi94, Tan99, Wat95f, Wat95j, Wat95l, Wat95h, Wat95k, Wat95d, Wat95g, Wat95i, Wat95e, Wat95a, Wat95b, Wat95c, Ano96e, Ano96f, B^{+}98, CKM98, Ell97a, RM97]. Releases [ABP02, Ano01]. Reliability [HDC94, Vân94d]. Remarks [JS00]. renewing [Gra93]. Rensselaer [Lop94a]. Repeated [BEEK94]. Report [Cha90b, Cha91, KMG96, Mad92, Cha90a]. reports [CVG94]. Representation [Wur96, Chi93b, Chi93b, Mic90]. Representations [Kal93, Roa96, Bre88, Die92, JKP03, JM98]. Represented [CM22]. Representing [MP14]. Reproduction [PM12]. Republic [Ska95]. Requires [Ano95f]. Rescheduling [BS02a]. Research [AHK93, EM94, Lav91, Par00]. researchers [BWB97]. résolus [AL98, Dur98, JLP99]. Resource [Wat02c, Wan06]. Respect [Wan94]. Response [CW05a]. Restore [Ano00b]. restrictions [JM98]. Result [Nes93]. resultant [LMJ93, WEM98, Min04]. resultant-based [WEM98]. resultants [Ose96]. résumés [MNP97]. retarded [QBH06]. Retreat [Mad92]. Review [Ano95b, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano97b, Ano97c, Ano98a, Ano98b, Ano99a, Ano99b, Ano03a, Ano04, Bag95, Bar92, Ber96, CT96, Gan03, Gre97, Ryb98, Sto03, SMC97, Was02, CVG94, PZ96, SMF^{+}90, Tan99, Wes95]. Reviews [Ano92b, Ano94c, Bra96, CW05c, CW05e, Mea96, Ste96, Wri96, CVG94]. Revised [CGG^{+}88a, HR92]. Revisiting [RLAH14]. rewriting [AB01, AB05b]. REX [McG84]. RHINE [Car96, Ano97d]. Ricci [Wil89]. Richard [Ano96c, Gre97]. Ridge [Ano02b]. Riemann [CM91, DP11, Liu19].

Ring [Pro00]. Rings [Koh94]. Rio [M^{+}94]. Risch [Bro91b, GS89]. Robertson [Koe91]. robot [KZ94]. Robotics [WWK92, MNEA95, Wan90]. robust [PG08]. Rogers [Sil04]. Rohr [EHR^{+}96]. Role [Mav93]. ROM [Ano98b, Ano99a, Ano99b, Ano03a]. Rome [Gia89]. Ronald [Ano96a]. Room [Dev93a]. root [GG12]. roots [BG11, GS12a]. Rosemont [Lum94]. Rostock [Glo98]. rotation [PFG06]. Rotational [LH93, Zir94]. Routine [Ron14]. Routines [Dub00, LS04]. RRtools [Sil04]. RSA [TT00]. Ruder [EHR^{+}96]. rule [FVGF98]. Rules [Wan94]. Runge [Cam06]. Russia [Mal01]. Russian [ZF91]. RWCA [Ano97d].

S [Ano96d, Koe91, Sto03]. SA [Ano00b]. SADE [FF11]. Sage [ES10, Ste13]. Samarkand [GMV00]. Sample [CGG83b, CGG84, CGG85, Ste94, SH10]. San [SE11]. Santander [Gut04]. Sasha [Gan03]. SAT [Bau94]. SAT-M0 [Bau94]. satisfied [AZ91]. SCAFI [CvGV92]. Scale [BH93]. Scales [San96, CR94a]. scans [Bra96]. scattering [BS99]. scenes [TT94]. Schenectady [GH84]. Schmidt-Rohr [EHR^{+}96]. School [LW88, M^{+}94, Sid94, BWB97]. Schritte [Bur94a]. Schrödinger [MG94]. Schroeder [vRFS^{+}00]. Schubert [Vei97]. Schur [CD01]. Scicos [CCN06]. Science [AHK93, Ano95a, Kol00, RW95, SvADS07, Ano98b, Bet98, Har14, Har15, ZJEP95]. Sciences [Ano95e, Bay94, Glo91, Lem97, Sco01]. Scientific [Cor95a, Cor95b, Dub00, GH93, GH95, GH97, GMV00, Hrě95, JS92, LP12, Ste96, Ano97b, Cor02, EM07, Fiu95, GH04, LP05, Lav91, Ryb98, SF90, SMFC92, Zac96]. scientifiques [ZSG5 ]. Scientists [EM97a, EM97c, Gre97, Rao02, BWB97, EM00, Ros96, Sch98b, Sch99c, TA96]. Scilab [CCN06]. Scilab/Scicos [CCN06]. scores [Fos01]. script [CS08]. Scripting [Ano01]. Search [Vil95, vRFS^{+}00, Koe93]. Second [KH97, Ano97a, CVG94, HK89, Mat89, VMRA14]. second-order [HK89, VMRA14]. sections [BS03]. Securities [Ano05, Pri00]. Security [Ano01]. see [Bar92]. Segmentation [MS94b]. Selected [HR92]. selectivity [CR94b]. Self [BCG^{+}05, BD93]. Self-consistent [BCG^{+}05]. Self-tutor [BD93]. Sellevag [vRFS^{+}00]. Semantics [Pin94]. Semigroups [Car97a, Car97b]. seminar [CvGV92]. semisimple [FJ93]. Sensitivity [CS93, WVG90]. Seoul [HY14]. Separated [Lie94, AGR95]. Separated-form [Lie94]. Separating [BLPR13]. September [Ano00a, Ano02a, BBST94, Bod01, Mal01]. Sequence [Roy93, SSTW94]. sequences [CB01, Gro91, LMJ93]. serial [GS98]. Series [Ano95f, CB02, Gru93, Yan93, BP92, BB92c, BG90, Bro91a, Déc93, Dom03, Gar99b, Gau99, Gro89, Gro91, HJOP03, Koe92a, Koe92b, Koo91, LL04a, Mat90, McL94a, Pak90, PN98, Rya99, Déc93, Déc93, IA92, Rod92]. Server [Ano00b, FKM95]. Servers [Ano00b]. Session [CGG83b, CGG84, CGG85]. sessions [Con97]. Set [Car97a, Car97b, Wan95b, Was04]. sets [Koe93, PG06, PG08]. Seven [MS94b, Wic95]. Seventh [KN94]. Several [Arn95, CV88, CVW93]. SFA [Pro00]. SFR [Ano04, Ano97a, Ano97b, Ano97c, Ano98a, Ano98b, Ano99a, Ano99b]. shallow [KBG01]. Share [GK98, GK98]. Share-Library [GK98]. Shear [Car90]. Shear-free [Car90]. Sheer [Gro96]. Shell [Ano00b, GSF05]. shift [NS05]. Shifting [NS05]. Ships [Ano01]. short [AR06]. Shpink [Ano00b]. sic [Kim97]. SIGCSE [Ano95a]. SIGMOD [SW94a]. Signal [IEE93a, Mol94]. SIGSAM [Gon89]. Simon [Doo99]. Simple [Ada94, CV88, Dom03, Lah92, Yan94].

Simplification [Zim95, HSW94, HSW98, McG84]. simplifies [Vol93]. Simulation [Ano01, CCN06, Cap93, JL94, Tal94]. Simulationen [JKM04]. simulations [Xu95]. Singer [DDdMS02]. Single [Ano95f, GSF05]. single-shell [GSF05]. singular [IA92]. singularities [Tar92, Vol93]. singularity [QBH06]. SIS [BD87]. Skeptic [HT98]. Skew [Faj19]. skewness [BS05b]. skin [SC94]. skin-effect [SC94]. Slater [PM12]. Slater-Type [PM12]. sliding [LS96, SRRB91]. sliding-mode [SRRB91]. small [Sad96]. smooth [Alu91]. SMT [SE22]. soap [Opr00]. Software [AV02, ABP02, Ano92b, Ano00b, Ano01, Bar92, Bee17, Bro95, CW92, Com96, EHR^{+}96, ES10, Fos06, GU93, Glo93, Hen94, HY14, SMW00b, Sim92, TRC94, Was04, Wat95h, Wat95i, Wes95, vRFS^{+}00, Ano96k, BFS93, CM01, CVG94, FH04, KRR12, Lav91, Ros96, Tal94, Ano00b, Ano01]. Software2Go [Ano00b]. Solar [DS94b]. Solid [Ano88, EHR^{+}96]. Solid-State [Ano88, EHR^{+}96]. solitary [bLpL02, pLbL03, LL04b]. soliton [FOW87, GI09]. Solution [AB02, CTE95, CJ90, CES94, DS94b, Hoh01, JKP03, Pet00a, BS99, CZL11, IA92, Ioa94, Rod92, SW90]. Solutions [Ano00b, Aue90, GC^{+}88, GMM^{+}89, Izu14, San96, dv96, AZ96, ABF14, ADdM12, BCEW12, BC94b, CDS95, CDL^{+}11, Die92, FH04, FM89, GI09, GKaM01, Ikl94, Li00b, bLpL02, LL04a, LSYL11, pLbL03, LL04b, MNP97, Rya99, SMF^{+}90, SMM^{+}90, XYXZ16, jZzFbL10]. Solvability [CG95]. solvable [SW94b]. solve [ADdM12, Bar92, Grä96]. solver [SE22, WEM98]. Solvers [Ber96, PZ96]. Solving [Ano95f, Bar96, B^{+}98, BCG^{+}05, Bro95, CDK94, GH93, GH95, GH97, GH04, GDS94a, MG94, Pet01a, SMW01, Sid94, Ste96, YC92, AY03, Ano97b, Bay94, CvGV92, DU97, DET07, GDS94b, HJR^{+}03, LP05, Roa94, Ryb98, SLC07, Shi09, VC89, YA07b, Zac96]. Some [BKR^{+}91, tHB91, Con92, GWV93, JKP96, Mic90, SF90, SMFC92, Wan94, AZ96, ACS91, McL94a, VC89, WVG90]. son [DG97]. Sons [Ano96a]. Soon [CW05e]. sortable [Zei92]. source [Ste08, Ste13]. Sources [TRC94]. South [Ano03b, HY14]. Sozopol [Tod07]. SP [Vei97]. Space [Mur07b, SMW00c, BV93, Car90, CM91]. Space-Time [Mur07b]. space-times [Car90]. spaces [Bog16]. spacetimes [Vul03]. Spain [Gut04]. Sparse [MP10, GG12, PM07]. SPE [MNP97]. Special [Mon96, Ron14, BC09]. species [ALL04, Chi93a, Chi93b]. Specification [KS12a, KS12b]. Spectral [Boy93, DJ89]. Speech [IEE93a]. Speed [FP89]. Spherical [CTE95, DJ89, IFF01]. spherically [VČ88]. Spiess [EHR^{+}96]. spin [AFK^{+}94, FF88, GSF05]. Spin-angular [GSF05]. spinor [CMC92]. Splice [Ano96k]. Spline [PQ23, Sei92, Sei93]. Sprache [Vet95]. Spreadsheets [CVG94]. Springer [Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano97b, Ano97c, Ano98a, Ano98b, Ano99a, Ano99b, Ano03a, Ano04, CT96, Ryb98]. Springer-Verlag [Ano96c, Ano96d, Ano96e, Ano97a, Ano97b, Ano97c, Ano99a, Ano99b, CT96]. square [Ber97b]. square-free [Ber97b]. Squares [PM12]. Stability [BC94a, GW20, GKLM92]. stack [Zei92]. Staff [Ful94]. Standard [GFF01]. Stanford [Wou86]. Stanley [vRFS^{+}00]. Stars [Mur07a]. Started [Map03a, Map04, Map05a, Map07a, Map08a, Put94, Wat01a, Wat02a, CKM98, CKM04, MMCK09]. State [Ano88, EHR^{+}96, WY93, Har94]. stationary [Pet94]. Statistical [AB02]. Statistics [RAB03, CB02, CS08, KT95, KT99, Sny95]. steady [CDS95]. Stehle [EHR^{+}96]. Steps [Bur94b]. stereogram [TT94]. STO [PM12]. Stochastic [Gan03, Res92, SZ94b, Sto03, VV13, Xu95, Cyg96, CGK01, CKO02, KOC02]. Stokes [HT85]. Stop [BS02b, BS02c, BS05a]. Storm [Ano00b]. Stormix [Ano00b]. story [Ken01]. Strain [Zir94]. Strange [BB92c]. STREED [Yua97]. Streifzüge [EHR^{+}96]. Strong [Arn95, EHR^{+}96]. structural [IOP21]. structure [GSF06, LRB95, MP12]. Structures [FS95, MP14, FP89, Fri97]. Stuck [Ano96a]. Student [EL89, May09, Tan99, Wat95f, Wat95j, Wat95l, Wat95h, Wat95k, Wat95d, Wat95g, Wat95i, Wat95e, Wat95a, Wat95b, Wat95c]. Students [Dur19, Roe94, BWB97]. studies [BF02, BF09, BWB97]. Study [BB92b, DHM95b, Roe94, CA07, KD90, Li00a, Sal93, YC07]. Styling [Ste25]. subdifferentials [BV97]. Subroutines [BFKS93]. subspaces [IS00]. substructures [KSK96]. subsums [ENS92]. Subway [ARL20]. Successful [Kun02]. Suche [vRFS^{+}00]. Sufficiency [Cor94]. Suitable [Com96]. Suite [Ano01]. Sum [FVGF98, MP14]. Sum-of-Products [MP14]. Summation [Ano95c, Str94a, ACGL04, PS95]. Summations [Wan94]. Summer [HA94, Lee93, Lop94a]. Sums [Le01, Pak90]. SUP [MNP97]. supersymmetries [AH97]. Supplement [Mal06, WG83]. support [DS94a]. Surf [PMS20]. Surface [GKLM91, Sha94, GKLM92]. Surfaces [DP11, FKM95, Ano98a, KK97, PMS20, Rov00a, Rov00b]. Surfing [PMS20]. Surprising [Ber93]. Switzerland [Lak96]. Symbolic [ACM94, AAFMJ01, Ano96g, Bac96, BCG^{+}05, BB92b, Ber96, Bro93, BKR^{+}91, Bro07, Car97a, Cha86, CV88, CVW93, Cor95b, Doo99, Fit89, Gar95, Glo98, Gon89, GDS94a, GDS94b, Gut04, HR92, HT85, Ioa94, Jef08, KH97, KMG96, Ken01, KZ94, Lak96, LP12, LMS07, LMRS08a, LS96, Mav93, MSW00, MCG13, Ngu93, Ngu94, Pie93, Roa94, SE11, SC94, Sim92, TTL93, TTLM94, TMH94, VAS06, Var95, Wan92, WWK92, Was02, WN90, Wat91, qXbL04, dJ91, dv96, vdHvH12, ALP96, ACGL04, AFK^{+}94, Ano03a, BD87, Bee05, BCLA83, CC92, CGG83a, Cha90a, CC98, Fit93, Fiu95, GF92, Gro90, Har14, Har15, LP05, Leo86, Lev95, Mat90, PC02, Rya99, SMF^{+}90, SMM^{+}90, SF90, Sie93, TP94, Van94a, Vân94d, VČ88, VC89, Vin91, Wan90, Xu95, de 93, de 95, Chi93b]. Symbolic [Gia89, KRR12, Tay98a]. Symbolic-Numeric [BCG^{+}05, GDS94a, GDS94b, Roa94, GF92]. Symbolic-Numerical [MSW00]. SymbolicC [SH10]. symbolique [Chi93b, KL96, Lem97]. Symbols [Lai94]. Symmetric [Pro00, Ste95, Koe94, VČ88]. Symmetries [LW88, FFB99, HK89, LRB95, VMRA14, VJC12]. Symmetry [LLLL94, vRFS^{+}00, FF11, Nag93, Vin92]. Symmetry-breaking [LLLL94]. Symposium [ACM82, ACM94, Ano88, Ano95a, BBST94, BGR91, Bro93, Bro07, Cha86, Doo99, Fit93, GW00, Gia89, Glo98, Gon89, Gut04, Jef08, KH97, KT91, Lak96, Lee93, Lop94a, MCG13, ÖSZ94, SE11, Wan92, WN90, Wat91, vdHvH12, CC91, Lev95]. Symsac [Cha86]. SyNRAC [AY03, YA07a, YA07b]. Synthesis [Shi05b, SRRB91, Shi05a]. Synthesizing [Tsa93]. System [AV02, Ano01, C^{+}84a, CFG^{+}84, C^{+}84b, HH95, JT91, JS92, KMG96, Mur07b, Pro00, Sch99a, Sch99b, Sch01a, SVV95, TTL93, TTLM94, TMH94, Vei97, Vod82, Wan92, BD87, BSS03, Bar93, Ber93, CHF94, CGGG83, CGG83a, Cha90a, CSS94, Die92, FP95, Grä96, JL94, LHJ92, Li00b, Mat89, OP95, PF95, RLHAL11, Sny95, Van94a, YZ07, jZzFbL06]. Systematic [KK91]. système [KL96, RM97]. Systemen [Ben98]. Systems [Ada94, Ano00a, Atz94, Bag95, Cap93, CS93, CES94, DST88, DST93, DPV14, HWH91, HR92, Mur94, Sch00a, SZ94b, WY93, BPS12, Ben98, BLPR13, CM11, CDL^{+}11, CSS94, DU97, DET07, FP89, Fat90, Fit93, Fit89, GK91, GLHR19, GWV93, GKW03, GKaM01, HSW94, HSW98, IA93, JKP96, KK91, Les94, Lie94, Lyn01, Lyn09, Ose96, Rod92, SRRB91, TRC94, TH94, Van94a, WGV90, WVG90, Wes99, YB00a, Yua97, dJ91]. systolic [Wou86]. syzygies [Stu89].

T [Ano95b, Ano96b]. Tables [WG83]. takes [CVG94]. taking [RLHAL11]. talk [Ada97, SMW00b]. Tallyman [Ano00b]. tangent [Alu91]. Tango [Ano00b]. Tape [CW05d]. Task [BS02a]. Tasks [SKB01a]. Taylor [IA92]. Taylor-series [IA92]. TDDS [GLHR19]. Teach [Mon94, Mat89, Mat90, PF95]. teachers [BWB97]. Teaching [FP95, Ful94, TH94]. Technical [Ano95a, Bra96, MAP87]. Technik [EHR^{+}96, Hör95b]. technique [Ben89]. Techniques [Pav94, Ano96j, GDS94b, Hol96, Sei92]. Technologies [Ano00b, Mal01, Ano01]. Technology [Lum93, Lum94, TRC94, Gra93, SH98, TH94, Kne94]. Telescoper [Le01]. Telescoping [ACGL04]. Template [BFKS93]. Template-Driven [BFKS93]. templates [VAS06]. Tensor [Bre88, KMG96, CM91, JM98, Liu19]. Terms [Le01, BP92, QBH06]. ternary [Bro91a]. Test [Mon92, Yan93, CS08, GH94, KBG01, qXbL04]. Testing [DG99, Yan95]. tetrad [CC98, Van88]. Textbook [Ano95f]. Texts [Sch93b]. Thailand [GW00]. Their [DK95, EZ23, Ben98, BGR91, Dev93a, GS12a, IS00, Wu92]. Them [DPV14, Mon94]. thèmes [Don97]. Theorem [ADG^{+}01, Rei96, VČ88, Wan95a]. Theorems [BHC95]. theoretic [Ano96j, Hol96]. Theoretical [Bay94, TP94]. Theoretische [vRFS^{+}00, JKM04]. Theories [Gar95]. Theory [Ada94, Ano88, Dur19, Koh94, ÖSZ94, Str94b, TW06, Bet01, Bry91, CCM91, CC91, Col05, Cos00, FP95, HSW98, Joy00, KD90, Mic90, PF95, Pos13, Wur96]. Thermal [vRFS^{+}00]. Thermodynamics [Tay98b, Tay98a, Lev92]. Third [Ano04, BGR91, GMV00, Top99]. Thomas [ABF14, GLHR19]. Three [CVG94, Pie93, CHF94, Rab96, VHC17]. Three-Body [Pie93]. three-dimensional [Rab96]. three-wave [CHF94]. threefold [ES92]. TikZ [Ste25]. Time [Ano00b, Ful94, Mur07b, SZ03, WY93, CR94a]. times [Car90]. tissues [YC07]. Titchmarsh [BEEK94]. Titchmarsh-Weyl [BEEK94]. Tits [GS12a]. Tokyo [WN90]. Tolerance [SKB01a, SKB01b]. Tolerating [BS02b, BS02c, BS05a]. Tome [Dou96, KL96]. Too [Com96]. Tool [Ano01, BF91, Krä07, Mur94, BWB97]. Toolbox [DB15, Ano90, YA07b]. Toolkit [Ano01, MB98]. Tools [CW92, Fos11, Was04]. topology [KRTZ20, Tan89]. torational [LSYL11]. tori [Nag93]. Torque [CP93]. torsion [Vul02, Vul03]. torus [KU03]. total [BS03]. tour [SH98]. Toys [Fos11]. traces [MNP97]. trajectory [SZ03]. transcendentals [KR03]. transfer [GKLM92]. transformation [Gro89, Vol93]. Transformations [Zim95, GF02, MWCY14, PN98, WW91, WyW93]. transforming [Gro91]. Transforms [DB15, Kau93]. Transition [WY93]. translation [GS12a]. Transmission [Hed92]. Transport [DS94b]. Trapani [BGR91]. Traveling [GS12b, LL04a]. Travelling [Li00b, bLpL02]. tree [RW92]. trees [Cam06]. trends [Fat90]. Triangular [Kal93, CM11]. triexcited [ČPV90]. Trigonometric [GHC92, Qua18, McL94a]. trigonometry [Par97c]. Trio [DPV14]. Tropical [AV11]. Troy [Lop94a].

Truncation [BH93, Cam06]. truss [IA94b]. Trust [DPV14]. Turkey [EM94]. turns [Shi98]. Turtle [LM94]. tutor [BD93]. Tutorial [C^{+}85a, C^{+}85b, C^{+}85c, CFG^{+}86, CGG^{+}88a, C^{+}89, CGG^{+}89, CGG^{+}90, CGG^{+}92, EL89, Lop94b, Gar99b]. Tutorials [E^{+}92, EJLS96, Ell97a]. Twenty [Gra94]. Twenty-Fourth [Gra94]. twisted [ES92]. twistor [Bry91]. Two [Bry91, ČPRV87, Pie93, Ron14, CVG94, Fit89, GR92, Koe91, Li00b, SC09, Zei92]. Two-Body [Ron14]. two-dimensional [Li00b]. two-part [CVG94]. two-phase [Fit89]. Two-Point [ČPRV87]. two-stack-sortable [Zei92]. Type [PM12, Car90, CTE95, DVL89, FJ93, Gec94, MWCY14, Sil04]. Typesetting [Pin94].

u [vRFS^{+}00]. U.K [BBST94]. überall [EHR^{+}96]. Udine [CC91]. UK [S^{+}01]. Ukraine [Bro93]. ultradiscrete [GI09]. ultraspherical [Koe94]. Umbral [RW92]. UML [Ano01]. uncertain [PG08, PG10, Yua97]. uncertainties [DU97]. undergraduate [CGG^{+}86, Run97]. Understanding [Tho17]. Unendlichen [vRFS^{+}00]. Unendlichen/Rebhan [vRFS^{+}00]. Unerkannte [vRFS^{+}00]. Unifying [ES10, RSL^{+}26]. Unipotent [GP92]. uniqueness [CDS95, Kwo90]. unitarizable [Bre88]. United [ACM94]. Univariate [Mon92, GG12]. University [Ano03b, Bro07, Doo99, Glo98, Gut04, Kne94, Lee93, Lum94, Tan89, Wou86, BWB97, FLC93, TR94]. Unix [RM97]. Unix/X [RM97]. Unix/X-Window [RM97]. unstable [YB00a]. Untergang [EHR^{+}96]. Untergang/Verhulst [EHR^{+}96]. up-date [Fri09]. Updates [Ano01]. Upgrade [Was04, Wat95f, Wat95l, Wat95g]. uravneniĭ [ZF91]. USA [Ano02b, Ano03b, MCG13, SE11]. usage [CT95, CT97, DG97, HF92]. Use [Cap93, CVW93, Dur19, de 93, de 95, Ben98, LM91a, LM91b, VC89]. User [C^{+}85a, C^{+}85b, GGC81a, GGC81b, GGC83, GH84, Map05b, Map07b, Map08b, Map09, Map10, Map11, Map12, Sch98a, WG83, LHJ92, Leo86, Was04]. user-friendly [Was04]. uses [BWB97]. Using [Abe93, AB94d, Bee17, Bil94, Bla5 , BR14, CS93, CES94, CM94b, Dav88, DG99, DS94b, DPV14, EP94, EG97, GH93, GH95, GKLM91, GHC92, JT91, KMG96, Kei01, MG94, MNEA95, Mat89, Mat90, May09, Mol94, PF95, Put94, PM12, Rhe93, Roy93, Sch5 , SSTW94, Ste96, SC09, SMC97, Wan90, WWK92, Wil98, AAFMJ01, Ano96b, Ano96j, Ano97b, Ano03a, BF02, BF09, BS03, Bay94, Bay05, Ber93, BR02, BS99, BS05b, Bra93, BDS96, Bug95, BD93, CB01, CDF92, CR94b, CB02, Col05, Cos00, DM05, DG95, FH04, Fee90, Fit89, GH97, GH04, GU93, Gro89, GKaM01, Har98a, Har00, Her01, Hol96, Hor95a, JM98, Jud96, Kim97, KBG01, LP05, LM91a, LTFN01, Lyn01, Lyn09, MS94a, Möl07, MBC92, OP95, OS95, Phi04, PC02, RLHAL11, RLAH14]. using [Ryb98, SIE98, Shi05a, Shi05b, SF10, Sie93, SH98, Ste94, SH10, Stu89, Sza02, TP94, Tsa93, LM91b, VČ88, VJC12, Vul02, Vul03, Xu95, qXbL04, YB00a, YB00b, Zac96]. Utilities [Red95, Red96b, Ano96f]. Uzbekistan [GMV00].

V [Ano95b, Ano96c, Ano96d, Ano96e, Ano96f, B^{+}98, Bra96, CT97, EHR^{+}96, EJLS94, Ell97b, Glo93, Kra97, Lop94a, Mea96, RM97, Spi94, Tan99, Wri96, Abe93, AB94a, AB94b, AB94c, AB99, AB00, Ano92a, Ano94b, Ano96i, Ano96j, Ano05, Art98, Atz94, Bar92, Bar96, Bay94, Bay96, Bia00, BM94a, BM94b, BS99, BBRM95, BDS96, Bro97b, Bro97c, Bro95, CL95, CGG^{+}90, CGG^{+}91a, CGG^{+}91b, CGG^{+}92, CH94, CP93, Cor96a, CT01, DG98, Dev93b, E^{+}92, Ell96, EJLS96, Ell97a, Ell97b, Fat92, Fat96, FMM94, FIBT01, Gar97, Har98a, Har00, HHR95, HHR96a, HHR96b, HHR98, Hol96, Joh93, Kof94a, Kof94b, KSK96, Lee93, Lem97, LBP98, MG94, MS94a, McL93, MB98, Mol94, MGH^{+}96, MD96, MP97, MGH^{+}98, OP95, Pet92, Pri00, Put94, Ram98, Red96a, Shi05a, Shi05b]. V [Shi98, TTL93, Wat96, Wat93, Wat95f, Wat95j, Wat95l, Wat95h, Wat95k, Wat95d, Wat95g, Wat95i, Wat95e, Wat95a, Wat95b, Wat95c, Wer93a, Wer93b, Wes95, Zir94, Ano99a, Ano99b]. V-Matrices [FMM94]. V.2 [LEGS93, LM94]. v.3.0 [Ano00b]. V.4 [AL98]. V0T0M [Bau94]. v1.4 [Ano00b]. v4.0.2 [Ano00b]. valid [Jef93]. Value [EP96b, EP04a, Hoh01, Pet01a, Art98]. Valued [Asl96, AH97, RLAH14]. values [FH04, Gec94]. Vancouver [Doo99]. Variable [Ano95f, SZ94a]. Variables [Wan94, VHC17]. Variance [BH93, BS05b]. Variational [Dav88, MS94b, Hel10]. variations [Cor94, LL95]. Varieties [Kal93, CLO92, CLO07, Rab96]. variety [GHL^{+}00]. VarInt [Hel10]. Various [Sch00a]. Varying [WY93]. Vassallo [CVG94]. vector [Ano98b, Bet98, CJ97, Wou86]. vectorial [BLM97]. vectorised [GS98]. Vedra [ÖSZ94]. velocity [Ikl94]. Verhulst [EHR^{+}96]. Verification [KS12b, Mav93, Nes93, KS13, Yan94]. Verified [DU97]. verifying [GI09]. Verlag [Ano96c, Ano96d, Ano96e, Ano97a, Ano97b, Ano97c, Ano99a, Ano99b, CT96]. versatile [Was04]. Version [Ano00b, C^{+}87, CGG^{+}88a, LC90, Sch99a, Sym88, Tan99, Bar95, BWB97, Bro95, ES92, Ell96, Pfl97]. versus [Sch00b, Sch01a]. vertex [AF92]. VI [GF02]. Via [Roe94, AF05, AB05b, ACS91, ADdM14, DB15, GKaM01, JKP98, JKP03, Lop94b, SRRB91]. Viable [ES10, Ste08, Ste13, de 93, de 95]. vibration [BLD96]. vibrational [VAS06]. VII [FIT03]. viii [Ano95b, GSF05]. Viosoft [Ano01]. Virasoro [Bre88]. Virtual [BG11]. Virtues [LMS07]. Visual [HPS05, IEE89]. Visualization [FKM95, Hrě95, Ska95, WA96]. Visualizing [Kei01, BWB97]. Vlasov [CTE95, ECT93]. Vogel [EHR^{+}96]. Vorkoetter [Ano96d]. Vorstoß [vRFS^{+}00].

Waals [Lev92, LH93]. Walter [Ste96]. Wannier [FMM94]. Washington [Gra93]. water [KBG01]. Waterloo [Ano00b, Bro07, Cha86, Ano01, FLC93, Tho94]. Wave [Dav88, CHF94, GS12b, HT85, Li00b, bLpL02, LL04a, pLbL03, LL04b]. wavelet [Kau93]. Wavelets [Kai94]. waves [GKLM92]. Way [Isr96]. ways [Boy15]. Web [Bra96, Bra96, Kun02, Man02]. Web-Based [Kun02]. WebEQ [Kun02]. Weierstrass [van95]. Weimar [Top99]. Weizsäcker [vRFS^{+}00]. Werner [EHR^{+}96, vRFS^{+}00]. West [Zei92]. Weyl [BEEK94]. Which [Arn95]. who [BWB97]. Wiemers [EHR^{+}96]. Wigner [Lai94, PFG06]. Wiley [Ano96a]. William [Lum94]. wind [CDS95]. Window [RM97]. Windows [Bro97c, Bar92, RM97]. Wissenschaftliche [EHR^{+}96]. without [CW05d, ENS92]. Wolfram [Bra96]. Words [EZ23]. Work [Pet00b, Pet00c]. Workbook [AB94d, C^{+}87, GC^{+}88, GMM^{+}89]. Worker [Sch00b, Sch01a]. Working [Pet01b, Pin93, TTLM94]. Worksheets [May09, Mon94]. Workshop [AHK93, Ano97d, Ano00a, Ano02a, EM94, GMV00, Lee93, Lop94a, Car96, DS94a, Wou86, Cha90b, Cha91]. Workstation [Ano00b]. World [Gra94, HR92]. Worlds [Ano00b]. Worskshop [IEE89]. worthy [Ano90]. Written [TTL93, TTLM94]. Wu [LL04b]. Wunner [EHR^{+}96]. Würzburg [SSKT91]. WZ [Ges95, Tef02].

X-Window [RM97]. XII [Sid94, WVG90]. XIII [GWV93].

Yau [Rab96]. Year [Roe94]. York [Ano96a, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano97b, Ano97c, Ano98a, Ano98b, Ano99a, Ano99b, Ano03a, Ano04, Lop94a, GH84].

Z [WW91]. Zachary [Ano97b, Ryb98]. Zee [vRFS^{+}00]. Zeilberger [Koo93]. zero [Bog16, dJ91]. zeros [FS92]. Zeuthen [KT91]. zum [GK98]. zur [tHB91, Mon93, Ger92]. Zurich [Lak96].

References

Alkahby:2001:SCA

[AAFMJ01] H. Alkahby, G. Ansong, P. Frempong-Mireku, and A. Jalbout. Symbolic computation of Appell polynomials using Maple. In Proceedings of the 16th Conference on Applied Mathematics (Edmond, OK, 2001), volume 7 of Electron. J. Differ. Equ. Conf., pages 1–13 (electronic). Southwest Texas State Univ., San Marcos, TX, 2001.

Abell:1994:DEM

[AB94a] Martha L. Abell and James P.

Braselton. Differential Equations with Maple V. Academic Press, New York, NY, USA, 1994. ISBN 0-12-041548-8. xiv + 684 pp. LCCN QA371.5.D37A23 1994. US$44.95.

Abell:1994:MVE

[AB94b] Martha L. Abell and James P.

Braselton. Maple V by Example. Academic Press, New York, NY, USA, 1994. ISBN 0-12-041545-3. xii + 500 pp. LCCN QA76.95 .A212 1994. US$39.95.

Abell:1994:MVH

[AB94c] Martha L. Abell and James P. Braselton. The Maple V Handbook. Academic Press, New York, NY, USA, 1994. ISBN 0-12-041542-9. LCCN QA76.95.A213 1994. US$39.95.

Argabright:1994:CWU

[AB94d] L. N. Argabright and R. C. Busby. Calculus Workbook Using Maple. Kendall/Hunt Publishing Company, Dubuque, IA, USA, 1994. ISBN 0-7872-0081-6. ???? pp.

Abell:1999:MVE

[AB99] Martha L. Abell and James P. Braselton. Maple V by example. Academic Press, New York, NY, USA, second edition, 1999. ISBN 0-12-041558-5 (book), 0-12-041559-3 (CD-ROM). xii + 644 pp. LCCN QA76.95 .A212 1999. US$.

Abell:2000:DEM

[AB00] Martha L. Abell and James P. Braselton. Differential equations with Maple V. Academic Press, New York, NY, USA, second edition, 2000. ISBN 0-12-041560-7. xii + 719 pp. LCCN QA371.5.D37 A33 2000.

Armando:2001:MEP

[AB01] Alessandro Armando and Clemens Ballarin. Maple’s evaluation process as constraint contextual rewriting. In Proceedings of the 2001 International Symposium on Symbolic and Algebraic Computation, pages 32–37 (electronic). ACM Press, New York, NY 10036, USA, 2001.

Aladjev:2002:IMP

[AB02] V. Z. Aladjev and M. A.

Bogdevicius. Interactive Maple: Programming, Solution of the Mathematical, Statistical and Engineering-Physical Problems. International Academy of Noosphere and Vilnius Technical University, Tallinn-Vilnius, Estonia, second edition, 2002. ISBN 9985-9277-1-0. ???? pp. LCCN ???? URL http:

//www.geocities.com/intl_
academy_noosphere/Information_

News.html. Includes CD-ROM.

Abell:2005:ME

[AB05a] Martha L. Abell and James P.

Braselton. Maple by example. Elsevier Academic Press, Amsterdam, The Netherlands, third edition, 2005. ISBN 0-12-088526-3. xi + 550 pp. LCCN QA76.95 .A2122 2005. Includes CD-ROM.

Armando:2005:REM

[AB05b] Alessandro Armando and Clemens Ballarin. A reconstruction and extension of Maple’s assume facility via constraint contextual rewriting. Journal of Symbolic Computation, 39(5):503–521, May 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Abe:1993:IQM

[Abe93] Y. Abe. Introductory quantum mechanics using Maple V. In Lee [Lee93], pages 3– 19. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Amore:2014:ACS

[ABF14] Paolo Amore, John P. Boyd, and Francisco M. Fernández. Accurate calculation of the solutions to the Thomas–Fermi equations. Applied Mathematics and Computation, 232(??): 929–943, April 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300314001829.

Ablamowicz:1996:CAC

[Abł96]  Rafał Abłamowicz.  Clif-

ford algebra computations with Maple. In Clifford (geometric) algebras (Banff, AB, 1995), pages 463–501. Birkhäuser Boston Inc., Cambridge, MA, USA, 1996.

Aladjev:2002:NSM

[ABP02] V. Aladjev, M. Bogdevicius, and O. Prentkovskis. New Software for Mathematical Package Maple of Releases 6, 7 and 8. International Academy of Noosphere and Vilnius Technical University, Tallinn-Vilnius, Estonia, 2002. ISBN 9985-9277-4-5, 9986-05-565-2. 404 (est.) pp. LCCN ???? US$. URL http:

//www.geocities.com/intl_
academy_noosphere/Information_
News.html.

Abramov:2004:TCS

[ACGL04] S. A. Abramov, J. J. Carette, K. O. Geddes, and H. Q. Le. Telescoping in the context of symbolic summation in Maple. Journal of Symbolic Computation, 38(4):1303–1326, 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

ACM:1982:CRN

[ACM82] Conference Record of the Ninth Annual ACM Symposium on Principles of Programming Languages: Papers Presented at the Symposium, Albuquerque, New Mexico, January 25–27, 1982. ACM Press, New York, NY 10036, USA, January 1982. ISBN 0-89791-065-6. LCCN QA76.7 .A15 1982. US$28.00.

ACM:1994:IPI

[ACM94] ACM, editor. ISSAC ’94: Proceedings of the 1994 International Symposium on Symbolic and Algebraic Computation: July 20–22, 1994, Oxford, England, United Kingdom. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

Augot:1991:MDS

[ACS91] D. Augot, P. Charpin, and N. Sendrier. The minimum distance of some binary codes via the Newton’s identities. In Cohen and Charpin [CC91], pages 65–73. ISBN 3-540-54303-1 (Berlin), 0-387-54303-1 (New York). LCCN QA268 .E95 1990.

Adams:1994:AAS

[Ada94] Barry G. Adams. Algebraic Approach to Simple Quantum Systems with Applications to Perturbation Theory. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57801-3. xvi + 451 pp. LCCN QC174.17.P45 A33 1994.

Adams:1997:MT

[Ada97]  Steven G. Adams.  Maple

talk. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-237306-8. xx + 236 pp. LCCN QA76.95.A28 1997.

Avellar:2012:PMI

[ADdM12] J. Avellar, L. G. S. Duarte, and L. A. C. P. da Mota. PSsolver: a Maple implementation to solve first order ordinary differential equations with Liouvillian solutions. Computer Physics Communications, 183 (10):2313–2314, October 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465512001518.

Avellar:2014:MPF

[ADdM14] J. Avellar, L. G. S. Duarte, and L. A. C. P. da Mota. A Maple package to find first order differential invariants of 2ODEs via a Darboux approach. Computer Physics Communications, 185(1):307–316, January 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465513002762.

Adams:2001:CAM

[ADG^{+}01] Andrew Adams, Martin Dunstan, Hanne Gottliebsen, Tom Kelsey, Ursula Martin, and Sam Owre. Computer algebra meets automated theorem proving: Integrating Maple and PVS. Lecture Notes in Computer Science, 2152:27–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
2152/21520027.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
2152/21520027.pdf.

Avis:1992:PAC

[AF92] David Avis and Komei Fukuda.

A pivoting algorithm for convex hulls and vertex enumeration of arrangements and polyhedra. Discrete and Computational Geometry, 8(3):295– 313, 1992. CODEN DCGEER. ISSN 0179-5376 (print), 1432-0444 (electronic). ACM Symposium on Computational Geometry (North Conway, NH, 1991).

Ablamowicz:2005:CGH

[AF05] Rafal Ablamowicz and Bertfried Fauser. Clifford and Graßmann Hopf algebras via the BIGEBRA package for Maple. Computer Physics Communications, 170(2):115– 130, August 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465505002602.

Allen:1994:CAS

[AFK^{+}94] S. Allen, G. J. Fee, A. T. Kachura, F. W. Letniowski, and R. G. McLenaghan. Comparison of algorithms for the symbolic computation of the NP spin coefficients and curvature components. General relativity and gravitation, 26(1):21– 40, 1994. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic).

Alonso:1994:FMP

[AGR94] C. Alonso, J. Gutierrez, and T. Recio. FRAC: a Maple package for computing in the rational function field K(X). In Lopez [Lop94a], pages 107–115. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Alonso:1995:RFD

[AGR95] Cesar Alonso, Jaime Gutierrez, and Tomas Recio. A rational function decomposition algorithm by near-separated polynomials. Journal of Symbolic Computation, 19(6):527– 544, June 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Anton:1992:CAG

[AH92] Howard Anton and Albert Herr.

Calculus with analytic geometry. Wiley, New York, NY, USA, fourth edition, 1992. ISBN 0-471-50901-9. various pp. LCCN QA303 .A53 1992.

Adams:1996:AC

[AH96] S. Adams and B. Horn. Anton’s Calculus. Wiley, New York, NY, USA, 1996. ISBN ???? ???? pp. LCCN ????

Ayari:1997:GMP

[AH97] M. A. Ayari and V. Hussin.

GLie: a MAPLE program for Lie supersymmetries of Grassmann-valued differential equations. Computer Physics Communications, 100(1-2):157– 176, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Angeles:1993:CKP

[AHK93] J. Angeles, G. Hommel, and P. Kovacs, editors. Computational kinematics: Proceedings of the Workshop on Computational Kinematics, held at the International Conference and Research Center for Computer Science (IBFI), of Germany, from October 11 to October 15, 1993, volume 28 of Solid Mechanics and its Applications. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. ISBN 0-7923-2585-0. LCCN QA841.C59 1993.

Artino:1994:ECM

[AKEJ94] C. Q. Artino, J. R. Kolod, B. Evans, and J. Johnson. Exploring Calculus with Maple. Wiley, New York, NY, USA, 1994. ISBN 0-471-09703-9. ???? pp.

Akritas:1989:ECA

[Akr89] Alkiviadis G. Akritas. Elements of computer algebra with applications. A Wiley-Interscience Publication. Wiley, New York, NY, USA, 1989. ISBN 0-471-61163-8. xvi + 425 pp. LCCN QA155.7.E4A37 1989.

Alame:1998:MVC

[AL98] Ibrahim Alame and Patrick Le Quere. Maple V.4: cours et problèmes résolus classes préparatoires et 1er cycle. Eyrolles, Paris, France, 1998. ISBN 2-212-08988-0. 323 pp. LCCN ????

Auger:2002:CME

[ALL04] Pierre Auger, Gilbert Labelle, and Pierre Leroux. Computing the molecular expansion of species with the Maple package Devmol. Sém. Lothar. Combin., 49:Art. B49z, 34 pp. (electronic), 2002/04.

Ablamowicz:1996:CAN

[ALP96] Rafaél Ablamowicz, Pertti Lounesto, and Josep M. Parra, editors. Clifford algebras with numeric and symbolic computations. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1996. ISBN 0-8176-3907-1, 3-7643-3907-1. xvii + 322 pp. LCCN QC20.7.C55C565 1996. US$59.50.

Aluffi:1991:HMS

[Alu91] Paolo Aluffi. How many smooth plane cubics with given jinvariant are tangent to 8 lines in general position? In Kleiman and Thorup [KT91], pages 15– 29. ISBN 0-8218-5131-4. LCCN QA564 .Z48 1989.

Andersson:1997:AMM

[And97]  Goran Andersson.  Applied

mathematics with Maple. Chartwell-Bratt, Bromley, Kent, UK, 1997. ISBN 91-44-00149-5 (paperback), 0-86238-490-7 (Chartwell-Bratt). 448 pp. LCCN ????

Anderson:2004:MPJ

[And04] I. M. Anderson. Maple packages and Java applets. London Mathematical Society Lecture Note Series, 312:193–206, 2004. CODEN ???? ISSN 0076-0552.

Anonymous:1988:ISQ

[Ano88] Anonymous, editor. International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods. Wiley, New York, NY, USA, 1988. CODEN IJQSDI. ISSN 0161-3642.

Anonymous:1990:MMT

[Ano90] Anonymous. Maple: $695 math toolbox is worthy competitor for Mathematica. PC Magazine, pages 48–49, February 27, 1990. CODEN PCMGEP. ISSN 0888-8507 (print), 1078-8085 (electronic).

Anonymous:1992:MV

[Ano92a] Anonymous. Maple V. Macworld, 9(5):190, May 1992. CO-DEN MACWEA. ISSN 0741-8647.

Anonymous:1992:SR

[Ano92b] Anonymous. Software reviews. Mathematics and computer education, 26(3):337–??, Fall 1992. CODEN MCEDDA. ISSN 0730-8639.

Anonymous:1994:IOM

[Ano94a]  Anonymous.  Introduction:

overview of MAPLE applications to general relativity. In MacCallum et al. [M^{+}94], pages 267–269. ISBN 0-19-853646-1. LCCN QC173.6.B73 1989.

Anonymous:1994:MV

[Ano94b] Anonymous. Maple V 3.0. Macworld, 11(10):73–??, October 1994. CODEN MACWEA. ISSN 0741-8647.

Anonymous:1994:PR

[Ano94c] Anonymous. Product reviews. Computer, 27(7):97–??, July 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:STS

[Ano95a]  Anonymous, editor.  26th

SIGCSE Technical Symposium on Computer Science Education, volume 27(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:1995:BRMb

[Ano95b]  Anonymous.  Book review:

Mathematical Computation with Maple V: Ideas and Applications, edited by T. Lee, Birkhäuser, Boston, 1993. viii + 199 pp. $34.50/£39.00. ISBN 0-8176-3724-9. Biometrics, 51 (4):1588, December 1995. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.
jstor.org/stable/2533303.

Anonymous:1995:E

[Ano95c] Anonymous. Erratum: “Summation by parts for finite difference approximations for d/dx”. Journal of Computational Physics, 116(1): 192, January 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0021999185710194.

See [Str94a].

Anonymous:1995:EPB

[Ano95d]  Anonymous.  Experiments

in project based calculus: the Manhattan College Maple project. American Society for Engineering Education Annual Conference. Proceedings, 1:316– 320, 1995. CODEN ACOPDW. ISSN 0190-1052.

Anonymous:1995:MMS

[Ano95e] Anonymous. Maple in the mathematical sciences. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(3): 20–??, ???? 1995. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1995:PNS

[Ano95f] Anonymous. The PWS Notebook Series Electronic Data Files for Solving Textbook Problems: Maple Notebook for Dick/ Patton—Calculus of a Single Variable (Requires Maple for PC compatibles). PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1995. ISBN 0-534-93943-0. ???? pp.

Anonymous:1996:BRDc

[Ano96a]  Anonymous.  Book review:

Differential equations with MAPLE^{T M}: By Kevin R. Coombes. Brian R. Hunt, Ronald L. Lipsman, John E. Osborn, and Garrett J. Stuck. John Wiley & Sons, New York. (1996). 232 pages. $20.95. Computers and Mathematics with Applications, 31 (2):129, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122196900838.

Anonymous:1996:BREj

[Ano96b] Anonymous. Book review: Experiments in mathematics using Maple: By C. T. J. Dodson and E. A. Gonzalez. Springer, Berlin. (1995). 465 pages. DM 48.00. Computers and Mathematics with Applications, 31 (11):138, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/0898122196900425.

Anonymous:1996:BRMi

[Ano96c]  Anonymous.  Book review:

Maple V: Learning guide: By K. M. Heal, M. L. Hansen and K. M. Richard. Springer-Verlag, New York. (1996). 269 pages. $24.00. Computers and Mathematics with Applications, 32 (6):130, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122196902035.

Anonymous:1996:BRMj

[Ano96d]  Anonymous.  Book review:

Maple V: Programming guide: By M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn and S. Vorkoetter. Springer-Verlag, New York. (1996). 379 pages. $34.00. Computers and Mathematics with Applications, 32 (6):130, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122196902011.

Anonymous:1996:BRMl

[Ano96e]  Anonymous.  Book review:

The Maple handbook, Maple V, release 4: By Darren Redfern. Springer-Verlag, New York. (1996). 495 pages. $29.00. Computers and Mathematics with Applications, 32(6): 131, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122196902138.

Anonymous:1996:BRPg

[Ano96f] Anonymous. Book review: The practical approach^{T M} utilities for Maple^{T M}: Maple V, release 3: By Darren Redfern. Springer, New York. (1995). 312 pages. $69.00 (including diskette). Computers and Mathematics with Applications, 32 (6):130, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122196901996.

Anonymous:1996:ICS

[Ano96g] Anonymous, editor. IEE Colloquium on Symbolic Computation for Control (Digest No.1996/078). IEE, London, UK, 1996.

Anonymous:1996:M

[Ano96h]  Anonymous.  Maple.  In

Ross [Ros96], pages 877–898. ISBN 0-8493-2530-7. LCCN QA76.754 .H35 1996. URL

http://www.loc.gov/catdir/
enhancements/fy0646/95006378-
d.html.

Anonymous:1996:MV

[Ano96i] Anonymous. Maple V. Waterloo Maple Software, Waterloo, Ontario, Canada, student version, release 4. Power edition, Windows and Macintosh CD-ROM edition, 1996. ISBN 0-387-14232-0. 1 learning guide, 1 disk pp. LCCN ????

Anonymous:1996:MVP

[Ano96j] Anonymous. Maple V package for field theoretic Poisson brackets using object-oriented techniques. Computer Physics Communications, 96(2-3):209– 216, August 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:1996:SAN

[Ano96k] Anonymous. Splice: An analytical network analysis software. IEEE transactions on education, 39(3):394–398, August 1996. CODEN IEEDAB. ISSN 0018-9359 (print), 1557-9638 (electronic).

Anonymous:1997:BRIc

[Ano97a]  Anonymous.  Book review:

Introduction to Maple, (second edition): By André Heck. Springer-Verlag, New York. (1996). 699 pages. DM 68.00; öS 496.40; sFr 60.00. Computers and Mathematics with Applications, 33(5):128, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0898122197829318.

Anonymous:1997:BRIe

[Ano97b] Anonymous. Book review: Introduction to scientific programming: Computational problem solving using Maple and C: By Joseph Zachary. Springer-Verlag, New York. (1996). 380 pages. DM 78.00; öS 569.40; sFr 69.00 (diskette included). Computers and Mathematics with Applications, 33 (5):131, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122197829446.

Anonymous:1997:BRMi

[Ano97c]  Anonymous.  Book review:

The Maple(R) O.D.E. lab book: By Darren Redfern and Edgar Chandler. Springer-Verlag, New York. (1996). 160 pages. DM 44.00, öS 321.20, sFr 39.50 (diskette included). Computers and Mathematics with Applications, 33(7):137, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0898122197846020.

Anonymous:1997:CAR

[Ano97d] Anonymous. Computer algebra. 5th Rhine Workshop (RWCA). Theoretical Computer Science, 187(1–2):??, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1998:BRDc

[Ano98a] Anonymous. Book review: Discovering curves and surfaces with Maple(R): By Grazyna Klimek and Maciej Klimek. Springer, New York. (1997). 217 pages. $49.00, DM 86.00, öS 627.80, sFr 78.50, GBP 28.00 (diskette included). Computers and Mathematics with Applications, 35(5):135, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S089812219890743X.

Anonymous:1998:BRPi

[Ano98b] Anonymous. Book review: Partial differential equations for computational science: With Maple(R) and vector analysis: By David Betounes. Springer, New York. (1998). $64.95, sFr 116.50, DM 128.00, öS 935.00 (CD-ROM included). Computers and Mathematics with Applications, 36(8): 126, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122198911816.

Anonymous:1999:BRId

[Ano99a] Anonymous. Book review: Interactive linear algebra with Maple V: By Elias Deeba and Ananda Gunawardena. Springer-Verlag, New York. (1998). 316 pages. $49.95, DM 98.00, öS 716.00, sFr 89.50, GBP 37.50 (CD-ROM included). Computers and Mathematics with Applications, 37 (3):136, February 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122199904315.

Anonymous:1999:BRIi

[Ano99b] Anonymous. Book review: Interactive linear algebra with Maple V: By Elias Deeba and Ananda Gunawardena. Springer-Verlag, New York. (1998). 316 pages. $49.95, DM 98.00, öS 716.00, sFr 89.50, GBP 37.50 (CD-ROM included). Computers and Mathematics with Applications, 37 (6):136, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S089812219990295X.

Anonymous:2000:DPS

[Ano00a] Anonymous, editor. Distributed and Parallel Systems – From Concepts to Architectures: Proceedings of DAPSYS’2000, 3rd Austrian-Hungarian Workshop on Distributed and Parallel Systems, Balatonfűred, Lake Balaton, Hungary September 10th–13th, 2000. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN ???? LCCN ????

Anonymous:2000:NPAb

[Ano00b] Anonymous. New products:

Active Worlds Servers, Activeworlds.com, Inc.; Atipa Monolith Firewall, Atipa Linux Solutions; Hard Hat Net CompactPCI, MontaVista Software Inc.; EZTerm v1.4, CSV Technologies, Inc.; Maple 6, Waterloo Maple; Motif 2.1.20, Software2Go, LLC; Network Shell v.3.0, Shpink Software; PC300, Cyclades Corporation; ESP Print Pro v4.0.2, Easy Software Products; Progress Version 8.3, Progress Software Corporation; Time Navigator, Quadratec SA; Quick Restore 2.6, Workstation Solutions, Inc.; Storm Linux 2000, Stormix Technologies, Inc.; Tallyman, Akopia; Tango 2000 Application Server, Pervasive Software. Linux Journal, 72:??, April 2000. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2001:PVL

[Ano01] Anonymous. Products: Viosoft’s Linux embedded development environment; Popkin Software releases development modeling suite; Iopsis Software’s Forte for Java IDE; NQL releases scripting language components; Ascend Software updates delivery management system; Excel Software ships UML design tool; Hyperformix’s discreteevent simulation modeler; In-Cert’s application fault management software; BioconX releases biometric security software; Waterloo Maple’s mathematical computation software [Maple 7]; Equator Technologies’ application development toolkit. Computer, 34(8):84– 86, August 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co2001/pdf/r8084.
                                        pdf;   http://www.computer.
                                        org/computer/co2001/r8084abs.
                                        htm.

Anonymous:2002:DAH

[Ano02a] Anonymous, editor. Dapsys 2002: 4th Austrian-Hungarian Workshop on Distributed and Parallel Processing, Linz, Austria, September 29–October 2, 2002. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN ???? LCCN ????

Anonymous:2002:MIC

[Ano02b] Anonymous, editor. MathML International Conference: Hickory Ridge Conference Center, Chicago, IL, USA, June 28–30, 2000. ????, ????, 2002. ISBN ???? LCCN ????

Anonymous:2003:BRFf

[Ano03a]  Anonymous.  Book review:

Finite elements using Maple: a symbolic programming approach: By A. Portela and A. Charafi. Springer, New York. (2002). 325 pages. $49.95. (CD-ROM included). Computers and Mathematics with Applications, 46(2–3):515–516, July/August 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122103900666.

Anonymous:2003:ECC

[Ano03b] Anonymous, editor. East Coast Computer Algebra Day, April 5, 2003, Clemson University, Clemson, South Carolina, USA. ????, ????, 2003. ISBN ???? LCCN ????

Anonymous:2004:BRIa

[Ano04] Anonymous. Book review: Introduction to Maple: Third Edition. By André Heck. Springer, New York, NY. (2003). 828 pages. $49.95, EUR49.95, SFR 86.00, GBP 38.50. Computers and Mathematics with Applications, 47(6–7):1149–1150, March/April 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122104901034.

Anonymous:2005:PDS

[Ano05] Anonymous. Pricing Derivative Securities with Matlab 2e: An Interactive, Dynamic Environment with Maple V and Matlab. Academic Press, New York, NY, USA, 2005. ISBN 0-12-565651-3. ???? pp. LCCN ????

Aratyn:2006:SCM

[AR06] Henrik Aratyn and Constantin Rasinariu. A short course in mathematical methods with Maple. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2006. ISBN 981-256-595-7. xv + 700 pp. LCCN QA297 .A73 2006.

Almech:2020:AGD

[ARL20] Alberto Almech and Eugenio Roanes-Lozano. Automatic generation of diagrammatic subway maps for any date with Maple. Mathematics in Computer Science, 14(2):193– 207, June 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Arnault:1995:CCN

[Arn95] François Arnault. Constructing Carmichael numbers which are strong pseudoprimes to several bases. Journal of Symbolic Computation, 20(2):151–162 (or 151–161??), August 1995. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Articolo:1998:PDE

[Art98] George A. Articolo. Partial differential equations and boundary value problems with Maple V. Academic Press, New York, NY, USA, 1998. ISBN 0-12-064475-4. xii + 628 pp. LCCN QA377.A82 1998.

Aslaksen:1996:MVC

[Asl96] Helmer Aslaksen. Multiplevalued complex functions and computer algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(2): 12–20, June 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Aschenwald:1996:MMP

[AST96] D. Aschenwald, T. Siegl, and R. F. Tichy. MAPinsure — a MAPLE package for life insurance. Journal of Symbolic Computation, 22(2):227–234, August 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Adams:2004:IMM

[ASV04] P. (Peter) Adams, K. (Ken)

Smith, and Rudolf Vyborny. Introduction to mathematics with Maple. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2004. ISBN 981-238-931-8, 981-256-009-2 (paperback). xx + 523 pp. LCCN QA76.95 .A26 2004.

Atzema:1994:MVG

[Atz94] E. Atzema. Maple V and geometrical optics: Aberration coefficients in arbitrary optical systems. In Lopez [Lop94a], pages 209–218. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Auer:1990:MSM

[Aue90] John W. Auer. Maple Solutions Manual for Linear Algebra with Applications. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1990. ISBN 0-13-535550-8. vii + 186 pp. LCCN ????

Auer:1991:LAA

[Aue91] John W. Auer. Linear Algebra with Applications. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1991. ISBN 0-13-538349-8. xv + 548 pp. LCCN QA 184 A84 1991.

Aladjev:2002:CAS

[AV02] V. Z. Aladjev and V. A. Vaganov. Computer Algebra System Maple: a New Software Library. International Academy of Noosphere and Vilnius Technical University, Tallinn-Vilnius, Estonia, 2002. ISBN 9985-9277-5-3. ???? pp. LCCN ???? URL http:

                                        //www.geocities.com/intl_
                                        academy_noosphere/Information_

News.html. Includes CD-ROM.

Adrovic:2011:TAG

[AV11] Danko Adrovic and Jan Verschelde. Tropical algebraic geometry in Maple: a preprocessing algorithm for finding common factors for multivariate polynomials with approximate coefficients. Journal of Symbolic Computation, 46 (7):755–772, July 2011. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0747717110001410.

Anai:2003:SMP

[AY03] Hirokazu Anai and Hitoshi Yanami. SyNRAC: a Maplepackage for solving real algebraic constraints. In Computational science—ICCS 2003. Part I, volume 2657 of Lecture Notes in Comput. Sci., pages 828–837. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003.

Almkvist:1991:MPF

[AZ91] G. Almkvist and D. Zeilberger.

A MAPLE program that finds, and proves, recurrences and differential equations satisfied by hyperexponential definite integrals. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(3):14–17, July 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Abramov:1996:DSI

[AZ96] Sergei A. Abramov and Eugene V. Zima. D’Alembertian solutions of inhomogeneous linear equations (differential, difference, and some other). In Lakshman Y. N. [Lak96], pages 232–240. ISBN 0-89791-796-0. LCCN QA 76.95 I59

                                        1996.  URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/issac/236869/
                                        p232-abramov/.

Barrow:1998:SDE

[B^{+}98] David Barrow et al. Solving differential equations with Maple V, release 4. Brooks/Cole symbolic computation series. Brooks/Cole, Pacific Grove, CA, USA, 1998. ISBN 0-534-34555-7. viii + 253 pp. LCCN QA371.5.D37S65 1998.

Bacher:1996:AGO

[Bac96] Rainer Bacher. Automatic generation of optimization code based on symbolic non-linear domain formulation. In Lakshman Y. N. [Lak96], pages 283–291. ISBN 0-89791-796-0. LCCN QA 76.95 I59

                                        1996.  URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/issac/236869/
                                        p283-bacher/.

Baglivo:1995:BRB

[Bag95]  Jenny Baglivo.  Book re-

view: Computer Algebra Systems: Maple and Mathematica. The American Statistician, 49 (1):86–92, February 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

                                        URL http://www.jstor.org/
                                        stable/2684820.

Banks:2001:DEG

[Ban01] Bernard W. Banks. Differential equations with graphical and numerical methods. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-084376-8. xiii + 462 pp. LCCN QA371 .B26 2001.

Baran:1992:WYS

[Bar92] N. Baran. What you see is what you solve: Mathematica [2.0] and MathCAD [3.0] for Windows. (includes related article on Maple V) (software review). BYTE Magazine, 17:263–265, May 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barkatou:1993:ACC

[Bar93] M. A. Barkatou. An algorithm for computing a companion block diagonal form for a system of linear differential equations. Applicable algebra in engineering, communication and computing, 4(3):185– 195, 1993. CODEN AAECEW. ISSN 0938-1279 (print), 1432-0622 (electronic).

Barkatou:1995:RVM

[Bar95] A. Barkatou. A rational version of Moser’s algorithm. In Levelt [Lev95], pages 297–302. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/220346/
p297-barkatou/.

Barrow:1996:SOM

[Bar96] David Barrow. Solving ODEs with Maple V. Brooks/Cole, Pacific Grove, CA, USA, preliminary edition, 1996. ISBN 0-534-34402-X. viii + 154 pp. LCCN ????

Barakat:2001:JMP

[Bar01] Mohamed Barakat. Jets. A MAPLE-package for formal differential geometry. In Computer algebra in scientific computing (Konstanz, 2001), pages 1–12. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001.

Bauldry:1994:SIM

[Bau94] W. C. Bauldry. (SAT-M0) an introduction to Maple V0T0M on the PC. In Lum [Lum94], pages 107–113. ISBN 0-201-54304-4. LCCN QA11.A1I454 1992.

Bauldry:1996:PMM

[Bau96] W. C. Bauldry. PreCalculus Math and Maple. William C. Brown Company Publishers, Dubuque, IA, USA, 1996. ISBN ???? ???? pp. LCCN ????

Baylis:1994:TMP

[Bay94] William E. Baylis. Theoretical Methods in the Physical Sciences: an introduction to problem solving using Maple V. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 0-8176-3715-X. xvii + 286 pp. LCCN Q183.9 .B39 1994. URL

                                        http://www.birkhauser.com/
                                        Books/baylis/summary.htm.

Baylis:1996:PAC

[Bay96] W. E. Baylis. Pauli-algebra calculations in MAPLE V. In Clifford algebras with numeric and symbolic computations, pages 69–82. Birkhäuser Boston Inc., Cambridge, MA, USA, 1996.

Bayram:2005:DCR

[Bay05]  Mustafa Bayram.  Deriva-

tion of conservation relationships for catalytic cycles using MAPLE. Applied Mathematics and Computation, 160(1): 189–195, January 5, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Borwein:1990:DRN

[BB90] Jonathan Borwein and Peter Borwein. A dictionary of real numbers. Wadsworth and Brooks/Cole Publishing Company, Pacific Grove, CA, USA, 1990. ISBN 0-534-12840-8. viii + 424 pp. LCCN QA47.B625 1990.

Baillon:1992:ORA

[BB92a] Jean-Bernard Baillon and Ronald E.

Bruck. Optimal rates of asymptotic regularity for averaged nonexpansive mappings, pages 27–66. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1992.

Bergeron:1992:SMS

[BB92b] François Bergeron and Nantel Bergeron. Symbolic manipulation for the study of the descent algebra of finite Coxeter groups. Journal of Symbolic Computation, 14(2–3):127–140 (or 127–139??), August–September 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Borwein:1992:SSH

[BB92c] J. M. Borwein and P. B. Borwein. Strange series and high precision fraud. Amer. Math. Monthly, 99(7):622–640, 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Boggess:1995:CMV

[BBRM95] Albert Boggess, David Barrow, Maurice Rahe, and Jeff Morgan. CalcLabs with Maple V, First Edition. Brooks/Cole, Pacific Grove, CA, USA, 1995. ISBN 0-534-25590-6. xii + 229 pp. LCCN 303.5.D37C34 1995. US$17.95.

Blake:1994:BDI

[BBST94] J. R. Blake, J. M. Boulton-Stone, and N. H. Thomas, editors. Bubble dynamics and interface phenomena: proceedings of an IUTAM Symposium held in Birmingham, U.K., 6–9 September, 1993, volume 23 of Fluid Mech. Appl. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-3008-0. LCCN TA357.5.C38 I88 1994.

Belair:1994:SBE

[BC94a] Jacques Bélair and Sue Ann Campbell. Stability and bifurcations of equilibria in a multiple-delayed differential equation. SIAM J. of Applied Math., 54(5):1402–1424, 1994. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bragg:1994:AAS

[BC94b] L. R. Bragg and B. Cahlon.

Analytic approximation of solutions of differential equations. Complex variables, theory and application, 26(1-2):11–40, 1994. CODEN CVTADV. ISSN 0278-1077.

Backeljauw:2009:ACF

[BC09] Franky Backeljauw and Annie Cuyt. Algorithm 895: a continued fractions package for special functions. ACM Transactions on Mathematical Software, 36 (3):15:1–15:20, July 2009. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bernardin:2011:MPG

[BCD^{+}11] L. Bernardin, P. Chin, P. De-Marco, D. E. G. Hare, K. M. Heal, G. Labahn, J. P. May, J. McCarron, M. B. Monagan, D. Ohashi, and S. M. Vorkoetter. Maple Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2011. ISBN 1-926902-08-4. xxvi + 678 pp. LCCN ????

Bernardin:2012:MPG

[BCD^{+}12] L. Bernardin, P. Chin, P. De-Marco, K. O. Geddes, D. E. G. Hare, K. M. Heal, G. Labahn, J. P. May, J. McCarron, M. B. Monagan, D. Ohashi, and S. M. Vorkoetter. Maple Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2012. ISBN 1-926902-25-4. xxviii + 676 pp. LCCN ????

Barkatou:2012:CCF

[BCEW12] M. A. Barkatou, T. Cluzeau, C. El Bacha, and J.-A. Weil. Computing closed form solutions of integrable connections. In van der Hoeven and van Hoeij [vdHvH12], pages 43–50. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Belyaeva:2005:MSN

[BCG^{+}05] I. N. Belyaeva, N. A. Chekanov, A. A. Gusev, V. A. Rostovtsev, Yu. A. Ukolov, Y. Uwano, and S. I. Vinitsky. A MAPLE symbolic-numeric program for solving the 2Deigenvalue problem by a selfconsistent basis method. Lecture Notes in Computer Science, 3718:32–39, 2005. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/content/pdf/
                                        10.1007/11555964_3.pdf.

Buchberger:1983:CAS

[BCLA83] B. Buchberger, G. E. Collins, R. Loos, and R. Albrecht. Computer algebra: symbolic and algebraic computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1983. ISBN 0-387-81776-X. vi + 283 pp. LCCN QA155.7.E4 C65 1983. US$24.50.

Bandyopadhyay:1987:SSI

[BD87] S. Bandyopadhyay and J. S. Devitt. SIS: a symbolic information management system. Journal of Symbolic Computation, 4(3):397–408, December 1987. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Burbulla:1993:STC

[BD93] Dietrich C. M. Burbulla and C. T. J. Dodson. Self-tutor for Computer Calculus using Maple. Prentice-Hall Canada, Scarborough, Ontario, Canada, 1993. ISBN 0-13-063926-5. xiv + 292 pp. LCCN ????

Bradford:2013:CAD

[BDE^{+}13] Russell Bradford, James H.

Davenport, Matthew England, Scott McCallum, and David Wilson. Cylindrical algebraic decompositions for boolean combinations. In Monagan et al. [MCG13], pages 125–132. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Braselton:1996:CLU

[BDS96] James P. Braselton, John W.

Davenport, and Arthur G. Sparks. Calculus labs using Maple V. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-673-99814-2. ???? pp. LCCN QA303.5.C65B73 1996.

Beebe:2004:MPQ

[Bee04] Nelson H. F. Beebe. Maple plot quality and PostScript conformance. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 11, 2004. 21 pp. URL

https://www.math.utah.edu/
~beebe/reports/maple-plot-

bugs.pdf. This report is updated frequently.

Beebe:2005:MSAa

[Bee05] Nelson H. F. Beebe. Maple (symbolic algebra) FAQ. World-Wide Web frequently-asked question document., Septem-

ber 2005. URL https://www.
math.utah.edu/faq/maple/.

This report is updated frequently.

Beebe:2017:MFC

[Bee17] Nelson H. F. Beebe. The Mathematical-Function Computation Handbook: Programming Using the MathCW Portable Software Library. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-64109-3 (hardcover), 3-319-64110-7 (ebook). xxxvi + 1114 pp. LCCN QA75.5-76.95. URL http:/

/www.springer.com/us/book/
9783319641096.

Brown:1994:RDN

[BEEK94] B. M. Brown, M. S. P. Eastham, W. D. Evans, and V. G. Kirby. Repeated diagonalization and the numerical computation of the Titchmarsh-Weyl m(λ)-function. Proc. Roy. Soc. London Ser. A, 445(1923):113– 126, 1994.

Bauldry:1995:LAM

[BEJ95] William C. Bauldry, Benny Evans, and Jerry Johnson. Linear Algebra with Maple. John Wiley, New York, NY, USA, 1995. ISBN 0-471-06368-1. xiii + 275 pp. LCCN QA185.D37B38 1995. US$20.95.

Beltzer:1995:EAM

[Bel95] Abraham I. Beltzer. Engineering analysis with Maple/ Mathematica. Academic Press, New York, NY, USA, 1995. ISBN 0-12-085570-4. xi + 282 pp. LCCN QA76.95 .B45 1995.

Benson:1989:ACT

[Ben89] Max Environments. Benson. An algebraic computing technique. In Tangora [Tan89], pages 63– 77. ISBN 0-8247-8031-0. LCCN QA448.D38 C67 1989.

Benker:1998:ICS

[Ben98] Hans Benker. Ingenieurmathematik mit Computeralgebra-Systemen: AXIOM, DE-RIVE, MACSYMA, MAPLE, MATHCAD, MATHEMAT-ICA, MATHLAP und MuPAD in der Anwendung. (German) [Engineering mathematics with computer algebra systems: AX-IOM, DERIVE, MACSYMA, MAPLE, MATHCAD, MATH-EMATICA, MATHLAP and MuPAD in their use]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1998. ISBN 3-528-05673-8. xiii + 439 pp. LCCN ????

Bergeron:1993:SMU

[Ber93] François Bergeron. Surprising mathematics using a computer algebra system. Journal of Symbolic Computation, 15(3):365–370, 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Bertrand:1994:INA

[Ber94] Laurent Bertrand. On the implementation of a new algorithm for the computation of hyperelliptic integrals. In ACM [ACM94], pages 211–215. ISBN 0-89791-638-7. LCCN QA76.95.I59

1994.  URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/190347/
p211-bertrand/.

Bernardin:1996:RSS

[Ber96] L. Bernardin. A review of symbolic solvers. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(1):9–20, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bernardin:1997:MMP

[Ber97a]  Laurent Bernardin.  Maple

on a massively parallel, distributed memory machine. In Kaltofen and Hitz [KH97], pages 217–222. ISBN 0-89791-951-3. LCCN QA76.58 .I5674 1997. URL http://www.

acm.org:80/pubs/citations/
proceedings/issac/266670/
p217-bernardin/.

Bernardin:1997:SFF

[Ber97b] Laurent Bernardin. On squarefree factorization of multivariate polynomials over a finite field. Theoretical Computer Science, 187(1–2):105– 116, November 15, 1997. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/
browse.cgi?year=1997&volume=
187&issue=1-2&aid=2595.

Betounes:1998:PDE

[Bet98] David Betounes. Partial differential equations for computational science: with Maple and vector analysis. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1998. ISBN 0-387-98300-7 (hardcover). xix + 517 pp. LCCN QA377.B53 1998.

Betounes:2001:DET

[Bet01] David Betounes. Differential equations: theory and applications with Maple. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 2001. ISBN 0-387-95140-7 (cloth). xiv + 680 pp. LCCN QA371.5.D37 B47 2001. With 1 CD-ROM (Windows, Macintosh and UNIX).

Bauldry:1991:CLM

[BF91] William C. Bauldry and Joseph R. Fiedler. Calculus Laboratories with Maple: a Tool, Not an Oracle. Brooks/Cole, Pacific Grove, CA, USA, 1991. ISBN 0-534-13788-1. xii + 144 pp. LCCN MLCM 90/01887 (Q).

Bauldry:1996:CPM

[BF96] W. C. Bauldry and Joseph R. Fiedler. Calculus Projects with Maple. Brooks/Cole, Pacific Grove, CA, USA, second edition, 1996. ISBN 0-534-23748-7. xx + 233 pp. LCCN QA303.5.D37B38 1996.

Barnes:2002:MMC

[BF02] Belinda Barnes and Glenn Fulford. Mathematical modelling with case studies: a differential equation approach using Maple. Taylor and Francis, New York, 2002. ISBN 0-415-29803-2 (hb), 0-415-29804-0 (pb). xiv + 428 pp. LCCN QA371.3 .B37 2002. Includes bibliographical references (p. 415-419) and index.

Barnes:2009:MMC

[BF09] Belinda Barnes and Glenn Fulford. Mathematical modelling with case studies: a differential equations approach using Maple and MATLAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2009. ISBN 1-4200-8348-1 (hardcover). ix + 349 pp. LCCN QA371.3 .B37 2009. URL http:

//www.loc.gov/catdir/toc/
fy0904/2008049372.html.

Bentley:1993:TDI

[BFKS93] Jon L. Bentley, Mary F.

Fernandez, Brian W. Kernighan, and Norman L. Schryer. Template-driven interfaces for numerical subroutines. ACM Transactions on Mathematical Software, 19(3):265–287, September 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/citations/journals/toms/
1993-19-3/p265-bentley/.

Beneke:1993:CSP

[BFS93] T. W. Beneke, S. Fingerling, and W. W. Schwippert. Computer software packages ‘Maple’, ‘Mathcad’ and ‘Mathematica’ compared. Elektronik, 42(13):50, 63–68, June 1993. CODEN EKRKAR. ISSN 0013-5658.

Brak:1990:AAN

[BG90] R. Brak and A. J. Guttmann.

Algebraic approximants: a new method of series analysis. J. Phys. A, 23(24):L1331–L1337, 1990. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Bembe:2011:VRR

[BG11] Daniel Bembé and André Galligo. Virtual roots of a real polynomial and fractional derivatives. In Schost and Emiris [SE11], pages 27–34. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Brezinski:1991:OPT

[BGR91] Claude Brezinski, Laura Gori, and André Ronveaux, editors. Orthogonal polynomials and their applications: proceedings of the Third International Symposium on Orthogonal Polynomials and their Applications, Erice (Trapani), June 1–8, 1990, volume 9 of IMACS Annals on Computing and Applied Mathematics. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, 1991. LCCN QA404.5 .I55 1990.

Bauldry:1993:TVS

[BH93] W. C. Bauldry and J. L. Hebert. Truncation and variance in scale mixtures. In Lee [Lee93], pages 67–76. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Ballarin:1995:TAI

[BHC95] Clemens Ballarin, Karsten Homann, and Jacques Calmet. Theorems and algorithms: An interface between Isabelle and Maple. In Levelt [Lev95], pages 150–157. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/220346/
p150-ballarin/.

Biazzo:2000:AEP

[Bia00] Veronica Biazzo. Algorithms for the extension of precise and imprecise conditional probability assessments: an implementation with Maple V. Matematiche (Catania), 55(1):91–101 (2001), 2000. CODEN MTM-TAA.

Bilonick:1994:UMD

[Bil94] R. A. Bilonick. Using Maple to design a complex mixture experiment. In Lopez [Lop94a], pages 228–234. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Bisoi:1991:MPG

[Bis91] A. K. Bisoi. A Maple program to generate orthonormal polynomials. Computers and Mathematics with Applications, 22(9): 1–5, 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Brown:1994:BCI

[BK94] R. C. Brown and Man Kam Kwong. On the best constant for the inequality ∫_{0}^{∞} y^{′2} ≤ K(∫_{0}^{∞} |y|) max_{(0,∞)} |y′′|. In Inequalities and applications, volume 3 of World Sci. Ser. Appl. Anal., pages 103–121. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1994.

Broughan:1991:SSC

[BKR^{+}91] K. A. Broughan, G. Keady, T. D. Robb, M. G. Richardson, and M. C. Dewar. Some symbolic computing links to the NAG numeric library. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(3): 28–37, July 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Blachman:1995:UM

[Bla5 ] N. Blachman. Using Maple. Brooks/Cole, Pacific Grove, CA, USA, 1995 (??). ???? pp.

Bardell:1996:FPV

[BLD96] N. S. Bardell, R. S. Langley, and J. M. Dunsdon. On the free in-plane vibration of isotropic rectangular plates. Journal of Sound and Vibration, 191 (3):459–467, April 1996. CO-DEN JSVIAG. ISSN 0022-460x (print), 1095-8568 (electronic).

Barrio:1997:CVP

[BLM97] Roberto Barrio, Carlos Lopez, and Eduardo Martinez. Calculo vectorial: practicas con Mathematica y Maple, volume 54 of Textos Docentes. Prensas Universitarias de Zaragoza, Zaragoza, Spain, 1997. ISBN 84-7733-488-9. viii + 229 pp. LCCN ????

Li:2002:RMP

[bLpL02] Zhi bin Li and Yin ping Liu. RATH: a Maple package for finding travelling solitary wave solutions to nonlinear evolution equations. Computer Physics Communications, 148(2):256– 266, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Bouzidi:2013:SLF

[BLPR13] Yacine Bouzidi, Sylvain Lazard, Marc Pouget, and Fabrice Rouillier. Separating linear forms for bivariate systems. In Monagan et al. [MCG13], pages 117–124. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

Blachman:1994:MVP

[BM94a] Nancy Blachman and Michael J.

Mossinghoff. Maple V, a Practical Approach. Brooks/Cole, Pacific Grove, CA, USA, March 1994. ISBN ???? ???? pp. LCCN ????

Blachman:1994:MVQ

[BM94b] Nancy Blachman and Michael J.

Mossinghoff. Maple V Quick Reference. Brooks/Cole, Pacific Grove, CA, USA, 1994. ISBN 0-534-20478-3. xiii + 522 pp. LCCN QA76.95 .B52 1994.

Bernardin:2002:MM

[BMH02] Laurent Bernardin, James Mc-Carron, and Douglas Harder. MathML in Maple. In Anonymous [Ano02b], page ?? ISBN ???? LCCN ????

Bekker:1994:IEM

[BMW94] Paul A. Bekker, Arjen Merckens, and Tom J. Wansbeek. Identification, Equivalent Models, and Computer Algebra. Academic Press, New York, NY, USA, 1994. ISBN 0-12-084775-2. ix + 211 pp. LCCN QA276.8 .B45 1994.

Bode:2001:EPP

[Bod01] Arndt Bode, editor. Euro-Par 2000 parallel processing: 6th International Euro-Par Conference, Munich, Germany, August 1–September 1, 2000: proceedings, volume 1900 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-67956-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E865 2000; QA76.58 .I553 2000.

Bogner:2016:MPC

[Bog16] Christian Bogner. MPL — a program for computations with iterated integrals on moduli spaces of curves of genus zero. Computer Physics Communications, 203(??):339–353, June 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465516300595.

Borwein:1989:PER

[Bor89] Jonathan M. Borwein. Pi, Euler, Ramanujan, and MAPLE. Colloquium, Department of Computer Science, University of Manitoba, Winnipeg, MB, Canada., November 9, 1989.

Boyd:1993:CLS

[Boy93] John P. Boyd. Chebyshev and Legendre spectral methods in algebraic manipulation languages. Journal of Symbolic Computation, 16(4):377– 399, October 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Boyd:2015:FWC

[Boy15] John P. Boyd. Four ways to compute the inverse of the complete elliptic integral of the first kind. Computer Physics Communications, 196(??):13– 18, November 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465515001733.

Bergeron:1992:CGF

[BP92] François Bergeron and Simon Plouffe. Computing the generating function of a series given its first few terms. Experimental mathematics, 1(4):307–312, 1992. ISSN 1058-6458 (print), 1944-950X (electronic).

Barkatou:2012:IMP

[BPS12] Moulay A. Barkatou, Eckhard Pflügel, and Flavia Stan. ISOLDE: a Maple package for systems of linear functional equations. ACM Communications in Computer Algebra, 46 (3/4):157–159, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Betounes:2002:MCI

[BR02] David Betounes and Mylan Redfern. Mathematical computing: an introduction to programming using Maple. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 2002. ISBN 0-387-95331-0. xii + 412 pp. LCCN QA76.6 .B4775 2002. Includes CD-ROM.

Barakat:2009:CCC

[BR09] Mohamed Barakat and Daniel Robertz. conley: Computing connection matrices in Maple. Journal of Symbolic Computation, 44(5):540–557, May 2009. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Botana:2014:UMP

[BR14] Francisco Botana and Tomás Recio. Using Maple’s RegularChains library to automatically classify plane geometric loci. In Hong and Yap [HY14], pages 500–503. ISBN 3-662-44198-5 (paperback), 3-662-44199-3 (ebook). LCCN QA76.9.M35.

Brandt:1993:DAP

[Bra93] Jeffrey Martin Brandt. The design and analysis of power electronics circuits using the MAPLE programming language. Thesis (M.S.E.E.), University of Kentucky, Lexington, KY, USA, 1993. ix + 119 pp.

Bramley:1996:TNR

[Bra96] Randall Bramley. Technical news & reviews: Maple V Release 4; Mathematica 3.0 from Wolfram; HotGAMS from NIST; Informant scans Web and e-mails you; more on the web. . .. IEEE Computational Science & Engineering, 3(3):80, Fall 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). Discusses a Java-based front end for the Guide to Available Mathematical Software, with a pointer to

URL http://math.nist.gov/
HotGAMS/.

Bremner:1988:TPU

[Bre88] Murray R. Bremner. Tensor products of unitarizable representations of the Virasoro algebra with central charge \frac{1}{2}. Communications in Algebra, 16 (7):1513–1523, 1988. CODEN COALDM. ISSN 0092-7872 (print), 1532-4125 (electronic).

Broer:1991:HST

[Bro91a] Bram Broer. Hilbert series for ternary forms. In Computational aspects of Lie group representations and related topics (Amsterdam, 1990), volume 84 of CWI Tract, pages 1–18. Math. Centrum, Centrum Wisk. Inform., Amsterdam, The Netherlands, 1991.

Bronstein:1991:RDE

[Bro91b] Manuel Bronstein. The Risch differential equation on an algebraic curve. In Watt [Wat91], pages 241–246. ISBN 0-89791-437-6. LCCN QA 76.95

                                        I59 1991. URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/issac/120694/
                                        p241-bronstein/.

Bronstein:1993:IPI

[Bro93] Manuel Bronstein, editor. IS-SAC ’93, proceedings of the 1993 International Symposium on Symbolic and Algebraic Computation, July 6–8, 1993, Kiev, Ukraine, ISSAC — Proceedings. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-604-2. LCCN QA76.5.I613 1993. ACM order number: 505930.

Brown:1995:MSP

[Bro95] Allen Brown. Model software for problem solving: Maple V is the latest version of a modelling package capable of solving equations analytically, and much more. Electronics world + wireless world, 101(5):379–??, May 1995. CODEN EWWWE6. ISSN 0959-8332, 0266-3244.

Brooks:1997:IMC

[Bro97a] David R. Brooks. An introduction to MAPLE for college algebra. Brooks/Cole, Pacific Grove, CA, USA, 1997. ISBN 0-534-34757-6. xii + 180 pp. LCCN ????

BrooksCole:1997:MVM

[Bro97b] Brooks and Cole Publishing Company. Maple V (Macintosh edition). Brooks/Cole, Pacific Grove, CA, USA, student, release 4. edition, 1997. ISBN 0-534-34025-3 (disks) 0-534-34023-7 (container) 0-534-26202-3 (manual). 9 computer disks pp. LCCN QA76.95 1996.

BrooksCole:1997:MVW

[Bro97c] Brooks and Cole Publishing Company. Maple V (Windows edition). Brooks/Cole, Pacific Grove, CA, USA, student, release 4. edition, 1997. ISBN 0-534-34024-5 (disks), 0-534-34022-9 (container), 0-534-26202-3 (manual). 9 computer disks pp. LCCN QA76.95 1996.

Brown:2007:PIS

[Bro07] C. W. Brown, editor. Proceedings of the 2007 International Symposium on Symbolic and Algebraic Computation, July 29–August 1, 2007, University of Waterloo, Waterloo, Ontario, Canada. ACM Press, New York, NY 10036, USA, 2007. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Bryant:1991:TEH

[Bry91] Robert L. Bryant. Two exotic holonomies in dimension four, path geometries, and twistor theory. In Carlson et al. [CCM91], pages 33–88. ISBN 0-8218-1492-3. LCCN QA641 .C6143 1991.

Bronstein:1993:FPF

[BS93] Manuel Bronstein and Bruno Salvy. Full partial fraction decomposition of rational functions. In Bronstein [Bro93], pages 157–160. ISBN 0-89791-604-2. LCCN QA76.5.I613

                                        1993.  URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/issac/164081/
                                        p157-bronstein/. ACM order

number: 505930.

Borgert:1995:MP

[BS95] J. Borgert and H. Schwarze. Maple in der Physik. ????, ????, 1995. ISBN ???? ???? pp. LCCN ????

Blackett:1999:AFB

[BS99] Anthony J. Blackett and Andris T. Stelbovics. Analytic first Born atomic scattering: a computer algebra solution using Maple V. Computer Physics Communications, 116 (1):78–94, January 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465598001854.

Bosa:2002:TLR

[BS02a] Karoly Bosa and Wolfgang Schreiner. Task logging, rescheduling, and peer checking in Distributed Maple. Technical Report 02-10, RISC-Linz, Johannes Kepler University, Linz, Austria, March 2002. URL

                                        http://www.risc.uni-linz.

ac.at/people/schreine/papers/
faultrep2002.ps.gz.

Bosa:2002:TSFa

[BS02b] Karoly Bosa and Wolfgang Schreiner. Tolerating stop failures in Distributed Maple. In Anonymous [Ano02a], page ?? ISBN ???? LCCN ???? URL

http://www.risc.uni-linz.
ac.at/people/schreine/papers/
dapsys2002.ps.gz.

Bosa:2002:TSFb

[BS02c] Karoly Bosa and Wolfgang Schreiner. Tolerating stop failures in Distributed Maple. PDCP: Parallel and Distributed Computing Practices, ??(??): ??, ???? 2002. CODEN ???? ISSN 1097-2803. URL http:

//www.risc.uni-linz.ac.at/
people/schreine/papers/pdcp2003.

ps.gz. To appear.

Bartlett:2003:BTI

[BS03] Philip L. Bartlett and Andris T.

Stelbovics. Born total ionisation cross sections: An algebraic computing program using Maple. Computer Physics Communications, 154(2):159– 174, August 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465503002868.

Bosa:2005:TSF

[BS05a] Károly Bósa and Wolfgang Schreiner. Tolerating stop failures in Distributed Maple.

Scalable Computing: Practice and Experience, 6(2):59– 70, June 2005. CODEN ???? ISSN 1895-1767. URL

http://www.scpe.org/vols/
vol06/no2/SCPE_6_2_06.pdf;
 http://www.scpe.org/vols/
vol06/no2/SCPE_6_2_06.zip.

Bowman:2005:AVS

[BS05b] K. O. Bowman and L. R. Shenton. The asymptotic variance and skewness of maximum likelihood estimators using Maple. Journal of Statistical Computation and Simulation, 75(12):975–986, 2005. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Borwein:2011:IMM

[BS11] Jonathan M. Borwein and Matthew P. Skerritt. An introduction to modern mathematical computing: with Maple. Springer Undergraduate Texts in Mathematics and Technology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4614-0121-6. xvi + 216 pp. LCCN QA76.6 .B665 2011.

Baranovskii:2003:CAG

[BSS03] S. P. Baranovskiĭ, I. V. Shirokov, and S. V. Shmakov. Computation of the automorphism group of a finitedimensional algebra in the Maple system. In Mathematical structures and modeling. No.

12 (Russian), pages 113–119. Omsk. Gos. Univ., Omsk, 2003.

Bugl:1995:EDE

[Bug95] Paul Bugl. Explorations in Differential Equations using Maple. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-374752-2. 149 pp. LCCN ????

Burkhardt:1994:ESM

[Bur94a]  Werner Burkhardt.  Erste

Schritte mit Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-56649-X. ???? pp.

Burkhardt:1994:FSM

[Bur94b] Werner Burkhardt. First Steps in Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-19874-1 (New York), 3-540-19874-1 (Berlin). ix + 118 pp. LCCN QA76.95 .B86713 1994. US$19.95.

Brackx:1993:CPK

[BV93] F. Brackx and N. Van Acker. A conjugate Poisson kernel in Euclidean space. Simon Stevin, 67 (1-2):3–14, 1993. CODEN SS-WNAX. ISSN 0037-5454.

Bauschke:1997:FCS

[BV97] H. H. Bauschke and Martin Von Mohrenschildt. Fenchel conjugates and subdifferentials in Maple. Research Report CORR 97-23, Dept. of Combinatorics and Optimization, Faculty of Mathematics, University of Waterloo, Waterloo, ON, Canada, 1997. 14 pp.

Borwein:1997:MIM

[BWB97] Jonathan M. Borwein, Carolyn Watters, and Ephraim J. Borowski, editors. The Math-Resource interactive math dictionary: version 1.0. The math visualizing program Newton and Leibniz didn’t have. [An essential tool for high school, college and university students, teachers, scientists, engineers, researchers and anyone else who uses or studies mathematics]. MathResources, Inc., Halifax, NS, Canada, 1997. ISBN 3-540-14650-4. ???? pp. LCCN

????  URL http://www.
mathresources.com.

Char:1984:DMC

[C^{+}84a] Bruce W. Char et al. The design of Maple: a compact, portable, and powerful computer algebra system. Technical Report CS-83-06, Institute for Computer Research, University of Waterloo, Waterloo, Ontario, Canada, 1984. 13 pp. Work appeared in Computer Algebra proceedings of Eurocal’83-European computer Algebra Conference, London, England, March 1983, pp. 101– 115, Springer-Verlag, Berlin, 1983.

Char:1984:DPMb

[C^{+}84b] Bruce W. Char et al. On the design and performance of the Maple system. Technical Report CS-84-13, Faculty of Mathematics, University of Waterloo, Waterloo, Ontario, Canada, 1984. 35 pp.

Char:1985:MUGa

[C^{+}85a] Bruce W. Char et al. Maple User’s Guide: First Leaves, a Tutorial Introduction to Maple and Maple Reference Manual. WATCOM Publications Limited, 415 Philip St, Waterloo, Ontario N2L 3X2, Canada, fourth edition, 1985. ISBN 0-919884-51-2. xiii + 310 pp. LCCN QA76.8.M46 M36 1985.

Char:1985:MUGb

[C^{+}85b] Bruce W. Char et al. Maple User’s Guide: First Leaves, a Tutorial Introduction to Maple and Maple Reference Manual. WATCOM Publications Limited, 415 Philip St, Waterloo, Ontario N2L 3X2, Canada, fourth edition, 1985. ISBN 0-919884-01-6. xiii + 310 pp. LCCN QA155.7.E4 M3 1985.

Char:1985:TIM

[C^{+}85c] Bruce W. Char et al. A tutorial introduction to Maple. Technical Report CS-85-56, Faculty of Mathematics, University of Waterloo, Waterloo, Ontario, Canada, 1985. 27 pp.

Char:1987:MWC

[C^{+}87] Bruce W. Char et al. Maple workbook of calculus problems: Preliminary version. Technical Report CS-87-27, Faculty of Mathematics, University of Waterloo, Waterloo, Ontario, Canada, 1987. 45 pp.

Char:1989:FLM

[C^{+}89] Bruce W. Char et al. First Leaves for the Macintosh: a Tutorial Introduction to Maple. Brooks/Cole, Pacific Grove, CA, USA, 1989. ISBN 0-534-10222-0. xv + 140 pp. LCCN QA155.7.E4 F57 1989.

Char:1990:MRM

[C^{+}90] Bruce W. Char et al. MAPLE Reference Manual. Waterloo Maple Publishing, 160 Columbia St, Waterloo, Ontario N2L 3L3, Canada, fifth edition, 1990. ISBN 0-9694580-1-0. LCCN ????

Char:1991:MLR

[C^{+}91] Bruce W. Char et al. Maple language reference manual. Technical report, University of Waterloo, Waterloo, Ontario, Canada, 1991. ii + 153 pp. Prepublication manuscript.

Coombes:1996:DEM

[C^{+}96] Kevin Robert Coombes et al. Differential equations with Maple. Wiley, New York, NY, USA, 1996. ISBN 0-471-10875-8. v + 232 pp. LCCN QA371.5.D37D53 1996.

Corless:2007:JHF

[CA07] Robert M. Corless and Dawit Assefa. Jeffery–Hamel flow with Maple: a case study of integration of elliptic functions in a CAS. In Brown [Bro07], pages 108–115. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Cameron:2006:MPA

[Cam06] Frank Cameron. A Matlab package for automatically generating Runge–Kutta trees, order conditions, and truncation error coefficients. ACM Transactions on Mathematical Software, 32(2):274–298, June 2006. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Capolsini:1993:UMM

[Cap93] P. Capolsini. The use of Maple for multibody systems modeling and simulation. In Lee [Lee93], pages 109–117. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Carminati:1990:SFP

[Car90] J. Carminati. Shear-free perfect fluids in general relativity. II. aligned, Petrov type III space-times. J. Math. Phys., 31(10):2434–2440, 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carriere:1996:PRW

[Car96] L. R. Oudin A. Carrière, editor. Proceedings of the 5th RHINE workshop on computer algebra. Institut Franco-Allemand de Recherches de Saint-Louis, 68300 Saint-Louis, France, April 1996. ISBN ???? LCCN ????

Caron:1997:ASMa

[Car97a]  P. Caron.  AG: a set of

Maple packages for symbolic computing of automata and semigroups. Lecture Notes in Computer Science, 1214:879– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caron:1997:ASMb

[Car97b] Pascal Caron. AG: a set of Maple packages for manipulating automata and semigroups. Software—Practice and Experience, 27(8):863–884, August 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=7323;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=7323&PLACEBO=IE.pdf.

Caron:1998:LMP

[Car98] P. Caron. LANGAGE: a Maple package for automaton characterization of regular languages. Lecture Notes in Computer Science, 1436:46–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caron:2000:LMP

[Car00] Pascal Caron. LANGAGE: a Maple package for automaton characterization of regular languages. Theoretical Computer Science, 231(1):5–15, 2000. CO-DEN TCSDIQ. ISSN 0304-3975 (print), 1879-2294 (electronic). Implementing automata (London, ON, 1997).

Camargo-Brunetto:2001:IGL

[CB01] M. A. O. Camargo-Brunetto.

Investigating the growth of length of intermediate expressions in polynomial sequences using a Maple program. In Computer mathematics (Matsuyama, 2001), volume 9 of Lecture Notes Ser. Comput., pages 28–35. World Sci. Publishing, River Edge, NJ, 2001.

Childs:2002:SAM

[CB02] Aaron Childs and N. Balakrishnan. Series approximations for moments of order statistics using MAPLE. Computational Statistics & Data Analysis, 38(3):331–347, 2002. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Cohen:1991:EIS

[CC91] G. Cohen and P. Charpin, editors. Eurocode ’90: international symposium on coding theory and applications, Udine, Italy, November 5–9, 1990, volume 514 of Lecture Notes in Comput. Sci. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-540-54303-1 (Berlin), 0-387-54303-1 (New York). LCCN QA268 .E95 1990.

Calmet:1992:AIS

[CC92] Jacques Calmet and John A. Campbell, editors. Artificial intelligence and symbolic mathematical computing: International Conference AISMC-1, Karlsruhe, Germany, August 1992: proceedings, volume 737 of Lecture Notes in Comput. Sci. Springer-Verlag, Berlin, Germany, 1992. ISBN 0-387-57322-4. LCCN Q334 .I5488 1992.

Connell:1993:EIA

[CC93] Amanda E. Connell and Robert. M. Corless. An experimental interval arithmetic package in Maple. Interval Computations = Interval’nye vychisleniia, 2:120–134, 1993. ISSN 0135-4868.

Cyganowski:1998:MPN

[CC98] Sasha Cyganowski and John Carminati. The Maple package NPTOOLS; a symbolic algebra package for tetrad formalisms in general relativity. Computer Physics Communications, 115(2–3):200–214, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0010465598001635.

Carlson:1991:CGL

[CCM91] James A. Carlson, C. Herbert Clemens, and David R. Morrison, editors. Complex geometry and Lie theory, volume 53 of Proc. Sympos. Pure Math.

Amer. Math. Soc., Providence, RI, USA, 1991. ISBN 0-8218-1492-3. LCCN QA641 .C6143 1991.

Chu:2002:HPM

[CCM02] K. C. Chu, S. R. Czapor, and R. G. McLenaghan. Huygens’ principle and MAPLE’s NPspinor package. In Recent developments in general relativity, Genoa 2000, pages 151–163. Springer Italia, Milan, 2002.

Cheung:1996:EMC

[CcMMK96] C. K. Cheung, William G. Multivariable calculus McCallum, Tim. Murdoch, and G. E. Keough. Exploring multivariable calculus with Maple to accompany Multivariable calculus, preliminary edition, [by] William G. McCallum, Deborah Hughes-Hallett, Andrew M. Gleason. Wiley, New York, NY, USA, 1996. ISBN 0-471-13753-7 (paperback). 228 pp. LCCN ????

Campbell:2006:MSS

[CCN06] S. L. (Stephen La Vern) Campbell, Jean-Philippe Chancelier, and Ramine Nikoukhah. Modeling and Simulation in Scilab/Scicos. Springer Science+Business Media, New York, NY, USA, 2006. ISBN 0-387-27802-8 (hardcover). xi + 313 pp. LCCN Q183.9 .C36

2006. URL http://www.loc.
gov/catdir/enhancements/fy0663/
2005930797-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0814/2005930797-b.html;

                                         http://www.loc.gov/catdir/
                                        toc/fy0705/2005930797.html.

Carvalho:2001:MPC

[CD01] M. J. Carvalho and S. D’Agostino. A Maple program for calculations with Schur functions. Computer Physics Communications, 141(3):282–295, November 30, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465501004027.

Campos:2002:IPI

[CDC^{+}02] M. A. Campos, G. P. Dimuro, A. C. R. Costa, J. F. F. Araújo, and A. M. Dias. Interval probability and interval Markov chains in Maple. TEMA Tend. Mat. Apl. Comput., 3(2): 53–62, 2002. ISSN 1677-1966. Selecta of the XXIV National Congress of Applied and Computational Mathematics, Part 2 (Portuguese) (Belo Horizonte, 2001).

Carli:2014:MPI

[CDdM14] H. Carli, L. G. S. Duarte, and L. A. C. P. da Mota. A Maple package for improved global mapping forecast. Computer Physics Communications, 185(3):1115–1129, March 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S001046551300413X.

Carminati:1992:IDE

[CDF92] John Carminati, John S. Devitt, and Greg J. Fee. Isogroups of differential equations using algebraic computing. Journal of Symbolic Computation, 14 (1):103–120, 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Chan:1994:DAP

[CDK94] K. C. Chan, A. Diaz, and E. Kaltofen. A distributed approach to problem solving in Maple. In Lopez [Lop94a], pages 13–21. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Chen:2011:CRS

[CDL^{+}11] Changbo Chen, James H.

Davenport, François Lemaire, Marc Moreno Maza, Bican Xia, Rong Xiao, and Yuzhen Xie. Computing the real solutions of polynomial systems with the RegularChains library in Maple. ACM Communications in Computer Algebra, 45(3–4): 166–168, September 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Carminati:1995:NUS

[CDS95] J. Carminati, J. S. Debitt, and W. D. Scott. On the nonuniqueness of solutions in the modelling of steady wind flows. Computers and Fluids, 24(2): 177–188, February 1995. CO-DEN CPFLBI. ISSN 0045-7930 (print), 1879-0747 (electronic).

Corless:1994:SBL

[CES94] R. M. Corless and K. El-Sawy. Solution of banded linear systems of equations in Maple using LU factorization. In Lopez [Lop94a], pages 219–227. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Chu:1997:GRC

[CFFM97] Kenneth C.-K. Chu, Cesar E. Farell, Greg J. Fee, and R. G. McLenaghan. General relativity calculations in Maple. In The Sixth Canadian Conference on General Relativity and Relativistic Astrophysics (Fredericton, NB, 1995), volume 15 of Fields Inst. Commun., pages 195–199. Amer. Math. Soc., Providence, RI, USA, 1997.

Char:1984:DPMa

[CFG^{+}84] B. Char, G. Fee, K. O. Geddes, G. H. Gonnet, M. B. Monagen, and S. M. Watt. On the design and performance of the Maple system. In Golden and Hussain [GH84], page 199.

Char:1986:TIM

[CFG^{+}86] Bruce W. Char, G. J. Fee, Keith O. Geddes, Gaston H. Gonnet, and Michael B. Monagan. A tutorial introduction to Maple. Journal of Symbolic Computation, 2(2):179– 200, June 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Czapor:1986:IBA

[CG86] S. R. Czapor and K. O. Geddes. On implementing Buchberger’s algorithm for Gröbner bases. In Char [Cha86], pages 233–238. ISBN 0-89791-199-7. LCCN QA 155 .7 E4A336 1986. US$21.

Campbell:1995:SGD

[CG95] Stephen L. Campbell and E. Griepentrog. Solvability of general differential algebraic equations. SIAM J. Scient. Stat. Comput., 16(2):257–270, 1995. CODEN SIJCD4. ISSN 0196-5204.

Char:1983:MSC

[CGG83a] Bruce Char, Keith Geddes, and Gaston Gonnet. The Maple symbolic computation system. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(3–4):31–42, August/November 1983. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Char:1983:IMS

[CGG83b] Bruce W. Char, Keith O. Geddes, and Gaston H. Gonnet. An introduction to Maple: Sample interactive session. Technical Report CS-83-16, Faculty of Mathematics, University of Waterloo, Waterloo, Ontario, Canada, 1983. 15 pp.

Char:1984:IMS

[CGG84] Bruce W. Char, Keith O. Geddes, and Gaston H. Gonnet. An introduction to Maple: Sample interactive session. Technical Report CS-84-04, Faculty of Mathematics, University of Waterloo, 1984. 15 pp.

Char:1985:MSI

[CGG85] Bruce W. Char, Keith O. Geddes, and Gaston H. Gonnet. Maple: a sample interactive session. Technical Report CS-85-01, Faculty of Mathematics, University of Waterloo, 1985. 19 pp.

Char:1986:CAU

[CGG^{+}86] B. W. Char, K. O. Geddes, G. H. Gonnet, B. J. Marshman, and P. J. Ponzo. Computer algebra in the undergraduate mathematics classroom. In Char [Cha86], pages 135–140. ISBN 0-89791-199-7. LCCN QA 155 .7 E4A336 1986. US$21.

Char:1988:FLT

[CGG^{+}88a] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton Leong, Michael B. Monagan, and Stephen M. Watt. First Leaves: a Tutorial Introduction to Maple, Revised for Maple Version 4.2. WATCOM Publications Limited, 415 Philip St, Waterloo, Ontario N2L 3X2, Canada, 1988. ISBN 0-921329-22-9. iii + 115 pp. LCCN ????

Char:1988:MRM

[CGG^{+}88b] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton Leong, Michael B. Monagan, and Stephen M. Watt.

Maple Reference Manual. WAT-COM Publications Limited, 415 Philip St, Waterloo, Ontario N2L 3X2, Canada, fifth edition, 1988. ISBN 0-921329-15-6. xxvi + 403 pp. LCCN ????

Char:1989:FLT

[CGG^{+}89] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton Leong, Michael B. Monagan, and Stephen M. Watt. First Leaves: a Tutorial Introduction to Maple. WAT-COM Publications Limited, 415 Philip St, Waterloo, Ontario N2L 3X2, Canada, third edition, 1989. ISBN 0-921329-68-7. 122 pp. LCCN ????

Char:1990:FLT

[CGG^{+}90] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton Leong, Michael B. Monagan, and Stephen M. Watt. First Leaves: a Tutorial Introduction to Maple V. Waterloo Maple Publishing, 160 Columbia St, Waterloo, Ontario N2L 3L3, Canada, third edition, 1990. ISBN 0-9694580-0-2. 122 pp. LCCN ????

Char:1991:MLVb

[CGG^{+}91a] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton Leong, Michael B. Monagan, and Stephen M. Watt. Maple Library V Reference Manual. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97592-6, 3-540-97592-6. xxv + 698 pp. LCCN QA155.7.E4 M353 1991.

Char:1991:MVLa

[CGG^{+}91b] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton Leong, Michael B. Monagan, and Stephen M. Watt. Maple V: Language Reference Manual. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97622-1 (New York), 3-540-97622-1 (Berlin). xv + 267 pp. LCCN QA155.7.E4 M36 1991. US$24.95, FF 193,00.

Char:1992:FLT

[CGG^{+}92] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton Leong, Michael B. Monagan, and Stephen M. Watt. First Leaves: a Tutorial Introduction to Maple V. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97621-3, 3-540-97621-3. xvii + 253 pp. LCCN QA155.7.E4 F56 1992. Also available in Japanese, ISBN 4-431-70651-8.

Char:1983:DMC

[CGGG83] B. W. Char, K. O. Geddes, W. M. Gentleman, and G. H. Gonnet. The design of Maple: a compact, portable and powerful computer algebra system. In van Hulzen [vH83], pages 101– 115. ISBN 0-387-12868-9 (New York), 3-540-12868-9 (Berlin). LCCN QA155.7.E4 E85 1983.

Cyganowski:2001:MSD

[CGK01] S. Cyganowski, L. Grüne, and P. E. Kloeden. Maple for stochastic differential equations. In Theory and numerics of differential equations (Durham, 2000), Universitext, pages 127–177. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001.

Cheung:1994:MCM

[CH94] C. K. Cheung and John Harer.

Multivariable Calculus with Maple V. John Wiley, New York, NY, USA, preliminary edition, 1994. ISBN 0-471-59835-6. xiv + 366 pp.

Char:1986:PSS

[Cha86] Bruce W. Char, editor. Proceedings of the 1986 Symposium on Symbolic and Algebraic Computation: Symsac ’86, July 21–23, 1986, Waterloo, Ontario. ACM Press, New York, NY 10036, USA, 1986. ISBN 0-89791-199-7. LCCN QA 155 .7 E4A336 1986. US$21.

Char:1990:PRS

[Cha90a] B. W. Char. Progress report on a system for general-purpose parallel symbolic algebraic computation. In Watanabe and Nagata [WN90], pages 96–103. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

URL http://www.acm.org:
80/pubs/citations/proceedings/
issac/96877/p96-char/.

Char:1990:RAM

[Cha90b] Bruce W. Char. Report on the 7th Annual Maple Workshop. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24(1):18, January 1990. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Char:1991:RAM

[Cha91] Bruce W. Char. Report on the 8th Annual Maple Workshop. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(1):24, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Chandler:1996:OM

[Cha96] E. Chandler. ODE’s and Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN ???? ???? pp. LCCN ????

Cairo:1994:ICT

[CHF94] L. Cairó, D. D. Hua, and M. R. Feix. Invariants for a cubic three-wave system. J. Phys. A, 27(15):L547–L551, 1994. CO-DEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Chiricota:1993:CEM

[Chi93a]  Yves Chiricota.  Classifica-

tion des espèces moléculaires de degré 6 et 7 (French) [Classification of molecular species of degrees 6 and 7 ]. Les Annales des sciences mathématiques du Québec, 17(1):11–37, 1993. ISSN 0707-9109.

Chiricota:1993:RSE

[Chi93b] Yves Chiricota. Représentation symbolique d’espèces moléculaires. (French) [Symbolic representation of molecular species]. Theoretical Computer Science, 117 (1–2):131–136, August 30, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/
browse.cgi?year=1993&volume=
117&issue=1-2&aid=1416. Con-

ference on Formal Power Series and Algebraic Combinatorics (Bordeaux, 1991).

Crow:1994:CMK

[CHM94] John Crow, Herb Holden, and Marvin Margolis. Computers and mathematics: Keith Devlin. Notices of the American Mathematical Society, 41 (3):195–201, March 1994. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Corless:1990:SHL

[CJ90] R. M. Corless and D. J. Jeffrey. Solution of a hydrodynamic lubrication problem with Maple. Journal of Symbolic Computation, 9(4):503–514 (or 503–513??), April 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Carlson:1997:MMM

[CJ97] James A. Carlson and Jennifer M. Johnson. Multivariable mathematics with Maple: linear algebra, vector calculus and differential equations. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-270315-7 (paperback). vi + 290 pp. LCCN QA185.D37 C37 1997.

Corless:1992:NEA

[CJR92a] R. M. Corless, D. J. Jeffrey, and H. Rasmussen. Numerical evaluation of Airy functions with complex arguments. Journal of Computational Physics, 99 (1):106–114, March 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0021999192902798.

Corless:1992:NEAb

[CJR92b] R. M. Corless, D. J. Jeffrey, and H. Rasmussen. Numerical evaluation of Airy functions with complex arguments. Journal of Computational Physics, 99 (1):106–114, March 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0021999192902798.

Carette:2011:PEM

[CK11] Jacques Carette and Michael Kucera. Partial evaluation of Maple. Science of Computer Programming, 76(6):469– 491, June 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cheung:1998:GSM

[CKM98] Chi Keung Cheung, G. E.

Keough, and Michael May, S. J. Getting started with Maple: (for release 3, 4, and 5). Wiley, New York, NY, USA, 1998. ISBN 0-471-25249-2. x + 172 pp. LCCN QA76.95.C46 1998.

Cheung:2004:GSM

[CKM04] Chi Keung Cheung, G. E.

Keough, and Michael May, S. J. Getting started with Maple. Wiley, New York, NY, USA, second edition, 2004. ISBN 0-471-47013-9. x + 206 pp. LCCN QA76.95 .C46 2004.

Cyganowski:2002:EPS

[CKO02] Sasha Cyganowski, Peter E.

Kloeden, and Jerzy Ombach. From elementary probability to stochastic differential equations with Maple. Universitext. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-42666-3 (paperback). ISSN 0172-5939 (print), 2191-6675 (electronic). xvi + 310 pp. LCCN QA273 .C94 2002. URL

http://www.loc.gov/catdir/
enhancements/fy0817/2001054970-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0817/
2001054970-t.html.

CarillodeAlbornoz:1995:MV

[CL95] Augustin Carillo de Albornoz and Immaculada Llamas. Maple V. RA-MA Editorial, ????, 1995. ISBN 84-7897-161-0. ???? pp. LCCN ????

Cox:1992:IVA

[CLO92] David Cox, John Little, and Donala O’Shea. Ideals, varieties, and algorithms: An introduction to computational algebraic geometry and commutative algebra. Undergraduate Texts in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97847-X. xii + 513 pp.

Cox:2007:IVA

[CLO07] David A. Cox, John B. Little, and Donal O’Shea. Ideals, varieties, and algorithms: an introduction to computational algebraic geometry and commutative algebra. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2007. ISBN 0-387-35650-9, 0-387-35651-7 (e-book). xv + 551 pp. LCCN QA564 .C688 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0814/2006930875-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0814/
2006930875-t.html;   http:
//www.springerlink.com/openurl.
asp?genre=book&isbn=978-0-
387-35650-1.

Czapor:1987:NMP

[CM87] S. R. Czapor and R. G. McLenaghan. NP: a Maple package for performing calculations in the NewmanPenrose formalism. General relativity and gravitation, 19(6):623–635, 1987. CO-DEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic).

Carminati:1991:AIR

[CM91] J. Carminati and R. G. McLenaghan. Algebraic invariants of the Riemann tensor in a fourdimensional Lorentzian space. J. Math. Phys., 32(11):3135– 3140, 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Campbell:1994:PGN

[CM94a] S. L. Campbell and E. Moore.

Progress on a general numerical method for nonlinear higher index DAEs. II. Circuits, systems, and signal processing: CSSP, 13(2-3):123–138, 1994. CO-DEN CSSPEH. ISSN 0278-081X (print), 1531-5878 (electronic).

Curran:1994:UMM

[CM94b] D. A. S. Curran and W. Middleton. Using Maple in a mathematical modelling course. In Ersoy and Moscardini [EM94], pages 33–58. ISBN 3-540-58010-7, 0-387-58010-7. ISSN 0258-1248. LCCN QA401.M3943 1994.

Chiang:1996:MP

[CM96] K. Chiang and H. T. Man. Maple for Physicists. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN ???? ???? pp. LCCN ????

Clark:1999:MM

[CM99] David Clark and James Mc-Carron. Modules for Maple. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):24–25, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Conde:2001:APS

[CM01] Antônio Conde and Mirian Percia Mendes. An algebraic problem and the software Maple. International Journal of Applied Mathematics, 6(2):129– 140, 2001. ISSN 1311-1728 (print), 1314-8060 (electronic).

Chen:2011:ACT

[CM11] Changbo Chen and Marc Moreno Maza. Algorithms for computing triangular decompositions of polynomial systems. In Schost and Emiris [SE11], pages 83–90. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Chen:2022:FMP

[CM22] Tian Chen and Michael Monagan. Factoring multivariate polynomials represented by black boxes: a Maple + C implementation. Mathematics in Computer Science, 16(2–3):??, September 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL https://

link.springer.com/article/
10.1007/s11786-022-00534-
7.

Chen:2024:MPF

[CM24] Tian Chen and Michael Monagan. A Maple program to factor multivariate polynomials given by black boxes. ACM Communications in Computer Algebra, 58(3):77–80, September 2024. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Czapor:1992:ACS

[CMC92] S. R. Czapor, R. G. McLenaghan, and J. Carminati. The automatic conversion of spinor equations to dyad form in MAPLE. General relativity and gravitation, 24(9):911–928, 1992. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic).

Chen:2008:CPM

[CML^{+}08] Changbo Chen, Marc Moreno Maza, François Lemaire, Wei Pan, Liyun Li, and Yuzhen Xie. The ConstructibleSet-Tools and ParametricSystem-Tools modules of the RegularChains library in Maple. ACM Communications in Computer Algebra, 42(3):182–184, September 2008. CODEN ????

ISSN 1932-2232 (print), 1932-2240 (electronic).

Colton:2002:MCA

[Col02] Simon Colton. Making conjectures about Maple functions. Lecture Notes in Computer Science, 2385:259–274, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/series/0558/bibs/
                                        2385/23850259.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/
                                        2385/23850259.pdf.

Colton:2005:ACM

[Col05] Simon Colton. Automated conjecture making in number theory using HR, Otter and Maple. Journal of Symbolic Computation, 39(5):593–615, May 2005. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Comello:1996:PST

[Com96] Vic Comello. Physics software that’s suitable for engineers, too. Research & Development, 38(13):36–38, 40, December 1996. CODEN RE-DEEA. ISSN 0746-9179. Reviews several software packages, including some from Mathematica and Maple.

Connell:1992:APK

[Con92] Ian Connell. Addendum to a paper of K. Harada and M.- L. Lang: “Some elliptic curves arising from the Leech lattice” [J. Algebra 125, 1989, 2, 298– 310, MR 90g:11072]. Journal of Algebra, 145(2):463–467, 1992. CODEN JALGA4. ISSN 0021-8693 (print), 1090-266x (electronic).

Connell:1996:DEM

[Con96] Howard J. (Howard John) Connell. Differential equations A: Maple laboratory handbook. RMIT lecture notes in mathematics. Mathematics Dept., RMIT, Melbourne, Victoria, Australia, February 1996. ISBN 0-86444-562-8. ISSN 1039-5040. 70 + xiv pp. LCCN ????

Connell:1997:MLS

[Con97] Howard J. (Howard John) Connell. Maple laboratory sessions. Part 1. RMIT lecture notes in mathematics. Mathematics Dept., RMIT, Melbourne, Victoria, Australia, February 1997. ISBN 0-86444-635-7. vii + 77 + xxxi pp. LCCN ????

Connell:1998:DEM

[Con98] Howard J. (Howard John) Connell. Differential equations: a Maple laboratory handbook. RMIT lecture notes in mathematics. Mathematics Dept., RMIT, Melbourne, Victoria, Australia, July 1998. ISBN 0-86444-741-8. [98] pp. LCCN ????

Coombes:1997:DEM

[Coo97] Kevin Robert Coombes. Differential equations with Maple.

Wiley, New York, NY, USA, second (revised for Maple V, release 4) edition, 1997. ISBN 0-471-17645-1. v + 240 pp. LCCN QA371.5.D37 D53 1997 Bar.

Corless:1994:SAC

[Cor94] Robert M. Corless. Sufficiency analysis for the calculus of variations. In ACM [ACM94], pages 197–204. ISBN 0-89791-638-7. LCCN QA76.95.I59

                                        1994.  URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/issac/190347/
                                        p197-corless/.

Corless:1995:EMI

[Cor95a] Robert M. Corless. Essential Maple: An Introduction for Scientific Programmers. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94209-2. xiv + 218 pp. LCCN QA76.95 .C68 1995. US$24.95.

Corless:1995:SRS

[Cor95b] Robert M. Corless. Symbolic Recipes: Scientific Computing with Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94210-6. 300 pp. US$39.00.

Corless:1996:ABD

[Cor96a] R. M. Corless. Abell and Braselton’s Differential Equations with Maple V. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(1):

57–??, ????  1996.  CO-

DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1996:ECG

[Cor96b] Robert M. Corless. Editor’s corner: Gröbner bases and matrix eigenproblems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(4): 26–32, December 1996. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corliss:19xx:IIA

[Corxx] George F. Corliss. INTPAK for interval arithmetic in Maple: Introduction and applications. Journal of Symbolic Computation, 19xx. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Submitted.

Corless:2002:EMI

[Cor02] Robert M. Corless. Essential Maple 7: an introduction for scientific programmers. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95352-3. xv + 282 pp. LCCN QA76.95 .C678 2002.

Cosgrave:2000:NTC

[Cos00] John Cosgrave. Number theory and cryptography (using Maple). In Coding theory and cryptography (Annapolis, MD, 1998), pages 124–143. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000.

Connolly:1993:FTC

[CP93] T. H. Connolly and F. Pfeiffer. Feedforward torque computations with the aid of Maple V. In Angeles et al. [AHK93], pages 109–118. ISBN 0-7923-2585-0. LCCN QA841.C59 1993.

Cizek:1987:TPP

[ČPRV87] J. Čižek, J. Paldus, U. W. Ramgulam, and F. Vinette. Two-point Padé approximants in electrochemical kinetic currents. Progress in Surface Science, 25(1-4):17–39, 1987. CO-DEN PSSFBP. ISSN 0079-6816 (print), 1878-4240 (electronic).

Cizek:1990:EAF

[ČPV90] J. Čížek, J. Paldus, and F. Vinette. Explicit algebraic form of coupled cluster equations for the PPP model of benzene with an approximate inclusion of triexcited clusters. International Journal of Quantum Chemistry, 38(6):831–851, December 1990. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Char:1994:AIT

[CR94a] Bruce W. Char and Mark F. Russo. Automatic identification of time scales in enzyme kinetics models. In ACM [ACM94], pages 74–83. ISBN 0-89791-638-7. LCCN QA76.95.I59

1994.  URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/190347/
p74-char/.

Chen:1994:ASE

[CR94b] Chungmin Melvin Chen and Nick Roussopoulos. Adaptive selectivity estimation using query feedback. In Snodgrass and Winslett [SW94a], pages 161–172. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2 1994.

URL http://www.acm.org/
pubs/articles/proceedings/
mod/191839/p161-chen/p161-
chen.pdf;   http://www.acm.
org/pubs/citations/proceedings/
mod/191839/p161-chen/.

Cushman:1988:CNF

[CS89] Richard Cushman and Jan A.

Sanders. The constrained normal form algorithm. Celestial mechanics, 45(1-3):181– 187, 1988–1989. CODEN CLM-CAV. ISSN 0008-8714.

Carr:1993:SAN

[CS93] S. Carr and G. J. Savage. Sensitivity analysis of nonlinear physical systems using Maple. In Lee [Lee93], pages 118–

127. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Cordeiro:2008:MSI

[CS08] Gauss M. Cordeiro and Borko D.

Stosić. Maple script for improving test statistics. Journal of Statistical Computation and Simulation, 78(11):1045– 1053, 2008. CODEN JSCSAJ.

ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cox:1994:ACA

[CSS94] Hazel Cox, Stephen J. Smith, and Brian T. Sutcliffe. Application of the computer algebra system Maple for calculations on atomic and molecular systems. Computer Physics Communications, 84(1-3):186– 200, November 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Cheb-Terrab:1994:MPP

[CT94a] E. S. Cheb-Terrab. Maple procedures for partial and functional derivatives. Computer Physics Communications, 79 (3):409–424, May 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Cowan:1994:FLM

[CT94b] Richard Cowan and Albert K. L. Tsang. The fallingleaves mosaic and its equilibrium properties. Adv. Appl. Prob., 26(1):54–62, 1994. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

Cornil:1995:MIR

[CT95] J. M. Cornil and P. Testud. Maple: Introduction raisonnée à l’usage de l’étudiant, de l’ingénieur et du chercheur. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-58923-6. ???? pp. LCCN ????

Cheb-Terrab:1996:BRB

[CT96] E. S. Cheb-Terrab. Book review: Classical mechanics with Maple: By R. L. Greene. Springer-Verlag, Berlin, 1995. Approx. 230 pages, 55 figures. Hardcover price DM 48,00. ISBN 0-387-94512-1. Computer Physics Communications, 97 (3):358, September 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/0010465596000483.

Cornil:1997:MVR

[CT97] Jack-Michel Cornil and Philippe Testud. Maple V Release 4: introduction raisonnèe à l’usage de l’étudiant, de l’ingénieur et du chercheur. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-63186-0. xxi + 471 pp. LCCN ????

Cornil:2001:IMV

[CT01] Jack-Michel Cornil and Philippe Testud. An introduction to Maple V. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-66442-4. xx + 470 pp. Translated from the 1997 French original by Thierry Van Effelterre.

Cheb-Terrab:1995:SVE

[CTE95] E. S. Cheb-Terrab and A. G. Elfimov. Solution of Vlasov’s equation for complicated plasma geometry. I. spherical type. Computer Physics Communications, 85(2):251–266, February 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Cizek:1988:SCQ

[CV88] J. Cizek and F. Vinette. Symbolic computation in quantum mechanics. several simple examples. In Anonymous [Ano88]. CODEN IJQSDI. ISSN 0161-3642.

Carminati:2001:GMP

[CV01] J. Carminati and K. T. Vu. GHP: a Maple package for performing calculations in the Geroch–Held–Penrose formalism. General relativity and gravitation, 33(2):295–308, 2001. CODEN GRHVA8. ISSN 0001-7701 (print), 1572-9532 (electronic).

Crow:1994:TSR

[CVG94] John Crow, Mario Vassallo, and Gustaf Gripenberg. Three software reviews comprise this month’s column. John Crow presents the second installment of his two-part comparative review of Maple and Macsyma. Mario Vassallo reports on his experiences with spreadsheets and mathematics. Gustaf Gripenberg takes a look at MAX. Notices of the American Mathematical Society, 41 (4):299–306, April 1994. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cohen:1992:CAI

[CvGV92] Arjeh M. Cohen, Leendert van Gastel, and Sjoerd Verduyn Lunel, editors. Computer algebra in industry 2. Problem solving in practice. Proceedings of the SCAFI seminar on computer algebra in industry, held in Amsterdam, Netherlands, Dec. 1992. Wiley, New York, NY, USA, 1992. ISBN 0-471-95529-9. LCCN TA345 .C64242 1995.

Cizek:1993:USL

[CVW93] J. Cizek, F. Vinette, and E. J.

Weniger. On the use of the symbolic language Maple in physics and chemistry: Several examples. In de Groot and Nadrchal [dN93], pages 31–44. ISBN 981-02-1245-3. LCCN QC20.I45 1992.

Cohen:1992:STM

[CW92] Jaques Cohen and Aline Weitzman. Software tools for microanalysis of programs. Software —Practice and Experience, 22 (9):777–808, September 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chonacky:2005:R

[CW05a] N. Chonacky and D. Winch.

3Ms: a response. Computing in Science and Engineering, 7(5):7–9, September/October 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/32219/01501733.pdf?isnumber=
32219&prod=JNL&arnumber=1501733&
arSt=+7&ared=+9&arAuthor=
+Chonacky%2C+N.%3B++Winch%
2C+D.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=32219&arnumber=1501733&
count=14&index=1.

Chonacky:2005:IPM

[CW05b] N. Chonacky and D. Winch. 3Ms for instruction, Part 2: Maple, Mathematica, and Matlab. Computing in Science and Engineering, 7(4): 14–23, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/31456/01463131.pdf?isnumber=
31456&prod=JNL&arnumber=1463131&
arSt=+14&ared=+23&arAuthor=
Chonacky%2C+N.%3B+Winch%2C+
D.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
31456&arnumber=1463131&count=
14&index=1.

Chonacky:2005:IRM

[CW05c] Norman Chonacky and David Winch. 3Ms for instruction: Reviews of Maple, Mathematica, and Matlab. Computing in Science and Engineering, 7 (3):7–13, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

URL http://csdl.computer.
org/comp/mags/cs/2005/03/
c3007.pdf;   http://csdl.
computer.org/comp/mags/cs/
2005/03/c3007abs.htm.

Chonacky:2005:MMM

[CW05d] Norman Chonacky and David Winch. Maple, Mathematica, and Matlab: The 3M’s without the tape. Computing in Science and Engineering, 7(1):8–16, January/February 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2005/01/c1008.pdf;
 http://csdl.computer.org/
dl/mags/cs/2005/01/c1008.
htm.

Chonacky:2005:RMM

[CW05e] Norman Chonacky and David Winch. Reviews of Maple, Mathematica, and Matlab: Coming soon to a publication near you. Computing in Science and Engineering, 7(2):9–10, March/April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2005/02/c2009.pdf;
 http://csdl.computer.org/
dl/mags/cs/2005/02/c2009.
htm.

Cyganowski:1996:MPS

[Cyg96] S. Cyganowski. A MAPLE package for stochastic differential equations. In Computational techniques and applications: CTAC95 (Melbourne, 1995), pages 223–230. World Sci. Publishing, River Edge, NJ, 1996.

Chu:2011:MPN

[CZL11] Hongmei Chu, Yinlong Zhao, and Yinping Liu. A MAPLE package of new ADM-Padé approximate solution for nonlinear problems. Applied Mathematics and Computation, 217(17): 7074–7091, May 1, 2011. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

David:1988:VCC

[Dav88] Carl W. David. Variational calculations on a correlated helium wave function using Maple. Computers and Chemistry, 12 (3):229–231, 1988. CODEN COCHDK. ISSN 0097-8485 (print), 1879-0763 (electronic).

Davis:2001:DEM

[Dav01] Jon H. Davis. Differential equations with Maple: an interactive approach. Birkhäuser Boston Inc., Cambridge, MA, USA, 2001. ISBN 0-8176-4181-5, 3-7643-4181-5. xiv + 409 pp. LCCN QA371.5.D37 D37 2001. An interactive approach, With 1 CD-ROM (Unix, Windows and Macintosh).

Dovlo:2015:TCF

[DB15] Edem Dovlo and Natalie Baddour. Toolbox for the computation of 2D Fourier transforms in polar coordinates via Maple.

Journal of Open Research Software, 3(1):e3–??, February 05, 2015. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
bo/.

Duarte:2002:EPS

[DDdMS02] L. G. S. Duarte, S. E. S. Duarte, L. A. C. P. da Mota, and J. E. F. Skea. An extension of the Prelle–Singer method and a Maple implementation. Computer Physics Communications, 144(1):46–62, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

deJager:1993:USC

[de 93] Bram de Jager. Use of symbolic computation in nonlinear control: Is it viable? Proceedings of the IEEE Conference on Decision and Control, 1:276–281, 1993. CODEN PCDCDZ. ISSN 0191-2216.

deJager:1995:USC

[de 95] Bram de Jager. Use of symbolic computation in nonlinear control: Is it viable? IEEE Transactions on Automatic Control, 40(1):84–89, January 1995. CO-DEN IETAA9. ISSN 0018-9286 (print), 1558-2523 (electronic).

Decoste:1993:SIS

[Déc93] Hélène Décoste. Séries indicatrices et q-séries. (French) [Indicator series and q-series]. Theoretical Computer Science, 117 (1–2):169–186, August 30, 1993.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/
browse.cgi?year=1993&volume=
117&issue=1-2&aid=1419. Con-

ference on Formal Power Series and Algebraic Combinatorics (Bordeaux, 1991).

Decker:1994:CM

[Dec94]  R. Decker.  Calculus and

Maple. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN ???? ???? pp. LCCN ????

Diochnos:2007:CRS

[DET07] Dimitrios I. Diochnos, Ioannis Z. Emiris, and Elias P. Tsigaridas. On the complexity of real solving bivariate systems. In Brown [Bro07], pages 127–134. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Devlin:1992:CM

[Dev92] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 39(8):825–843, October 1992. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devitt:1993:COT

[Dev93a] J. S. Devitt. Combinatorial objects and their generating functions: a Maple class room environment. In Lee [Lee93], pages 20–26. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Devitt:1993:CMV

[Dev93b] John S. Devitt. Calculus with Maple V. Brooks/Cole, Pacific Grove, CA, USA, 1993. ISBN 0-534-16362-9. xxii + 502 pp. LCCN QA303.5.D37 D48 1993.

Devlin:1993:CM

[Dev93c] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 40(3):236–241, March 1993. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devlin:1994:CMa

[Dev94a] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 41(3):195–??, March 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devlin:1994:CMb

[Dev94b] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 41(6):611–??, July 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dodson:1995:EMU

[DG95] C. T. J. Dodson and E. A.

(Elizabeth A.) Gonzalez. Experiments in mathematics using Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-59284-9. xix + 465 pp. LCCN QA76.95 .D63 1995.

Dumas:1997:MSB

[DG97] Philippe Dumas and Xavier Gourdon. Maple: son bon usage en mathématiques. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-63140-2. xv + 460 pp. LCCN ????

Deeba:1998:ILA

[DG98] Elias Deeba and Ananda Gunawardena. Interactive linear algebra with Maple V. Textbooks in mathematical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98240-X (softcover). xiii + 317 pp. LCCN QA185.C65D44 1998.

DiDonato:1999:TFC

[DG99] Armido R. DiDonato and Russ Gnoffo. Testing a Fortran 90 compiler using the NSWC Fortran 77 Mathematics Library. Technical Report NSWCDD/TR-98/75, Naval Surface Warfare Center, Dahlgren, VA 22448-5100, USA, February 1999. v + 64 pp. URL https:

//apps.dtic.mil/sti/pdfs/
ADA360604.pdf.

DiNardo:2008:MAP

[DGS08] E. Di Nardo, G. Guarino, and D. Senato. A Maple algorithm for polykays and multivariate polykays. Advances and Applications in Statistics, 8(1): 19–36, February 2008. CO-DEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/
abstract/2836.htm.

Davis:1995:CGM

[DHM95a] Phillip J. Davis, Reuben Hersh, and Elena Anne Marchisotto. The Companion Guide to the Mathematical Experience. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1995. ISBN 0-8176-3849-0. vi + 120 pp. LCCN ?QA8.4 .D37 1995. US$14.95. URL http://

www.birkhauser.com/Books/
Davis/companion.htm.

Davis:1995:MES

[DHM95b] Phillip J. Davis, Reuben Hersh, and Elena Anne Marchisotto. The Mathematical Experience: Study Edition. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1995. ISBN 0-8176-3739-7. xxi + 487 pp. LCCN QA8.4.D37 1995. US$38.50. URL http:/

/www.birkhauser.com/Books/
Davis/SUMMARY.HTM.

DiBenedetto:1995:PDE

[DiB95] Emmanuele DiBenedetto. Partial Differential Equations. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1995. ISBN 0-8176-3708-7. xiv + 416 pp. LCCN QA377.D624 1995. US$42.50. URL http://

                                        www.birkhauser.com/Books/
                                        dibenedetto/summary.htm.

Dietrich:1992:EAC

[Die92] Volker Dietrich. ELISE, an algorithm to compute asymptotic representations for solutions of linear differential equations, realized with the computer algebra system MAPLE. Journal of Symbolic Computation, 14(1): 85–92, 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Dudley:1989:CAD

[DJ89] M. L. Dudley and R. W. James. Computer-aided derivation of spherical harmonic spectral equations in astro-geophysics. Journal of Symbolic Computation, 8(4):423–427, October 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

deJager:1991:SCZ

[dJ91] Bram de Jager. Symbolic calculation of zero dynamics for nonlinear control systems. In Watt [Wat91], pages 321–322. ISBN 0-89791-437-6. LCCN QA 76.95

                                        I59 1991. URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/issac/120694/
                                        p321-de_jager/.

Diaz:1995:CGC

[DK95] Angel Díaz and Erich Kaltofen.

On computing greatest common divisors with polynomials given by black boxes for their evaluations. In Levelt [Lev95], pages 232–239. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/220346/
p232-diaz/.

Devitt:1986:RCF

[DM86] J. S. Devitt and R. A. Mollin.

The rational canonical form of a matrix. International journal of mathematics and mathematical sciences, 9(2):253–260, 1986. ISSN 0161-1712 (print), 1687-0425 (electronic).

Delahaye:2005:DAE

[DM05] David Delahaye and Micaela Mayero. Dealing with algebraic expressions over a field in Coq using Maple. Journal of Symbolic Computation, 39(5):569–592, May 2005. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

deGroot:1993:PCP

[dN93] Robert A. de Groot and Jaroslav Nadrchal, editors. Physics computing ’92: proceedings of the 4th international conference, Prague, Czechoslovakia, August 24–28, 1992, Physics Computing — International Conference. World Scientific Publishing Co. Pte. Ltd., P.

O. Box 128, Farrer Road, Singapore 9128, 1993. ISBN 981-02-1245-3. LCCN QC20.I45 1992.

Dominici:2003:NDS

[Dom03] Diego Dominici. Nested derivatives: a simple method for computing series expansions of inverse functions. International journal of mathematics and mathematical sciences, 58:3699–3715, 2003. CODEN ???? ISSN 0161-1712 (print), 1687-0425 (electronic). URL

                                        http://www2.newpaltz.edu/~
                                        dominicd/NESTED14.pdf.

Donato:1997:MQT

[Don97]  P. (Paul) Donato.  Maple:

quinze thèmes mathématiques. Bibliothèque pratique: premier cycle and classes préparatoires. Diderot multimedia, Paris, France, 1997. ISBN 2-84352-043-6. vi + 181 pp. LCCN ????

Dooley:1999:IJS

[Doo99] Sam Dooley, editor. ISSAC 99: July 29–31, 1999, Simon Fraser University, Vancouver, BC, Canada: proceedings of the 1999 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 1999. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999.

Douglas:1993:EEA

[Dou93] S. C. Douglas. Exact expectation analysis of the LMS adaptive filter for correlated Gaussian input data. In IEEE [IEE93a], pages 519–522. CO-DEN IPRODJ. ISBN 0-7803-0948-0 (microfiche), 0-7803-0947-2 (casebound), 0-7803-0946-4 (softbound), 0-7803-0949-9 (CD ROM). ISSN 0736-7791. LCCN TK 7882 S65 I16 1993. Five volumes. IEEE catalog number 93CH3252-4.

Douillet:1996:MAM

[Dou96] Pierre Douillet. Maths avec MAPLE. Tome 2. Ellipses, Paris, France, 1996. ISBN 2-7298-5631-5. 239 pp. LCCN ????

Douglas:1995:EEA

[DP95] Scott C. Douglas and Weimin Pan. Exact expectation analysis of the LMS adaptive filter. IEEE Transactions on Signal Processing, 43(12):2863– 2871, December 1995. CO-DEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic).

Deconinck:2011:CPA

[DP11] Bernard Deconinck and Matthew S.

Patterson. Computing with plane algebraic curves and Riemann surfaces: The algorithms of the Maple package “Algcurves”. Lecture Notes in Mathematics, 2013:67–123, 2011. CODEN LNMAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/
10.1007/978-3-642-17413-1_

Duran:2014:MTM

[DPV14] Antonio J. Durán, Mario Pérez, and Juan L. Varona. The misfortunes of a trio of mathematicians using computer algebra systems. Can we trust in them? Notices of the American Mathematical Society, 61(10):1249– 1252, November 2014. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL http://www.ams.
                                        org/notices/201410/rnoti-
                                        p1249.pdf.

Dean:1994:CSD

[DS94a] Nathaniel Dean and Gregory E. Shannon, editors. Computational support for discrete mathematics: DIMACS workshop, March 12–14, 1992, Piscataway, NJ, volume 15 of DI-MACS Series in Discrete Mathematics and Theoretical Computer Science. Amer. Math. Soc., Providence, RI, USA, 1994. ISBN 0-8218-6605-2. LCCN QA76.9.M35C64 1994.

Dorn:1994:SSC

[DS94b] R. T. Dorn and R. J. Soukup. Solution of the solar cell transport equations using Maple. In Lopez [Lop94a], pages 90–96. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Davenport:1988:CAS

[DST88] J. H. Davenport, Y. Siret, and E. Tournier. Computer Algebra: Systems and Algorithms for Algebraic Computation. Academic Press, New York, NY, USA, 1988. ISBN 0-12-204230-1. xix + 267 pp. LCCN QA155.7.E4 D38 1988.

Davenport:1993:CAS

[DST93] J. H. Davenport, Y. Siret, and E. Tournier. Computer algebra: Systems and algorithms for algebraic computation. Academic Press, New York, NY, USA, second edition, 1993. ISBN 0-12-204232-8. xx + 298 pp. LCCN QA155.7.E4 D38 1993. With a preface by Daniel Lazard. Translated from the French by A. Davenport and J. H. Davenport. With a foreword by Anthony C. Hearn.

Dimitrova:1997:VSL

[DU97] Nelli S. Dimitrova and Christian P. Ullrich. Verified solving of linear systems with uncertainties in Maple. Lecture Notes in Computer Science, 1196:142– 149, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:2000:SPE

[Dub00] Paul Dubois. Scientific programming: Extending Maple with compiled routines. Computing in Science and Engineering, 2(4):82–86, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://dlib.computer.org/
cs/books/cs2000/pdf/c4082.
pdf.

Durand:1998:PMR

[Dur98] Robert Durand. Problèmes de mathématiques résolus avec Maple et Mathematica. Ellipses, Paris, France, 1998. ISBN 2-7298-4821-5. x + 180 pp. LCCN ????

Durcheva:2019:HUC

[Dur19] Mariana Durcheva. How to use CAS (Maple) to help students learn number theory. Mathematics in Computer Science, 13 (1–2):69–78, June 2019. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

deJager:1996:SSC

[dv96] Bram de Jager and Bram van Asch. Symbolic solutions for a class of partial differential equations. Journal of Symbolic Computation, 22(4):459–468, October 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Debever:1989:DEM

[DVL89] R. Debever, N. Van den Bergh, and J. Leroy. Diverging Einstein–Maxwell null fields of Petrov type D. Class. Quant. Grav., 6(10):1373–1378, 1989. CODEN CQGRDG. ISSN 0264-9381 (print), 1361-6382 (electronic).

Yang:2013:HMP

[dYyR13] Xu dong Yang and Hang yu Ruan. HBFGen: a Maple package to construct the Hirota bilinear form for nonlinear equations. Applied Mathematics and Computation, 219(15): 8018–8025, April 1, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0096300313001598.

Ellis:1992:MVF

[E^{+}92] Wade Ellis, Jr. et al. Maple V Flight Manual: Tutorials for Calculus, Linear Algebra, and Differential Equations. Brooks/ Cole, Pacific Grove, CA, USA, 1992. ISBN 0-534-17338-1. xxi + 183 pp. LCCN QA155.7.E4 M35 1992.

Elfimov:1993:VME

[ECT93] A. G. Elfimov and E. S. Cheb-Terrab. Vlasov–Maxwell equations for complicated plasma geometry. Computer Physics Communications, 77(3):357– 373, November 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/001046559390182C.

England:2014:BCM

[ECTB^{+}14] M. England, E. Cheb-Terrab, R. Bradford, J. H. Davenport, and D. Wilson. Branch cuts in Maple 17. ACM Communications in Computer Algebra, 48 (1/2):24–27, March 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Edwards:1995:LA

[Edw95] Harold M. Edwards. Linear Algebra. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1995. ISBN 0-8176-3731-1. xiii + 184 pp. LCCN QA184.E355 1995. US$36.50. URL http:/

/www.birkhauser.com/Books/
edwards/summary.htm.

ElHamidi:1997:UMA

[EG97] A. El Hamidi and M. Garbey. Using MAPLE for the analysis of bifurcation phenomena in gas combustion. Theoretical Computer Science, 187(1– 2):249–262, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/
browse.cgi?year=1997&volume=
187&issue=1-2&aid=2604.

Eckhardt:1996:BWS

[EHR^{+}96] B. Eckhardt, G. Hellbardt, H. Rechenberg, I. Heber, R. Haag, M. Carrier, S. Hüfner, M. Pillin, R. R. Ernst, H. Friedrich, C. Kleint, G. Baumann, and A. Ruider. Buchbesprechungen —Wissenschaftliche Software: Stehle: Order, Chaos, Order/Kafka: Gegen den Untergang/Verhulst: Der Glanz von Kopenhagen/Luchner: Physik ist überall: Streifzüge durch Natur, Alltag, Technik und Forschung/Omnès: The Interpretation of Quantum Mechanics/Neuser: Natur und Begriff/Vogel: Gerthsen Physik/ Kassel: Quantum Groups/ Schmidt-Rohr, Spiess: Multidimensional Solid-State NMR and Polymers/Ruder, Wunner, Herold, Geyer: Atoms in Strong Magnetic Fields/ Herrmann: Der Photoeffekt/ Kleint, Wiemers: Werner Heisenberg in Leipzig/Kofler: Maple V, Release 2: Einführung und Leitfaden für den Praktiker/Mathcad Plus 6.0. Physikalische Blätter, 52(1):55–60, January 1996. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:

//onlinelibrary.wiley.com/
doi/10.1002/phbl.19960520116/
abstract.

Eilers:2017:IEM

[EJ17] Søren Eilers and Rune Johansen. Introduction to Experimental Mathematics. Cambridge mathematical textbooks. Cambridge University Press, Cambridge, UK, 2017. ISBN 1-107-15613-0 (hardcover), 1-108-13375-4 (PDF e-book). xv + 303 pp. LCCN QA9 .E45 2017.

Ellis:1994:MVM

[EJLS94] W. Ellis, E. Johnson, E. Lodi, and D. Schwalbe. Maple V in der mathematischen Anwendung (für Release 2 und 3). International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1994. ISBN 3-929821-21-4. ???? pp.

Ellis:1996:MVF

[EJLS96] Wade Ellis, Jr., E. Johnson, E. Lodi, and D. Schwalbe. Maple V Flight Manual: Tutorials for Calculus, Linear Algebra, and Differential Equations. Brooks/Cole, Pacific Grove, CA, USA, second edition, 1996. ISBN ???? ???? pp. LCCN ????

Ellis:1989:MCS

[EL89] Wade Ellis, Jr. and Ed Lodi. Maple for the Calculus Student: a Tutorial. Brooks/Cole, Pacific Grove, CA, USA, 1989. ISBN 0-534-11874-7. vii + 67 pp. LCCN QA155.7.E4 E44 1989.

Ellis:1996:MVE

[Ell96] Wade Ellis. Maple V edition étudiant: version 4. International Thomson Publishing France, Paris, France, 1996. ISBN 2-84180-032-6. xiv + 172 pp. LCCN ????

Ellis:1997:MVF

[Ell97a] Wade Ellis. Maple V flight manual release 4: tutorials for calculus, linear algebra, and differential equations. Brooks/Cole symbolic computation series. Brooks/Cole, Pacific Grove, CA, USA, 1997. ISBN 0-534-26202-3. xiv + 168 pp. LCCN QA76.95 1996.

Ellis:1997:MVM

[Ell97b] Wade Ellis. Maple V in der mathematischen Anwendung. (German) [Maple V in mathematical application]. International Thomson Publishing, Bonn, Germany, 2. aktualisierte und überarbeitete Aufl. edition, 1997. ISBN 3-8266-0353-2. 186 pp. LCCN ????

Ersoy:1994:MMC

[EM94] Yasar Ersoy and Alfredo O.

Moscardini, editors. Mathematical modelling courses for engineering education: Proceedings of the NATO Advanced Research Workshop on the Design of Mathematical Modelling Courses for Engineering Education, held in Izmir, Turkey, July 12–16, 1993, volume 132 of NATO ASI Series F Computer and Systems Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58010-7, 0-387-58010-7. ISSN 0258-1248. LCCN QA401.M3943 1994.

Enns:1997:LMN

[EM97a] Richard Enns and George McGuire. Laboratory Manual for Nonlinear Physics with Maple for Scientists and Engineers. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1997. ISBN 0-8176-3841-5 (lab manual), 0-8176-3977-2 (text + lab manual). xii + 136 pp. LCCN QC20.7.N6E56 1997. US$54.50.

Enns:1997:NPMb

[EM97b] Richard H. Enns and George McGuire. Nonlinear Physics with Maple Files and Experiments. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1997. ISBN 0-8176-3838-5 (paperback), 0-8176-3841-5 (lab manual), 0-8176-3977-2 (set). x + 389 pp. LCCN QC20.7.N6 E56 1997.

Enns:1997:NPMa

[EM97c] Richard H. Enns and George McGuire. Nonlinear Physics with Maple for Scientists and Engineers. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1997. ISBN 0-8176-3841-5 (paperback), 0-8176-3841-5 (lab manual), 0-8176-3977-2 (set). x + 389 pp. LCCN QC20.7.N6 E56 1997. US$54.50.

Enns:2000:NPM

[EM00] Richard H. Enns and George McGuire. Nonlinear physics with Maple for scientists and engineers. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, second edition, 2000. ISBN 0-8176-4119-X, 3-7643-4119-X. xiii + 661 pp. LCCN QC20.7.N6 E56 2000.

Enns:2007:CAR

[EM07] Richard H. Enns and George McGuire. Computer algebra recipes: an advanced guide to scientific modeling. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-25768-3 (softcover), 0-387-49333-6 (ebook). x + 372 pp. LCCN QA371 .E56 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0824/2006936017-

d.html;   http://www.loc.
gov/catdir/toc/fy0710/2006936017.
html.

Erdos:1992:NPP

[ENS92] P. Erdős, J.-L. Nicolas, and A. Sárközy. On the number of pairs of partitions of n without common subsums. Colloq. Math., 63(1):61–83, 1992. CO-DEN CQMAAQ. ISSN 0010-1354 (print), 1730-6302 (electronic).

Edwards:1994:CUM

[EP94] C. H. Edwards, Jr. and David E. Penney. Calculus Using Maple. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-458282-9. ???? pp.

Edwards:1996:CPD

[EP96a] C. H. Edwards, Jr. and David E. Penney. Computing Projects for Differential Equations: Computing and Modeling. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-504465-0. 304 pp.

URL http://www.prenhall.
com/books/esm_0135044650.

html. Lab supplement to [EP96b].

Edwards:1996:DEB

[EP96b] C. H. Edwards, Jr. and David E. Penney. Differential Equations and Boundary Value Problems: Computing and Modeling. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-382094-7. 718 pp. LCCN QA371.E28 1996.

                                        URL http://www.prenhall.
                                        com/books/esm_0133820947.

html. See also lab supplement [EP96a].

Edwards:1996:DEC

[EP96c] C. H. Edwards, Jr. and David E. Penney. Differential Equations: Computing and Modeling. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-382102-1. ???? pp. LCCN QA371.E282 1996.

Edwards:2003:DEC

[EP03] C. H. (Charles Henry) Edwards and David E. Penney. Differential equations: computing and modeling. Pearson Education, Upper Saddle River, NJ, USA, third edition, 2003. ISBN 0-13-067337-4. various pp. LCCN QA371 .E29 2003.

Edwards:2004:DEB

[EP04a] C. H. (Charles Henry) Edwards and David E. Penney. Differential Equations and Boundary Value Problems: Computing and Modeling. Pearson Education, Upper Saddle River, NJ, USA, 2004. ISBN 0-13-065245-8. xv + 787 + 7 pp. LCCN QA371 .E28 2004.

Edwards:2004:EDE

[EP04b] C. H. (Charles Henry) Edwards and David E. Penney. Elementary differential equations. Pearson Education, Upper Saddle River, NJ, USA, fifth edition, 2004. ISBN 0-13-145773-X. xi + 613 + 6 pp. LCCN QA371 .E33 2004b.

Ellingsrud:1992:NTC

[ES92] Geir Ellingsrud and Stein Arild Strømme. The number of twisted cubic curves on the general quintic threefold (preliminary version). In Yau [Yau92], pages 181–222. ISBN 962-7670-01-4. LCCN QC20.7.M24 E83 1992.

Erocal:2010:SPU

[ES10] Burçin Eröcal and William Stein. The Sage Project: Unifying free mathematical software to create a viable alternative to Magma, Maple, Mathematica and MATLAB. Lecture Notes in Computer Science, 6327:12–27, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/content/pdf/
10.1007/978-3-642-15582-6_
4.pdf.

Ekhad:2023:CCW

[EZ23] Shalosh B. Ekhad and Doron Zeilberger. Counting clean words according to the number of their clean neighbors. ACM Communications in Computer Algebra, 57(1):5–9, March 2023. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3610377.3610379.

Fajardo:2019:CML

[Faj19]  William Fajardo.  A com-

putational Maple library for skew PBW extensions. Fundamenta Informaticae, 167(3): 159–191, ???? 2019. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fateman:1990:ATD

[Fat90] R. J. Fateman. Advances and trends in the design and construction of algebraic manipulation systems. In Watanabe and Nagata [WN90], pages 60–67. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

                                        URL http://www.acm.org:
                                        80/pubs/citations/proceedings/
                                        issac/96877/p60-fateman/.

Fattahi:1992:MVC

[Fat92] Abi Fattahi. Maple V Calculus Labs. Brooks/Cole, Pacific Grove, CA, USA, 1992. ISBN 0-534-19272-6. 92 + 3 pp. LCCN QA155.7.E4 F37 1992.

Fattahi:1996:MVC

[Fat96] Abi Fattahi. Maple V Calculus Labs. Brooks/Cole, Pacific Grove, CA, USA, second edition, 1996. ISBN 0-534-26208-2. vii + 100 pp. LCCN QA303.5.D37F38 1996. US$13.95.

Fee:1990:CCC

[Fee90] G. J. Fee. Computation of Catalan’s constant using Ramanujan’s formula. In Watanabe and Nagata [WN90], pages 157–160. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

URL http://www.acm.org:
80/pubs/citations/proceedings/
issac/96877/p157-fee/.

Ferrard:1998:MM

[Fer98] Jean-Michel Ferrard. Maths et Maple. Sciences Sup. Mathématiques. Dunod, Paris, France, 1998. ISBN 2-10-003651-3. xii + 465 pp. LCCN ????

Fuchssteiner:1988:CAD

[FF88] Benno Fuchssteiner and Uwe Falck. Computer algorithms for the detection of completely integrable quantum spin chains. In Levi and Winternitz [LW88], pages 22–50. ISBN 9971-5-0663-7. LCCN QC20.7.N6 I54 1988.

Filho:2011:SMP

[FF11] Tarcísio M. Rocha Filho and Annibal Figueiredo. [SADE]: a Maple package for the symmetry analysis of differential equations. Computer Physics Communications, 182(2):467– 476, February 2011. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465510003814.

Filho:1999:QMP

[FFB99] Tarcísio M. Rocha Filho, Annibal Figueiredo, and Léon Brenig. [QPSI] a MAPLE package for the determination of quasi-polynomial symmetries and invariants. Computer Physics Communications, 117 (3):263–272, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Fitz-Gerald:1994:EMC

[FGH94] G. F. Fitz-Gerald and W. P. Healy. Enlivening the mathematics curriculum with Maple. In Lopez [Lop94a], pages 71–80. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Faghihi:2004:NSC

[FH04] F. Faghihi and K. Hadad. Numerical solutions of coupled differential equations and initial values using Maple software. Applied Mathematics and Computation, 155(2):563– 572, August 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fritzsche:2001:MPC

[FIBT01] S. Fritzsche, T. Inghoff, T. Bastug, and M. Tomaselli. Maple procedures for the coupling of angular momenta. V. Recoupling coefficients. Computer Physics Communications, 139 (3):314–326, October 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465501002193.

Fitt:1989:SCH

[Fit89] A. D. Fitt. Symbolic computation of hyperbolicity regions for systems of two-phase flow conservation laws using MAPLE. Journal of Symbolic Computation, 8(3):305–308, 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Fitch:1993:DIS

[Fit93]  John Fitch, editor.  De-

sign and implementation of symbolic computation systems. Proceedings of the International Symposium (DISCO ’92) held in Bath, April 13–15, 1992, volume 721 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-57272-4. LCCN QA76.9.S88I576 1992.

Fritzsche:2003:MPC

[FIT03] S. Fritzsche, T. Inghoff, and M. Tomaselli. Maple procedures for the coupling of angular momenta. VII. Extended and accelerated computations. Computer Physics Communications, 153(3):424–444, July 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465503002273.

Fiume:1995:ISS

[Fiu95] Eugene L. Fiume. An introduction to scientific, symbolic, and graphical computation. A. K. Peters, Ltd., Wellesley, MA, USA, 1995. ISBN 1-56881-051-2. xvi + 306 pp. LCCN QA76.F54 1995.

Fleischmann:1993:SCC

[FJ93] Peter Fleischmann and Ingo Janiszczak. The semisimple conjugacy classes of finite groups of Lie type E_{6} and E_{7}. Communications in Algebra, 21(1):93–161, 1993. CO-DEN COALDM. ISSN 0092-7872 (print), 1532-4125 (electronic).

Fournier:1995:VMS

[FKM95] Robert Fournier, Norbert Kajler, and Bernard Mourrain. Visualization of mathematical surfaces: the IZIC server approach. Journal of Symbolic Computation, 19(1/2/3): 159–174 (or 159–173??), January, February, March 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Fofana:1993:EME

[FLC93] M. Fofana, I. LeGrow, and S. Carr. Experiences with Maple in engineering education at the University of Waterloo. In Lee [Lee93], pages 27– 33. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Faugeras:1989:MPM

[FM89] Olivier D. Faugeras and Steve Maybank. Motion from point matches: multiplicity of solutions. In IEEE [IEE89], pages 248–255. ISBN 0-8186-1903-1, 0-8186-5903-3 (microfiche), 0-8186-8903-X (case). LCCN TJ211.4 .W67 1989. IEEE Computer Society order number 1903; IEEE catalog number 89CH2716-9.

Fowkes:1994:IMM

[FM94] Neville D. Fowkes and John J.

Mahony. An Introduction to Mathematical Modelling. Wiley, New York, NY, USA, 1994. ISBN 0-471-93422-4 (hardcover), 0-471-94309-6 (softcover). xvi + 447 pp. LCCN QA401 .F68 1994.

Frempong-Mireku:1994:KWV

[FMM94] P. Frempong-Mireku and K. J. K.

Moriarty. Kramers and Wannier V-Matrices for the partition functions of the Ising model. In Lopez [Lop94a], pages 97–104. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Foster:1991:PM

[Fos91] Kenneth R. Foster. Prepackaged math. IEEE Spectrum, 28(11):44–50, November 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1999:MI

[Fos99] Kenneth R. Foster. Math on the Internet. IEEE Spectrum, 36(4):36–40, April 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2001:MSA

[Fos01] K. R. Foster. Maple scores again. IEEE Spectrum, 38 (12):63, December 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2006:SMG

[Fos06] Kenneth R. Foster. Software: Maple goes graphical. IEEE Spectrum, 43(2):64, February 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2011:MMT

[Fos11] Kenneth R. Foster. Mathematica 8 and Maple 15 [tools toys]. IEEE Spectrum, 48 (12):23, December 2011. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fuchssteiner:1987:CAM

[FOW87] Benno Fuchssteiner, Walter Oevel, and Waldemar Wiwianka. Computer-algebra methods for investigation of hereditary operators of higher order soliton equations. Computer Physics Communications, 44(1-2):47–55, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Fateman:1989:SDS

[FP89] R. J. Fateman and C. G. Ponder. Speed and data structures in computer algebra systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23(2):8–11, April 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fitzgerald:1995:TEQ

[FP95] S. Fitzgerald and J. Place.

Teaching elementary queueing theory with a computer algebra system. In Anonymous [Ano95a], pages 350–354. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fritzsche:1997:MPC

[Fri97] S. Fritzsche. Maple procedures for the coupling of angular momenta I. data structures and numerical computations. Computer Physics Communications, 103(1):51–73, June 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465597000325.

Fritzsche:2009:MPC

[Fri09] S. Fritzsche. Maple procedures for the coupling of angular momenta. An up-date of the Racah module. Computer Physics Communications, 180 (10):2021–2023, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465509001969.

Feinsilver:1992:BFR

[FS92] P. Feinsilver and R. Schott. On Bessel functions and rate of convergence of zeros of Lommel polynomials. Mathematics of Computation, 59(199):153–156, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Flajolet:1995:CAL

[FS95] Philippe Flajolet and Bruno Salvy. Computer algebra libraries for combinatorial structures. Journal of Symbolic Computation, 20(5–6):653–672 (or 653–671??), November– December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Fulton:1994:CMH

[Ful94] S. R. Fulton. Can Maple help in teaching calculus with limited time and staff? In Lopez [Lop94a], pages 63–70. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Farr:1994:MLP

[FV94] W. W. Farr and M. VanGeel. Maple labs and programs for calculus. In Lopez [Lop94a], pages 159–166. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Fritzsche:1998:MPC

[FVGF98] S. Fritzsche, S. Varga, D. Geschke, and B. Fricke. Maple procedures for the coupling of angular momenta II. sum rule evaluation. Computer Physics Communications, 111 (1–3):167–184, June 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465598000216.

GomezPardo:2013:ICM

[G´ 13] José Luis Gómez Pardo. Introduction to Cryptography with Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-32165-8, 3-642-32166-6. xxx + 705 pp. LCCN QA76.9.A25 G66

2013.  URL http://www.
springerlink.com/content/
978-3-642-32166-5.

Gander:2003:BRB

[Gan03]  Walter Gander.  Book re-

view: From Elementary Probability to Stochastic Differential Equations with MAPLE, by Sasha Cyganowski; Peter Kloeden; Jerzy Ombach. Journal of the American Statistical Association, 98(461):254–255, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/30045222.

Garvan:1995:RTE

[Gar95] Frank G. Garvan. Ramanujan’s theories of elliptic functions to alternative bases — a symbolic excursion. Journal of Symbolic Computation, 20(5–6): 517–536, November–December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Garvan:1997:MVP

[Gar97] Frank Garvan. The Maple V Primer: Release 4. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1997. ISBN 0-8493-2681-8. 143 pp. LCCN QA155.7.E4G37 1997. US$15.95.

Garvan:1999:MFM

[Gar99a] Frank Garvan. Modular functions, Maple and Andrews’ 10th problem. In Topics in number theory (University Park, PA, 1997), volume 467 of Math. Appl., pages 163–179. Kluwer Acad. Publ., Dordrecht, 1999.

Garvan:1999:PTS

[Gar99b] Frank Garvan. A q-product tutorial for a q-series MAPLE package. Sém. Lothar. Combin., 42:Art. B42d, 27 pp. (electronic), 1999. The Andrews Festschrift (Maratea, 1998).

Garvan:2002:MB

[Gar02] Frank (Frank G.) Garvan. The Maple book. Chapman and Hall/CRC, Boca Raton, FL, USA, 2002. ISBN 1-58488-232-8 (paperback). 479 pp. LCCN QA155.7.E4 G36 2002.

Gauthier:1999:HMP

[Gau99] Bruno Gauthier. HYPERG: a Maple package for manipulating hypergeometric series. Sém. Lothar. Combin., 43:Art. S43a, 10 pp. (electronic), 1999.

Geddes:1988:MCW

[GC^{+}88] Keith O. Geddes, Bruce W.

Char, et al. Maple: Calculus workbook problems and solutions. Technical report, Faculty of Mathematics, University of Waterloo, 1988. xxiii + 211 pp.

Geddes:1992:ACA

[GCL92] Keith O. Geddes, S. R. Czapor, and G. Labahn. Algorithms for Computer Algebra. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-9259-0. xviii + 585 pp. LCCN QA155.7.E4 G43 1992.

Grotendorst:1994:SCP

[GDS94a] J. Grotendorst, J. Dornseiffer, and S. M. Schoberth. Symbolic-numeric computations for problem-solving in physical chemistry and biochemistry. In Lopez [Lop94a], pages 131–140. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Grotendorst:1994:SCT

[GDS94b] Johannes Grotendorst, Juergen Dornseiffer, and Siegfried M. Schoberth. Symbolic-numeric computation techniques for problem-solving in physical chemistry and biochemistry. Technical Report KFA-ZAM-IB-9414, KFA Research Centre, Jülich, Jülich, Germany, 1994.

9 pp. URL ftp://ftp.zam.
kfa-juelich.de/pub/zamdoc/
ib/ib-94/ib-9414.ps.

Geck:1994:CVI

[Gec94] Meinolf Geck. On the character values of Iwahori-Hecke algebras of exceptional type. Proc. Roy. Soc. London Ser. A, 68(1): 51–76, 1994.

Gerbracht:1992:AED

[Ger92] Eberhard H. A. Gerbracht. Zur Algorithmischen und Exakten Diskussion Elementarer Funktionen. Technical report, Technischen Universität Carolo-Wilhelmina zu Braunschweig, Braunschweig, Germany, 1992.

Gessel:1995:FIW

[Ges95] Ira M. Gessel. Finding identities with the WZ method. Journal of Symbolic Computation, 20(5-6):537–566, November– December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Geddes:1992:HSI

[GF92] K. O. Geddes and G. J. Fee.

Hybrid symbolic-numeric integration in MAPLE. In Wang [Wan92], pages 36–41. ISBN 0-89791-489-9. LCCN QA 76.95 I59 1992.

Gaigalas:2002:MPC

[GF02] G. Gaigalas and S. Fritzsche.

Maple procedures for the coupling of angular momenta. VI. LS–jj transformations. Computer Physics Communications, 149(1):39–60, November 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S001046550200591X.

Gaigalas:2001:MPC

[GFF01] G. Gaigalas, S. Fritzsche, and B. Fricke. Maple procedures for the coupling of angular momenta. III. Standard quantities for evaluating many-particle matrix elements. Computer Physics Communications, 135 (2):219–237, April 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465500002344.

Garcia:2012:RIA

[GG12] Maria Emilia Alonso Garcia and André Galligo. A root isolation algorithm for sparse univariate polynomials. In van der Hoeven and van Hoeij [vdHvH12], pages 35–42. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Geddes:1981:MUMa

[GGC81a] Keith O. Geddes, Gaston H. Gonnet, and Bruce W. Char. Maple user’s manual. Technical Report CS-81-25, Computer Science Department, University of Waterloo, Waterloo, Ontario, Canada, 1981. 40 pp.

Geddes:1981:MUMb

[GGC81b] Keith O. Geddes, Gaston H. Gonnet, and Bruce W. Char. Maple user’s manual. Technical Report CS-82-40, Computer Science Department, University of Waterloo, Waterloo, Ontario, Canada, 1981. iv + 95 pp.

Geddes:1983:MUM

[GGC83] Keith O. Geddes, Gaston H. Gonnet, and Bruce W. Char. MAPLE: User’s manual. Technical Report CS-83-41, Faculty of Mathematics, University of Waterloo, 1983. 130 pp.

Golden:1984:PMU

[GH84] V. Ellen Golden and M. A. Hussain, editors. Proceedings of the 1984 MACSYMA Users’ Conference: Schenectady, New York, July 23–25, 1984. General Electric, Schenectady, NY, USA, 1984.

Gander:1993:SPS

[GH93] Walter Gander and Jiřı Hřebíček. Solving Problems in Scientific Computing Using MAPLE and MATLAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-57329-1, 3-540-57329-1. xiii + 268 pp. LCCN Q183.9 .G36 1993.

Govindarajulu:1994:AEA

[GH94] Z. Govindarajulu and H. C.

Howard. Asymptotic expansions applied to F -test criteria. Australian journal of statistics, 36(1):101–113, 1994. CODEN AUJSA3. ISSN 0004-9581.

Gander:1995:SPS

[GH95] Walter Gander and Jiřı Hřebíček.

Solving Problems in Scientific Computing Using Maple and MATLAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1995. ISBN 3-540-58746-2. xv + 315 pp. LCCN Q183.9 .G36 1995.

Gander:1997:SPS

[GH97] W. Gander and J. Hrěbíček.

Solving Problems in Scientific Computing using Maple and MATLAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1997. ISBN 3-540-61793-0. xvii + 408 pp. LCCN Q183.9 .G36 1997. US$49.95.

Gilbert:2003:MPD

[GH03] Robert Gilbert and George Hsiao. Maple Projects for Differential Equations. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-047974-8. 236 pp. LCCN QA371.5.D37 G54 2003.

Gander:2004:SPS

[GH04] Walter Gander and Jiří Hrěbíček. Solving problems in scientific computing using Maple and MATLAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2004. ISBN 3-540-21127-6. xxii + 476 pp. LCCN Q183.9 .G36 2004 UCD.

Guimaraes:1992:DRT

[GHC92] J. E. F. Guimaraes, G. R. Heppler, and S. R. Czapor. On the derivation and reduction of C^{1} trigonometric basis functions using Maple. Journal of Symbolic Computation, 13(6): 643–676 (or 643–675??), June 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Giusti:2000:PNM

[GHL^{+}00] Marc Giusti, Klemens Hägele, Grégoire Lecerf, Joël Marchand, and Bruno Salvy. The projective Noether Maple package: computing the dimension of a projective variety. Journal of Symbolic Computation, 30(3):291–307, 2000. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gao:2009:MPV

[GI09] Min Gao and Masaaki Ito. A Maple package for verifying ultradiscrete soliton solutions. Computer Physics Communications, 180(7):1196–1205, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S001046550900023X.

Gianni:1989:SAC

[Gia89] P. (Patrizia) Gianni, editor.

Symbolic and Algebraic Computation. International Symposium ISSAC ’88, Rome, Italy, July 4–8, 1988. Proceedings, volume 358 of Lecture Notes In Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51084-2. LCCN QA76.95 .I57 1988. Conference held jointly with AAECC-6.

Gaubert:1991:RCD

[GK91] Stéphane Gaubert and Carlos Klimann. Rational computation in dioid algebra and its application to performance evaluation of discrete event systems. In Algebraic computing in control (Paris, 1991), volume 165 of Lecture Notes in Control and Inform. Sci., pages 241–252. Springer-Verlag, Berlin, Germany, 1991.

Geulig:1998:IME

[GK98] I. Geulig and W. Krämer. Intervallrechnung in Maple –0 Die Erweiterung intpakX zum Paket intpak der Share-Library. (German) [Interval arithmetic in Maple — the extension intpakX to the share library package intpak]. Report, IWRMM, University of Karlsruhe, Karlsruhe, Germany, 1998.

Gerhard:2020:MC

[GK20] Jürgen Gerhard and Ilias S. Kotsireas. Maple Conference 2019. ACM Communications in Computer Algebra, 53(4):182, March 2020. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3388974.3388976.

Guyer:2001:NMC

[GKaM01] Tolga Güyer, Onur Kıymaz, and Göksal Bilgici andŞeref Mirasyedioğlu. A new method for computing the solutions of differential equation systems using generalized inverse via Maple. Applied Mathematics and Computation, 121(2–3):291–299, June 15, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/9/12/105/25/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/9/12/105/25/33/article.
                                        pdf.

Garbey:1991:UMA

[GKLM91] Marc Garbey, Hans G. Kaper, Gary K. Leaf, and Bernard J. Matkowsky. Using MAPLE for the analysis of bifurcation phenomena in condensed-phase surface combustion. Journal of Symbolic Computation, 12 (1):89–114 (or 89–113??), July 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Garbey:1992:QPW

[GKLM92] Marc Garbey, Hans G. Kaper, Gary K. Leaf, and Bernard J. Matkowsky. Quasi-periodic waves and the transfer of stability in condensed-phase surface combustion. SIAM J. of Applied Math., 52(2):384–395, 1992. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gray:1994:MPE

[GKW94] Simon Gray, Norbert Kajler, and Paul Wang. MP: a protocol for efficient exchange of mathematical expressions. In ACM [ACM94], pages 330–335. ISBN 0-89791-638-7. LCCN QA76.95.I59

1994.  URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/190347/
p330-gray/.

Grabmeier:2003:CAH

[GKW03] Johannes Grabmeier, Erich Kaltofen, and Volker Weispfenning, editors. Computer algebra handbook: foundations, applications, systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-65466-6. xx + 637 pp. LCCN QA155.7.E4 C64954 2003. URL http:/

/www.springer.com/sgw/cda/
frontpage/0,11855,1-102-22-
1477871-0,00.html. Includes

CD-ROM.

Gerdt:2019:MPT

[GLHR19] Vladimir P. Gerdt, Markus Lange-Hegermann, and Daniel Robertz. The MAPLE package TDDS for computing Thomas decompositions of systems of nonlinear PDEs. Computer Physics Communications, 234 (??):202–215, January 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465518302820.

Glowinski:1991:PIC

[Glo91] R. Glowinski, editor. Proceedings of the 10th International Conference on Computing Methods in Applied Sciences and Engineering. INRIA. Nova Science, New York, NY, USA, 1991.

Gloggengiesser:1993:MVS

[Glo93] H. Gloggengießer. Maple V Software für Mathematiker. Markt & Technik, ??, 1993. ISBN 3-87791-439-X. ???? pp.

Gloor:1998:IPI

[Glo98] Oliver Gloor, editor. ISSAC 98: Proceedings of the 1998 International Symposium on Symbolic and Algebraic Computation, August 13–15, 1998, University of Rostock, Germany. ACM Press, New York, NY 10036, USA, 1998. ISBN 1-58113-002-3. LCCN ????

Gruntz:1994:IG

[GM94] D. Gruntz and M. Monagan. Introduction to Gauss. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 28(2): 3–19, August 1994. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Geddes:1989:MCW

[GMM^{+}89] Keith O. Geddes, B. J. Marshman, I. McGee, P. J. Ponzo, and Bruce W. Char. Maple Calculus Workbook: Problems and Solutions. University of Waterloo, Waterloo, Ontario, Canada, 1989. ISBN ???? ???? pp. LCCN ???? No longer in print.

Ganzha:2000:CTI

[GMV00] V. G. (Victor Grigor’evich)

Ganzha, Ernst Mayr, and E. V. (Evgenii Vasil’evich) Vorozhtsov, editors. CASC-2000: Third International Workshop on Computer Algebra in Scientific Computing, October 5–9, 2000, Samarkand, Uzbekistan. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-41040-6. LCCN ????

Gomez:1990:MFC

[Gom90] Claude Gomez. MACROFORT:

a FORTRAN code generator in MAPLE = MACROFORT: un générateur de code FORTRAN dans MAPLE. Rapports techniques 119, INRIA (Institut National de Recherche en Informatique et en Automatique), Rocquencourt, France, 1990. 14 pp.

Gonnet:1988:EMA

[Gon88] Gaston H. Gonnet. Examples of Maple applied to problems from the American Mathematical Monthly. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22(2):8–16, April 1988. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gonnet:1989:PAS

[Gon89] Gaston H. Gonnet, editor. Proceedings of the ACM-SIGSAM 1989 International Symposium on Symbolic and Algebraic Computation: ISSAC ’89 / July 17–19, 1989, Portland, Oregon. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

Geck:1992:UCC

[GP92] Meinolf Geck and Götz Pfeiffer. Unipotent characters of the Chevalley groups D_{4}(q), q odd. Manuscripta Mathematica, 76(3-4):281–304, 1992. CO-DEN MSMHB2. ISSN 0025-2611 (print), 1432-1785 (electronic).

Grant:1993:CBF

[GQ93] I. P. Grant and H. M. Quiney. A class of Bessel function integrals with application in particle physics. J. Phys. A, 26(24):7547–7562, 1993. CO-DEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Gutierrez:1992:PIT

[GR92] Jaime Gutierrez and Tomas Recio. A practical implementation of two rational function decomposition algorithms. In Wang [Wan92], pages 152–157. ISBN 0-89791-489-9. LCCN QA 76.95

I59 1992. URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/143242/
p152-gutierrez/.

Grayson:1993:EER

[Gra93] L. P. Grayson, editor. Engineering education: renewing America’s technology: 23rd Annual conference on frontiers in education — November 1993, Washington, DC, Frontiers in Education Conference. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1482-4, 0-7803-1483-2, 0-7803-1484-0. ISSN 0190-5848. LCCN ????

Grayson:1994:PFE

[Gra94] L. P. Grayson, editor. Proceedings. Frontiers in Education. Twenty-Fourth Annual Conference. Educating Engineers for World Competition (Cat. No.94CH35723). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994.

ISBN 0-7803-2413-7. LCCN ????

Grabe:1996:APS

[Grä96] Hans-Gert Gräbe. About the polynomial system solve facility of Axiom, Macsyma, Maple, Mathematica, MuPAD, and Reduce. Technical report 96-11, Institut für Informatik, Universität Leipzig, Leipzig, Germany, 1996. 28 pp.

Greene:1993:ICP

[Gre93] R. L. Greene. On integrating computers into the physics curriculum. In Lee [Lee93], pages 34–42. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Greene:1995:CMM

[Gre95] Ronald L. Greene. Classical mechanics with Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94512-1. ix + 173 pp. LCCN QC125.2 .G72 1995.

Greene:1997:BRR

[Gre97] Ronald L. Greene. Book review: Richard H. Enns and George C. McGuire, Nonlinear Physics with Maple for Scientists and Engineers. Computers in Physics, 11(5):451– ??, September 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

                                        tronic). URL https://aip.
                                        scitation.org/doi/10.1063/
                                        1.4822590.

Gresser:1999:MAC

[Gre99] John T. Gresser. A Maple approach to calculus. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-010583-X. xiv + 292 pp. LCCN QA303.5.C65G74 1999.

Grotendorst:1989:MPC

[Gro89] J. Grotendorst. MAPLE programs for converting series expansions to rational functions using the Levin transformation. automatic generation of FORTRAN functions for numerical applications. Computer Physics Communications, 55(3):325–335, October 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Grotendorst:1990:AFM

[Gro90] J. Grotendorst. Approximating functions by means of symbolic computation and a general extrapolation method. Computer Physics Communications, 59(2):289–301, June 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Grotendorst:1991:MPT

[Gro91] J. Grotendorst. Maple package for transforming series, sequences and functions. Computer Physics Communications, 67(2):325–342, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Grossman:1996:SJC

[Gro96]  Nathaniel Grossman.  The

Sheer Joy of Celestial Mechanics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1996. ISBN 0-8176-3832-6, 3-7643-3832-6. 159 pp. LCCN QB351.G69 1995. US$34.50. URL http:/

                                        /www.birkhauser.com/Books/
                                        grossman/summary.htm.

Gruntz:1993:NAC

[Gru93] Dominik Gruntz. A new algorithm for computing asymptotic series. In Bronstein [Bro93], pages 239–244. ISBN 0-89791-604-2. LCCN QA76.5.I613

                                        1993.  URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/issac/164081/
                                        p239-gruntz/. ACM order

number: 505930.

Geddes:1989:RIM

[GS89] K. O. Geddes and L. Y. Stefanus. On the Risch–Norman integration method and its implementation in MAPLE. In Gonnet [Gon89], pages 212–217. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. URL http://www.

                                        acm.org:80/pubs/citations/
                                        proceedings/issac/74540/p212-

geddes/. ACM order number: 505890. English and French.

Gomez:1998:MPG

[GS98] Claude Gomez and Tony Scott. Maple programs for generating efficient FORTRAN code for serial and vectorised machines. Computer Physics Communications, 115(2–3):548–562, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465598001143.

Gasiorek:2012:OPP

[GS12a] Marcin Gaşiorek and Daniel Simson. One-peak posets with positive quadratic Tits form, their mesh translation quivers of roots, and programming in Maple and Python. Linear Algebra and its Applications, 436 (7):2240–2272, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0024379511007555.

Griffiths:2012:TWA

[GS12b] Graham W. Griffiths and W. E.

Schiesser. Traveling wave analysis of partial differential equations: numerical and analytical methods with MATLAB and Maple. Academic Press, New York, NY, USA, 2012. ISBN 0-12-384652-8 (hardcover). xiii + 447 pp. LCCN QA374 .G75 2012.

Gaigalas:2005:MPC

[GSF05] G. Gaigalas, O. Scharf, and S. Fritzsche. Maple procedures for the coupling of angular momenta. VIII. Spin-angular coefficients for single-shell configurations. Computer Physics Communications, 166(2):141– 169, March 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465504005582.

Gaigalas:2006:HSP

[GSF06] G. Gaigalas, O. Scharf, and S. Fritzsche. Hyperfine structure parametrisation in Maple. Computer Physics Communications, 174(3):202– 221, February 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465505005709.

Gomez:1995:CFM

[GSZ95] C. Gomez, B. Salvy, and P. Zimmermann. Calcul Formel: Mode d’Emploi. Exemples en Maple, volume 13 of Logique mathématiques informatique. Masson Editeur, Paris, France, 2e tirage avec mise à jour Maple V.4 edition, 1995. ISBN 2-225-84780-0. xii + 327 pp. LCCN ????

Gillespie:1993:ECK

[GU93] B. Gillespie and J. Uro. Exact calculation of the Kaplan– Meier bias using Maple software. In Lee [Lee93], pages 128– 136. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Gutierrez:2004:IJU

[Gut04] Jaime Gutierrez, editor. ISAAC 2004: July 4–7, 2004, University of Cantabria, Santander, Spain: proceedings of the 2004 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 2004. ISBN 1-58113-827-X. LCCN ????

Gerald:1989:ANA

[GW89] Curtis F. Gerald and Patrick O.

Wheatley. Applied numerical analysis. Addison-Wesley, Reading, MA, USA, fourth edition, 1989. ISBN 0-201-11583-2. xi + 679 + 36 + 7 pp. LCCN QA297 .G47 1989.

Gerald:1994:ANA

[GW94] Curtis F. Gerald and Patrick O.

Wheatley. Applied numerical analysis. Addison-Wesley, Reading, MA, USA, fifth edition, 1994. ISBN 0-201-56553-6. various pp. LCCN QA 297 G47 1994.

Gao:2000:CMP

[GW00] Xiao-Shan Gao and Dongming Wang, editors. Computer mathematics: proceedings of the fourth Asian Symposium (ASCM 2000), Chiang Mai, Thailand, 17-21 December 2000. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2000. ISBN 981-02-4498-3. LCCN QA76.9.M35.

Gander:2020:EBS

[GW20] Martin J. Gander and Gerhard Wanner. Exact BDF stability angles with Maple. BIT Numerical Mathematics, 60(3): ??, September 2020. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https://

                                        link.springer.com/article/
                                        10.1007/s10543-019-00796-
                                        x.

Gille:1993:SMD

[GWV93] Jean-Charles Gille, Stefan Wȩgrzyn, and Pierre Vidal. Some models for developmental systems. XIII. multilevel development by parallel programming. International Journal of Systems Science, 24(3):441– 450, 1993. CODEN IJSYA9. ISSN 0020-7721 (print), 1464-5319 (electronic).

Huston:1994:SCJ

[HA94] R. Huston and D. Armel, editors. 27th Summer conference: — June 1994, North Myrtle Beach, SC, Educational Resources Information Center — Microfiche Reports — ED 1994; ED 372752. EDRS, ????, 1994.

Harris:1992:DCM

[Har92] Kent Harris. Discovering Calculus with Maple. Wiley, New York, NY, USA, 1992. ISBN 0-471-55156-2. 192 pp. LCCN QA303.5.D37 H37 1992.

Hardy:1994:MLC

[Har94] D. W. Hardy. Maple labs and calculus at Colorado State. In Lum [Lum94], pages 773–776. ISBN 0-201-54304-4. LCCN QA11.A1I454 1992.

Hare:1997:CPB

[Har97] D. E. G. Hare. Computing the principal branch of log-Gamma. Journal of Algorithms, 25(2): 221–236, November 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0196677497908816.

Harris:1998:ARD

[Har98a] Frank E. Harris. Algebraic reduction in discrete light-cone quantized electrodynamics using Maple V. Technical report, Department of Physics, University of Utah and Quantum Theory Project, Departments of Physics and Chemistry, University of Florida, Salt Lake City, UT 84112, USA and Gainesville, FL 32611, USA, November 1998. Submitted to Computer Physics Communications.

Harris:1998:CGC

[Har98b] Frank E. Harris. Computer generation of coupled-cluster equations. Technical report, Department of Physics, University of Utah and Quantum Theory Project, Departments of Physics and Chemistry, University of Florida, Salt Lake City, UT 84112, USA and Gainesville, FL 32611, USA, November 1998. To appear in International Journal of Quantum Chemistry. A Maple V program is avail-

able at http://courses.chem.
utah.edu/facultyandstaff/
harrisf/harrisf.html#software.

Harris:2000:ARD

[Har00] Frank E. Harris. Algebraic reduction in discrete light-cone quantized electrodynamics using Maple V. Computer Physics Communications, 132(1–2):21– 29, October 15, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465500001338.

Harris:2014:MMU

[Har14] Frank E. Harris. Mathematics for physical science and engineering: symbolic computing applications in Maple and Mathematica. Academic Press, New York, NY, USA, 2014. ISBN 0-12-801000-2 (hardcover), 0-12-801049-5 (e-book). xiv + 768 pp. LCCN QA37.3 .H36 2014. URL http://www.

sciencedirect.com/science/
book/9780128010006.

Harris:2015:IMBb

[Har15] Frank E. Harris. Instructor’s Manual for Mathematics for physical science and engineering: symbolic computing applications in Maple and Mathematica. Elsevier Academic Press, Amsterdam, The Netherlands, 2015. ???? pp.

Harper:1992:IM

[HC92] David Harper and Rachid Chalabi. An introduction to Maple. Report, Computer Algebra Support Project, Astronomy Unit, School of Mathematical Sciences, Queen Mary and Westfield College, University of London, London, E1 4NS, November 1992. ???? pp. Third edition. Available by anonymous

ftp from maths.qmw.ac.uk
(138.37.80.15) in pub/mapledoc.

Hagin:1995:CEM

[HC95] Frank G. Hagin and Jack K.

Cohen. Calculus Explorations with Maple. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-328576-6. ???? pp. LCCN ????

Hagin:1996:CM

[HC96] Frank G. Hagin and Jack K. Cohen. Calculus with Maple. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-518994-2. x + 125 pp. LCCN QA303.5.C65H34 1996.

Harms:1994:NRM

[HDC94] D. D. Harms, J. S. Devitt, and C. J. Colbourn. Networks and reliability in Maple. In Dean and Shannon [DS94a], pages 223–244. ISBN 0-8218-6605-2. LCCN QA76.9.M35C64 1994.

Healy:1996:DEB

[Hea96] W. P. Healy. Differential equations B: Maple laboratory handbook. RMIT lecture notes in mathematics. Mathematics Department, RMIT, Melbourne, Victoria, Australia, second edition, 1996. ISBN 0-86444-588-1. vi + 67 + xix pp. LCCN ????

Holmes:1993:ECM

[HEBS93] Mark H. Holmes, Joseph G. Ecker, William E. Boyce, and William L. Siegmann. Exploring Calculus with Maple. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-52616-6. vii + 258 pp. LCCN QA303.5.D37 E97 1993.

Heck:1993:IM

[Hec93]  André Heck.  Introduction

to Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97662-0 (New York), 3-540-97662-0 (Berlin). xiii + 497 pp. LCCN QA155.7.E4H43 1993.

Heck:1996:IM

[Hec96] André Heck. Introduction to Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1996. ISBN 0-387-94535-0 (hardcover), 1-4684-0484-9, 1-4684-0486-5. xx + 699 pp. LCCN QA155.7.E4H43 1993. US$39.00.

Heck:2003:IM

[Hec03] André Heck. Introduction to Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2003. ISBN 0-387-00230-8. xvi + 828 pp.

Hedman:1992:TAA

[Hed92] Anders Hedman. Transmission analysis: Automatic derivation of relationships. American Society of Mechanical Engineers, Design Engineering Division (Publication) DE, 43 pt 1:259–266, 1992. CODEN AMEDEH.

Hellstrom:2010:VVI

[Hel10] Christian Hellström. VarInt:

variational integrator design with Maple. ACM Communications in Computer Algebra, 44 (3):115, September 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hennequin:1989:CAQ

[Hen89] P. Hennequin. Combinatorial analysis of Quicksort algorithm. RAIRO. Informatique théorique et applications := Theoretical informatics and applications, 23(3):317–333, ???? 1989. CODEN RSITD7, RI-TAE4. ISSN 0988-3754 (print), 1290-385X (electronic). URL

http://archive.numdam.org/
article/ITA_1989__23_3_317_
0.pdf;   http://www.rairo-
ita.org/action/displayIssue?
jid=ITA&tab=currentissue.

Hench:1994:SMA

[Hen94] D. L. Hench. Software for mathematical analysis. WESCON Conference Record, pages 258– 263, 1994. CODEN WCREDI. ISSN 1044-6036, 0083-8837.

Herman:2001:LAM

[Her01] Eugene A. Herman. Linear algebra: modules for interactive learning using MAPLE 6. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-44135-7. xviii + 496 pp. LCCN QA185.C65 L56 2001. Includes CD-ROM.

Hoshi:1992:PUQ

[HF92] Mamoru Hoshi and Philippe Flajolet. Page usage in a quadtree index. BIT, 32(3): 384–402, 1992. CODEN BIT-TEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Hutton:1995:MCA

[HH95] J. Hutton and J. Hutton. The Maple computer algebra system. Journal of Applied Econometrics, 10(3):329–??, 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Heal:1995:MVM

[HHR95] K. M. Heal, M. L. Hansen, and K. M. Rickard. Maple V Mathematics Learning Guide. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94536-9. ix + 269 pp.

LCCN QA155.7.E4 H42 1996. US$24.00. With the editorial assistance of J. S. Devitt; based in part on the work of B. W. Char.

Heal:1996:EMV

[HHR96a] K. M. Heal, M. L. Hansen, and K. M. Rickard. Einführung in Maple V. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60545-2. ix + 280 + 4 pp. LCCN ????

Heal:1996:MVL

[HHR96b] K. M. Heal, M. L. Hansen, and K. M. Rickard. Maple V Learning Guide. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94536-9, 0-387-94575-X (software boxed version). ix + 269 pp. LCCN QA155.7.E4H42 1996. US$24.00. With the editorial assistance of J. S. Devitt. Based in part on the work of B. W. Char.

Heal:1998:MVL

[HHR98] K. M. Heal, M. L. Hansen, and K. M. Rickard. Maple V Learning Guide. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98399-6, 0-387-98397-X. ix + 284 pp. LCCN QA155.7.E4 M362 1998. With the editorial assistance of J. S. Devitt. Based in part on the work of B. W. Char.

Houseaux:2003:CMP

[HJOP03] V. Houseaux, G. Jacob, N. E. Oussous, and M. Petitot. A complete Maple package for noncommutative rational power series. In Computer mathematics, volume 10 of Lecture Notes Ser. Comput., pages 174–188. World Sci. Publishing, River Edge, NJ, 2003.

Hazaveh:2003:EHS

[HJR^{+}03] K. Hazaveh, D. J. Jeffrey, G. J. Reid, S. M. Watt, and A. D. Wittkopf. An exploration of homotopy solving in Maple. In Computer mathematics, volume 10 of Lecture Notes Ser. Comput., pages 145–162. World Sci. Publishing, River Edge, NJ, 2003.

Hsu:1989:CSO

[HK89] L. Hsu and N. Kamran. Classification of second-order ordinary differential equations admitting Lie groups of fibre-preserving point symmetries. Proceedings of the London Mathematical Society. Third Series, 58(2):387– 416, 1989. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244x (electronic).

Heck:1995:IMb

[HK95] Andre Heck and Wolfram Koepf. Introduction to Maple. SIAM Review, 37(1):132–??, 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hostetler:1994:IC

[HL94] D. Hostetler and E. Larson. Interactive Calculus. D. C. Heath and Company, Lexington, MA, USA, 1994. ISBN ???? ???? pp. LCCN ????

Harris:1995:DCM

[HL95] Kent Harris and Robert J.

Lopez. Discovering calculus with Maple. Wiley, New York, NY, USA, second edition, 1995. ISBN 0-471-00973-3. 344 pp. LCCN QA303.5.D37 H37 1995.

Hogben:2007:HLA

[Hog07] Leslie Hogben, editor. Handbook of Linear Algebra. Discrete Mathematics and its Applications (Boca Raton). Chapman and Hall/CRC, Boca Raton, FL, USA, 2007. ISBN 1-58488-510-6 (hardcover), 1-4200-1057-3 (e-book). xxx + 1370 pp. LCCN QA184.2 .H36 2007. URL http:/

/www.crcnetbase.com/isbn/
9781420010572;   http://www.
crcnetbase.com/isbn/9781584885108;
 http://www.loc.gov/catdir/
enhancements/fy0647/2006045491-

d.html. Associate editors: Richard Brualdi, Anne Greenbaum and Roy Mathias.

Hohn:2001:SMB

[Hoh01] Michael Hohn. On the Solution of Mixed Boundary Value Problems in Elasticity. Ph.D. thesis, Department of Mathematics, University of Utah, Salt Lake City, UT, USA, December 2001. x + 221 pp.

Hollmann:1996:MVP

[Hol96] Helia Hollmann. A Maple V package for field theoretic Poisson brackets using objectoriented techniques. Computer Physics Communications, 96(2– 3):209–216, August 1, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0010465596000768.

Horbatsch:1995:QMU

[Hor95a] Marko Horbatsch. Quantum mechanics using Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-58875-2. x + 331 pp. LCCN QC174.17.D37 H67 199.

Horhager:1995:MTN

[Hör95b] M. Hörhager. Maple in Technik und Naturwissenschaft. ????, ????, December 1995. ISBN ???? ???? pp. LCCN ????

Herman:2005:VLA

[HPS05] Eugene A. Herman, Michael D. Pepe, and Eric P. Schulz. Visual linear algebra. Wiley, New York, NY, USA, 2005. ISBN 0-471-68299-3 , 0-471-70628-0 (CD-ROM). xix + 550 pp. LCCN QA185.C65 H47 2005. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        toc/fy053/2004059088.html.

Houstis:1992:AIE

[HR92] E. N. (Elias N.) Houstis and John Rischard Rice, editors. Artificial Intelligence, Expert Systems and Symbolic Computing. Selected and Revised Papers from the IMACS 13th World Congress, Dublin, Ireland, July 22–26, 1991. North-Holland Publishing Co., Amsterdam, The Netherlands, 1992. ISBN 0-444-89703-8. LCCN Q334 .I45 1991.

Hrebicek:1995:SVM

[Hrě95] J. Hrěbíček. Scientific visualization with MAPLE. In Skala [Ska95], pages 366–375.

Hiss:1992:BMC

[HS92] Gerhard Hiss and Josephine Shamash. 2-blocks and 2modular characters of the Chevalley groups G_{2}(q). Mathematics of Computation, 59 (200):645–672, 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Helton:1994:CSE

[HSW94] J. W. Helton, M. Stankus, and J. Wavrik. Computer simplification of engineering systems formulas. In IEEE, editor, Proceedings of the 33rd IEEE Conference on Decision and Control, 14–16 December 1994, volume 2, pages 1893–1898. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1968-0. LCCN TJ217 .I11c 1994. IEEE catalog number 94CH3460-3.

Helton:1998:CSF

[HSW98] J. W. Helton, M. Stankus, and J. J. Wavrik. Computer simplification of formulas in linear systems theory. IEEE Transactions on Automatic Control, 43 (3):302–314, March 1998. CO-DEN IETAA9. ISSN 0018-9286 (print), 1558-2523 (electronic).

Hui:1985:SCS

[HT85] W. H. Hui and G. Tenti. Symbolic computation of the Stokes wave. In Pavelle [Pav85], pages 329–337. ISBN 0-89838-173-8. LCCN QA155.7.E4 A67 1985.

Harrison:1993:RAR

[HT93] John Harrison and Laurent Thery. Reasoning about the reals: The marriage of HOL and Maple. Lecture Notes in Computer Science, 698:351– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1998:SAC

[HT98] John Harrison and L. Théry. A skeptic’s approach to combining HOL and Maple. Journal of Automated Reasoning, 21(3):279–294, December 1998. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A%3A1006023127567.

Harper:1991:GCA

[HWH91] David Harper, Chris Wooff, and David Hodgkinson. A Guide to Computer Algebra Systems. John Wiley, New York, NY, USA, 1991. ISBN 0-471-92910-7. xii + 148 pp. LCCN QA155.7.E4 H37 1991.

Hong:2014:MSI

[HY14] Hoon Hong and Chee Yap, editors. Mathematical Software — ICMS 2014: 4th International Conference, Seoul, South Korea, August 5–9, 2014, Proceedings, volume 8592. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-662-44198-5 (paperback), 3-662-44199-3 (e-book). LCCN QA76.9.M35.

Ioakimidis:1992:DTS

[IA92] N. I. Ioakimidis and G. T. Anastasselos. Direct Taylor-series solution of singular integral equations with MAPLE. Computers and structures, 45(3):613–617, October 1992. CODEN CM-STCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Ioakimidis:1993:CBM

[IA93] N. I. Ioakimidis and E. G. Anastasselou. Computer-based manipulation of systems of equations in elasticity problems with Gröbner bases. Computer Methods in Applied Mechanics and Engineering, 110(1-2):103– 111, 1993. CODEN CMMECC. ISSN 0045-7825, 0374-2830.

Ioakimidis:1994:AGB

[IA94a] N. I. Ioakimidis and E. G. Anastasselou. Application of Gröbner bases to problems of movement of a particle. Computers and Mathematics with Applications, 27 (3):51–57, February 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122194900469.

Ioakimidis:1994:GBT

[IA94b] N. I. Ioakimidis and E. G. Anastasselou. Gröbner bases in truss problems with Maple. Computers and structures, 52(5):1093– 1096, September 1994. CO-DEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Ioakimidis:1994:DEC

[IA94c] Nikolaos I. Ioakimidis and Eleni G. Anastasselou. Derivation of the equation of caustics in Cartesian coordinates with Maple. Engineering fracture mechanics, 48(1):147–149 (or 147–150??), May 1994. CO-DEN EFMEAH. ISSN 0013-7944 (print), 1873-7315 (electronic).

IEEE:1989:PAM

[IEE89]  IEEE, editor.  Proceedings:

analysis, motion / Worskshop on Visual Motion, March 20– 22, 1989, Irvine, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-1903-1, 0-8186-5903-3 (microfiche), 0-8186-8903-X (case). LCCN TJ211.4 .W67 1989. IEEE Computer Society order number 1903; IEEE catalog number 89CH2716-9.

IEEE:1993:III

[IEE93a] IEEE, editor. ICASSP 93:

1993 IEEE International Conference on Acoustics, Speech, and Signal Processing, April 27–30, 1993, Minneapolis Convention Center, Minneapolis, Minnesota. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. CODEN IPRODJ. ISBN 0-7803-0948-0 (microfiche), 0-7803-0947-2 (casebound), 0-7803-0946-4 (softbound), 0-7803-0949-9 (CD ROM). ISSN 0736-7791. LCCN TK 7882 S65 I16 1993. Five volumes. IEEE catalog number 93CH3252-4.

IEEE:1993:JNP

[IEE93b]  IEEE, editor.  JCNN ’93

— Nagoya: proceedings of 1993 International Joint Conference on Neural Networks, Nagoya Congress Center, October 25–29, 1993, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1421-2. LCCN QA76.87 .I57 1993. Three volumes.

Inghoff:2001:MPC

[IFF01] T. Inghoff, S. Fritzsche, and B. Fricke. Maple procedures for the coupling of angular momenta. IV. Spherical harmonics. Computer Physics Communications, 139(3):297– 313, October 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465501002181.

Ikle:1994:EDS

[Ikl94]  M. Iklé.  Exact (1 + 1)-

dimensional solutions to a discrete velocity model for coagulation-fragmentation. Mathematical models and methods in applied sciences, 4(6):857– 869, 1994. CODEN MMMSEU. ISSN 0218-2025 (print), 1793-6314 (electronic).

Ioakimidis:1994:SCS

[Ioa94] N. I. Ioakimidis. Symbolic computations for the solution of inverse/design problems with Maple. Computers and structures, 53(1):63–68, October 1994. CODEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Ilmer:2021:MAS

[IOP21] Ilia Ilmer, Alexey Ovchinnikov, and Gleb Pogudin. Maple application for structural identifiability analysis of ODE models. ACM Communications in Computer Algebra, 55(2):

49–53, June 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL

https://dl.acm.org/doi/10.
1145/3493492.3493497.

Ikramov:2000:ISM

[IS00] Kh. D. Ikramov and N. V.

Savel’eva. Invariant subspaces of matrix algebras, and MAPLE procedures for detecting their presence. Programmirovanie, 2:63–66, 2000. CO-DEN PROGD3. ISSN 0132-3474, 0361-7688.

Israel:1996:CMW

[Isr96] Robert B. Israel. Calculus the Maple Way. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-82829-4. ???? pp. LCCN ????

Izumi:2014:MPI

[Izu14] Hideaki Izumi. Maple package itsol for formal solutions of iterative functional equations. ACM Communications in Computer Algebra, 48(3/4):192–196, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jeffrey:1993:IOE

[Jef93]  D. J. Jeffrey.  Integration

to obtain expressions valid on domains of maximum extent. In Bronstein [Bro93], pages 34–41. ISBN 0-89791-604-2. LCCN QA76.5.I613

1993.  URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/164081/

                                        p34-jeffrey/. ACM order

number: 505930.

Jeffrey:2008:PAM

[Jef08] David Jeffrey, editor. Proceedings of the 21st annual meeting of the International Symposium on Symbolic Computation, ISSAC 2008, July 20–23, 2008, Hagenberg, Austria. ACM Press, New York, NY 10036, USA, 2008. ISBN 1-59593-904-0. LCCN ????

Janssen:2004:MTG

[JKM04] Klaus Janssen, Hanns Klinger, and Reinhold Meise. Markovketten: theoretische Grundlagen, Beispiele und Simulationen mit MAPLE. Mathematische Semesterberichte, 51(1):69–93, 2004. ISSN 0720-728x (print), 1432-1815 (electronic).

Jones:1996:SAM

[JKP96] J. Jones, N. P. Karampetakis, and A. C. Pugh. Some applications of Maple in linear systems analysis. In Anonymous [Ano96g], pages 7/1–5.

Jones:1998:CAG

[JKP98] Jon Jones, N. P. Karampetakis, and A. C. Pugh. The computation and application of the generalized inverse via ∥MAPLE∥. Journal of Symbolic Computation, 25(1):99–124, January 1998. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Jones:2003:SDA

[JKP03] J. Jones, N. P. Karampetakis, and A. C. Pugh. Solution of discrete ARMA-representations via MAPLE. Applied Mathematics and Computation, 139 (2–3):437–489, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jerosolimski:1994:NMF

[JL94] M. Jerosolimski and L. Levacher. New method for fast calculation of Jacobian matrices: Automatic differentiation for power system simulation. IEEE Transactions on Power Systems, 9(2):700–706, May 1994. CODEN ITPSEG. ISSN 0885-8950 (print), 1558-0679 (electronic).

Jacob:1991:ACC

[JLL91] G. Jacob and F. Lamnabhi-Lagarrigue, editors. Algebraic computing in control: proceedings of the first European conference, Paris, March 13–15, 1991, volume 165 of Lecture Notes in Control and Information Sciences. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-540-54408-9. LCCN QA402.3 .A4535 1991.

Jeanneau:1999:EMR

[JLP99] Xavier Jeanneau, Daniel Lignon, and Jean-Louis Poss. Exercices de mathématiques résolus à l’aide de Maple et Mathematica. Ellipses-Marketing, Paris, France, 1999. ISBN 2-7298-9909-X. viii + 230 pp. LCCN ????

Joyner:1998:MPD

[JM98] David Joyner and Roland Martin. A Maple package for the decomposition of certain tensor products and restrictions of representations using crystal graphs. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(2):19–22, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jing:2017:PLM

[JM17] Rui-Juan Jing and Marc Moreno Maza. The polyhedra library in Maple. ACM Communications in Computer Algebra, 51(3):86– 88, September 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Johnson:1993:LAM

[Joh93] Eugene W. Johnson. Linear Algebra with Maple V. Brooks/ Cole, Pacific Grove, CA, USA, 1993. ISBN 0-534-13069-0. xii + 142 pp. LCCN QA185.D37 J64 1993.

Jones:1996:FMC

[Jon96]  Kevin Rowe Jones.  A

fuzzy mathematical computation package for Maple. Thesis (M.Eng.), Department of Engineering Mathematics and Computer Science, University of Louisville, Louisville, KY, USA, 1996. x + 168 pp.

Joyner:2000:CTC

[Joy00] David Joyner, editor. Coding theory and cryptography: from Enigma and Geheimschreiber to quantum theory. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-66336-3 (softcover), 3-642-59663-0 (ebook). LCCN QA268 .C67 1999. UK£44.50. URL http:/

/frode.home.cern.ch/frode/
pubs/cryptoday.pdf.  Pro-

ceedings of the Conference on Coding Theory, Cryptography and Number Theory held at the U.S. Naval Academy during October 25–26, 1998.

Jenks:1992:ASC

[JS92] Richard D. Jenks and Robert S.

Sutor. AXIOM: The Scientific Computation System. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97855-0 (New York), 3-540-97855-0 (Berlin). xxiv + 742 pp. LCCN QA76.95.J46 1992. URL

http://link.springer.com/
10.1007/978-1-4612-2940-7.

Joyner:2000:RCM

[JS00] David Joyner and Salahoddin Shokranian. Remarks on codes from modular curves: MAPLE applications. In Coding theory and cryptography (Annapolis, MD, 1998), pages 229–250. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000.

Jank:1991:CMU

[JT91] G. Jank and L. H. Tack. Conformal mapping using Bergman’s method and the MAPLE system. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(2):18–23, April 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Judson:1996:CEU

[Jud96] Phoebe Judson. Calculus explorations using Maple. Saunders College Publishing, Ft. Worth, TX, USA, 1996. ISBN 0-03-017434-1. 148 pp. LCCN QA303.5.C65J83 1996.

Jakubczyk:2021:MPC

[JWK^{+}21] Paweł Jakubczyk, Andrzej Wal, Michał Kaczor, Dorota Jakubczyk, Mirosław Łabuz, and Jan Milewski. A Maple package for combinatorial aspects of Bethe Ansatz. Computer Physics Communications, 261(??):Article 107720, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0010465520303568.

Zhou:2006:IAH

[jZzFbL06] Zhen jiang Zhou, Jing zhi Fu, and Zhi bin Li. An implementation for the algorithm of Hirota bilinear form of PDE in the Maple system. Applied Mathematics and Computation, 183(2):872–877, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhou:2010:MPC

[jZzFbL10] Zhen jiang Zhou, Jing zhi Fu, and Zhi bin Li. Maple packages for computing Hirota’s bilinear equation and multisoliton solutions of nonlinear evolution equations. Applied Mathematics and Computation, 217(1): 92–104, September 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kaiser:1994:FGW

[Kai94] Gerald Kaiser. A Friendly Guide to Wavelets. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 0-8176-3711-7. xiv + 300 pp. LCCN QA403.3.K35 1994. US$39.50. URL http:/

/www.birkhauser.com/Books/
kaiser/summary.htm.

Kalkbrener:1993:GEA

[Kal93] Michael Kalkbrener. A generalized Euclidean algorithm for computing triangular representations of algebraic varieties. Journal of Symbolic Computation, 15(2):143–168 (or 143– 167??), February 1993. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Kamminga-vanHulsen:1993:CMM

[Kam93] M. Kamminga-van Hulsen.

Computeralgebra met Maple.

Academic Service, ????, 1993. ISBN 90-6233-923-9. ???? pp.

Kamerich:1995:GM

[Kam95] Ernic Kamerich. A Guide to Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94116-9. 335 pp. LCCN QA76.95 .K355 1995. US$29.00.

Kamerich:1999:GM

[Kam99] Ernic Kamerich. A guide to Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-94116-9. xxiii + 325 pp. LCCN QA76.95.K355 1999.

Kautsky:1993:ACD

[Kau93] Jaroslav Kautský. An algebraic construction of discrete wavelet transforms. Applications of Mathematics, 38(3): 169–193, 1993. CODEN APM-TEO. ISSN 0862-7940 (print), 1572-9109 (electronic).

Klimov:2001:KPT

[KBG01] D. M. Klimov, V. G. Baĭdulov, and V. A. Gorodtsov. The Kovalevskaya–Painlevé test for shallow water equations using the MAPLE package. Doklady Akademii nauk SSSR, 376 (5):600–604, 2001. CODEN DANKAS. ISSN 0869-5652.

Kwong:1990:CAS

[KD90] Man Kam Kwong and Mary Beth Dever. Computer-aided study of a problem in Hermitian matrix theory. Journal of Symbolic Computation, 9(1):87–112, 1990. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Keith:2001:VLA

[Kei01] Sandra Keith. Visualizing Linear Algebra Using Maple. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-041816-1. 250 (est.) pp. LCCN ????

Kendall:2001:SIC

[Ken01] Wilfrid S. Kendall. Symbolic Itô calculus in AXIOM: an ongoing story. Statistics and Computing, 11(1):25–35, 2001. CO-DEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Kaltofen:1997:PPS

[KH97] Erich Kaltofen and Markus Hitz, editors. PASCO ’97. Proceedings of the Second International Symposium on Parallel Symbolic Computation, July 20–22, 1997, Maui, HI. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-951-3. LCCN QA76.58 .I5674 1997.

Kimura:1997:MDU

[Kim97] Ikuo Kimura. Multi-link dynamics using Lagrange’s equations, Maple, and a kinamatic [sic] recursive formula. Thesis (M.S.), California State University, Long Beach, Long Beach, CA, USA, 1997. xii + 130 pp.

Kirkwood:1995:AIE

[Kir95] J. R. Kirkwood. Annotated Instructor’s Edition to accompany Calculus Projects for Maple. William C. Brown Company Publishers, Dubuque, IA, USA, 1995. ISBN 0-697-25488-7. ???? pp.

Kleczka:1991:SCA

[KK91] W. Kleczka and E. Kreuzer. Systematic computer-aided analysis of dynamic systems. In Watt [Wat91], pages 429– 430. ISBN 0-89791-437-6. LCCN QA 76.95 I59

1991.  URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/120694/
p429-kleczka/.

Klimek:1997:DCS

[KK97] Grazyna Klimek and Maciej Klimek. Discovering curves and surfaces with Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94890-2 (hardcover). xi + 217 pp. LCCN QA 567 K58 1997.

Krob:1996:IAC

[KL96] Daniel Krob and Stephane Legros. Introduction au calcul symbolique et aux mathématiques expérimentales. Tome 1, Le système Maple. Passeport pour l’informatique. International Thomson Publishing France, Paris, France, 1996. ISBN 2-84180-038-5. xxi + 354 pp. LCCN ????

Klasa:2010:FPD

[Kla10] J. Klasa. A few pedagogical designs in linear algebra with Cabri and Maple. Linear Algebra and its Applications, 432(8): 2100–2111, April 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kavian:1996:MPR

[KMG96] Masoud Kavian, R. G. McLenaghan, and K. O. Geddes. MapleTensor: Progress report on a new system for performing indicial and component tensor calculations using symbolic computation. In Lakshman Y. N. [Lak96], pages 204–211. ISBN 0-89791-796-0. LCCN QA 76.95 I59

1996.  URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/236869/
p204-kavian/.

Kreyszig:1994:MCM

[KN94] E. Kreyszig and E. J. Normington. Maple Computer Manual for Seventh Edition Advanced Engineering Mathematics: Erwin Kreyszig. Wiley, New York, NY, USA, 1994. ISBN 0-471-31126-X. xii + 506 pp. LCCN QA401 K74 1993. There is a companion Instructor’s Maple Manual, but I have not yet found an exact citation for it.

Kneppers:1994:IML

[Kne94] H. A. W. M. Kneppers. The influence of Maple on a linear algebra course at the Delft University of Technology. In Lopez [Lop94a], pages 57–62. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Kloeden:2002:EPS

[KOC02] Peter E. Kloeden, Jerzy Ombach, and Sasha Cyganowski. From elementary probability to stochastic differential equations with MAPLE. Universitext. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-42666-3 (paperback). xvi + 310 pp. LCCN QA273 .K56 2002.

Koepf:1991:TCM

[Koe91] Wolfram Koepf. On two conjectures of M. S. Robertson. Complex variables, theory and application, 16(2-3):127–130, 1991. CODEN CVTADV. ISSN 0278-1077.

Koepf:1992:ADP

[Koe92a] Wolfram Koepf. Algorithmic development of power series. In Calmet and Campbell [CC92], pages 195–213. ISBN 0-387-57322-4. LCCN Q334 .I5488 1992.

Koepf:1992:PSC

[Koe92b] Wolfram Koepf. Power series in computer algebra. Journal of Symbolic Computation, 13(6):581–603, 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Koekoek:1993:SDE

[Koe93] Roelof Koekoek. The search for differential equations for certain sets of orthogonal polynomials. Journal of Computational and Applied Mathematics, 49(1–3):111–119, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Proceedings of the Seventh Spanish Symposium on Orthogonal Polynomials and Applications (VII SPOA) (Granada, 1991).

Koekoek:1994:DES

[Koe94] Roelof Koekoek. Differential equations for symmetric generalized ultraspherical polynomials. Trans. Amer. Math. Soc., 345(1):47–72, 1994.

Kofler:1994:MVRa

[Kof94a] M. Kofler. Maple V Release 2.

Addison-Wesley, Reading, MA, USA, 1994. ISBN 3-89319-635-8. ???? pp.

Kofler:1994:MVRb

[Kof94b] M. Kofler. Maple V Release 3. Addison-Wesley, Reading, MA, USA, 1994. ISBN 3-89319-765-6. ???? pp. Contains a CD-ROM with examples and a Maple demo disk.

Kofler:1997:MIR

[Kof97] Michael Kofler. Maple: an introduction and reference. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-17899-0. xx + 522 pp. LCCN QA76.95.K64 1997.

Kohl:1994:GRH

[Koh94] T. Kohl. Group rings and Hopf-Galois theory in Maple. In Lopez [Lop94a], pages 116–121. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Kolev:2000:EIC

[Kol00] Dimitar Kolev, editor. Eighth International Colloquium on Numerical Analysis and Computer Science with Applications, Plovdiv, Bulgaria, August 13– 17, 1999, volume 1A(4) of International journal of differential equations. Academic Publications, Sofia, Bulgaria, 2000. ISBN ???? LCCN ????

Komma:1995:MPM

[Kom95]  Michael Komma.  Moderne

Physik mit Maple. International Thomson Publishing, Bonn, Germany, November 1995. ISBN 3-929821-81-8. ???? pp. LCCN ????

Koornwinder:1991:HHS

[Koo91] Tom H. Koornwinder. Handling hypergeometric series in Maple. In Orthogonal polynomials and their applications (Erice, 1990), volume 9 of IMACS Ann. Comput. Appl. Math., pages 73–80. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, 1991.

Koornwinder:1993:ZAA

[Koo93] Tom H. Koornwinder. On Zeilberger’s algorithm and its qanalogue. Journal of Computational and Applied Mathematics, 48(1–2):91–111, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Proceedings of the Seventh Spanish Symposium on Orthogonal Polynomials and Applications (VII SPOA) (Granada, 1991).

Kinderlehrer:1993:EDO

[KOW93] David Kinderlehrer, Biao Ou, and Noel Walkington. The elementary defects of the Oseen-Frank energy for a liquid crystal. C. R. Acad. Sci. Paris Sér. I Math., 316(5):465–470, 1993. CODEN CHASAP. ISSN 0764-4442 (print), 1778-3577 (electronic).

Krawcewicz:2003:CML

[KR03] Wieslaw Krawcewicz and B. (Bindhyachal) Rai. Calculus with Maple Labs: early transcendentals. Alpha Science, Pangbourne, UK, 2003. ISBN 1-84265-074-2. xi + 697 pp. LCCN QA303.5.C65 K7 2003 UCD.

Krawietz:1997:MVI

[Kra97] Arnold Krawietz. Maple V für Ingenieurstudium. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-60223-2. ix + 358 pp. LCCN ????

Kramer:2007:IMP

[Krä07] W. Krämer. Introduction to the Maple power tool intpakX.

Serdica Journal of Computing, Bulgarian Academy of Sciences, 1(4):469–504, 2007. CODEN ???? ISSN 1312-6555.

Korporal:2010:MPI

[KRR10] Anja Korporal, Georg Regensburger, and Markus Rosenkranz. A Maple package for integrodifferential operators and boundary problems. ACM Communications in Computer Algebra, 44(3):120–122, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Korporal:2012:ISD

[KRR12] Anja Korporal, Georg Regensburger, and Markus Rosenkranz. ISSAC 2012 software demonstrations: Symbolic computation for ordinary boundary problems in Maple. ACM Communications in Computer Algebra, 46(3/4):154–156, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Katsamaki:2020:PMP

[KRTZ20] Christina Katsamaki, Fabrice Rouillier, Elias Tsigaridas, and Zafeirakis Zafeirakopoulos. PTOPO: a MAPLE package for the topology of parametric curves. ACM Communications in Computer Algebra, 54(2):49– 52, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL

https://dl.acm.org/doi/10.
1145/3427218.3427223.

Khan:2012:FSM

[KS12a] Muhammad Taimoor Khan and Wolfgang Schreiner. On formal specification of Maple programs. Lecture Notes in Computer Science, 7362:443– 447, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/
content/pdf/10.1007/978-3-
642-31374-5_33.

Khan:2012:TFS

[KS12b] Muhammad Taimoor Khan and Wolfgang Schreiner. Towards the formal specification and verification of Maple programs. Lecture Notes in Computer Science, 7362:231–247, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.
com/chapter/10.1007/978-3-
642-31374-5_16/;   http://
link.springer.com/content/
pdf/10.1007/978-3-642-31374-
5_16.

Khan:2013:VFM

[KS13] Muhammad Taimoor Khan and Wolfgang Schreiner. A verification framework for MiniMaple programs. ACM Communications in Computer Algebra, 47 (3–4):98–99, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kugi:1996:OOA

[KSK96] A. Kugi, K. Schlacher, and M. Kaltenbacher. Object oriented approach for large circuits with substructures in the computer algebra program Maple V. In Proceedings of the 1996 3rd International Conference on Software for Electrical Engineering Analysis and Design, ELECTROSOFT’96, pages 491–500. ACM Press, New York, NY 10036, USA, 1996.

Klima:2000:AAA

[KSS00] Richard E. Klima, Neil Sigmon, and Ernest Stitzinger. Applications of abstract algebra with Maple. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2000. ISBN 0-8493-8170-3. xii + 256 pp. LCCN QA162.K65 2000.

Klima:2007:AAA

[KSS07] Richard E. Klima, Neil P. Sigmon, and Ernest L. Stitzinger. Applications of abstract algebra with Maple and MAT-LAB. Discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2007. ISBN 1-58488-610-2, 1-4200-1119-7. xii + 505 pp. LCCN QA162 .K65 2007. Includes CD-ROM.

Kleiman:1991:EAG

[KT91] Steven L. Kleiman and Anders Thorup, editors. Enumerative algebraic geometry: proceedings of the 1989 Zeuthen Symposium, Copenhagen, Denmark, volume 123 of Contemp Math. Amer. Math. Soc., Providence, RI, USA, 1991. ISBN 0-8218-5131-4. LCCN QA564 .Z48 1989.

Karian:1995:PSE

[KT95] Zaven A. Karian and Elliot A.

Tanis. Probability and statistics explorations with Maple. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-436759-6. xi + 226 pp. LCCN QA273.19.E4 K37 1995.

Karian:1999:PSE

[KT99] Zaven A. Karian and Elliot A.

Tanis. Probability and statistics, explorations with Maple. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1999. ISBN 0-13-021536-8. 270 pp. LCCN QA273.19.E4 K37 1999 Stacks.

Kopuzlu:2003:MMG

[KU03] Abdullah Kopuzlu and Tamer Ugur. A method on the matrix of graph of torus knot k(2, q) and its Maple application. Applied Mathematics and Computation, 141(2– 3):331–338, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kunkle:2002:WBI

[Kun02] Wanda Kunkle. A Web-based integral evaluator: a demonstration of the successful integration of WebEQ, Maple, and Java. In Anonymous [Ano02b], page ?? ISBN ???? LCCN ????

Kwong:1990:KCM

[Kwo90] Man Kam Kwong. On the Kolodner-Coffman method for the uniqueness problem of Emden–Fowler BVP. Zeitschrift für Angewandte Mathematik und Physik, 41(1):79–104, 1990. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Kwong:1991:NIP

[KZ91] Man Kam Kwong and A. Zettl. Norm inequalities for the powers of a matrix. Amer. Math. Monthly, 98(6):533–538, 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kouchakpour:1994:SCR

[KZ94] Peyman Kouchakpour and Albert Y. Zomaya. Symbolic computation of robot dynamic parameters. Cybernetics and Systems, 25(5):697–727, September–October 1994. CO-DEN CYSYDH. ISSN 0196-9722 (print), 1087-6553 (electronic).

Lahtonen:1992:LBM

[Lah92] Jyrki Lahtonen. Lower bounds for the multiplicities of simple composition factors of the induced modules. Communications in Algebra, 20(7): 1873–1895, 1992. CODEN COALDM. ISSN 0092-7872 (print), 1532-4125 (electronic).

Lai:1994:CAF

[Lai94] Shan-Tao Lai. Computation of algebraic formulas for Wigner 3-j, 6-j, and 9-j symbols by Maple. International Journal of Quantum Chemistry, 52(3): 593–607, October 15, 1994. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

LakshmanYN:1996:IPI

[Lak96] Lakshman Y. N., editor. ISSAC ’96: Proceedings of the 1996 International Symposium on Symbolic and Algebraic Computation, July 24–26, 1996, Zurich, Switzerland. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996.

Lassalle:1991:PJG

[Las91] Michel Lassalle. Polynômes de Jacobi généralisés (French) [Generalized Jacobi polynomials]. C. R. Acad. Sci. Paris Sér. I Math., 312(6):425–428, 1991. CODEN CHASAP. ISSN 0764-4442 (print), 1778-3577 (electronic).

Lavaud:1991:ASE

[Lav91] M. Lavaud. A^{S} TEX: a software environment on PC adapted to scientific research. In Glowinski [Glo91], pages 779–788.

Lemelin:1998:MVP

[LBP98] Mario Lemelin, Sylvie Bourque Picotte, and Eric Paquette. Maple V pour l’algèbre linéaire. Editions Le Griffon d’argile, Sainte-Foy, PQ, Canada, 1998. ISBN 2-89443-060-4. 143 pp. LCCN ????

Leicknam:1997:LM

[LBQTV97] Jean-Claude Leicknam, Geneviève Bidaud, Michel Quaggetto, and François-Xavier Testard-Vaillant. 15 leçons de Maple. Enseignement de l’informatique. Masson Editeur, Paris, France, 1997. ISBN 2-225-85625-7. ISSN 1270-9069. xii + 216 pp. LCCN ????

Levine:1990:MV

[LC90] Minna Levine and Linda Custer. Mathematica, version 1.2. MacUser, pages 148–154, 156, 158, 160, 162, 164, November 1990. CODEN MCUSEY. ISSN 0884-0997. Requires IBM PC or; compatibles; 4MB RAM, DOS 3.0 or later. Discusses Mathematica, Theorist, Maple, and Milo.

Le:2001:CMT

[Le01] H. Q. Le. Computing the minimal telescoper for sums of hypergeometric terms. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35(3): 2–10, September 2001. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lee:1993:MCM

[Lee93] Thomas Lee, editor. Mathematical Computation with Maple V: Ideas and Applications: Proceedings of the Maple Summer Workshop and Symposium, University of Michigan, Ann Arbor, June 28–30, 1993. Birkhäuser Boston Inc., Cambridge, MA, USA, 1993. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Lim:1993:ACA

[LEGS93] C. T. Lim, M. T. Ensz, M. A.

Ganter, and D. W. Storti. Algebraic computer aided-design with Maple V.2. In Lee [Lee93], pages 169–175. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Lemelin:1997:MVP

[Lem97] Mario Lemelin. Maple V pour les sciences: introduction sur les fonctionnalités du logiciel de calcul symbolique. Editions Le Griffon d’argile, Sainte-Foy, PQ, Canada, ed. provisoire. edition, 1997. ISBN 2-89443-052-3. 221 pp. LCCN ????

Leong:1986:IDU

[Leo86] B. L. Leong. Iris: design of an user interface program for symbolic algebra. In Char [Cha86], pages 1–6. ISBN 0-89791-199-7. LCCN QA 155 .7 E4A336 1986. US$21.

Leroux:1995:AM

[Ler95] Alain Leroux. Applications de Maple. Librairie Vuibert, ????, 1995. ISBN 2-7117-8859-8. ???? pp. LCCN ????

Lesser:1994:NMC

[Les94] Martin Lesser. Nonlinear modeling of complex mechanical systems. International journal of bifurcation and chaos in applied sciences and engineering, 4(3):521–551, 1994. CODEN IJBEE4. ISSN 0218-1274.

Levelt:1992:VWK

[Lev92] A. H. M. Levelt. Van der Waals, Korteweg, van Laar: a MAPLE excursion into the thermodynamics of binary mixtures. In Cohen et al. [CvGV92], pages 227–238. ISBN 0-471-95529-9. LCCN TA345 .C64242 1995.

Levelt:1995:IIS

[Lev95] A. H. M. Levelt, editor. IS-SAC ’95: International symposium on symbolic and algebraic computation: July 10–12, 1995, Montréal, Canada, ISSAC — Proceedings. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995.

Levine:1997:AMM

[Lev97] Albert Levine. Aide-mémoire Maple. Ellipse-Edition Marketing, Paris, France, 1997. ISBN 2-7298-6797-X. 127 pp. LCCN ????

Levine:1998:IM

[Lev98] Albert Levine. Introduction à Maple. Ellipses, Paris, France, deuxième edition, 1998. ISBN 2-7298-9804-2. 269 pp. LCCN ????

Liao:1995:EHP

[LF95] Hsin-Chao Liao and Richard J. Fateman. Evaluation of the heuristic polynomial GCD. In Levelt [Lev95], pages 240–247. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/220346/
p240-liao/.

Lohr:1993:RED

[LH93] L. L. Lohr and C. H. Huben.

Rotational energy dispersions for van der Waals molecular clusters. In Lee [Lee93], pages 137–143. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Lamagna:1992:DUI

[LHJ92] Edmund A. Lamagna, Michael B.

Hayden, and Catherine W. Johnson. The design of a user interface to a computer algebra system for introductory calculus. In Wang [Wan92], pages 358–368. ISBN 0-89791-489-9. LCCN QA 76.95

I59 1992. URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/143242/
p358-lamagna/.

Li:2000:AMS

[Li00a] Shi Qi Li. An application of Maple to the study of quaternions and the quaternion group. J. Chongqing Norm. Univ. Nat. Sci. Ed., 17(2):34–40, 69, 2000.

CODEN CSXKER. ISSN 1001-8905.

Li:2000:TWS

[Li00b] Zhi-Bin Li. Travelling wave solutions to the two-dimensional dissipative-dispersive nonlinear system with Maple. In Computer mathematics (Chiang Mai, 2000), volume 8 of Lecture Notes Ser. Comput., pages 224–228. World Sci. Publishing, River Edge, NJ, 2000.

Lieh:1994:SFE

[Lie94] Junghsen Lieh. Separated-form equations of motion of controlled flexible multibody systems. Journal of Dynamic Systems, Measurement, and Control, 116(4):702–712, December 1994. CODEN JDSMAA. ISSN 0022-0434.

Liu:2019:BDI

[Liu19] Jiang Liu. Block distance invariant method for monoterm canonicalization of Riemann tensor polynomials. ACM Communications in Computer Algebra, 53(3):134–137, September 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Labelle:1995:CVM

[LL95] Gilbert Labelle and Louise Laforest. Combinatorial variations on multidimensional quadtrees. Journal of Combinatorial Theory (Series A), 69(1): 1–16, 1995. CODEN JCBTA7. ISSN 0097-3165 (print), 1096-0899 (electronic).

Lawson:1996:MLL

[LL96] Terry Lawson and Robert J.

Lopez. Maple labs for linear algebra. Wiley, New York, NY, USA, 1996. ISBN 0-471-13594-1 (paperback). vi + 112 pp. LCCN ????

Li:2004:RMP

[LL04a] Zhi-Bin Li and Yin-Ping Liu.

RAEEM: a Maple package for finding a series of exact traveling wave solutions for nonlinear evolution equations. Computer Physics Communications, 163(3):191–201, November 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0010465504004370.

Liu:2004:MPF

[LL04b] Yin Ping Liu and Zhi Bin Li. A Maple package for finding exact solitary wave solutions based on Wu’s elimination method and its application. J. Systems Sci. Math. Sci., 24(1):118–124, 2004. CODEN XKSHEW. ISSN 1000-0577.

LeBlanc:1994:SBL

[LLLL94] Victor G. LeBlanc, Ali Lari-Lavassani, and William F. Langford. Symmetry-breaking for leading order gradient maps in R^{2} with applications to O(3). Nonlinearity (Bristol), 7(2):577–594, 1994. CODEN NONLE5. ISSN 0951-7715 (print), 1361-6544 (electronic).

Letniowski:1988:IAQ

[LM88] F. W. Letniowski and R. G. McLenaghan. An improved algorithm for quartic equation classification and Petrov classification. General relativity and gravitation, 20(5):463–483, 1988. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic).

LeVey:1991:JUM

[LM91a] G. Le Vey and J. Masse. Joint use of Maple and Basile: using computer algebra in a CACSD environment. In Algebraic computing in control (Paris, 1991), volume 165 of Lecture Notes in Control and Inform. Sci., pages 156–164. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991.

Vey:1991:JUM

[LM91b] G. Le Vey and J. Masse. Joint use of Maple and Basile: using computer algebra in a CACSD environment. In Algebraic computing in control (Paris, 1991), volume 165 of Lecture Notes in Control and Inform. Sci., pages 156–164. Springer-Verlag, Berlin, Germany, 1991.

Lozano:1994:TGM

[LM94] E. R. Lozano and E. R. Macias. “Turtle graphics” in Maple V.2. In Lopez [Lop94a], pages 3–12. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Lescure:1998:EAMa

[LM98a] Nicolas Lescure and Bruno Mombelli. Électrocinétique avec Maple et Pspice: 68 exercices et problèmes corrigés, rappels de cours. J’integre. Prepas scientifiques. Dunod, Paris, France, 1998. ISBN 2-10-003723-4. xiv + 170 pp. LCCN ????

Lescure:1998:EAMb

[LM98b] Nicolas Lescure and Bruno Mombelli. Électronique avec Maple et Pspice: 66 exercices et problèmes corrigés, rappels de cours. J’integre, prepas scientifiques. Dunod, Paris, France, 1998. ISBN 2-10-003847-8. x + 182 pp. LCCN ????

Llovet:1993:LRS

[LMJ93] J. Llovet, R. Martínez, and J. A. Jaén. Linear recurring sequences for computing the resultant of multivariate polynomials. Journal of Computational and Applied Mathematics, 49(1–3):145–152, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Proceedings of the Seventh Spanish Symposium on Orthogonal Polynomials and Applications (VII SPOA) (Granada, 1991).

Li:2008:HPS

[LMRS08a] Xin Li, M. M. Maza, R. Rasheed, and E. Schost. High-performance symbolic computation in a hybrid compiled-interpreted programming environment. In ICCSA ’08. International Conference on Computational Sciences and Its Applications, 2008, pages 331–341. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 2008.

Li:2008:MLB

[LMRS08b] Xin Li, Marc Moreno Maza, Raqeeb Rasheed, and Éric Schost. The modpn library: bringing fast polynomial arithmetic into MAPLE. ACM Communications in Computer Algebra, 42(3):172–174, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Li:2011:MLB

[LMRS11] Xin Li, Marc Moreno Maza, Raqeeb Rasheed, and Éric Schost. The modpn library: Bringing fast polynomial arithmetic into Maple. Journal of Symbolic Computation, 46 (7):841–858, July 2011. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S074771711000146X.

Li:2007:VGP

[LMS07] Xin Li, Marc Moreno Maza, and Éric Schost. On the virtues of generic programming for symbolic computation. In Shi et al. [SvADS07], pages 251–258. CO-DEN LNCSD9. ISBN 3-540-72585-7 (print), 3-540-72586-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.
springerlink.com/content/
978-3-540-72586-2.

Lemaire:2005:RLM

[LMX05] F. Lemaire, M. Moreno Maza, and Y. Xie. The RegularChains library in MAPLE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(3):96– 97, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Lopez:1993:MNA

[Lop93] R. J. Lopez. Maple: The new apprenticeship in applied mathematics. In Grayson [Gra93], pages 559–564. ISBN 0-7803-1482-4, 0-7803-1483-2, 0-7803-1484-0. ISSN 0190-5848. LCCN ????

Lopez:1994:MVM

[Lop94a] R. J. Lopez, editor. Maple V:

mathematics and its application: proceedings of the Maple Summer Workshop and Symposium, Rensselaer Polytechnic Institute, Troy, New York, August 9–13, 1994. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Lopez:1994:MCT

[Lop94b] Robert J. Lopez. Maple via Calculus: a Tutorial Approach. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 0-8176-3771-0. 166 pp. LCCN QA303.5.D37 L66 1994. URL

                                        http://www.birkhauser.com/
                                        Books/lopez/summary.htm.

Li:1994:IMP

[LP94] P. Li and R. L. Peskin. Implementation of a mapping procedure for the domain decomposition for PDEs in Maple. In Lopez [Lop94a], pages 46–54. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Landau:2005:FCS

[LP05] Rubin H. Landau and M. J. Páez. A first course in scientific computing: symbolic, graphical, and numeric problem solving using Maple, Java, Mathematica, and Fortran. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-12183-4 (hardcover). xxiv + 481 pp. LCCN Q183.9 .L36 2005.

Langer:2012:NSS

[LP12] Ulrich Langer and Peter Paule, editors. Numerical and Symbolic Scientific Computing: Progress and Prospects. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. viii + 358 pp.

Lisle:1995:ADS

[LRB95] I. G. Lisle, G. J. Reid, and A. Boulton. Algorithmic determination of structure of infinite Lie pseudogroups of symmetries of PDEs. In Levelt [Lev95], pages 1–6. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/220346/
p1-lisle/.

Lu:1996:SCD

[LS96] X. Y. Lu and S. K. Spurgeon. Symbolic computation for dynamic sliding mode controller design. In IEE, editor, IEE Colloquium on Symbolic Computation for Control (Digest No: 1996/078), 2 April 1996, pages 4/1–4/5. IEE, London, UK, 1996. ISBN ???? LCCN ????

Lee:2004:OPD

[LS04] H. J. (Hyun Jin) Lee and W. E.

Schiesser. Ordinary and partial differential equation routines in C, C++, Fortran, Java^{(}R), Maple^{(}R), and MATLAB^{(}R). Chapman and Hall/CRC, Boca Raton, FL, USA, 2004. ISBN 1-58488-423-1. xiii + 519 pp. LCCN QA371.5.D37 L44 2004.

Li:2011:ARS

[LSYL11] Hongbo Li, Ruiyong Sun, Shoubin Yao, and Ge Li. Approximate rational solutions torational ODEs defined on discrete differentiable curves. In Schost and Emiris [SE11], pages 217–224. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Loreto:2001:CEC

[LTFN01] A. B. Loreto, L. V. Toscani, M. P. Fachin, and M. M.

Negrón. Computing the exact complexity of divide-and-conquer algorithm using Maple. TEMA Tend. Mat. Apl. Comput., 2:125–134, 2001. ISSN 1677-1966. Selecta of the XXIII National Congress of Applied and Computational Mathematics (Portuguese) (Santos, 2000).

Lum:1993:TCM

[Lum93]  L. Lum, editor.  Technol-

ogy in collegiate mathematics: 4th International conference — November 1991, Portland, OR, Proceedings of the International Conference on Technology in Collegiate Mathematics. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-50013-2. LCCN QA11.A1I454 1991.

Lum:1994:PFA

[Lum94] Lewis Lum, editor. Proceedings of the Fifth Annual International Conference on Technology in Collegiate Mathematics: William Rainey Harper College and Northern Illinois University, Rosemont, Illinois, November 12–15, 1992, Proceedings of the Annual International Conference on Technology in Collegiate Mathematics. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-54304-4. LCCN QA11.A1I454 1992.

Levi:1988:SNP

[LW88] D. Levi and P. Winternitz, editors. Symmetries and nonlinear phenomena: proceedings of the International School on Applied Mathematics, Paipa (Colombia), Feb. 22–26, 1988, volume 9 of CIF Ser. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1988. ISBN 9971-5-0663-7. LCCN QC20.7.N6 I54 1988.

Lopez:1994:LER

[LW94] Robert J. Lopez and Lyle P.

Wiedeman. Laplace’s equation on a rectangle: a Maple-AVS connection. IFIP Transactions. A. Computer Science and Technology, A-48(????):155– 165, 1994. CODEN ITATEC. ISSN 0926-5473.

Lynch:2001:DSA

[Lyn01] Stephen Lynch. Dynamical systems with applications using MAPLE. Birkhäuser Boston Inc., Cambridge, MA, USA, 2001. ISBN 0-8176-4150-5, 3-7643-4150-5. xiii + 398 pp. LCCN QA614.8.L96 2001.

Lynch:2009:DSA

[Lyn09] Stephen Lynch. Dynamical systems with applications using Maple. Springer, New York, NY, USA, second edition, 2009. ISBN 0-8176-4389-3. ???? pp. LCCN ????

MacCallum:1994:ACG

[M^{+}94] M. A. H. MacCallum et al., editors. Algebraic computing in general relativity: lecture notes from the First Brazilian School on Computer Algebra, July 1989, Rio de Janeiro, Brazil, volume 2 of Lecture Notes — Brazilian School on Computer Algebra. Clarendon Press, New York, NY, USA, 1994. ISBN 0-19-853646-1. LCCN QC173.6.B73 1989.

Madore:1992:RAM

[Mad92] Blair F. Madore. Report on the 10th Annual Maple Retreat. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26(3):7–8, September 1992. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Malle:1993:PGG

[Mal93]  Gunter Malle.  Polynome

mit Galoisgruppen PGL_{2}(p) und PSL_{2}(p) über Q(t) (German) [Polynomials with Galois groups PGL_{2}(p) and PSL_{2}(p) over Q(t) ]. Communications in Algebra, 21(2):511–526, 1993. CODEN COALDM. ISSN 0092-7872 (print), 1532-4125 (electronic).

Malyshkin:2001:PCT

[Mal01] Victor E. Malyshkin, editor. Parallel computing technologies: 6th International Conference, PaCT 2001, Novosibirsk, Russia, September 3–7, 2001: proceedings, volume 2127 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42522-5 (paperback). LCCN QA267.A1 L43 no.2127; QA76.58 .I547 2001.

URL http://link.springer-
ny.com/link/service/series/
0558/tocs/t2127.htm.

Maloney:2006:MSL

[Mal06] John Maloney. A Maple Supplement for Linear Algebra. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2006. ISBN 0-13-145337-8. 180 (est.) pp. LCCN ????

Mansfield:2002:EIM

[Man02] Paul Mansfield. Essential ingredients for mathematical content deployment on the Web: The MapleNet experience. In Anonymous [Ano02b], page ?? ISBN ???? LCCN ????

Scott:1987:MTN

[MAP87]  Tony Scott, editor.  The

Maple Technical Newsletter, 1987. ISSN 1061-5733. URL

http://www.birkhauser.com/
Journals/MapleTech/SUMMARY.

HTM. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland. Published twice annually.

Maplesoft:2003:MGS

[Map03a] Maplesoft. Maple 9 Getting Started Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2003. ISBN 1-894511-41-7. iv + 35 pp. LCCN ???? URL

http://www.maplesoft.com/
products/maple/manuals/GettingStartedGuide.
pdf.

Maplesoft:2003:MLG

[Map03b] Maplesoft. Maple 9 Learning Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2003. ISBN 1-894511-42-5. viii + 324 pp. LCCN ???? URL

http://www.maplesoft.com/
products/maple/manuals/LearningGuide.

pdf. Based in part on the work of B. W. Char.

Maplesoft:2004:MGS

[Map04] Maplesoft. Maple 9.5 Getting Started Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2004. ISBN 1-894511-56-5. iv + 38 pp. LCCN ???? URL

http://www.maplesoft.com/
products/maple/manuals/GettingStartedGuide.
pdf.

Maplesoft:2005:MGS

[Map05a]  Maplesoft.  Maple Getting

Started Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2005. ISBN 1-894511-74-3. viii + 88 pp. LCCN ???? URL

http://www.maplesoft.com/
products/maple/manuals/GettingStartedGuide.
pdf.

Maplesoft:2005:MUM

[Map05b] Maplesoft. Maple User Manual. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2005. ISBN 1-894511-75-1. xvi + 398 pp. LCCN ????

Maplesoft:2007:MGS

[Map07a]  Maplesoft.  Maple Getting

Started Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2007. ISBN 1-894511-74-3. viii + 88 pp. LCCN ???? URL

http://www.maplesoft.com/
products/maple/manuals/GettingStartedGuide.
pdf.

Maplesoft:2007:MUM

[Map07b] Maplesoft. Maple User Manual. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2007. ISBN 1-897310-20-X. xvi + 396 pp. LCCN ????

Maplesoft:2008:MGS

[Map08a]  Maplesoft.  Maple Getting

Started Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2008. ISBN 1-897310-49-8. xvi + 116 pp. LCCN ???? URL

http://www.maplesoft.com/
products/maple/manuals/GettingStartedGuide.
pdf.

Maplesoft:2008:MUM

[Map08b] Maplesoft. Maple User Manual. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2008. ISBN 1-897310-48-X. xviii + 400 pp. LCCN ????

Maplesoft:2009:MUM

[Map09] Maplesoft. Maple User Manual. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2009. ISBN 1-897310-69-2. xvi + 524 pp. LCCN ????

Maplesoft:2010:MUM

[Map10] Maplesoft. Maple User Manual. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2010. ISBN 1-897310-89-7. xvi + 442 pp. LCCN ????

Maplesoft:2011:MUM

[Map11] Maplesoft. Maple User Manual. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2011. ISBN 1-926902-07-6. xx + 442 pp. LCCN ????

Maplesoft:2012:MUM

[Map12] Maplesoft. Maple User Manual. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2012. ISBN 1-926902-23-8. xix + 438 pp. LCCN ????

Mathews:1989:UCA

[Mat89]  J. H. Mathews.  Using a

computer algebra system to teach second order differential equations. CoED, 9(4):7–10, October–December 1989. CO-DEN CWLJDP. ISSN 0736-8607.

Mathews:1990:UCS

[Mat90] J. Mathews. Using computer symbolic algebra in applied engineering mathematics to teach Fourier series. CoED, 10(1):47– 52, January–March 1990. CO-DEN CWLJDP. ISSN 0736-8607.

Mavaddat:1993:RSP

[Mav93] F. Mavaddat. The role of a symbolic programming language in hardware verification: The case of Maple. In Lee [Lee93], pages 176–187. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

May:2009:UMW

[May09]  Mike May, S. J.  Using

Maple worksheets to enable student explorations of cryptography. Cryptologia, 33(2):151– 157, 2009. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Meade:1998:ETM

[MB98] Douglas B. Meade and Etan Bourkoff. Engineer’s toolkit: Maple V for engineers. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-8053-6445-5. vi + 154 pp. LCCN ????

Mutrie:1992:AFE

[MBC92] Mark P. W. Mutrie, Richard H.

Bartels, and Bruce W. Char. An approach for floating-point error analysis using computer algebra. In Wang [Wan92], pages 284–293. ISBN 0-89791-489-9. LCCN QA 76.95

I59 1992. URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/143242/
p284-mutrie/.

McDowell:1997:CGM

[McD97] Sean A. C. McDowell. Computation of general multipole moment expansions for N atoms by MAPLE. International Journal of Quantum Chemistry, 62(4):343–351, April 15, 1997. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=42521;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=42521&PLACEBO=IE.pdf.

McGeer:1984:DIB

[McG84] Patrick C. McGeer. A discussion and implementation of Brown’s REX simplification algorithm. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(1):10–19, February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Monagan:2013:IPI

[MCG13] Michael Monagan, Gene Cooperman, and Mark Giesbrecht, editors. ISSAC 2013: Proceedings of the 2013 International Symposium on Symbolic and Algebraic Computation, June 26– 29, 2013, Boston, MA, USA. ACM Press, New York, NY 10036, USA, 2013. ISBN 1-4503-2059-7. LCCN QA76.95 .I59 2013.

McLaughlin:1993:CMV

[McL93] R. McLaughlin. Calculus and Maple V. Saunders College Publishing, Ft. Worth, TX, USA, 1993. ISBN 0-03-096778-3. ???? pp.

McLean:1994:NES

[McL94a] W. McLean. Numerical evaluation of some trigonometric series. Mathematics of Computation, 63(207):271–275, 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McLenaghan:1994:MAG

[McL94b] R. G. McLenaghan. MAPLE applications to general relativity. In MacCallum et al. [M^{+}94], pages 265–266. ISBN 0-19-853646-1. LCCN QC173.6.B73 1989.

Monagan:1996:PMV

[MD96] Michael B. (Michael Burnett)

Monagan and John Stanley Devitt. Programmieren mit Maple V. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60544-4. xi + 400 pp. LCCN ????

Magron:2018:RMP

[ME18] Victor Magron and Mohab Safey

El Din.  RealCertify: a

Maple package for certifying non-negativity. ACM Communications in Computer Algebra, 52(2):34–37, June 2018. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Meade:1996:BRM

[Mea96] Douglas B. Meade. Book reviews: The Maple V handbook, by Martha L. Abell and James P. Braselton. Mathematics of Computation, 65(216): ??, October 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/
jourcgi/jour-pbprocess?fn=
110&arg1=S0025-5718-96-00768-
5&u=/mcom/1996-65-216/.

Mahmood:1994:AAS

[MG94] M. F. Mahmood and T. L.

Gill. Analytical approaches to solving coupled nonlinear Schrödinger equations using Maple V. In Lopez [Lop94a], pages 83–89. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Monagan:1996:MVP

[MGH^{+}96] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, and S. Vorkoetter. Maple V Programming Guide. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94537-7, 0-387-94576-8. xii + 379 pp. LCCN QA76.95.M363 199. US$34.00. With the editorial assistance of J. S. Devitt, M. L. Hansen, D. Redfern, and K. M. Rickard.

Monagan:1998:MVP

[MGH^{+}98] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, and S. M. Vorkoetter. Maple V Programming Guide. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98398-8 (softcover), 0-387-98400-3 (boxed). xi + 379 pp. LCCN QA155.7.E4M363 1998. With the assistance of J. S. Devitt, M. L. Hansen, D. Redfern, and K. M. Rickard.

Monagan:2001:MPG

[MGH^{+}01] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M. Vorkoetter, J. McCarron, and P. DeMarco. Maple 7 Programming Guide. Waterloo Maple Inc., Waterloo, ON, Canada N2L 6C2, 2001. xii + 630 pp. LCCN ????

Monagan:2002:MAP

[MGH^{+}02a] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M.

Vorkoetter, J. McCarron, and P. DeMarco. Maple 8 Advanced Programming Guide. Waterloo Maple Inc., Waterloo, ON, Canada N2L 6C2, 2002. ISBN 1-894511-28-X. x + 382 pp. LCCN ????

Monagan:2002:MIP

[MGH^{+}02b] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M. Vorkoetter, J. McCarron, and P. DeMarco. Maple 8 Introductory Programming Guide. Waterloo Maple Inc., Waterloo, ON, Canada N2L 6C2, 2002. ISBN 1-894511-27-1. xii + 380 pp. LCCN ????

Monagan:2003:MAP

[MGH^{+}03a] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M. Vorkoetter, J. McCarron, and P. DeMarco. Maple 9 Advanced Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2003. ISBN 1-894511-44-1. x + 444 pp. LCCN ????

Monagan:2003:MIP

[MGH^{+}03b] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M. Vorkoetter, J. McCarron, and P. DeMarco. Maple 9 Introductory Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2003. ISBN 1-894511-43-3. x + 388 pp. LCCN ????

Monagan:2005:MAP

[MGH^{+}05a] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M.

Vorkoetter, J. McCarron, and P. DeMarco. Maple Advanced Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2005. ISBN 1-894511-77-8. x + 442 pp. LCCN ????

Monagan:2005:MIP

[MGH^{+}05b] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M. Vorkoetter, J. McCarron, and P. DeMarco. Maple Introductory Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2005. ISBN 1-894511-76-X. x + 388 pp. LCCN ????

Monagan:2007:MAP

[MGH^{+}07a] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M. Vorkoetter, J. McCarron, and P. DeMarco. Maple Advanced Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2007. ISBN 1-894511-77-8. x + 442 pp. LCCN ????

Monagan:2007:MIP

[MGH^{+}07b] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M. Vorkoetter, J. McCarron, and P. DeMarco. Maple Introductory Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2007. ISBN 1-897310-17-X. x + 388 pp. LCCN ????

Monagan:2008:MAP

[MGH^{+}08a] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M. Vorkoetter, J. McCarron, and P. DeMarco. Maple Advanced Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2008. ISBN 1-897310-47-1. x + 442 pp. LCCN ????

Monagan:2008:MIP

[MGH^{+}08b] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M. Vorkoetter, J. McCarron, and P. DeMarco. Maple Introductory Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2008. ISBN 1-897310-46-3. x + 388 pp. LCCN ????

Monagan:2009:MAP

[MGH^{+}09a] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M. Vorkoetter, J. McCarron, and P. DeMarco. Maple Advanced Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2009. ISBN 1-897310-72-2. x + 442 pp. LCCN ????

Monagan:2009:MIP

[MGH^{+}09b] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M. Vorkoetter, J. McCarron, and P. DeMarco. Maple Introductory Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2009. ISBN 1-897310-73-0. x + 388 pp. LCCN ????

Monagan:2010:MAP

[MGH^{+}10a] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M. Vorkoetter, J. McCarron, and P. DeMarco. Maple Advanced Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2010. ISBN 1-897310-96-X. x + 442 pp. LCCN ????

Monagan:2010:MIP

[MGH^{+}10b] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M. Vorkoetter, J. McCarron, and P. DeMarco. Maple Introductory Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2010. ISBN 1-897310-95-1. x + 388 pp. LCCN ????

Murray:1993:EPD

[MH93] Carl D. Murray and David Harper. Expansion of the Planetary Disturbing Function to Eighth Order in the Individual Orbital Elements, volume 15 of QMW Maths Notes. School of Mathematical Sciences, Queen Mary and Westfield College (University of London), Mile End Road, London E1 4NS, England, 1993. ISBN 0-902480-14-6. vii + 436 pp. LCCN QB603.M6 M87 1993. UD£13.00. The machinereadable version is available via anonymous ftp from ei-

ther galois.maths.qmw.ac.uk
(138.37.80.15) or euclid.
maths.qmw.ac.uk (138.37.80.

16) in the directory pub/expan. The book was generated automatically from the Maple/ Mathematica output to produce a Plain TEX file.

Michler:1990:SPC

[Mic90] Gerhard O. Michler. Some problems in computational representation theory. computational group theory, part I. Journal of Symbolic Computation, 9(5-6):571–582, 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Minimair:2004:MMR

[Min04]  Manfred Minimair.  MR:

Macaulay Resultant package for Maple. ACM SIGPLAN Notices, 39(4):26–29, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mittermaier:2000:PAC

[Mit00] Christian Mittermaier. Parallel algorithms in constructive algebraic geometry. Diploma thesis, RISC-Linz, Johannes Kepler University, Linz, Austria, March 2000. vi + 107 pp.

URL ftp://ftp.risc.uni-
linz.ac.at/pub/techreports/
2000/00-07.ps.gz.

Meade:2009:GSM

[MMCK09] Douglas B. Meade, Michael May, S. J., C.-K. Cheung, and G. E. Keogh. Getting started with Maple. Wiley, New York, NY, USA, third edition, 2009. ISBN 0-470-45554-3. x + 208 pp. LCCN QA76.95 .C46 2009. URL

http://www.loc.gov/catdir/
enhancements/fy1117/2011290786-
d.html;   http://www.loc.
gov/catdir/enhancements/fy1117/
2011290786-t.html.

Monagan:1993:GAD

[MN93] M. B. Monagan and W. M. Neuenschwander. GRADIENT, algorithmic differentiation in Maple. In Bronstein [Bro93], pages 68–76. ISBN 0-89791-604-2. LCCN QA76.5.I613 1993. ACM order number: 505930.

Mahyat:1995:UMD

[MNEA95] M. Mahyat, A. Naifi, D. Elhajji, and M. Ankrim. Using Maple to drive the closed-form identification model of robotics manipulator. Modelling, Measurement & Control B: Solid & Fluid Mechanics & Thermics, Mechanical Systems, 58 (1-2):53–63, 1995. CODEN MMBTEC. ISSN 0761-2516.

Merceille:1997:EMC

[MNP97] Jean-Bernard Merceille, Alain-F Nizard, and Marc Peronnet. Exercices de Maple: calculs, traces, programmation SUP et SPE: résumés de cours, 81 exercices et problèmes avec solutions. Flash. Prepa. Masson Editeur, Paris, France, 1997. ISBN 2-225-85444-0. 188 pp. LCCN ????

Molinder:1994:SPU

[Mol94] J. I. Molinder. Signal processing using Maple V. In Lopez [Lop94a], pages 39–45. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Moller:2007:OLC

[Möl07] Karl Dieter Möller. Optics: learning by computing with examples using MathCAD, Matlab, Mathematica, and Maple. Undergraduate texts in contemporary physics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2007. ISBN 0-387-69492-7, 0-387-26168-0. xvi + 453 pp. LCCN QC381 .M66 2007eb.

Monagan:1992:HIT

[Mon92] Michael B. Monagan. A heuristic irreducibility test for univariate polynomials. Journal of Symbolic Computation, 13 (1):47–58 (or 47–57??), January 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Monagan:1993:GFC

[Mon93] Michael B. Monagan. von zur Gathen’s factorization challenge. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27(2):13–18, April 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Monagan:1994:WCW

[Mon94]  M. Monagan.  Worksheets:

Can we teach mathematical algorithms with them? In Lopez [Lop94a], pages 187–196. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Monagan:1996:MSI

[Mon96] M. Monagan. MapleTech special issue. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(1):63, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Monagan:1997:PAM

[MP97] M. B. Monagan and Nicolas Puech. Programmer avec Maple V. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-60734-X. 395 pp. LCCN ????

Monagan:2010:SPM

[MP10] Michael Monagan and Roman Pearce. Sparse polynomial multiplication and division in Maple 14. ACM Communications in Computer Algebra, 44(4):205–209, December 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Monagan:2012:PNP

[MP12] Michael Monagan and Roman Pearce. POLY: a new polynomial data structure for Maple 17. ACM Communications in Computer Algebra, 46(3/4): 164–167, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Monagan:2014:DMS

[MP14] Michael Monagan and Roman Pearce. The design of Maple’s sum-of-products and POLY data structures for representing mathematical objects. ACM Communications in Computer Algebra, 48(3/4):166–186, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Monagan:2016:FBR

[MP16] Michael Monagan and Roman Pearce. Fermat benchmarks for rational expressionals in Maple. ACM Communications in Computer Algebra, 50(4):188–190, December 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Mathews:1994:PIU

[MS94a] David M. Mathews and Keith Schwingendorf. Precalculus Investigations using Maple V. HarperCollins College Publishers, New York, NY, USA, 1994. ISBN 0-673-99410-4. ???? pp. LCCN QA331 .M4317 1994.

Morel:1994:VMI

[MS94b] Jean-Michel Morel and Sergio Solimini. Variational Methods in Image Segmentation: with Seven Image Processing Experiments, volume 14 of Progress in Nonlinear Differential Equations and Their Applications. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 0-8176-3720-6, 3-7643-3720-6. xvi + 245 pp. LCCN TA1637.M67 1995. US$59.50. URL http:/

/www.birkhauser.com/Books/
morel/summary.htm.

Mittermaier:2000:PSN

[MSW00] Christian Mittermaier, Wolfgang Schreiner, and Franz Winkler. A parallel symbolicnumerical approach to algebraic curve plotting. In Ganzha et al. [GMV00], page ?? ISBN 3-540-41040-6. LCCN ???? URL

http://www.risc.uni-linz.
ac.at/people/schreine/papers/
casc2000.ps.gz;   http://
www14.informatik.tu-muenchen.
de/CASC2000.

Muller:1997:EFA

[Mul97] Jean-Michel Muller. Elementary Functions: Algorithms and Implementation. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1997. ISBN 0-8176-3990-X. xv + 204 pp. LCCN QA331.M866

1997.  US$59.95.  URL

http://www.birkhauser.com/
cgi-win/ISBN/0-8176-3990-
X;   http://www.ens-lyon.
fr/~jmmuller/book_functions.
html.

Muller:2006:EFA

[Mul06] Jean-Michel Muller. Elementary Functions: Algorithms and Implementation. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, second edition, 2006. ISBN 0-8176-4372-9. xxii + 266 pp. LCCN QA331 .M866 2006. US$59.95. URL http:

//perso.ens-lyon.fr/jean-
michel.muller/SecondEdition.
html;   http://www.springer.
com/sgw/cda/frontpage/0,,
4-40109-22-72377986-0,00.
html.

Muller:2016:EFA

[Mul16] Jean-Michel Muller. Elementary Functions: Algorithms and Implementation. Birkhäuser Boston Inc., Cambridge, MA, USA, third edition, 2016. ISBN 1-4899-7981-6 (print), 1-4899-7983-2 (e-book). xxv + 283 pp. LCCN QA331 .M866 2016.

Murray:1994:CAT

[Mur94] D. Murray. Computer algebra as a tool for analyzing nonlinear systems. In Lopez [Lop94a], pages 199–208. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Murariu:2007:MPBb

[Mur07a]  Gabriel Murariu.  MAPLE

package for bosons stars on curved manifolds. In Cristiana Dumitrache, Nedelia A. Popescu, Marian Doru Suran, and Vasile Mioc, editors, Fifty Years of Romanian Astrophysics, volume 895 of AIP Conference Proceedings, pages 333–336. American Institute of Physics, Woodbury, NY, USA, 2007. ISBN 0-7354-0400-3. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. LCCN ???? URL http://

scitation.aip.org/dbt/dbt.
jsp?KEY=APCPCS&Volume=895&
Issue=1.

Murariu:2007:MPBa

[Mur07b] Gabriel Murariu. MAPLE procedures for boson fields system on curved space-time. In Serkant Ali Cetin and Iskender Hikmet, editors, BPU’6: Sixth International Conference of the Balkan Physical Union, Istanbul, Turkey, 22–26 August 2006, volume 899 of AIP Conference Proceedings, pages 719–?? American Institute of Physics, Woodbury, NY, USA, 2007. ISBN 0-7354-0404-6. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

LCCN ???? URL http://www.
springer.com/physics/book/
978-0-7354-0404-5.

Miao:2014:PMP

[MWCY14] Qian Miao, Yunhu Wang, Yong Chen, and Yunqing Yang. PDE-BellII: a Maple package for finding bilinear forms, bilinear Bäcklund transformations, Lax pairs and conservation laws of the KdV-type equations. Computer Physics Communications, 185(1):357–367, January 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465513003068.

Nagata:1993:NIT

[Nag93] Wayne Nagata. Nonhyperbolic invariant tori in a double Hopf bifurcation with reflection symmetry. Nonlinearity (Bristol), 6(2):259–275, 1993. CO-DEN NONLE5. ISSN 0951-7715 (print), 1361-6544 (electronic).

Naylor:2002:IBO

[Nay02] Bill Naylor. Interactions between OpenMath and MathML under the Maple environment. In Anonymous [Ano02b], page ?? ISBN ???? LCCN ????

Nesterov:1993:PIC

[Nes93] V. M. Nesterov, editor. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification. Held in Lafayette, Louisiana, February 25–March 1, 1993, volume 2 of Interval Comput./Interval. Vychisl. Moscow Institute for New Technologies in Education, Moscow, Russia, 1993.

Nguyen:1993:SCC

[Ngu93] T. T. Nguyen. Symbolic computation in computable general equilibrium modeling. In Lee [Lee93], pages 144–156. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Nguyen:1994:SCM

[Ngu94] T. Nguyen. Symbolic Computation with Maple in Economic Modeling. ????, ????, 1994. ISBN ???? ???? pp. LCCN ????

Najafi:2005:SAM

[NS05] H. Saberi Najafi and M. Shams.

Shifting algorithms with Maple and implicit shift in the QR algorithm. Applied Mathematics and Computation, 161(3):947– 962, February 25, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nicolaides:1996:MCI

[NW96] Roy A. Nicolaides and Noel Walkington. Maple: a comprehensive introduction. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56230-9. xix + 466 pp. LCCN QA76.95.N53 1996. US$39.95. Covers Maple V.4.

Nicolaides:1997:MCI

[NWF97] Roy Nicolaides, Noel Walkington, and Patrick Fitzpatrick. Maple: a comprehensive introduction. SIAM Review, 39(4): 792–??, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Okano:1992:NRA

[Oka92] Takeshi Okano. A note on the rational approximations to e. Tokyo journal of mathematics, 15(1):129–133, 1992. ISSN 0387-3870.

Ogunye:1995:CSG

[OP95] Ayowale B. Ogunye and Alexander Penlidis. Computation of system Gramians and balanced realizations using Maple V. International Journal of Systems Science, 26(4):899–926, April 1995. CODEN IJSYA9. ISSN 0020-7721 (print), 1464-5319 (electronic).

Oprea:2000:MSF

[Opr00] John Oprea. The mathematics of soap films: explorations with Maple, volume 10 of Student mathematical library. Amer. Math. Soc., Providence, RI, USA, 2000. ISBN 0-8218-2118-0 (softcover). xiii + 266 pp. LCCN QA644 .O67 2000.

Ozaki:1995:CFD

[OS95] Hideshi Ozaki and Kiyoshi Shirayanagi. Calculation of floating decimal Gröbner bases using Maple’s interval package. Sūrikaisekikenkyūsho Kōkyūroku, 920:38–52, 1995. Research on the theory and applications of computer algebra (Japanese) (Kyoto, 1994).

Osetrova:1996:MPF

[Ose96] T. A. Osetrova. MAPLE— procedures for finding the resultants of systems of nonlinear algebraic equations. In Complex analysis and differential equations (Russian), pages 163–174. Krasnoyarsk. Gos. Univ., Krasnoyarsk, 1996.

Ohrn:1994:PIS

[ÖSZ94] N. Y. Öhrn, J. R. Sabin, and M. C. Zerner, editors. Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory and Computational Methods: held at Ponte Vedra Beach, Florida, February 12– 19, 1994, volume 28 of International Journal of Quantum Chemistry Symposium. Wiley, New York, NY, USA, 1994. CODEN IJQSAF. ISBN 0-471-12125-8. ISSN 0538-821X. LCCN QD453.A1 I58 S95, no.28.

Pakes:1990:SRS

[Pak90] Anthony G. Pakes. On series of reciprocal sums of powers. Austral. Math. Soc. Gaz., 17(3): 61–68, 1990. ISSN 0311-0729 (print), 1326-2297 (electronic).

Parker:1997:MA

[Par97a] Richard Parker. Maple for Algebra. Delmar Publishers, Albany, NY, USA, 1997. ISBN 0-8273-7407-0. xi + 308 pp. LCCN QA155.7.E4P35 1997.

Parker:1997:MBC

[Par97b] Richard Parker. Maple for basic calculus. Delmar Publishers, Albany, NY, USA, 1997. ISBN 0-8273-7408-9. 375 pp. LCCN QA303.5.C65P37 1997.

Parker:1997:MT

[Par97c] Richard Parker. Maple for trigonometry. Delmar, Albany, NY, USA, 1997. ISBN 0-8273-7409-7 (paperback). viii + 235 pp. LCCN QA538.P37 1997.

Parlar:2000:IOR

[Par00]  Mahmut Parlar.  Interac-

tive operations research with Maple: methods and models. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2000. ISBN 0-8176-4165-3, 3-7643-4165-3. xiv + 468 pp. LCCN T57.6. P38 2000. Methods and models.

Pavelle:1985:ACA

[Pav85] Richard Pavelle, editor. Applications of computer algebra (Philadelphia, PA, 1984). Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1985. ISBN 0-89838-173-8. LCCN QA155.7.E4 A67 1985.

Pavelcak:1994:ALA

[Pav94] J. A. Pavelcak. Another look at learning the techniques of elementary integration. In Lopez [Lop94a], pages 179–186. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Polking:2006:DE

[PBA06] John Polking, Albert Boggess, and David Arnold. Differential Equations. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 2006. ISBN 0-13-143738-0. LCCN ????

Portela:2002:FEU

[PC02] A. (Artur) Portela and A. (Abdel) Charafi. Finite elements using Maple: a symbolic programming approach. Engineering online library. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-42986-7. xv + 325 pp. LCCN TA347.F5 P665 2002. Includes CD-ROM.

Pecelli:1991:FMD

[Pec91]  Giampiero Pecelli.  Formal

methods in delay-differential equations. In Watt [Wat91], pages 317–318. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/120694/
p317-pecelli/.

Peterson:1992:MV

[Pet92] N. C. Peterson. Maple V. Laboratory microcomputer, 11(1): 37–38, 1992. CODEN LMICDI. ISSN 0262-2955.

Petitjean:1994:ACS

[Pet94] Sylvain Petitjean. Automating the construction of stationary multiple-point classes. In ACM [ACM94], pages 9–14. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

URL http://www.acm.org:
80/pubs/citations/proceedings/
issac/190347/p9-petitjean/

.

Petcu:2000:NSO

[Pet00a] Dana Petcu. Numerical solution of ODEs with Distributed Maple. Technical Report 00-09, RISC-Linz, Johannes Kepler University, Linz, Austria, February 2000. URL ftp:

//ftp.risc.uni-linz.ac.at/
pub/techreports/2000/00-09.
ps.gz.

Petcu:2000:PDAa

[Pet00b] Dana Petcu. PVMaple: a distributed approach to cooperative work of Maple processes. Technical report, Westers University of Timisoara, Timisoara, Romania, May 2000. URL

http://www.risc.uni-linz.
ac.at/software/distmaple/
misc/PVMaple.ps.gz.

Petcu:2000:PDAb

[Pet00c] Dana Petcu. PVMaple: a distributed approach to cooperative work of Maple processes. Lecture Notes in Computer Science, 1908:216–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
1908/19080216.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
1908/19080216.pdf.

Petcu:2001:SIV

[Pet01a] Dana Petcu. Solving initial value problems with parallel Maple processes. Lecture Notes in Computer Science, 2150:926– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
2150/21500926.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
2150/21500926.pdf.

Petcu:2001:WMM

[Pet01b] Dana Petcu. Working with multiple Maple kernels connected by Distributed Maple or PVMaple. Technical report, Westers University of Timisoara, Timisoara, Romania, March 2001. URL

                                        http://www.risc.uni-linz.
                                        ac.at/software/distmaple/
                                        misc/petcu2001.ps.gz.

Place:1995:UCA

[PF95] Jerry Place and Sue Fitzgerald. Using a computer algebra system (Maple) to teach elementary queueing theory. Computer Applications in Engineering Education, 3(1):65–73, 1995. CO-DEN CAPEED. ISSN 1061-3773 (print), 1099-0542 (electronic).

Pagaran:2006:MPC

[PFG06] J. Pagaran, S. Fritzsche, and G. Gaigalas. Maple procedures for the coupling of angular momenta. IX. Wigner D-functions and rotation matrices. Computer Physics Communications, 174(8):616–630, April 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S001046550600004X.

Pflugel:1997:LVD

[Pfl97] E. Pflügel. On the latest version of DESIR-II. Theoretical Computer Science, 187(1– 2):81–86, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        187&issue=1-2&aid=2593.

Pineda:2003:GOB

[PG03] M. Delgado Pineda and E. A.

Galperin. Global optimization in R^{n} with box constraints and applications: a MAPLE code. Mathematical and computer modelling, 38(1-2):77–97, 2003. CODEN MCMOEG. ISSN 0895-7177 (print), 1872-9479 (electronic).

Pineda:2006:GOG

[PG06] M. Delgado Pineda and E. A.

Galperin. Global optimization over general compact sets by the beta algorithm: a MAPLE code. Computers and Mathematics with Applications, 52 (1–2):33–54, July 2006. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122106001805.

Pineda:2008:MCG

[PG08] M. Delgado Pineda and E. A.

Galperin. Maple code of gamma algorithm for global optimization of uncertain functions over compact robust sets. Computers and Mathematics with Applications, 56(1):90–103, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0898122107008292.

Pineda:2010:MCG

[PG10] M. Delgado Pineda and E. A.

Galperin. MAPLE code for the gamma algorithm for global optimization of uncertain functions in economy and finance. Computers and Mathematics with Applications, 59 (8):2951–2963, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122110001227.

Pineda:2013:MCC

[PGG13] M. Delgado Pineda, E. A. Galperin, and P. Jiménez Guerra. MAPLE code of the cubic algorithm for multiobjective optimization with box constraints. Numerical Algebra, Control and Optimization, 3(3):407–424, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:/

/aimsciences.org/article/
doi/10.3934/naco.2013.3.407.

Phillips:2004:HPL

[Phi04] Andrew T. Phillips. Highlighting programming language issues using mixed language programming in Maple and C. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):85–89, March 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Piecuch:1993:MSC

[Pie93] Piotr Piecuch. MAPLE symbolic computation of the longrange many-body intermolecular potentials: Three-body induction forces between two atoms and a linear molecule. International Journal of Quantum Chemistry, 47(4):261–305, August 15, 1993. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pinchback:1993:WLM

[Pin93] R. Pinchback. Working with large matrices in Maple. In Lee [Lee93], pages 77–86. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Pinchback:1994:DSA

[Pin94] R. M. Pinchback. Denotational semantics applied to the typesetting of Maple expressions. In Lopez [Lop94a], pages 22–27. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Liu:2003:MPF

[pLbL03] Yin ping Liu and Zhi bin Li.

A Maple package for finding exact solitary wave solutions of coupled nonlinear evolution equations. Computer Physics Communications, 155(1):65–76, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Pearce:2007:MLH

[PM07] Roman Pearce and Michael Monagan. A Maple library for high performance sparse polynomial arithmetic. ACM Communications in Computer Algebra, 41(3):110–111, September 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Pye:2012:LSF

[PM12] Cory C. Pye and Colin J. Mercer. On the least-squares fitting of Slater-type orbitals with Gaussians: Reproduction of the STO–NG fits using Microsoft Excel and Maple. Journal of Chemical Education, 89(11): 1405–1410, September 2012. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Preiser:1995:CM

[PM5 ] U. Preiser and N. Mondel. Campusline Maple. tewi-Verlag, ????, 1995 (??). ???? pp.

Piropo:2020:SCS

[PMS20] Marcelo Pirôpo, Fernando Moraes, and Fernando A. N. Santos. Surfing on curved surfaces — the Maple package Surf. Computer Physics Communications, 249(??):Article 107002, April 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S001046551930339X.

Paramonova:1998:ATH

[PN98] O. S. Paramonova and A. V. Niukkanen. Analytic transformations of hypergeometric series by means of a new extension of MAPLE. Programmirovanie, 6:25–26, 1998. CO-DEN PROGD3. ISSN 0132-3474, 0361-7688.

Ponce-Nunez:1991:GHB

[PNG91] E. J. Ponce-Núñez and E. Gamero.

Generating Hopf bifurcation formulae with MAPLE. In Seydel et al. [SSKT91], pages 295–299. ISBN 3-7643-2593-3 (Basel), 0-8176-2593-3 (Boston). LCCN QA372.B532 1991.

Postma:2013:NGT

[Pos13] E. J. Postma. The new group theory package in Maple 17. ACM Communications in Computer Algebra, 47(3–4):174–178, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Petcu:2005:DIG

[PPD05] Dana Petcu, Marcin Paprzycki, and Diana Dubu. Design and implementation of a Grid extension for Maple. Scientific Programming, 13(2):137–149, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pijls:2023:CMM

[PQ23] Henk Pijls and Le Phuong Quan. A computational method with Maple for finding the maximum curvature of a Bézierspline curve. Mathematical and Computational Applications, 28 (2):56–??, April 2023. CO-DEN ???? ISSN 2297-8747.

URL https://www.mdpi.com/
2297-8747/28/2/56.

Perrin-Riou:1998:MAA

[PR98] Bernadette Perrin-Riou. Maple: arithmétique et algèbre. Paris Onze édition, Paris, France, 1998. ISBN 2-87800-142-7. 146 pp. LCCN ????

Prisman:2000:PDS

[Pri00] Eli Prisman. Pricing derivative securities: an interactive, dynamic environment with Maple V and Matlab. Academic Press, New York, NY, USA, 2000. ISBN 0-12-564915-0, 0-12-564916-9 (CD-ROM). xxviii + 754 pp. LCCN HG6024.A3 E44 2000.

Prosper:2000:SPS

[Pro00] Vincent Prosper. SFA, a package on symmetric functions considered as operators over the ring of polynomials for the computer algebra system Maple. Journal of Symbolic Computation, 29(1):83–94, February 2000. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http://www.

idealibrary.com/links/doi/
10.1006/jsco.1999.0291;
http://www.idealibrary.com/
links/doi/10.1006/jsco.1999.
0291/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jsco.
1999.0291/ref.

Pirastu:1995:PCI

[PS95] Roberto Pirastu and Kurt Siegl. Parallel computation and indefinite summation: a ∥MAPLE∥ application for the rational case. Journal of Symbolic Computation, 20(5-6):603– 616, November–December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Portugal:1997:AMG

[PS97] R. Portugal and S. L. Sautú.

Applications of Maple to general relativity. Computer Physics Communications, 105 (2–3):233–253, October 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465597000787.

Putz:1994:UMV

[Put94] J. Putz. Using Maple V in calculus: Getting started affordably at a Liberal Arts College. In Lum [Lum94], pages 803–811. ISBN 0-201-54304-4. LCCN QA11.A1I454 1992.

Putz:2003:MA

[Put03] John F. Putz. Maple animation.

Chapman and Hall/CRC, Boca Raton, FL, USA, 2003. ISBN 1-58488-378-2. xiv + 216 pp. LCCN QA155.7.E4 P88 2003. Includes CD-ROM.

Pfalzgraf:1995:APR

[PW95] Jochen Pfalzgraf and Dongming Wang, editors. Automated practical reasoning: algebraic approaches, Texts Monogr. Symbol. Comput. Springer, Vienna, Austria, 1995. ISBN 3-211-82600-9 (Wien), 0-387-82600-9 (New York). ISSN 0943-853X. LCCN QA76.9.A96 A9 1995.

Postel:1996:ROS

[PZ96] F. Postel and P. Zimmermann. A review of the ODE solvers of Maple, Mathematica, Macsyma and MuPAD. In Carrière [Car96], page ???? ISBN ???? LCCN ???? updated version includes Axiom, Derive and Reduce.

Qesmi:2006:MPC

[QBH06] R. Qesmi, M. Ait Babram, and M. L. Hbid. A Maple program for computing a terms of a center manifold, and element of bifurcations for a class of retarded functional differential equations with Hopf singularity. Applied Mathematics and Computation, 175(2):932–968, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Quan:2018:CMM

[Qua18] Le Phuong Quan. A computational method with MAPLE for a piecewise polynomial approximation to the trigonometric functions. Mathematical and Computational Applications, 23 (4):??, December 2018. CO-DEN ???? ISSN 2297-8747.

                                        URL https://www.mdpi.com/
                                        2297-8747/23/4/63.

Xu:2004:SCP

[qXbL04] Gui qiong Xu and Zhi bin Li. Symbolic computation of the Painlevé test for nonlinear partial differential equations using Maple. Computer Physics Communications, 161(1-2):65– 75, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Rabotin:1996:MPD

[Rab96] V. V. Rabotin. MAPLE—a program for determining the Hodge number of three-dimensional Calabi–Yau varieties. In Complex analysis and differential equations (Russian), pages 175– 183. Krasnoyarsk. Gos. Univ., Krasnoyarsk, 1996.

Rafter:2003:SM

[RAB03] John A. (John Arthur) Rafter, Martha L. Abell, and James P. Braselton. Statistics with Maple. Academic Press, New York, NY, USA, 2003. ISBN 0-12-041556-9. xxv + 822 pp. LCCN QA276.4 .R34 2003. Includes CD-ROM.

Rambach:1998:MVC

[Ram98] Philippe Rambach. Maple V en classes prepas. Ellipses, Paris, France, 1998. ISBN 2-7298-9835-2. 142 pp. LCCN ????

Rao:2002:ANM

[Rao02] Singiresu S. Rao. Applied Numerical Methods for Engineers and Scientists. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-089480-X. xx + 1080 pp. LCCN TA345 .R36 2002.

Ravenscroft:1994:RGF

[Rav94] R. A. Ravenscroft. Rational generating function applications in Maple. In Lopez [Lop94a], pages 122–128. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Redfern:1996:MDL

[RC96] Darren Redfern and Ed Chandler. Maple O.D.E. lab book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94733-7 (softcover). ix + 160 pp. LCCN QA371.5.D37R43 1996.

Redfern:1997:MDL

[RC97] Darren Redfern and Ed Chandler. Maple O.D.E. lab book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94733-7. ix + 160 pp. LCCN QA371.5.D37R43 1996. US$29.95.

Reckziegel:1998:EDM

[Rec98]  Helmut Reckziegel.  Ele-

mentare Differentialgeometrie mit Maple. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1998. ISBN 3-528-06991-0. x + 188 pp. LCCN ????

Redfern:1993:MH

[Red93] Darren Redfern. The Maple Handbook. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-94054-5 (New York), 3-540-94054-5 (Berlin). 497 pp. LCCN QA76.95 .R43 1993. FF 181,00.

Redfern:1994:MH

[Red94] Darren Redfern. The Maple Handbook. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1994. ISBN 0-387-94331-5 (New York), 3-540-94331-5 (Berlin). 531 pp. LCCN QA76.95 .R43 1994.

Redfern:1995:PAU

[Red95] Darren Redfern. The Practical Approach Utilities for Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-14221-5. 400 pp. LCCN QA76.95.R434 1995. US$69.00.

Redfern:1996:MH

[Red96a] Darren Redfern. The Maple Handbook: Maple V Release 4. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94538-5, 0-387-94622-5. 495 pp. LCCN QA76.95.R43 1996. US$29.00.

Redfern:1996:PAU

[Red96b] Darren Redfern. The Practical Approach, Utilities for Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-14225-8. ???? pp. LCCN ????

Reitberger:1996:MAP

[Rei96] Heinrich Reitberger. A MAPLE assisted proof of an assertion in Misiurewicz’s theorem on the Lozi attractor. In Iteration theory (Batschuns, 1992), pages 281–283. World Sci. Publishing, River Edge, NJ, 1996.

Resnick:1992:ASP

[Res92] Sidney I. Resnick. Adventures in Stochastic Processes. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1992. ISBN 0-8176-3591-2. xii + 626 pp. LCCN QA274.R46 1992. US$64.50. URL http://

                                        www.birkhauser.com/Books/
                                        resnick/summary.htm.

Rincon:1995:CCC

[RGM95] Felix Rincon, Alfonsa Garcia, and Angeles Martinez. Calcul cientifico con Maple. RA-MA Editorial, ????, 1995. ISBN 84-7897-199-8. ???? pp. LCCN ????

Rhee:1993:UMA

[Rhe93] N. H. Rhee. Using Maple for asymptotic convergence analysis. In Lee [Lee93], pages 87– 94. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Richards:2002:AMM

[Ric02]  Derek Richards.  Advanced

mathematical methods with Maple. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-77010-6, 0-521-77981-2 (paperback). xv + 862 pp. LCCN QA297 .R495 2002.

Riesel:1994:PNC

[Rie94] Hans Riesel. Prime Numbers and Computer Methods for Factorization, volume 126 of Progress in Mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 0-8176-3743-5. xvi + 464 pp. LCCN QA246.R54 1994. US$69.50. URL http://

www.birkhauser.com/Books/
riesel/summary.htm.

Roanes-Lozano:2014:RFV

[RLAH14] Eugenio Roanes-Lozano, Jose Antonio Alonso, and Antonio Hernando. Revisiting fourvalued logics from Maple using the Logics Explorer package. Mathematics and Computers in Simulation, 104: 31–42, October 2014. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (elec-

tronic). URL https://www.
sciencedirect.com/science/
article/pii/S0378475412001486.

Roanes-Lozano:2011:LAD

[RLHAL11] Eugenio Roanes-Lozano, Antonio Hernando, Jose Antonio Alonso, and Luis M. Laita. A logic approach to decision taking in a railway interlocking system using Maple. Mathematics and Computers in Simulation, 82(1):15–28, September 2011.

CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (elec-

                                        tronic). URL https://www.
                                        sciencedirect.com/science/
                                        article/pii/S0378475410001874.

Roanes-Lozano:2010:CDC

[RLvLRM10] Eugenio Roanes-Lozano, Nicolas van Labeke, and Eugenio Roanes-Macías. Connecting the 3D DGS Calques3D with the CAS Maple. Mathematics and Computers in Simulation, 80 (6):1153–1176, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (elec-

                                        tronic). URL https://www.
                                        sciencedirect.com/science/
                                        article/pii/S0378475409002961.

Rybowicz:1997:MVR

[RM97] Marc Rybowicz and Jean-Pierre Massias. Maple V release 4: un système de calcul formel PC/ Windows — Unix/X-Window. Eyrolles, Paris, France, 1997. ISBN 2-212-08875-2. 207 pp. LCCN ????

Roach:1994:SNE

[Roa94]  Kelly Roach.  Symbolic-

numeric nonlinear equation solving. In ACM [ACM94], pages 278–284. ISBN 0-89791-638-7. LCCN QA76.95.I59

                                        1994.  URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/issac/190347/
                                        p278-roach/.

Roach:1996:HFR

[Roa96] Kelly Roach. Hypergeometric function representations.

In Lakshman Y. N. [Lak96], pages 301–308. ISBN 0-89791-796-0. LCCN QA 76.95

I59 1996. URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/236869/
p301-roach/.

Robertson:1996:EMM

[Rob96] John S. Robertson. Engineering Mathematics with Maple. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-053120-X. xiv + 278 pp. LCCN TA345 .R58 1996.

RodriguezAzara:1992:MPG

[Rod92] Jose Luis Rodriguez Azara. A MAPLE program for the generation of the Lie-series solution of systems of non-linear ordinary differential equations. Computer Physics Communications, 67(3):537–542, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Roennow:1994:IMF

[Roe94]  U. Roennow.  Introducing

Maple to first year engineering students via project study. In Lopez [Lop94a], pages 173–178. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Ronnow:2014:ESC

[Ron14] Troels F. Ronnow. Evaluating a special class of two-body integrals through Maple’s integral routine extension. Computing in Science and Engineering, 16(1):18–23, January/February 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Ross:1996:HSE

[Ros96] Paul W. Ross, editor. The handbook of software for engineers and scientists. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1996. ISBN 0-8493-2530-7. xix + 1557 pp. LCCN QA76.754 .H35 1996. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0646/95006378-
                                        d.html.

Rosen:1997:EDM

[Ros97] Kenneth H. Rosen. Exploring discrete mathematics with Maple. McGraw-Hill, New York, NY, USA, third edition, 1997. ISBN 0-07-054128-0. viii + 387 + 4 pp. LCCN QA76.95.E96 1997.

Rotaru:1997:MAM

[Rot97] Paul Rotaru. Mathématiques avec Maple. Bibliothèque des sciences. Diderot Editeur, Arts et sciences, Paris, France, 1997. ISBN 2-84134-052-X. x + 734 pp. LCCN ????

Rovenski:2000:GCS

[Rov00a] Vladimir Rovenski. Geometry of curves and surfaces with MAPLE. Birkhäuser Boston Inc., Cambridge, MA, USA, 2000. ISBN 0-8176-4074-6. x + 310 pp.

Rovenskii:2000:GCS

[Rov00b] Vladimir Y. Rovenskii. Geometry of curves and surfaces with MAPLE. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2000. ISBN 0-8176-4074-6, 3-7643-4074-6 (Basel). x + 310 pp. LCCN QA448.D38 R68 2000.

Royster:1993:UMC

[Roy93] D. C. Royster. Using Maple and the calculus reform material in the calculus sequence. In Lee [Lee93], pages 43–50. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Raz:2026:EAU

[RSL^{+}26] Tomer Raz, Michael Shalyt, Elyasheev Leibtag, Rotem Kalisch, Shachar Weinbaum, Yaron Hadad, and Ido Kaminer. From Euler to AI: Unifying formulas for mathematical constants. arXiv.org, ??(??):1–82, February 24, 2026. URL https:

//arxiv.org/abs/2502.17533;
https://arxiv.org/pdf/2502.
17533.

Runge:1997:IMT

[Run97] Alan P. Runge. Impact of Maple on the design, instruction and performance in an undergraduate physics mathematical methods course. Thesis (Ph.D.), University of Nebraska, Lincoln, NE, USA, 1997. v + 334 pp.

Ray:1992:UIA

[RW92] Nigel Ray and Colin Wright. Umbral interpolation and the addition/contraction tree for graphs. Discrete Math., 103 (1):67–74, 1992. CODEN DSMHA4. ISSN 0012-365X (print), 1872-681X (electronic).

Raade:1995:MHS

[RW95] Lennart Råde and Bertil Westergren. Mathematics Handbook for Science and Engineering. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1995. ISBN 0-8176-3858-X. 539 pp. LCCN QA41 .R33 1995. US$45.00. URL http://www.

birkhauser.com/Books/rade/
summary.htm.

Ryabenko:1999:MPS

[Rya99] A. A. Ryabenko. The MAPLE package of symbolic constructions of solutions of linear ordinary differential equations as power series. Programmirovanie, 5:71–80, 1999. CO-DEN PROGD3. ISSN 0132-3474, 0361-7688.

Rybowicz:1998:BRB

[Ryb98]  M. Rybowicz.  Book re-

view: Introduction to scientific programming. Computational problem solving using Maple and C by Joseph L. Zachary (Springer, Berlin, 1996) ISBN 0-387-94630-6. Computer Physics Communications, 109 (1):90–92, March 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465598000046.

Sakellariou:2001:EPP

[S^{+}01] Rizos Sakellariou et al., editors. Euro-Par 2001 parallel processing: 7th International Euro-Par Conference, Manchester, UK, August 28– 31, 2001: proceedings, volume 2150 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42495-4 (paperback). LCCN QA76.58 .I553 2001; QA267.A1 L43 no.2150. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t2150.htm.

Sadr:1996:MPS

[Sad96] Amir H. Sadr. A Maple package for a small pattern graph enumerations. Thesis (M.S.), University of Kentucky, Lexington, KY, USA, 1996. vii + 77 pp.

Salvy:1989:EAA

[Sal89] B. Salvy. Examples of automatic asymptotic expansions. Technical Report 114, Inst. Nat. Recherche Inf. Autom., Le Chesnay, France, December 1989. 18 pp.

Salvy:1991:EAA

[Sal91] B. Salvy. Examples of automatic asymptotic expansions.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(2):4–17, April 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Salvy:1993:EPM

[Sal93] Bruno Salvy. Efficient programming in Maple: a case study. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27(2):1–12, April 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sanchez:1996:MMS

[San96] Nestor E. Sanchez. The method of multiple scales: Asymptotic solutions and normal forms for nonlinear oscillatory problems. Journal of Symbolic Computation, 21(2):245–252, February 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Silvester:1994:SGF

[SC94] Peter P. Silvester and Simon V. Chamlian. Symbolic generation of finite elements for skin-effect integral equations. IEEE Transactions on Magnetics, 30(5 pt 2):3594–3596, September 1994. CODEN IEMGAQ. ISSN 0018-9464 (print), 1941-0069 (electronic).

Stosic:2009:UMM

[SC09] Borko D. Stosić and Gauss M.

Cordeiro. Using Maple and Mathematica to derive bias corrections for two parameter distributions. Journal of Statistical Computation and Simulation, 79(6):751–767, 2009. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Scheer:1993:KBF

[Sch93a] Axel Scheer. Kurvendiskussion bestimmter funktionen. Studienarbeit, 1993.

Scheftic:1993:IMT

[Sch93b] C. Scheftic. Interactive mathematics texts: Ideas for developers. In Lee [Lee93], pages 51–64. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Schreiner:1998:DMU

[Sch98a]  Wolfgang Schreiner.  Dis-

tributed Maple — user and reference manual. Technical report 98-05, RISC-Linz, Johannes Kepler University, Linz, Austria, May 1998. URL http:

//www.risc.uni-linz.ac.at/
software/distmaple/report;
http://www.risc.uni-linz.
ac.at/software/distmaple/
report/report-main.ps.gz.

Schwartz:1998:IME

[Sch98b] David I. Schwartz. Introduction to Maple for engineers and scientists. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-095133-1. ???? pp. LCCN QA76.95.S33 1998.

Schreiner:1999:DDS

[Sch99a] Wolfgang Schreiner. Developing a distributed system for algebraic geometry (preliminary version). In Topping [Top99], page ?? ISBN 0-948749-59-8. LCCN ???? URL http:

//www.risc.uni-linz.ac.at/
people/schreine/papers/euro99;
 http://www.risc.uni-linz.
ac.at/people/schreine/papers/
euro99.ps.gz;   http://www.
risc.uni-linz.ac.at/people/
schreine/papers/euro99/slides/
slides-main.ps.gz.

Schreiner:1999:PCA

[Sch99b] Wolfgang Schreiner. A parallel computer algebra system based on Maple and Java. Technical report 99-08, RISC-Linz, Johannes Kepler University, Linz, Austria, March 1999.

URL http://www.geocities.
com/%7Etechnionges/Pages/
Sympos.html;   http://www.
risc.uni-linz.ac.at/people/
schreine/papers/swcomm99/
distmaple.rtf;   http://www.
risc.uni-linz.ac.at/people/
schreine/papers/swcomm99/
swcomm99.ps.gz. Presented

at Software for Communication Technologies: Third International Austrian-Israeli Technion Symposium with Industrial Forum, April 26–27, 1999, Hagenberg, Austria.

Schwartz:1999:IME

[Sch99c] David I. Schwartz. Introduction to Maple for engineers and scientists. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-095133-1 (paperback). xvi + 240 pp. LCCN QA76.95 .S33 1999.

Schreiner:2000:APD

[Sch00a] Wolfgang Schreiner. Analyzing the performance of Distributed Maple on various parallel and distributed systems. Technical Report ??, RISC-Linz, Johannes Kepler University, Linz, Austria, 2000. URL http:

//www.risc.uni-linz.ac.at/
people/schreine/papers/distperf.
ps.gz.

Schreiner:2000:MWP

[Sch00b] Wolfgang Schreiner. Managerworker parallelism versus dataflow in Distributed Maple. Technical Report 00-34, RISC-Linz, Johannes Kepler University, Linz, Austria, December 2000. URL http:

//www.risc.uni-linz.ac.at/
people/schreine/papers/distflow2000.
ps.gz.

Schreiner:2001:MWP

[Sch01a] Wolfgang Schreiner. Managerworker parallelism versus dataflow in a distributed computer algebra system. In Malyshkin [Mal01], pages 329–?? ISBN 3-540-42522-5 (paperback). LCCN QA267.A1 L43 no.2127; QA76.58 .I547 2001. URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
2127/21270329.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
2127/21270329.pdf; http://
www.pact.sscc.ru/conference/
pact2001;   http://www.risc.
uni-linz.ac.at/people/schreine/
papers/pact2001.ps.gz.

Schreiner:2001:PBP

[Sch01b] Wolfgang Schreiner. Parallelizing the Big Prime Berlekamp Algorithm with Distributed Maple. Technical Report 01-15, RISC-Linz, Johannes Kepler University, Linz, Austria, July 2001. URL http:

//www.risc.uni-linz.ac.at/
people/schreine/papers/berlekamp2001.
ps.gz.

Schreiner:2003:DML

[Sch03a]  Wolfgang Schreiner.  Dis-

tributed Maple — lessons learned from parallel computeralgebra in distributed environments. In Anonymous [Ano03b], page ?? ISBN ????

LCCN ???? URL http://www.
math.clemson.edu/%7Esgao/
ECCAD03;   http://www.risc.
uni-linz.ac.at/people/schreine/
papers/eccad2003.pdf.

Schwartz:2003:IM

[Sch03b] David I. Schwartz. Introduction to Maple 8. Pearson Education, Upper Saddle River, NJ, USA, 2003. ISBN 0-13-032844-8. xix + 252 pp. LCCN QA155.7.E4 S39 2003.

Schwalbe:1995:DEU

[Sch5 ] D. Schwalbe. Differential Equations Using Maple. Brooks/ Cole, Pacific Grove, CA, USA, 1995 (??). ???? pp.

SotoPrieto:1995:ALC

[SCL95] Manuel Jesus Soto Prieto, Vicente Cordoba, and Jose Luis, aut. Algebra lineal: con Matlab y Maple. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-526815-X. viii + 301 pp. LCCN ????

Schori:1996:MLM

[SCMS96] Richard M. Schori, Todd S.

Coffey, Lawrence C. Moore, and David A. (David Alexander) Smith. Maple laboratory manual for Calculus: modeling and application. D. C. Heath and Company, Lexington, MA, USA, 1996. ISBN 0-669-32794-8. vi + 380 pp. LCCN ????

Scott:2001:MES

[Sco01] Bill Scott. Maple for environmental sciences: a helping hand. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-65826-2 (paperback). 406 pp. LCCN GE45.D37 S25 2001.

Scott:1995:CAM

[Sco5 ] T. Scott. Computer Algebra and the Magic of Maple. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1995 (??). ???? pp.

Shizgal:2010:MCC

[SD10] Bernie D. Shizgal and Raouf Dridi. Maple code for the calculation of the matrix elements of the Boltzmann collision operators for mixtures. Computer Physics Communications, 181 (9):1633–1640, September 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465510001657.

Schost:2011:IPI

[SE11] Éric Schost and Ioannis Z. Emiris, editors. ISSAC 2011: Proceedings of the 2011 International Symposium on Symbolic and Algebraic Computation, June 7–11, 2011, San Jose, CA, USA. ACM Press, New York, NY 10036, USA, 2011. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Sadeghimanesh:2022:SSN

[SE22] AmirHosein Sadeghimanesh and Matthew England. An SMT solver for non-linear real arithmetic inside Maple. ACM Communications in Computer Algebra, 56(2):76–79, June 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3572867.3572880.

Seidel:1992:NAT

[Sei92] H.-P. Seidel. New algorithms and techniques for computing with geometrically continuous spline curves of arbitrary degree. Mathematical modelling and numerical analysis = Modelisation mathématique et analyse numérique: M^{2}AN, 26 (1):149–176, 1992. CODEN RMMAEV. ISSN 0764-583X (print), 1290-3841 (electronic). Topics in computer aided geometric design (Erice, 1990). RAIRO Modél.

Seidel:1993:PFG

[Sei93] Hans-Peter Seidel. Polar forms for geometrically continuous spline curves of arbitrary degree. ACM Transactions on Graphics, 12(1):1–34, January 1993. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Scott:1990:SAM

[SF90] Tony C. Scott and Greg J.

Fee. Some applications of Maple symbolic computation to scientific and engineering problems. In Watanabe and Nagata [WN90], pages 302–303. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Shiskowski:2010:LAU

[SF10] Kenneth Shiskowski and Karl Frinkle. Linear algebra using Maple. Pure and applied mathematics. Wiley, New York, NY, USA, 2010. ISBN 0-470-63759-5 (hardcover). ???? pp. LCCN QA185.D37 S45 2010.

Skidmore:1998:GTD

[SH98] Alexandra Skidmore and Margie Hale. A guided tour of differential equations using computer technology. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-592767-6. 100 (est.) pp.

Shonkwiler:2009:MBI

[SH09] Ronald W. Shonkwiler and J. V. Herod. Mathematical biology: an introduction with Maple and Matlab. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2009. ISBN 0-387-70984-3 (paperback), 0-387-70983-5 (hardcover). ISSN 0172-6056. xiii + 551 pp. LCCN QH323.5 .S56 2009.

Steeb:2010:QMU

[SH10] Willi-Hans Steeb and Yorick Hardy. Quantum mechanics using computer algebra: including sample programs in C++, SymbolicC++, Maxima, Maple, and Mathematica. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, second edition, 2010. ISBN 981-4307-16-5 (hardcover). x + 234 pp. LCCN QC174.17.D37 S74 2010.

Shaw:1994:SMD

[Sha94] Stephen J. Shaw. Surface mode deformations on an oscillating bubble. In Blake et al. [BBST94], pages 355–363.

ISBN 0-7923-3008-0. LCCN TA357.5.C38 I88 1994.

Shirayanagi:1993:ACF

[Shi93] Kiyoshi Shirayanagi. An algorithm to compute floating point Gröbner bases. In Lee [Lee93], pages 95–106. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Shirer:1998:MVT

[Shi98] Donald L. Shirer. Maple V turns over a new leaf. Computers in Physics, 12(3):267–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/10.
1063/1.168673.

Shirazi:2005:FPA

[Shi05a] Kourosh H. Shirazi. Five points of accuracy synthesis of motion generator linkages using Maple-V. Applied Mathematics and Computation, 169(2):1142– 1171, October 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shirazi:2005:SLF

[Shi05b] Kourosh H. Shirazi. Synthesis of linkages with four points of accuracy using Maple-V. Applied Mathematics and Computation, 164(3):731–755, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shingareva:2009:MMP

[Shi09] Inna K. Shingareva. Maple and Mathematica: a problem solving approach for mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2009. ISBN 3-211-99431-9 (softcover). xviii + 483 pp. LCCN ????

Sidki:1994:SCG

[Sid94] Said Sidki. Solving certain group equations in PGL(2, k)— a computational approach. XII School of Algebra, Part II (Portuguese) (Diamantina, 1992). Matemática Contemporânea, 7: 59–70, 1994. ISSN 0103-9059.

Siegl:1993:PAS

[Sie93] Kurt Siegl. Parallelizing algorithms for symbolic computation using ∥Maple∥. ACM SIG-PLAN Notices, 28(7):179–186, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Searles:1998:NEM

[SIE98] Debra J. Searles, Dennis J. Isbister, and Denis J. Evans. Nonequilibrium molecular dynamics integrators using Maple. Mathematics and Computers in Simulation, 45(1-2):147–162, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). Non-standard applications of computer algebra (Linz/Wailea, HI, 1996/1997).

Sills:2004:RMP

[Sil04] Andrew V. Sills. RRtools—a Maple package for aiding the discovery and proof of finite Rogers–Ramanujan type identities. Journal of Symbolic Computation, 37(4):415–448, 2004. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Simon:1992:SMS

[Sim92] Barry Simon. Symbolic math software: It’s not just for mainframes anymore. PC Magazine, 11(14):405–??, August 1992. CODEN PCMGEP. ISSN 0888-8507 (print), 1078-8085 (electronic).

Skala:1995:CGV

[Ska95] V. Skala, editor. Computer graphics and visualization: 3rd International conference in Central Europe — February 1995, Pilsen, Czech Republic, volume 2 of International Conference in Central Europe on Computer Graphics and Visualization. University of West Bohemia, Plzen, Czech Republic, 1995.

Schreiner:2001:FTC

[SKB01a] Wolfgang Schreiner, Gabor Kusper, and Karoly Bosa. Fault tolerance for cluster computing based on functional tasks. In Sakellariou et al. [S^{+}01], pages 712–716. ISBN 3-540-42495-4 (paperback). LCCN QA76.58 .I553 2001; QA267.A1 L43 no.2150.

                                        URL http://europar.man.
                                        ac.uk/;   http://www.risc.
                                        uni-linz.ac.at/people/schreine/
                                        papers/europar2001.ps.gz.

Schreiner:2001:IFT

[SKB01b] Wolfgang Schreiner, Gabor Kusper, and Karoly Bosa. Introducing fault tolerance to Distributed Maple. Technical Report 01-03, RISC-Linz, Johannes Kepler University, Linz, Austria, January 2001.

Stroeker:1999:DMM

[SKH99] Roelof J. Stroeker, Johan F. Kaashoek, and L. F. Hoogerheide. Discovering mathematics with Maple: an interactive exploration for mathematicians, engineers, and econometricians. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1999. ISBN 3-7643-6091-7 (Basel), 0-8176-6091-7 (Boston). xvii + 227 pp. LCCN QA76.95.S76 1999.

Shingareva:2007:MMP

[SLC07] Inna Shingareva and Carlos Lizárraga-Celaya. Maple and Mathematica: a problem solving approach for mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-211-73265-9, 3-211-73264-0. xiii + 263 pp. LCCN QA155.7.E4 S55 2007eb.

Schreiner:2003:DMP

[SMB03] Wolfgang Schreiner, Christian Mittermaier, and Karoly Bosa. Distributed Maple: parallel computer algebra in networked environments. Journal of Symbolic Computation, 35(3):305–347, March 2003. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). URL http:/

/www.risc.uni-linz.ac.at/
people/schreine/papers/jsc2002.
ps.gz.

Styer:1997:BRM

[SMC97] Daniel Styer, Susan R. McKay, and Wolfgang Christian. Book review: Marko Horbatsch, Quantum Mechanics Using Maple. Computers in Physics, 11(2):165–??, March 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.
scitation.org/doi/10.1063/
1.4822533.

Scott:1990:PSQa

[SMF^{+}90] T. C. Scott, R. A. Moore, G. J. Fee, M. B. Monagan, G. Labahn, and K. O. Geddes. Perturbative solutions of quantum mechanical problems by symbolic computation: a review. International Journal of Modern Physics C [Physics and Computers], 1(1):53–76, 1990. CODEN IJMPEO. ISSN 0129-1831 (print), 1793-6586 (electronic).

Scott:1992:SAM

[SMFC92] T. C. Scott, M. B. Monagan, G. J. Fee, and R. M. Corless. Some applications of Maple to mathematical, scientific and engineering problems. In Houstis and Rice [HR92], pages 165–176. ISBN 0-444-89703-8. LCCN Q334 .I45 1991.

Scott:1990:PSQb

[SMM^{+}90] T. C. Scott, R. A. Moore, M. B. Monagan, G. J. Fee, and E. R. Vrscay. Perturbative solutions of quantum mechanical problems by symbolic computation. Journal of Computational Physics, 87(2):366–395, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0021999190902583.

Schreiner:2000:AAC

[SMW00a] Wolfgang Schreiner, Christian Mittermaier, and Franz Winkler. Analyzing algebraic curves by cluster computing. In Anonymous [Ano00a], pages 179–188. ISBN ???? LCCN ???? URL http:

                                        //www.risc.uni-linz.ac.at/
                                        people/schreine/papers/dapsys2000.
                                        ps.gz.

Schreiner:2000:PSL

[SMW00b] Wolfgang Schreiner, Christian Mittermaier, and Franz Winkler. On parallelizing a software library for algebraic geometry (invited talk). In Kolev [Kol00], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.risc.uni-linz.ac.at/
                                        people/schreine/papers/numerical99.
                                        ps.gz.

Schreiner:2000:PAS

[SMW00c] Wolfgang Schreiner, Christian Mittermaier, and Franz Winkler. Plotting algebraic space curves by cluster computing. In Gao and Wang [GW00], page ?? ISBN 981-02-4498-3. LCCN QA76.9.M35. URL http:/

/www.risc.uni-linz.ac.at/
people/schreine/papers/ascm2000.
ps.gz.

Schreiner:2001:SPA

[SMW01] Wolfgang Schreiner, Christian Mittermaier, and Franz Winkler. On solving a problem in algebraic geometry by cluster computing. In Bode [Bod01], pages 1196– ?? ISBN 3-540-67956-

1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E865 2000; QA76.58 .I553 2000. URL http:/

/www.risc.uni-linz.ac.at/
people/schreine/papers/europar2000-
long.ps.gz;   http://www.
risc.uni-linz.ac.at/people/
schreine/papers/europar2000.
ps.gz.

Snyder:1995:ACA

[Sny95] R. Snyder. Applications of a computer algebra system in introductory statistics. Behavior research methods, instruments, and computers, 27(2):139–143, May 1995. CODEN BRMCEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Spieth:1994:MVR

[Spi94]  A. Spieth.  Maple V Re-

lease 2 Referenzhandbuch. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1994. ISBN 3-929821-73-7. ???? pp.

Sira-Ramirez:1991:SSM

[SRRB91] Hebertt Sira-Ramirez and Miguel Rios-Bolívar. Synthesis of sliding-mode controllers for nonlinear systems via extended linearization. Dynamics Control, 1(4):379–403, 1991.

Seydel:1991:BCA

[SSKT91] R. Seydel, F. W. Schneider, T. Küpper, and H. Troger, editors. Bifurcation and chaos: analysis, algorithms, applications. Proceedings of the conference held in Würzburg, August 20–24, 1990, volume 97 of International Series of Numerical Mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1991. ISBN 3-7643-2593-3 (Basel), 0-8176-2593-3 (Boston). LCCN QA372.B532 1991.

Smith:1994:UMC

[SSTW94] T. Smith, C. Stolze, P. Tiffany, and K. Weld. Using Maple in the calculus sequence. In Huston and Armel [HA94], pages 220–227.

Stavroulakis:1999:PDE

[ST99] Ioannis P. Stavroulakis and Stepan A. Tersian. Partial Differential Equations: An Introduction With Mathematica and Maple. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1999. ISBN 981-02-3891-6. x + 297 pp. LCCN QA377 .S79 1999.

Stavroulakis:2004:PDE

[ST04] Ioannis P. Stavroulakis and Stepan A. Tersian. Partial differential equations an introduction with Mathematica and MAPLE. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, second edition, 2004. ISBN 981-256-244-3 (electronic book). xii + 306 pp. LCCN QA377 .S79 2004.

Stancheva:2007:DEM

[Sta07]  T. Stancheva.  Differential

equations with MAPLE. In Todorov [Tod07], pages 176–

187. ISBN 0-7354-0460-7. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. LCCN QA299.6.A63 2007.

URL http://proceedings.
aip.org/proceedings/confproceed/
946.jsp.

Steeb:1994:QMU

[Ste94] Willi-Hans Steeb. Quantum mechanics using computer algebra: includes sample programs for REDUCE, MAPLE, MATHE-MATICA and C++. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1994. ISBN 981-02-1770-6. viii + 189 pp. LCCN QC174.17.D37 S74 1994.

Stembridge:1995:MPS

[Ste95]  John R. Stembridge.  A

Maple package for symmetric functions. Journal of Symbolic Computation, 20(5-6):755– 768, November–December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Stenger:1996:BRS

[Ste96] Frank Stenger. Book reviews: Solving Problems in Scientific Computing Using MAPLE and MATLAB, by Walter Gander and Jírí Hrebícek. Mathematics of Computation, 65(214): 880–882, April 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                        URL http://www.ams.org/
                                        jourcgi/jour-pbprocess?fn=
                                        110&arg1=S0025-5718-96-00700-
                                        4&u=/mcom/1996-65-214/.

Stein:2008:CWC

[Ste08] William A. Stein. Can we create a viable free open source alternative to Magma, Maple, Mathematica and Matlab? In Jeffrey [Jef08], pages 5–6. ISBN 1-59593-904-0. LCCN ????

Stein:2013:SCV

[Ste13] William Stein. Sage: creating a viable free open source alternative to Magma, Maple, Mathematica, and MATLAB. In Felipe Cucker, Teresa Krick, Allan Pinkus, and Agnes Szanto, editors, Foundations of computational mathematics, Budapest 2011, volume 403 of London Mathematical Society lecture note series, pages 230–238. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-60407-9 (paperback), 1-139-09540-4 (e-book), 1-139-61690-0. LCCN QA297 .F635 2011.

Stenborg:2025:SMG

[Ste25]  Travis Stenborg.  Styling

Maple graphics with TikZ and LaTEX. TUGboat, 46(3):379– 380, 2025. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb46-3/
tb144stenborg-maple.pdf.

Stoyanov:2003:BRBc

[Sto03] Jordan Stoyanov. Book review: From Elementary Probability to Stochastic Differential Equations with MAPLE(R), by S. Cyganowski; P. Kloeden; J. Ombach. Journal of the Royal Statistical Society. Series D (The Statistician), 52 (3):405–406, ???? 2003. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/
stable/4128217.

Strand:1994:SPF

[Str94a] Bo Strand. Summation by parts for finite difference approximations for d/dx. Journal of Computational Physics, 110 (1):47–67, January 1994. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0021999184710059.

See erratum [Ano95c].

Stroock:1994:CIT

[Str94b]  Daniel W. Stroock.  Con-

cise Introduction to the Theory of Integration. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, second edition, 1994. ISBN 0-8176-3759-1. viii + 184 pp. LCCN QA312.S78 1994. US$24.50. URL http://

                                        www.birkhauser.com/Books/
                                        stroock/summary.htm.

Stroyan:1999:PCL

[Str99] K. D. Stroyan. Projects for calculus: the language of change. Academic Press, New York, NY, USA, second edition, 1999. ISBN 0-12-673031-8. xii + 326 pp. LCCN QA303.5.C65S77 1998.

Sturmfels:1989:CFP

[Stu89] Bernd Sturmfels. Computing final polynomials and final syzygies using Buchberger’s Gröbner bases method. Results in Mathematics = Resultate der Mathematik, 15(3-4): 351–360, 1989. ISSN 0378-6218.

Soto:1996:MCM

[SV96] Manuel J. Soto and Jose Luis Vicente Cordoba. Matematicas con Maple. Addison-Wes-ley Iberoamericana, Wilmington, DE, USA, 1996. ISBN 0-201-65397-4, 0-201-65398-2 (disk). viii + 281 pp. LCCN QA76.95.S67 1996.

Shi:2007:CSIb

[SvADS07] Yong Shi, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational Science — ICCS 2007: 7th International Conference, Beijing, China, May 27 — 30, 2007, Proceedings, Part II, volume 4488 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CO-DEN LNCSD9. ISBN 3-540-72585-7 (print), 3-540-72586-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.
springerlink.com/content/
978-3-540-72586-2.

Sokolsky:1995:GNR

[SVV95] Andrej G. Sokolsky, Akmal A.

Vakhidov, and Nickolay N. Vasiliev. Generating a new recurrence relations system for elliptic Hansen coefficients by means of computer algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(2S (special issue)): 16–18, June 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Stephenson:1990:NSF

[SW90] John Stephenson and Yong Wang. Numerical solution of Feigenbaum’s equation. Applied Mathematics Notes, 15:68–78, 1990. ISSN 0700-9224.

Snodgrass:1994:PAS

[SW94a] Richard T. Snodgrass and Marianne Winslett, editors. Proceedings of the 1994 ACM SIG-MOD International Conference on Management of Data / SIG-MOD ’94, Minneapolis, Minnesota, May 24–27, 1994, volume 23(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2 1994.

Spearman:1994:CSQ

[SW94b] Blair K. Spearman and Kenneth S. Williams. Characterization of solvable quintics x^{5} + ax + b. Amer. Math. Monthly, 101(10):986–992, 1994. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sanders:1998:CMF

[SW98] Jan A. Sanders and Jing Ping Wang. Combining Maple and Form to decide on integrability questions. Computer Physics Communications, 115(2-3):447– 459, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

SymbolicComputation:1988:MV

[Sym88] Symbolic Computation Group Staff. Maple, Version 4.2. Brooks/Cole, Pacific Grove, CA, USA, 1988. ISBN 0-534-10218-2. LCCN ???? US$350.00.

Salvy:1994:GMP

[SZ94a] Bruno Salvy and Paul Zimmermann. GFUN: a Maple package for the manipulation of generating and holonomic functions in one variable. ACM Transactions on Mathematical Software, 20(2):163–177, June 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sethi:1994:HDM

[SZ94b] Suresh P. Sethi and Qing Zhang. Hierarchical Decision Making in Stochastic Manufacturing Systems: Systems & Control: Foundations & Applications. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 0-8176-3735-4. xvi + 419 pp. LCCN TS155.8.S45 1994. US$49.50. URL http:/

/www.birkhauser.com/Books/
sethi/summary.htm.

Stefani:2003:TOB

[SZ03] Gianna Stefani and PierLuigi Zezza. Time optimality of a bang-bang trajectory with Maple. In Lagrangian and Hamiltonian methods for nonlinear control 2003, pages 135– 140. IFAC, Laxenburg, 2003.

Szabo:2002:LAI

[Sza02] Fred Szabo. Linear algebra: an introduction using Maple. Harcourt/Academic Press, San Diego, 2002. ISBN 0-12-680140-1. various pp. LCCN QA184.2 .S93 2002 Stacks SEMSTAX-NBKS.

Taylor:1994:CEM

[TA94] R. Taylor and K. Atherley. Chemical engineering with Maple. In Lopez [Lop94a], pages 149–155. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Taylor:1995:CEM

[TA95] R. Taylor and K. Atherley. Chemical engineering with Maple. CEE. Chemical engineering education, 29(1):56–??, 1995. CODEN CHEDAY. ISSN 0009-2479.

Tocci:1996:AME

[TA96] Christopher Tocci and Steven G. Adams. Applied Maple for engineers and scientists. The Artech House computer science library series. Artech House Inc., Norwood, MA, USA, 1996. ISBN 0-89006-853-4. xvii + 406 pp. LCCN TA345.T63 1996.

Talay:1994:PSP

[Tal94] D. Talay. Presto: a software package for the simulation of diffusion processes. Statistics and Computing, 4(4):247– 252, December 1994. CO-DEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Tangora:1989:CGT

[Tan89] Martin C. Tangora, editor.

Computers in geometry and topology. Proceedings of the Conference on Computers in Geometry and Topology, held Mar. 24–28, l986, at the University of Illinois at Chicago, volume 114 of Lecture Notes in Pure and Appl. Math. Marcel Dekker, Inc., New York, NY, USA, 1989. ISBN 0-8247-8031-0. LCCN QA448.D38 C67 1989.

Tanis:1999:BRM

[Tan99] Elliot A. Tanis. Book review: Maple V Student Version: Release 5 . The American Statistician, 53(4):389–??, November 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tari:1992:RSF

[Tar92] Farid Tari. Recognition of Ksingularities of functions. Experimental mathematics, 1(3): 225–229, 1992. ISSN 1058-6458 (print), 1944-950X (electronic).

Taylor:1998:TMS

[Tay98a] R. Taylor. Thermodynamics with Maple. I. Symbolic computation. Mathematics and Computers in Simulation, 45(1-2): 101–119, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). Nonstandard applications of computer algebra (Linz/Wailea, HI, 1996/1997).

Taylor:1998:TMI

[Tay98b] Ross Taylor. Thermodynamics with Maple. II. Numerical and graphical applications. Mathematics and Computers in Simulation, 45(1-2):121–146, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). Non-standard applications of computer algebra (Linz/Wailea, HI, 1996/1997).

Tefera:2002:MMP

[Tef02]  Akalu Tefera.  MultInt, a

MAPLE package for multiple integration by the WZ method. Journal of Symbolic Computation, 34(5):329–353, November 2002. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tomovic:1994:CAS

[TH94] M. M. Tomovic and V. S. Hillman. Computer algebra systems enhance teaching engineering technology courses. In Grayson [Gra94], pages 161– 164. ISBN 0-7803-2413-7. LCCN ????

Bui:1991:ABE

[tHB91] Thi to Hong Bui. Algorithmen zur Bestimmung Einiger Eigenschaften von Reellen Funktionen. (German) [Algorithms for determining some properties of real functions]. Master’s thesis, Technischen Universität Carolo-Wilhelmina zu Braunschweig, Braunschweig, Germany, 1991.

Thompson:1994:MMC

[Tho94] W. J. Thompson. MATHE-MATICA and MAPLE: Has Champaign met its Waterloo? Computers in Physics, 8(3):269–278, 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Thompson:2013:AIG

[Tho13]  Ian Thompson.  Algorithm

926: Incomplete Gamma functions with negative arguments. ACM Transactions on Mathematical Software, 39(2):14:1– 14:9, February 2013. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Thompson:2017:UM

[Tho17] Ian Thompson. Understanding Maple. Cambridge University Press, Cambridge, UK, 2017. ISBN 1-316-62814-0 (paperback). viii + 228 pp. LCCN QA76.95 .T52 2017.

Todd:1994:SEM

[TMH94] P. Todd, R. McLeod, and M. Harris. A symbolic engineering mechanics system built on OEM Maple. In Lopez [Lop94a], pages 31–38. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Todorov:2007:AME

[Tod07] Michail Dmytrovych Todorov, editor. Applications of mathematics in engineering and economics 33: 33rd International Conference: Sozopol, Bulgaria, 8–14 June 2007, volume 946 of AIP conference proceedings. American Institute of Physics, Woodbury, NY, USA, 2007. ISBN 0-7354-0460-7. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. LCCN QA299.6.A63 2007.

URL http://proceedings.
aip.org/proceedings/confproceed/
946.jsp.

Topping:1999:TEC

[Top99] Barry H. V. Topping, editor. Third Euro-conference on Parallel and Distributed Computing for Computational Mechanics, Weimar, Germany, March 20–25, 1999. Civil-Comp Press, Edinburgh, Scotland, 1999. ISBN 0-948749-59-8. LCCN ????

Taneri:1994:EGT

[TP94] U. Taneri and J. Paldus. Evaluation of group theoretical characteristics using the symbolic manipulation language MAPLE. International Journal of Quantum Chemistry, 52 (S28):139–154, February 12– 19, 1994. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory and Computational Methods.

Taylor-Russell:1994:MAU

[TR94] G. Taylor-Russell. Maple at the University of North London. In Lopez [Lop94a], pages 167–172.

ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Tanik:1994:SSE

[TRC94] Murat M. Tanik, Wilhelm Rossak, and Daniel E. Cooke, editors. Software systems in engineering: presented at the Energy-Sources Technology Conference, New Orleans, Louisiana, January 23–26, 1994, volume 59 of ASME — Publications – PD. ASME, New York, NY, USA, 1994. ISBN 0-7918-1186-7. LCCN QA76.758 .S68 1994.

Tsang:1993:SNN

[Tsa93] C. P. Tsang. Synthesizing neural network applications using computer algebra. In IEEE [IEE93b], pages 919–922. ISBN 0-7803-1421-2. LCCN QA76.87 .I57 1993. Three volumes.

Terrell:1994:BSR

[TT94] Maria S. Terrell and Robert E.

Terrell. Behind the scenes of a random dot stereogram. Amer. Math. Monthly, 101(8): 715–724, 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Tahir:2000:RCM

[TT00] Farasat Tahir and Muhammad Tahir. RSA ciphers with Maple. Punjab Univ. J. Math. (Lahore), 33:145–152, 2000. ISSN 1016-2526.

Thompson:1993:SCS

[TTL93] D. Thompson, T. Trias, and L. Leff. A symbolic CSG system written in Maple V. In Lee [Lee93], pages 188– 199. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Thompson:1994:WSC

[TTLM94] D. Thompson, T. Tries, L. Leff, and Z. Malik. A working 2-D symbolic CSG system written in Maple. In Tanik et al. [TRC94], pages 171–180. ISBN 0-7918-1186-7. LCCN QA76.758 .S68 1994.

Trappe:2006:ICC

[TW06] Wade Trappe and Lawrence C. Washington. Introduction to Cryptography: with Coding Theory. Pearson Prentice Hall, Upper Saddle River, NJ, USA, second edition, 2006. ISBN 0-13-186239-1, 0-13-198199-4 (paperback). xiv + 577 pp. LCCN QA268 .T73 2006.

Ulmer:1995:NKA

[UW95] Felix Ulmer and Jacques-Arthur Weil. Note on Kovacic’s algorithm. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(2):10–11, April 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

VandenBergh:1988:OMP

[Van88]  N. Van den Bergh.  OR-

THOFRAME: a MAPLE package for performing calculations in the orthonormal tetrad formalism. Class. Quant. Grav., 5(10):L169–L179, 1988. CO-DEN CQGRDG. ISSN 0264-9381 (print), 1361-6382 (electronic).

VanEssen:1994:ADN

[Van94a] H. Van Essen. Analysis and design of nonlinear control systems with the symbolic computation system MAPLE. Journal A. Revue A. Tijdschrift A. Zeitschrift A, 35(2):3–??, 1994. CODEN JRNAAD. ISSN 0771-1107.

vanHoeij:1994:ACI

[van94b] Mark van Hoeij. An algorithm for computing an integral basis in an algebraic function field. Journal of Symbolic Computation, 18(4):353–364 (or 353–363??), October 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

vanHoeij:1994:CPR

[van94c] Mark van Hoeij. Computing parameterizations of rational algebraic curves. In ACM [ACM94], pages 187–190. ISBN 0-89791-638-7. LCCN QA76.95.I59

1994.  URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/190347/
p187-van_hoeij/.

Vantu:1994:MPC

[Vân94d] M. Vântu. A Maple program for computing the symbolic reliability polynomial in a directed network. Bull. Transilv. Univ. Braşov Ser. B (N.S.), 1(36):53– 57 (1995), 1994. ISSN 1223-964X.

vanHoeij:1995:ACW

[van95] Mark van Hoeij. An algorithm for computing the Weierstrass normal form. In Levelt [Lev95], pages 90–95. ISBN 0-89791-699-9. LCCN QA 76.95

I59 1995. URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/220346/
p90-van_hoeij/.

Vardi:1995:ISC

[Var95] Ilan Vardi. Introduction to Symbolic Computation. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1995. ISBN 0-387-94147-9. 320 pp. US$49.95. Covers C, Fortran, Maple, and Mathematica.

Vail:2006:SCE

[VAS06] Benjamin Vail, Damian Aris, and Mihai Scarlete. Symbolic computation engines and molecular modeling templates: Maple-assisted point group analysis of the vibrational activity of molecules. International Journal of Quantum Chemistry, 106(1):167–175, ???? 2006. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vinette:1988:PEE

[VČ88] F. Vinette and J. Čížek. Perturbation energy expansion using hypervirial theorem and symbolic computation for the Ndimensional hydrogen atom in an external spherically symmetric field. Computer Physics Communications, 52(1):35–41, 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Vinette:1989:USC

[VC89] F. Vinette and J. Cizek. The use of symbolic computation in solving some nonrelativistic quantum mechanical problems. In Gianni [Gia89], pages 85– 95. ISBN 3-540-51084-2. LCCN QA76.95 .I57 1988. Conference held jointly with AAECC-6.

vanderHoeven:2012:IPI

[vdHvH12] Joris van der Hoeven and Mark van Hoeij, editors. ISSAC 2012: Proceedings of the 2012 International Symposium on Symbolic and Algebraic Computation, July 22–25, 2012, Grenoble, France. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4503-1269-1. LCCN QA76.95 .I59 2012.

Veigneau:1997:SPS

[Vei97] Sébastien Veigneau. SP, a package for Schubert polynomials realized with the computer algebra system MAPLE. Journal of Symbolic Computation, 23(4):413–425, April 1997. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Vetsch:1995:SMP

[Vet95] M. Vetsch. Die Sprache Maple. Probleme, Beispiele, Lösungen. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1995. ISBN 3-929821-22-2. ???? pp.

vanHulzen:1983:CAE

[vH83] J. A. van Hulzen, editor. Computer algebra: EUROCAL ’83, European Computer Algebra Conference, London, England, March 28-30, 1983, volume 162 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 0-387-12868-9 (New York), 3-540-12868-9 (Berlin). LCCN QA155.7.E4 E85 1983.

Velez:2017:LQP

[VHC17] Juan D. Velez, Juan P. Hernandez, and Carlos A. Cadavid. Limits of quotients of polynomial functions in three variables. ACM Communications in Computer Algebra, 51(2):42– 56, June 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Vilenkin:1995:SI

[Vil95] N. Ya. Vilenkin. In Search of Infinity. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1995. ISBN 0-8176-3819-9, 3-7643-3819-9. 145 pp. LCCN QA9.V5513 1995. US$24.50. URL http://

www.birkhauser.com/Books/
vilenkin/summary.htm. Trans-

lated from Russian and edited by Abe Shenitzer, York University, Ontario, Canada.

Vinette:1991:FSC

[Vin91] F. Vinette. Features of symbolic computation exploited in the calculation of lower energy bounds of cyclic polyene models. In Watt [Wat91], pages 458–459. ISBN 0-89791-437-6. LCCN QA 76.95

I59 1991. URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/120694/
p458-vinette/.

Vinette:1992:QSD

[Vin92] F. Vinette. Quasispin symmetry for the derivation of coupled cluster equations for the Hubbard model of benzene. International Journal of Quantum Chemistry, 42(6): 1737–1746, June 20, 1992. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vivaldi:2001:EMM

[Viv01] Franco Vivaldi. Experimental mathematics with Maple. Chapman and Hall/CRC mathematics. Chapman and Hall/CRC, Boca Raton, FL, USA, 2001. ISBN 1-58488-233-6. xi + 223 pp. LCCN QA76.95 .V58 2001.

Vu:2012:FHS

[VJC12] K. T. Vu, G. F. Jefferson, and J. Carminati. Finding higher symmetries of differential equations using the MAPLE package DESOLVII. Computer Physics Communications, 183 (4):1044–1054, April 2012. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465512000070.

Villard:1999:AIA

[VM99] Dominique Villard and Michael B. Monagan. ADrien: an implementation of automatic differentiation in Maple. In Dooley [Doo99], pages 221–228 (electronic). ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999.

                                        URL http://www.acm.org/
                                        pubs/contents/proceedings/
                                        issac/309831/.

Vidal:2014:MPB

[VMRA14] J. Vidal, C. Muriel, J. L. Romero, and J. J. Alonso. A Maple procedure based on λ-symmetries for secondorder ordinary differential equations. Applied Mathematics and Computation, 249(??):147– 163, December 15, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300314013800.

Voda:1982:MPL

[Vod82] P. J. Voda. Maple: a programming language and operating system. In ACM:POPL’82 [ACM82], pages 157–168. ISBN 0-89791-065-6. LCCN QA76.7 .A15 1982. US$28.00.

Volcheck:1993:NTS

[Vol93] Emil J. Volcheck. Noether’s Stransformation simplifies curve singularities rationally: a local analysis. In Bronstein [Bro93], pages 164–172. ISBN 0-89791-604-2. LCCN QA76.5.I613

1993.  URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/164081/
p164-volcheck/. ACM order

number: 505930.

vonRauchhaupt:2000:BSK

[vRFS^{+}00] Ulf von Rauchhaupt, Harald Fritzsch, Klaus Stierstadt, Rainer Kayser, Klaus Stierstadt, Rainer Kayser, Norbert Dragon, Tilman Pfau, Christoph Bennemann, and Holger Perlt. Bücher und Software: Krafft: Vorstoß ins Unerkannte. Lexikon großer Naturwissenschaftler/Weizsäcker: Große Physiker. Von Aristoteles bis Werner Heisenberg/Zee: Fearful Symmetry. The Search for Beauty in Modern Physics/Schroeder: An Introduction to Thermal Physics/Fraser, Lillestol u. Sellevag: Auf der Suche nach dem Unendlichen/Rebhan: Theoretische Physik Band 1: Mechanik, Elektrodynamik, Relativitätstheorie, Kosmologie/Meschede: Optik, Licht und Laser/Mantegna u. Stanley: An Introduction to Econophysics. Correlations and Complexity in Finance/Maple 6. Physikalische Blätter, 56(9):75–79, September 2000. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:

                                        //onlinelibrary.wiley.com/
                                        doi/10.1002/phbl.20000560918/
                                        abstract.

Vulcanov:2002:GTD

[Vul02] Dumitru N. Vulcanov. Gravity, torsion, Dirac field and computer algebra using MAPLE and REDUCE. arxiv.org, page 20, September 25, 2002.

                                        URL http://arxiv.org/abs/
                                        gr-qc/0209096.

Vulcanov:2003:CDE

[Vul03] Dumitru N. Vulcanov. Calculation of the Dirac equation in curved spacetimes with possible torsion using MAPLE and REDUCE. Computer Physics Communications, 154(3):205– 218, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Vrbik:2013:IIS

[VV13] Jan Vrbík and Paul Vrbík. Informal Introduction to Stochastic Processes with Maple. Universitext, 0172-5939. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4056-4. x + 287 + 54 pp. LCCN QA273.A1-274.9; QA274-274.9.

Welfert:1996:ANM

[WA96] Bruno D. Welfert and Ricardo Aguilar. Applied numerical methods and graphical visualization. Computer Applications in Engineering Education, 4(2):127–143, 1996. CODEN CAPEED. ISSN 1061-3773 (print), 1099-0542 (electronic).

Wang:1990:UMS

[Wan90]  Rongdong Wang.  Using

MAPLE for symbolic computations in robotics. Thesis (M.S.), Northern Illinois University, DeKalb, IL, USA, 1990.

Wang:1992:ISS

[Wan92] Paul S. Wang, editor. International System Symposium on Symbolic and Algebraic Computation 92, July 27–29, Berkeley, California. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-489-9. LCCN QA 76.95 I59 1992.

Wang:1994:DII

[Wan94] Dong Ming Wang. Differentiation and integration of indefinite summations with respect to indexed variables — some rules and applications. Journal of Symbolic Computation, 18(3):249–264 (or 249– 263??), September 1994. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wang:1995:EPM

[Wan95a] Dong Ming Wang. Elimination procedures for mechanical theorem proving in geometry. Annals of mathematics and artificial intelligence, 13(1-2):1– 24, 1995. CODEN AMAIEC. ISSN 1012-2443 (print), 1573-7470 (electronic).

Wang:1995:ICS

[Wan95b] Dong Ming Wang. An implementation of the characteristic set method in Maple. In Pfalzgraf and Wang [PW95], pages 187–201. ISBN 3-211-82600-9 (Wien), 0-387-82600-9 (New York). ISSN 0943-853X. LCCN QA76.9.A96 A9 1995.

Wang:2006:PMC

[Wan06] Frank Y. Wang. Physics with Maple: the computer algebra resource for mathematical methods in physics. Wiley-VCH, Weinheim, Germany, 2006. ISBN 3-527-40640-9. xx + 605 pp. LCCN cs QC20.7.E4 W36 2006 UCB.

Wass:2002:SCG

[Was02] John A. Wass. Symbolic computation gets more power: a review of Maple 7.0. Scientific Computing and Instrumentation, 19(5):18, 21–22, 24, April 2002. ISSN 1524-

2560.  URL http://www.
scimag.com/scripts/ShowPR.
asp?PUBCODE=030&ACCT=3000039877&
ISSUE=0204&RELTYPE=PR&PRODCODE=
00000000&PRODLETT=A.

Wass:2004:MSC

[Was04] John A. Wass. Mathematical software: Not a cosmetic upgrade: Maple 9 delivers a versatile set of user-friendly tools. Scientific Computing and Instrumentation, 21(3):28– 29, 32, February 2004. ISSN 1524-2560. URL http://www.

scimag.com/scripts/ShowPR.
asp?PUBCODE=030&ACCT=3000039877&
ISSUE=0402&RELTYPE=FE&PRODCODE=
00000000&PRODLETT=A&.

Wass:2006:GOM

[Was06] John A. Wass. Global optimization with Maple. Scientific Computing and Instrumentation, 23(7):16, 41, June 2006. ISSN 1524-2560.

Watt:1991:IPI

[Wat91] Stephen M. Watt, editor.

ISSAC ’91: proceedings of the 1991 International Symposium on Symbolic and Algebraic Computation, July 15–17, 1991, Bonn, Germany. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-89791-437-

6. LCCN QA 76.95 I59 1991.

Waterloo:1993:MV

[Wat93] Waterloo Maple Software.

Maple V. Waterloo Maple Software, Waterloo, Ontario, Canada, student version, release 2 edition, 1993. ISBN 0-387-14211-8. LCCN QA76.95. Six computer disks.

Waterloo:1995:MVRj

[Wat95a] Waterloo Maple Software.

Maple V Release 2: Student Edition, 386/486 Computers, Release 2. Brooks/Cole, Pacific Grove, CA, USA, 1995. ISBN 0-534-21222-0. ???? pp. LCCN ????

Waterloo:1995:MVRk

[Wat95b] Waterloo Maple Software. Maple V Release 2: Student Edition, 386/486 Computers, Release 2. Brooks/Cole, Pacific Grove, CA, USA, 1995. ISBN 0-534-21222-0. ???? pp. LCCN ???? US$79.95.

Waterloo:1995:MVRl

[Wat95c] Waterloo Maple Software. Maple V Release 2: Student Edition, Macintosh, Release 2. Brooks/Cole, Pacific Grove, CA, USA, 1995. ISBN 0-534-21228-X. ???? pp. LCCN ???? US$79.95.

Waterloo:1995:MVRf

[Wat95d] Waterloo Maple Software. Maple V Release 3 Student Edition Lab Pak: Student Maple Lab Pak, Release 3. Brooks/ Cole, Pacific Grove, CA, USA, 1995. ISBN 0-534-25564-7. ???? pp. LCCN ???? US$670.00.

Waterloo:1995:MVRi

[Wat95e] Waterloo Maple Software. Maple V Release 3 Student Edition Lab Pak: Student Maple Lab Pak, Release 3. Brooks/ Cole, Pacific Grove, CA, USA, 1995. ISBN 0-534-25565-5. ???? pp. LCCN ???? US$670.00.

Waterloo:1995:MVRa

[Wat95f] Waterloo Maple Software. Maple V Release 3 Student Edition Lab Pak Upgrade: Student Maple Lab Pak Upgrade, Release 3. Brooks/Cole, Pacific Grove, CA, USA, 1995. ISBN 0-534-25567-0 (invalid checksum). ???? pp. LCCN ????

Waterloo:1995:MVRg

[Wat95g] Waterloo Maple Software.

Maple V Release 3 Student Edition Lab Pak Upgrade: Student Maple Lab Pak Upgrade, Release 3. Brooks/Cole, Pacific Grove, CA, USA, 1995. ISBN 0-534-25575-2. ???? pp. US$250.00.

Waterloo:1995:MVRd

[Wat95h] Waterloo Maple Software.

Maple V Release 3 Student Edition Software Only: Student Maple Software Only, Release 3. Brooks/Cole, Pacific Grove, CA, USA, 1995. ISBN 0-534-25562-0. ???? pp. LCCN ???? US$77.95.

Waterloo:1995:MVRh

[Wat95i] Waterloo Maple Software.

Maple V Release 3 Student Edition Software Only: Student Maple Software Only, Release 3. Brooks/Cole, Pacific Grove, CA, USA, 1995. ISBN 0-534-25563-9. ???? pp. LCCN ???? US$77.95.

Waterloo:1995:MVRb

[Wat95j] Waterloo Maple Software.

Maple V Release 3: Student Edition: Student Maple, Release 3. Brooks/Cole, Pacific Grove, CA, USA, 1995. ISBN 0-534-25560-4. ???? pp. LCCN ???? US$90.95.

Waterloo:1995:MVRe

[Wat95k] Waterloo Maple Software. Maple V Student Edition Release 3: Student Maple, Release 3. Brooks/Cole, Pacific Grove, CA, USA, 1995. ISBN 0-534-25561-2. LCCN ???? US$90.95.

Waterloo:1995:MVRc

[Wat95l] Waterloo Maple Software. Maple V Student Edition Upgrade Release 3: Student Maple Upgrade, Release 3. Brooks/ Cole, Pacific Grove, CA, USA, 1995. ISBN 0-534-25567-1. ???? pp. LCCN ???? US$25.00.

WMS:1996:MV

[Wat96] Waterloo Maple Software. Maple V, 1996. ISBN 0-387-14229-0 (disks) 0-387-94536-9, 0-387-94575-X (boxed version).

WMI:1997:Ma

[Wat97a] Waterloo Maple, Inc. Math-View, 1997.

WMI:1997:Mb

[Wat97b] Waterloo Maple, Inc. Math-View, 1997.

WMI:2001:MGS

[Wat01a] Waterloo Maple Inc. Maple 7 Getting Started Guide. Waterloo Maple Inc., Waterloo, ON, Canada N2L 6C2, 2001. iv + 34 pp.

WMI:2001:MLG

[Wat01b] Waterloo Maple Inc. Maple 7 Learning Guide. Waterloo Maple Inc., Waterloo, ON, Canada N2L 6C2, 2001. viii + 292 pp. LCCN ???? Based in part on the work of B. W. Char.

WMI:2002:MGS

[Wat02a] Waterloo Maple Inc. Maple 8 Getting Started Guide. Waterloo Maple Inc., Waterloo, ON, Canada N2L 6C2, 2002. ISBN 1-894511-25-5. xiv + 33 pp.

WMI:2002:MLG

[Wat02b] Waterloo Maple Inc. Maple 8 Learning Guide. Waterloo Maple Inc., Waterloo, ON, Canada N2L 6C2, 2002. ISBN 1-894511-26-3. viii + 308 pp. LCCN ???? Based in part on the work of B. W. Char.

WMI:2002:MRG

[Wat02c] Waterloo Maple Inc. Maple 8 Resource Guide. Waterloo Maple Inc., Waterloo, ON, Canada N2L 6C2, 2002. 6 pp.

Wallack:1998:MMM

[WEM98] Aaron Wallack, Ioannis Z.

Emiris, and Dinesh Manocha. MARS: a MAPLE/MATLAB/ C resultant-based solver. In Gloor [Glo98], pages 244–251. ISBN 1-58113-002-3. LCCN

???? URL http://www.acm.
org/pubs/citations/proceedings/
issac/281508/p244-wallack/

.

Werner:1993:MLMa

[Wer93a] W. Werner. Mathematik lernen mit Maple V. ELBI-Verlag GmbH, ????, 1993. ISBN 3-929694-03-4. ???? pp.

Werner:1993:MLMb

[Wer93b] W. Werner. Mathematikaufgaben lösen mit Maple V. ELBI-Verlag GmbH, ????, 1993. ISBN 3-929694-04-2. ???? pp. Includes diskette.

Werner:1996:MLM

[Wer96] Wilhelm Werner. Mathematik lernen mit Maple: ein Lehrund Arbeitsbuch für das Grundstudium: Grundlagen, Differentialrechnung, numerische Methoden. dpunkt-Verlag, Heidelberg, Germany, 2. überarb. und erw. Aufl. mit CD-ROM edition, 1996. ISBN 3-920993-43-8. 688 pp. LCCN ????

Wessels:1995:SRM

[Wes95] P. Wessels. Software review: Maple V. SciTech Journal, 5(10):15, 17–19, November– December 1995. CODEN SITJE8. ISSN 1072-0995.

Westermann:1997:MIM

[Wes97] T. (Thomas) Westermann. Mathematik für Ingenieure mit Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-61248-3. v. pp. LCCN ????

Wester:1999:CAS

[Wes99] Michael J. Wester, editor. Computer algebra systems: a practical guide. Wiley, New York, NY, USA, 1999. ISBN 0-471-98353-5. xvi + 436 pp. LCCN QA155.7.E4 W48 1999. URL

http://www.loc.gov/catdir/
description/wiley033/99021149.
html;   http://www.loc.gov/
catdir/toc/onix03/99021149.
html.

Wiwianka:1990:ADC

[WF90] W. Wiwianka and B. Fuchssteiner. Algorithms to detect complete integrability in 1 + 1 dimension. In Nonlinear evolution equations and dynamical systems (Kolymbari, 1989), Res. Rep. Phys., pages 131–135. Springer-Verlag, Berlin, Germany, 1990.

Watt:1983:ATM

[WG83] Stephen Michael Watt and Keith O. Geddes. Arrays and tables in Maple. supplement to the Maple user’s manual. Technical Report CS-83-10, Computer Science Department, University of Waterloo, Waterloo, Ontario, Canada, 1983. 27 pp.

Wegrzyn:1990:DS

[WGV90] S. Wȩgrzyn, J.-C. Gille, and P. Vidal. Developmental systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-97340-0. viii + 122 pp. LCCN TA168.W42 1990.

Wang:1992:MAN

[WH92] S. Wang and K. Huseyin.

“Maple” analysis of nonlinear oscillations. Mathematical and computer modelling, 16(11):49– 57, November 1992. CODEN MCMOEG. ISSN 0895-7177 (print), 1872-9479 (electronic).

White:2010:CMC

[Whi10] Ralph E. White. Computational methods in chemical engineering with Maple. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-04310-0. ???? pp. LCCN ????

Wick:1995:IBS

[Wic95] David Wick. The Infamous Boundary: Seven Decades of Controversy in Quantum Physics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1995. ISBN 0-8176-3785-0, 3-7643-3785-0. xii + 244 pp. LCCN QC173.98.W53 1995. US$49.50. URL http://www.

                                        birkhauser.com/Books/wick/
                                        summary.htm.

Wils:1989:HCR

[Wil89] Patrick Wils. Homogeneous and conformally Ricci flat pure radiation fields. Class. Quant. Grav., 6(9):1243–1251, 1989. CODEN CQGRDG. ISSN 0264-9381 (print), 1361-6382 (electronic).

Wilsdon:1998:UMI

[Wil98] Colin Wilsdon. Using Maple to investigate Bernoulli numbers. Mathematics Today, 34(3):74– 76, 1998. ISSN 1361-2042.

Wilf:2006:G

[Wil06] Herbert S. Wilf. Generatingfunctionology. A. K. Peters, Ltd., Wellesley, MA, USA, third edition, 2006. ISBN 1-56881-279-5 (hardcover). x + 245 pp. LCCN QA353.G44 W55 2006.

Wills:1995:LFI

[WL95] Randall Wills and Steve Ligh.

On linear functions II. Mathematics and computer education, 29(1):32–??, Winter 1995. CO-DEN MCEDDA. ISSN 0730-8639.

Watanabe:1990:IPI

[WN90] Shunro Watanabe and Morio Nagata, editors. ISSAC ’90: proceedings of the International Symposium on Symbolic and Algebraic Computation: August 20–24, 1990, Tokyo, Japan. ACM Press and Addison-Wes-ley, New York, NY 10036, USA and Reading, MA, USA, 1990. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Wouk:1986:NCE

[Wou86] Arthur Wouk, editor. New computing environments: parallel, vector and systolic. Papers from a workshop held at Stanford University, Stanford, CA, November 7–9, 1984. Society for Industrial and Applied Mathematics (SIAM), Philadelphia, PA, USA, 1986. ISBN 0-89871-201-7. LCCN QA76.9.A73N49

Wright:1996:BRM

[Wri96] Francis J. Wright. Book reviews: Maple V by example, by Martha L. Abell and James P. Braselton. Mathematics of Computation, 65(214): 881–884, April 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                        URL http://www.ams.org/
                                        jourcgi/jour-pbprocess?fn=
                                        110&arg1=S0025-5718-96-00700-
                                        4&u=/mcom/1996-65-214/.

Wright:2001:CM

[Wri01]  Francis Wright.  Comput-

ing with Maple. Chapman and Hall/CRC, Boca Raton, FL, USA, 2001. ISBN 1-58488-236-0. xvii + 532 pp. LCCN QA155.7.E4 W75 2002. UK£24.99. URL http://www.

                                        maths.qmw.ac.uk/~fjw/CwM/.

Wu:1992:CCD

[Wu92] Xian Wu. Chern classes and degenerations of hypersurfaces and their lines. Duke Math. J., 67(3):633–652, 1992. CO-DEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic).

Wursthorn:1996:RTM

[Wur96] Martin Wursthorn. Representation theory with MAPLE: an example. An. Ştiinţ. Univ. Ovidius Constanţa Ser. Mat., 4 (1):197–203, 1996. ISSN 1223-723X. Representation theory of groups, algebras, and orders (Constanţa, 1995).

Wegrzyn:1990:SMD

[WVG90] Stefan Wȩgrzyn, Pierre Vidal, and Jean-Charles Gille. On some models for developmental systems. XII. more on internal context sensitivity. International Journal of Systems Science, 21(8):1441–1449, 1990. CODEN IJSYA9. ISSN 0020-7721 (print), 1464-5319 (electronic).

Wang:1991:ACZ

[WW91] Rongdong Wang and Peng-Yung Woo. Automatic computation of Z and inverse Z transformations by MAPLE. IECON Proceedings (Industrial Electronics Conference), 2:1092–1094, 1991. CODEN IEPREA. IEEE catalog number 91CH2976-9.

Wang:1992:UMS

[WWK92] R. Wang, P. Woo, and M. K. Kwong. Using Maple for symbolic computations in robotics. International Journal of Robotics and Automation, 7(2):41–49, 1992. CO-DEN IJAUED. ISSN 0826-8185 (print), 1925-7090 (electronic).

Watkins:1993:CST

[WY93] J. Watkins and S. Yurkovich.

Calculation of the state transition matrix for linear time varying systems. In Lee [Lee93], pages 157–166. ISBN 0-8176-3724-9, 3-7643-3724-9. LCCN QA76.95 .M36 1993. US$34.50. Also known as “MSWS 93”.

Wang:1993:ACI

[WyW93] Rongdong Wang and Peng yung Woo. Automatic computation of z- & inverse z-transformations by Maple. Journal of Symbolic Computation, 15(3):349–363, March 1993. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Xie:2010:DEE

[Xie10] Wei-Chau Xie. Differential equations for engineers. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19424-5. ???? pp. LCCN TA347.D45 X54 2010. URL

http://www.loc.gov/catdir/
enhancements/fy1005/2010001101-
b.html;   http://www.loc.
gov/catdir/enhancements/fy1005/
2010001101-t.html;   http:
//www.loc.gov/catdir/enhancements/
fy1006/2010001101-d.html.

Xu:1995:SPB

[Xu95] Kedai Xu. Stochastic pitchfork bifurcation: numerical simulations and symbolic calculations using MAPLE. Mathematics and Computers in Simulation, 38(1-3):199–209, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). Probabilités numériques (Paris, 1992).

Xiazhi:2016:MPF

[XYXZ16] Hao Xiazhi, Liu Yinping, Tang Xiaoyan, and Li Zhibin. A Maple package for finding interaction solutions of nonlinear evolution equations. Computers and Mathematics with Applications, 72(9):2450– 2461, November 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122116305089.

Yanami:2007:MPS

[YA07a] Hitoshi Yanami and Hirokazu Anai. The Maple package SyN-RAC. Future Generation Computer Systems, 23(5):721–726, June 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yanami:2007:SMT

[YA07b] Hitoshi Yanami and Hirokazu Anai. SyNRAC: a Maple toolbox for solving real algebraic constraints. ACM Communications in Computer Algebra, 41(3):112–113, September 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Yang:1993:MIT

[Yan93] W.-C. Yang. Maple and the integral test for series. In Lum [Lum93], pages 465–467. ISBN 0-201-50013-2. LCCN QA11.A1I454 1991.

Yan:1994:SVM

[Yan94] Song Y. Yan. A simple verification method for the Goldbach conjecture. International journal of mathematical education in science and technology, 25(5):681–688, 1994. CO-DEN IJMEBM. ISSN 0020-739X (print), 1464-5211 (electronic).

Yan:1995:PTL

[Yan95] S. Y. Yan. Primality testing of large numbers in Maple. Computers and Mathematics with Applications, 29 (12):1–8, June 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812219500052Z.

Yau:1992:EMM

[Yau92] Shing-Tung Yau, editor. Essays on mirror manifolds. International Press, Hong Kong, 1992. ISBN 962-7670-01-4. vi + 502 pp. LCCN QC20.7.M24 E83 1992.

Yildirim:2000:AKU

[YB00a] Necmettin Yildirim and Mustafa Bayram. An analysis of the kinetics of unstable enzymatic systems using MAPLE. Applied Mathematics and Computation, 112(1):41–48, June 1, 2000. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/29/17/
                                        20/86/21/24/abstract.html;
                                         http://www.elsevier.nl/gej-
                                        ng/29/17/20/86/21/24/article.
                                        pdf.

Yildirim:2000:DCR

[YB00b] Necmettin Yildirim and Mustafa Bayram. Derivation of conservation relationships for metabolic networks using MAPLE. Applied Mathematics and Computation, 112(2–3):255–263, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/17/20/86/23/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/17/20/86/23/28/article.
pdf.

Ye:1992:SLI

[YC92] Honglin Ye and R. M. Corless.

Solving linear integral equations in Maple. In Wang [Wan92], pages 95–102. ISBN 0-89791-489-9. LCCN QA 76.95 I59 1992.

Yip:2007:MCM

[YC07] Man Ho Yip and M. J. Carvalho. A Monte–Carlo maplet for the study of the optical properties of biological tissues. Computer Physics Communications, 177(12):965–975, December 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465507003712.

Yang:2004:EME

[YF04] Michael Yang and Richard Fateman. Extracting mathematical expressions from Post-Script documents. In Gutierrez [Gut04], pages 305–311. ISBN 1-58113-827-X. LCCN

????  URL http://www.

                                        ocf.berkeley.edu/~mlyang/
                                        papers/MichaelYangPsmath.
                                        pdf.

Yan:1999:FGO

[YJ99] S. Y. Yan and G. James. Fast group operations on elliptic curves in Maple. Computers and Mathematics with Applications, 37(8):129–138, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/S0898122199001066.

Yao:2006:CMP

[YL06] Ruo-Xia Yao and Zhi-Bin Li. CONSLAW: a Maple package to construct the conservation laws for nonlinear evolution equations. Applied Mathematics and Computation, 173(1):616– 635, February 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yeargers:1996:IMB

[YSH96] Edward K. Yeargers, Ronald W. Shonkwiler, and James V. Herod. An introduction to the mathematics of biology: with computer algebra models. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1996. ISBN 0-8176-3809-1, 3-7643-3809-1. x + 417 pp. LCCN QH323.5.Y435 1996. US$64.50.

Yuan:1997:SMP

[Yua97] Dongsheng Yuan. STREED:

a Maple package of modeling of uncertain systems. Thesis (M.S.), Florida State University, Tallahassee, FL, USA, 1997. x + 103 pp.

Yu:2007:CMF

[YZ07] Fajun Yu and Hongqing Zhang.

Constructing a multi-function equation hierarchy and its integrable coupling system with MAPLE. Applied Mathematics and Computation, 187(2): 1208–1218, April 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zachary:1996:ISP

[Zac96] Joseph Zachary. Introduction to scientific programming: computational problem solving using Maple and C. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94630-6, 1-4612-7518-0 (print), 1-4612-2366-0 (electronic). xxiv + 380 pp. LCCN QA76.6 .Z32 1996.

Zeilberger:1991:MPP

[Zei91] Doron Zeilberger. A Maple program for proving hypergeometric identities. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(3):4–13, July 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zeilberger:1992:PJW

[Zei92] Doron Zeilberger. A proof of Julian West’s conjecture that the number of two-stack-sortable permutations of length n is 2(3n)!/((n + 1)!(2n + 1)!). Discrete Math., 102(1):85–93, 1992. CODEN DSMHA4. ISSN 0012-365X (print), 1872-681X (electronic).

Zaitsev:1991:CDG

[ZF91] V. F. Zaĭtsev and A. V. Flegontov. Diskretno-gruppovye metody integrirovaniya obyknovennykh differentsial’nykh uravneniĭ. (Russian) [Discretegroup methods for the integration of ordinary differential equations]. Akad. Nauk SSSR, Leningrad. Inst. Inform. Avtomat., Leningrad, Russia, 1991. ISBN 5-201-10181-X. 240 pp. LCCN QA155.7.E4 Z3 1991.

Zietz:2007:EID

[Zie07] Joachim Zietz. Economic instruction — dynamic programming: An introduction by example. The Journal of Economic Education, 38(2):165–??, 2007. CODEN ???? ISSN 0022-0485.

Zima:1995:SOT

[Zim95] Eugene V. Zima. Simplification and optimization transformations of chains of recurrences. In Levelt [Lev95], pages 42–50. ISBN 0-89791-699-9. LCCN QA 76.95 I59

1995.  URL http://www.
acm.org:80/pubs/citations/
proceedings/issac/220346/
p42-zima/.

Zirwas:1994:MVA

[Zir94] G. Zirwas. Maple V analytically calculating plane strain F Ematrices with rotational degrees of freedom. In Lopez [Lop94a], pages 141–148. ISBN 0-8176-3791-5, 3-7643-3791-5. LCCN QA76.95.M36 1994.

Zizi:1993:MIE

[Ziz93] Jacqueline Zizi. Mathématiques, informatique et enseignement. Ed. Archimède, Ed. du Choix, Argenteuil, France, 1993. ISBN 2-909028-13-5. xxv + 303 pp. LCCN ????

Zachary:1995:ELC

[ZJEP95] J. L. Zachary, C. R. Johnson, E. N. Eide, and K. W. Parker. An entry-level course in computational engineering and science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27 (1):209–213, March 1995. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zotos:2007:PCM

[Zot07] Kostas Zotos. Performance comparison of Maple and Mathematica. Applied Mathematics and Computation, 188(2): 1426–1429, May 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zimmermann:1995:CFP

[ZSG5 ] P. Zimmermann, B. Salvy, and C. Gomez. Calcul formel pour scientifiques (Exemples en Maple). Masson Editeur, Paris, France, 1995 (??). ???? pp.